A Complete Bibliography of Publications in Fundamenta Informaticae: (2010–2019)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

02 May 2023 Version 1.10

Title word cross-reference

#2SAT [320].

0 [1018]. 1 [1018, 232, 1113]. 122^{+}3 [1441]. 1^{j+1}0^{j} [412]. 2 [529, 46, 715]. 3 [1297, 1298, 602, 861, 728]. 4 [574]. 5^{→}3′ [250, 125]. 6 [614]. ????_{∞} [1131]. ^{0} [910]. ^{2} [1144, 220]. ^{d} [1230]. ^{±} [1361]. _{Lingua} [980]. _{n} [723]. α^{0} [705]. β [117]. C [561, 456]. C [1132]. M [431]. S⟨⟩⨿ [1362]. X [499]. D_{0} [350]. ⋄ [725]. ℓ [1419]. ϵL [1395, 1325, 1326]. ϵL^{+} [1325, 1326]. ϵ · log n [124]. H [110]. hp [344, 345]. hv [859, 858, 1020]. J [1111]. K [179, 1276, 1260, 1120, 1522, 292, 406, 1248]. L [25, 1127, 380, 272]. l^{p} [911]. λ [1418, 502, 117]. λP [720]. ⟨ωωλ; ×⟩ [1463]. m [772]. A_{n} [584]. D_{4} [569]. K [216]. Z_{n} [1547]. μ [576, 435, 1214]. n [1547, 1092, 1237, 1325, 1326]. O(m log n) [158]. ω [276, 915]. P [244, 831, 956, 529, 833, 889, 897, 898, 899, 765, 837, 901, 980, 902, 766, 292, 222, 903, 262, 840, 1254, 263, 944, 1077, 1312, 1257, 1198, 1081, 841, 1210, 842, 1128, 351, 716]. p_{4} ± p_{3} + p_{2} ± p + 1 [351]. π [720, 506]. P ROEL [1399]. Q [977, 351, 423, 1221]. r [415, 416]. S [984]. σ [856, 271]. SL(2, Z) [1420]. T [971, 215, 1042]. θ [271]. V [976]. v^{2} = u^{p} − au − b [716]. Z [529, 563].

-Abelian [1276]. -additivity [1111]. -Adhesive [431]. -Approach [25]. -approximation [984]. -ary [1325, 1326]. -Automata [276]. -block [292]. -bonacci [772]. -Bonds [380]. -Calculus [576, 435, 502, 720, 1214, 1221, 506]. -Chu [380]. -circuit [715]. -Codes [725]. -colored [728]. -Comma [179]. -complexity [971]. -connected [1120]. -Convex [858, 859, 1020]. -DAGs [117]. -dimensional [529]. -Expressions [216]. -Fuzzy [271, 272, 215]. -Gauss [423]. -Grammars [276]. -level [574]. -Means [1522, 1260, 561, 456]. -Metric [1419]. -number [563]. -order [976]. -Orthogonal [423]. -overlap-free [406]. -palindromic [856]. -Petri [915]. -quasi-optimal [1248]. -rationality [977]. -reachability [529, 1127]. -reduction [117]. -Rough [1042]. -Safe [232]. -scrambling [1418]. -Similarity [1237]. -state [614]. -tag [46]. -trees [110].

1 [349].

2 [431]. 2009 [185]. 2017 [1497].

3 [1550].

60th [1250, 19, 105, 386].

Abducible [1293]. Abduction [1188]. Abductive [571, 1397, 1360, 1361, 99]. Abelian [1276, 1451, 1195]. Abnormality [1529]. ABox [88]. Absorbed [417]. Absorption [417]. Absorptive [1157]. Abstract [1074, 1353, 170, 1003, 849, 1291, 484, 1504, 173]. Abstraction [1078]. AC [910]. ACBC [1168]. ACBC-Adequate [1168]. Accelerated [1025]. Accelerations [332]. Acceptance [1266]. Accepted [305, 533]. Accepting [758, 887, 130]. Access [464]. Accommodate [1400]. Account [1543]. Accuracy [748, 136, 799]. Accurate [730]. Action [188, 730, 641]. Actions [230, 641]. Activation [559]. Active [1543, 1363, 837, 1257, 1366]. Activities [1405]. Activity [1380, 1444, 953, 1531, 787]. Actual [1327, 1328]. Acyclic [286, 1564]. Acyclicity [1145]. Ad [160]. Adaptation [98, 7]. Adaptive [446, 1435, 1427, 662, 453, 1377, 13, 300, 344, 345, 1449]. Adding [1067]. Addition [529, 1541]. Additional [1151, 1365]. additivity [1111]. Adequate [1168]. Adhesive [431]. Adhesivity [430]. Adjacency [1314]. Adjusting [662]. Admissible [624]. Advanced [1464, 187]. Advances [682, 1091]. Advertising [5]. Advice [1266]. AES [348]. Affine [1119]. After [1554]. against [349, 1236]. Agent [1359, 85, 148, 1240, 1007, 1347, 1401, 301, 69, 65, 370]. Agents [1385, 1346, 657]. Aggregate [75, 164]. Aggregated [1133]. Aggregates [1183, 197]. Aggregation [1372, 1466, 780, 517, 1159, 536, 1332, 1494, 883]. Akt [1375]. Akt-Wnt-mTOR-p70S6K [1375]. Algebra [686, 1252, 500, 584, 863, 86, 180, 1458, 1520, 864, 606, 89, 679, 272]. Algebraic [44, 172, 848, 1213, 33, 229, 362, 340, 1389, 24, 853, 312, 1512, 1025]. Algebraical [494]. Algebraically [1231]. Algebras [1517, 629, 905, 1012, 566, 176, 1076, 1080, 1238, 1516, 208, 215, 1446, 750]. Algorithm [965, 243, 1130, 1496, 665, 1304, 595, 1298, 201, 1408, 287, 45, 107, 1498, 439, 351, 54, 618, 896, 454, 1261, 907, 876, 1232, 1265, 1476, 737, 900, 561, 1267, 26, 139, 544, 1108, 1022, 826, 728, 1478, 1356, 1157, 1133, 511, 586, 1512, 536, 1449, 587, 1525, 1020, 442, 804, 346]. Algorithmic [1190, 966, 986, 863, 906, 949, 1524, 903, 1095]. Algorithmic-type [863]. Algorithms [2, 1116, 194, 487, 482, 1430, 288, 531, 185, 915, 591, 725, 777, 982, 978, 979, 458, 202, 87, 517, 13, 1237, 798, 878, 313, 1461, 239, 551, 569, 1093, 1094, 742, 1387, 1239]. Algorithms* [987]. Aligning [854]. All-or-nothing [1345]. Almost [946, 21, 267]. Almost-Sure [267]. Alphabet [952]. Alphabets [23]. AlPiNA [329]. Alternating [296, 152]. Alternating-Time [152]. Alternative [1548, 330]. Alvis [705]. Ambiguity [947, 1420, 522, 1525]. Among [646]. Amplification [1433]. Analogues [1112]. Analyses [1428]. Analysing [1251, 1376]. Analysis [1088, 2, 1353, 1, 1496, 1002, 1309, 961, 790, 60, 297, 1380, 786, 1537, 1426, 198, 465, 914, 188, 491, 516, 180, 1367, 1228, 181, 979, 689, 1044, 1132, 866, 738, 815, 115, 312, 165, 1425, 1085, 568, 1255, 479, 793, 648, 1135, 183, 803, 274]. Analysis* [1172]. Analytic [684, 1473]. Analytical [1296, 1226]. Analyzed [1381]. Analyzing [673]. Anatomy [1331]. Angular [865]. Animation [116]. Annealing [1156]. Annotated [102]. Announcement [436]. Anonymous [475, 927]. Answer [1143, 1183, 1145, 318, 1004, 486, 321, 1292, 1188]. Ant [452, 674]. Ant-Based [674]. Antichains [1538]. Antimatter [835]. Antimicrobial [787]. Antiport [1434, 980]. Antirandom [459]. Application [1495, 1298, 1435, 1380, 431, 1097, 979, 1389, 401, 309, 787, 1014, 624, 675, 598, 523, 313, 8, 27, 167, 1332, 1169]. Applications [1417, 1353, 574, 576, 53, 518, 1021, 561, 1443, 1108, 1109, 177, 1235, 42, 1085, 1527, 1035, 1314, 335, 685, 1096, 989]. Applied [1497, 1138, 1247]. Applying [1488]. Approach [132, 1149, 462, 572, 855, 1231, 317, 392, 697, 1344, 1371, 602, 860, 1278, 453, 229, 362, 585, 25, 823, 270, 371, 1245, 949, 1400, 650, 28, 29, 563, 643, 425, 738, 562, 853, 921, 1428, 1440, 818, 448, 984, 1100, 68, 1502, 1357, 1048, 679, 1490, 1197, 240, 1169]. Approaches [1456, 1511, 1471, 1034, 683, 165, 376]. Appropriate [1377]. Approximate [1130, 487, 1226, 595, 1537, 1518, 182, 280, 1061]. Approximating [71, 251, 853]. Approximation [1116, 635, 589, 1355, 269, 1034, 864, 1012, 963, 1039, 564, 827, 1461, 56, 234, 374, 1173, 272, 1036, 1096, 1480, 271, 1042, 639, 984]. Approximation-oriented [1034]. Approximations [638, 790, 452, 1168, 1038, 1039, 311, 1041, 655, 1373, 799, 1503, 1049, 792]. Apriori [734]. Apriori-Based [734]. Araucaria [550]. Arbiters [203]. Arbitrary [1070, 934]. Arc [1131, 995, 1070]. Arc-search [1131]. Architectural [1427]. Architecture [464]. Architectures [28, 29]. Archive [965]. Area [4]. Argumentation [1508, 1201, 1291, 319, 177, 1293, 1292]. Argumentation* [1003]. Arguments [572]. Arithmetic [535, 841]. Arithmetical [1055, 689]. Armin [321]. Array [1456, 1006, 1203, 1289, 559]. Arrow [37]. Artifacts [631]. Artificial [1435, 900]. ary [1325, 1326]. ASP [1395, 575, 197]. Aspects [209, 1165]. Assembling [248]. Assembly [1054, 1286, 951]. Assessing [937]. Assessment [697]. Assigns [587]. Associated [423, 1221]. Association [2, 1149, 1491, 1168, 921, 1165, 1169]. Assumption [1509, 560]. Assumptions [1066, 88]. Astrocytes [898, 899]. Asymmetric [729]. Asymptotic [136, 115]. Asymptotically [134]. Asynchronous [760, 667, 483, 901, 262, 204, 894, 610]. Atkin [586]. Attack [1253, 353]. Attacker [492]. Attacking [354]. Attacks [162, 237]. Attention [868]. Attitudes [1352]. Attractor [364]. Attribute [1371, 1372, 384, 270, 623, 383, 622, 1462, 374, 788, 1332, 1494]. Attribute-based [1462]. Attributes [523]. Audience [5]. Audio [739]. Augmentation [323, 1391]. Autarkies [217]. Authentication [748, 1549]. Author [293, 315, 336, 358, 385, 407, 428, 450, 474, 497, 509, 526, 552, 582, 603, 625, 663, 692, 717, 740, 768, 800, 828, 851, 885, 912, 939, 970, 990, 1010, 1031, 1050, 1073, 1129, 1141, 1163, 1181, 1199, 1218, 1249, 1269, 1302, 1329, 1330]. Automata [1334, 735, 609, 246, 226, 488, 296, 528, 1065, 122, 303, 847, 123, 276, 848, 304, 774, 305, 365, 611, 364, 1028, 762, 1284, 888, 250, 743, 889, 775, 1068, 386, 1176, 455, 398, 1177, 307, 23, 1178, 259, 1179, 1282, 1283, 125, 126, 941, 1092, 1421, 127, 137, 138, 1306, 260, 732, 825, 128, 943, 469, 238, 613, 893, 1180, 1422, 533, 129, 894, 610, 884, 1248, 890]. Automata-based [743]. Automated [961, 635, 669, 578]. Automatic [1044, 560, 484, 321]. Automation [94]. Automaton [1274, 614, 399, 689, 612, 1479, 190, 471]. Automorphism [607]. Automorphissm [128]. Autonomous [85, 369]. Autonomy [35]. Autonomy-Oriented [35]. Auxiliary [1065]. Avalanches [366]. Average [695, 1454, 1449]. Avoidance [339]. Avoiding [412, 1303, 418, 1441, 411]. Aware [1483, 682, 1427, 1511]. Axes [1230]. Axes-parallel [1230]. Axiom [1036]. Axiomatic [1080, 441, 240, 480]. Axiomatisation [171, 1340]. Axiomatizability [1368]. Axiomatization [1211, 796]. Axioms [578]. Axis [40]. Axis-Parallel [40].

B [519]. Back [1201, 288, 197]. Back-edges [288]. Backprojection [853]. Backward [804]. Bacterial [1254]. Balances [772]. Balancing [1133]. Ball [372]. Ballot [427]. Banach [1225]. Banacha [649]. Bang [618]. Barabasz [143]. Base [977]. Based [965, 636, 372, 241, 1026, 1052, 1485, 961, 782, 1395, 855, 150, 1435, 981, 284, 1432, 1336, 1464, 273, 1371, 602, 936, 1487, 1471, 1207, 1433, 1097, 896, 698, 1530, 73, 1458, 1170, 647, 421, 931, 4, 900, 651, 87, 1245, 980, 1044, 1172, 925, 650, 1338, 484, 222, 865, 1436, 674, 1246, 1478, 879, 752, 477, 1041, 1076, 146, 355, 559, 734, 313, 1425, 1535, 536, 537, 473, 793, 166, 819, 130, 15, 640, 1189, 1313, 16, 690, 789, 788, 1301, 1357, 1503, 911, 883, 1410, 1438, 184, 804, 617, 1243, 271, 630, 80, 805, 281, 341, 1529, 1042, 1545, 1366, 571, 2]. based [1359, 1382, 1078, 1351, 1510, 1397, 862, 1006, 1398, 743, 577, 362, 1253, 930, 908, 644, 1012, 973, 1038, 578, 232, 1400, 606, 1166, 470, 738, 1501, 205, 1512, 622, 1015, 103, 333, 580, 1255, 1101, 1462, 1449, 555, 642, 691, 868, 1314, 1062, 1248, 357, 1162]. Bases [34, 135, 908, 476, 910, 142, 1166]. Basic [1079]. Basis [272]. Baxter [419]. Bayes [454]. Bayesian [1298, 107, 589, 198, 637, 162, 690, 1301]. Bazilevic [1221]. BDD [136, 232]. BDD-based [232]. BDI [657]. Be [334, 1274]. Bee [900]. Beeps [1259]. Behavior [1382, 1058]. Behavioral [1053, 172]. Behaviors [809, 12, 565]. Behavioural [335]. Behaviours [481, 122, 90]. Belief [32, 908]. Beliefs [84, 52]. Belitskii [439]. Benefits [1348]. Bent [350, 289, 1132]. Better [531, 728]. Between [694, 1304, 196, 808, 1305, 522, 963, 1241, 238, 695, 231, 378, 781]. Bi [880, 867, 1221]. bi-Bazilevic [1221]. Bi-connected [867]. Bi-objective [880]. Biautomata [891]. Biclustering [1404, 1514]. Bifurcation [862]. Big [618, 112]. Big-step [112]. Bigraphs [846]. Bijective [419]. Bimanual [300]. Binary [1295, 71, 1019, 711, 412, 854, 388, 46, 1018, 352, 406, 1255, 934, 272]. Bio [1377, 264]. Bio-semantics [1377]. Bio-Turing [264]. Biochemical [698]. Biography [959]. Bioinformatics [682]. Biological [1251, 83, 559]. Biology [447, 116]. Biomodel [392]. Biomolecular [689]. Bipartite [771, 1314, 584, 907, 982, 978, 979, 458, 1267, 1108, 1109, 568, 1093, 1094, 1239]. Bipolar [362, 17]. Birthday [1250, 105, 386, 19]. Birthdays [414]. Bis [787]. Bis-Quaternary [787]. Bisimilarity [158]. Bisimulation [872, 1194, 1338, 819]. Bisimulation-Based [1338, 819]. BIST [459]. Bit [966, 45]. Bit-Parallel [45]. BITES [233]. Bits [124]. Bivariate [323, 427]. Blackboard [1244]. Blackhole [304]. Blending [700]. Block [1097, 292]. Blood [1151]. Bloom [236]. Blum [1247]. BMC [470]. BNF [817, 1527]. Bodies [857, 1020]. bonacci [772]. Bonds [380]. Bongartz [544]. Boolean [711, 392, 520, 289, 864, 1404, 1514, 42, 357, 989]. Border [936]. Borders [704]. Borel [20]. Boruta [77]. Bottom [238, 117]. Bottom-Up [238, 117]. Bound [736, 289, 74, 1057, 1460, 733]. Boundaries [704, 816]. Boundary [1224, 1130, 198]. Bounded [1335, 528, 993, 1111, 531, 306, 61, 47, 232, 879, 999, 884, 1062]. Boundedness [490]. Bounds [1284, 1117, 1236, 136, 1448]. Box [1520]. Boys [414]. Boys-and-girls [414]. BRACID [1167]. Brain [1442, 1531, 1044, 1444]. Branches [20, 36]. Branching [872, 1216, 512, 1069]. Branching-Time [1069]. Breaking [143]. Bridging [19, 633]. Brief [945]. Broadcast [475, 1067]. Brouwer [362]. Brownian [1479]. BTC [219]. BTC-compressed [219]. Buchanan [1525]. Büchi [743]. Buffer [1247]. Bug [994]. Building [540]. Bulk [1029]. Bursting [967]. Bus [442]. Business [1397, 1483, 99]. Butterfly [1449].

CA [483]. Caching [1155]. Cake [1345]. Calculi [1204]. Calculus [588, 1227, 832, 150, 576, 435, 502, 93, 1504, 427, 720, 1214, 1221, 506]. Camassa [1223]. Can [764, 334, 1274]. Cancellation [813]. Cancerous [645]. Candidate [1498]. Canonical [1140, 1172]. Canonicity [711]. Capabilities [76]. Capacitated [453]. Capacity [18]. Capacity-Raising [18]. Cardinal [781]. Cardinality [49, 1278, 1391]. Cardiovascular [1381]. Care [923]. Carlo [644, 579]. Caron [1525]. Carpets [839]. Cartan [1190, 1356]. Cartesian [1013]. Case [381, 1442, 277, 1362, 1153, 35, 127, 1505, 684, 479, 1189]. Case-Based [1189]. Cases [803]. Catalyst [1081]. Catalysts [956]. Catalytic [889]. Catching [994]. Categorial [173]. Categorical [90]. Categories [850]. Categorization [1026]. Category [380, 371]. Catenation [672, 702]. Catenations [1384]. Cattaneo [361]. Causal [992, 335]. Causality [844, 515, 1410]. Cause [1146]. Cause-effect [1146]. Caused [1151]. Causes [1086]. CCS [220]. CD [196]. Cell [1297, 1434, 944]. Cell-like [1434]. Cells [1077]. Cellular [609, 246, 488, 365, 611, 364, 614, 399, 612, 941, 1092, 128, 129, 610, 1479]. Center [1544, 1468]. Centrality [1123, 733]. Centralized [445]. Cerebral [1530]. Certificate [1015]. Certificate-based [1015]. Certificateless [1402, 1390, 314, 714]. Certificates [773]. Certifying [1559]. Chain [195, 400]. Chained [9]. Chameleon [607]. Changes [1151, 834, 473]. Changing [57]. Channel [1259, 1521]. Channels [808]. Chaotic [1437, 80]. Character [822]. Characterisation [494]. Characterising [972, 501]. Characteristic [716]. Characterization [848, 962, 936, 397, 1349, 1288, 943, 715]. Characterizations [1190, 389, 891, 330, 265]. Characterize [1036]. Characterizing [1105, 1238, 547]. Characters [413]. Charts [1258]. Chase [1331]. Chebyshev [395, 1223]. Check [340]. Check-only [340]. Checker [1240]. Checking [296, 1055, 743, 329, 87, 65, 1069, 232, 88, 879, 1341, 1489, 1462, 1410, 525, 884, 1062, 746, 1157]. Chemical [832]. Child [830]. CHIRA [536]. Chlorides [787]. Choice [977, 37, 17]. Chomsky [943]. Choosing [1377]. Chosen [1015]. Chosen-key [1015]. CHR [571]. CHR-based [571]. Chu [380]. Church [111]. CICT [1017]. Ciliate [245]. Ciliates [951]. Ciobanu [1250]. Cipher [349, 353]. Ciphers [1387]. Circuit [133, 437, 1432, 1436, 715]. Circuits [134, 1236, 841]. Circulant [1097]. Circular [844, 1467]. Circulation [1433]. Circumscription [1508]. CK [723]. Class [848, 907, 1458, 1132, 677]. Classes [773, 476, 910, 1539, 519]. Classical [1555, 356]. Classification [133, 922, 621, 863, 73, 399, 978, 979, 371, 78, 128, 1093, 1094, 642, 70, 1197, 346]. Classifier [1052, 1485, 78]. Classifiers [636, 1125, 658, 1053, 162, 937]. Classifying [1532]. Clausal [217, 218]. Clearing [123]. Clickstream [5]. Client [228]. Client-server [228]. Clinical [1442, 659]. Clock [700, 204].

Clocks [735, 1065]. Closed [821, 560]. Closeness [1411, 567, 1515]. Closest [728]. Closure [154, 518, 1399, 1362, 231, 1037]. Closures [227, 463, 666, 1175]. Cloud [682, 967, 651, 1405, 1462, 804]. CLP [149, 148]. Clues [1118]. Cluster [830, 97, 1460, 274]. Clustering [965, 367, 1260, 34, 618, 375, 1531, 727, 3, 561, 1467, 13, 925, 674, 545, 617, 369]. Clusters [3]. Co [720, 1017, 1541, 1217, 1473, 1266]. Co- [1217]. Co-domain [1017]. Co-Emptiness [1541]. Co-Finiteness [1541]. Co-intuitionistic [720]. co-RE [1266]. co-Semi-analytic [1473]. Coalgebraic [850]. Coalgebraically [1216]. Coalition [709]. Coalitions [572]. Coarse [446]. Coarsest [1208]. CoCaml [1215]. Code [14]. Codes [71, 179, 725, 824]. Coding [1027, 561, 1124, 1096]. Coefficient [1221, 1103]. Cognitive [275]. Coherent [1553]. Coinductive [1001, 1215]. Collaboration [1405, 1425]. Collaborative [698]. Collective [1352, 354]. Colonies [833]. Colony [452, 900]. Color [1467, 40, 1357, 1102]. Colored [1379, 1081, 728]. Coloring [342]. Colorings [291]. Colour [456]. Coloured [747, 738]. Column [1295, 379, 1018]. Combinatorial [185, 100, 798, 1436, 425, 1124]. Combinators [807]. Combinatory [279]. Combined [822, 394, 330]. Combines [1059]. Combining [1304, 151, 1156, 1506]. Comics [1534]. Comixify [1534]. Comma [179]. Commentary [1091]. Commercial [149]. Commitment [1359]. Commitment-based [1359]. Common [712, 45, 1451, 1454, 903, 753]. Communes [114]. Communicating [808, 894]. Communication [9, 132, 196, 696, 1240, 927, 1521, 897, 1122, 916, 1070]. Communication-Free [916, 1070]. Community [449, 35]. Commutative [161, 1157]. Compact [918, 202]. Companion [1262]. Comparative [1496, 665, 1003, 1040, 1306, 546]. Comparing [554]. Comparison [1088, 976, 1491, 349, 638, 1090, 1457]. Comparisons [1091, 796, 930, 1089, 823, 988, 1338, 1083, 1085, 1135]. Compatibility [673]. Competing [708]. Competition [317]. Complement [1432]. Complementary [283]. Complementation [388]. Complemented [362]. Complementer [1438]. Complete [666, 962, 985, 269, 566, 586, 601]. Completeness [170, 361, 889, 1287, 1524, 1070, 610, 499]. Completion [257, 934]. Complex [782, 1432, 1466, 812, 865, 1084, 781, 648]. Complexity [133, 134, 387, 1407, 1192, 367, 1463, 447, 20, 1384, 196, 548, 904, 611, 108, 1079, 366, 914, 927, 252, 394, 915, 253, 1176, 455, 1117, 221, 701, 1127, 1244, 259, 522, 401, 745, 916, 476, 910, 24, 275, 544, 142, 41, 404, 1072, 405, 1247, 938, 118, 971]. Compliance [1397, 1360]. Component [205]. Component-based [205]. Components [43, 489, 1018]. Composable [1334]. Composing [1092]. Composite [433]. Compositeness [1311]. Composition [298, 484, 1339, 321]. Compositionality [490]. Compositions [410, 917]. Compound [857, 1437, 713]. Comprehensive [1220]. compressed [219]. Compression [1046, 284, 1548, 597]. Compromises [671]. Computable [1517, 770]. Computation [1453, 797, 1460, 335, 1479, 975, 1103]. Computational [387, 360, 447, 1513, 366, 889, 253, 1176, 1287, 1434, 1244, 578, 723, 544, 404, 1257, 1465]. Computations [832, 228, 897, 864, 55]. Compute [1456, 1356]. Computed [1023, 1298, 1301]. Computer [19, 584, 863, 558, 1560, 1403, 753, 619, 120]. Computerized [1017]. Computing [1, 966, 1497, 712, 783, 483, 835, 1029, 896, 1454, 699, 38, 1563, 76, 836, 384, 653, 1364, 651, 1044, 1092, 382, 798, 1288, 650, 1338, 292, 839, 66, 1242, 818, 1093, 1094, 56, 1048, 1173, 376, 1049]. Concatenation [604]. Concept [1309, 605, 33, 1412, 556, 1476, 1047, 819, 1516, 628, 271]. Concept-Drift [556]. Concepts [39, 384, 378, 1557, 1506]. Conceptual [558, 566]. Conclusions [1146]. Concrete [1008, 1291, 878, 1156]. Concretisation [484]. Concurrency [720, 154, 92, 760, 969, 1256, 1410]. Concurrent [1250, 282, 845, 486, 972, 204, 619, 1196]. Condensing [167]. Conditional [523, 950]. Conditionals [1499]. Conditions [1224, 1296, 1130, 489, 431, 1442, 1300, 537, 27]. Configuration [149, 573, 1498, 1121]. Confirmation [637]. Conflict [1107, 218]. Conflicts [821, 648]. Confluence [431]. Conformance [673, 746]. Congestion [87]. Congruences [1108, 1109, 1093, 1094]. Conjugation [1254]. Conjunctive [38, 1414]. Connected [1241, 1421, 138, 1120, 867]. Connection [696, 815]. Connection-Oriented [696]. Connections [33, 1323, 1324]. Connectivity [1391]. Connector [28, 29]. Connectors [719]. Consensus [927, 1305]. Consequence [801, 532]. Conservation [596]. Conservative [493]. Considerations [1505]. Consistency [1296, 1152, 1300, 908, 100, 88]. Consistency-based [908]. Consistent [1087, 1440, 869]. Constant [520, 837]. Constant-Space [837]. Constrained [367, 1193, 45, 744]. Constraint [31, 1394, 282, 572, 845, 97, 1203, 190, 217, 218, 75, 164, 1245, 1069, 460, 617]. Constraint-Based [1245]. Constraints [1295, 1394, 1397, 575, 621, 635, 1543, 729]. Constructed [80]. Constructing [1071, 1268]. Construction [1149, 487, 665, 1272, 782, 92, 33, 1007, 395, 906, 1097, 1559, 139, 471, 533, 550, 479, 271]. Constructions [1306, 770, 42, 615]. Constructive [1395, 1007]. Constructivism [1393]. Consumption [1312]. Contact [1517, 1327, 1328]. Containing [417, 427]. Containment [433]. Content [1313, 868]. Content-based [868]. Context [1150, 887, 1043, 1211, 950, 1476, 1178, 1350, 47, 542, 1233, 925, 275, 382, 893]. Context-Free [1211, 893]. Contexts [140, 86, 39, 270, 622]. Contextual [1289, 522, 723, 642]. Continuation [760, 310, 1471, 1256]. Continuation-Passing [1256]. Continuous [914, 1377, 995, 1514, 1083, 1260]. Continuous-Time [995]. Contract [285]. Contraction [1002]. Contradiction [1252]. Contradiction-Tolerant [1252]. Contrast [1114]. Contributions [957]. Control [1289, 889, 87, 300]. Controllability [1483]. Controlled [1549, 1287, 765, 1075]. Controllers [719]. Controlling [577, 1058]. Conventional [836]. Convergence [930, 159]. Convergent [173]. Converting [817]. Convex [1183, 858, 1468, 536, 859, 1020]. Convexity [1019]. Convolution [425, 1234]. Convolutional [1535]. Cooperation [196, 1257]. Cooperative [867, 938, 897]. Coordinate [1020]. Coordinated [162]. Cop [617]. Corank [986, 1239]. Core [149, 1154]. Corpora [726]. Correct [92]. Correct-by-Construction [92]. Correcting [968, 442]. Correction [1357]. corrector [1232, 1265]. Correlated [6, 3]. Correlating [375]. Correlation [697, 602, 861, 1172]. Correlations [609]. Correspondences [380]. Cosets [827]. Cosine [646, 1313]. Cosparse [857]. Cost [682, 987, 1268]. Cost-aware [682]. Cost-Optimal [987]. Countable [615]. Counter [1068, 945, 1283, 1072]. Counterexamples [1412]. Counterpart [435]. Counterparts [32, 1185]. Counters [1177]. Counting [904, 419, 421, 948, 151, 427, 767]. Coupled [1224, 707]. Coupling [556]. Covariance [7]. Cover [286, 1241]. Coverability [1071, 511, 332, 742]. Coverage [694, 4, 1169]. Covering [1389, 963, 270, 374, 546, 691, 1314, 1480, 1042, 1161, 989]. Covering-based [691, 1314]. Coverings [210, 929]. Covers [1455, 624]. Cox [982, 978, 979, 1267]. Cox-regular [982, 978, 979, 1267]. Coxeter [584, 986, 978, 979, 1109, 551, 568, 1093, 1094]. cP [1431]. CPDL [1059]. CPU [880]. CPU/GPU [880]. CREAMS [1527]. Credibility [1426]. Crick [1285, 250, 125]. Crisp [518]. Crisply [1047, 378]. Criteria [243, 1478]. Criterion [482]. Critical [431, 598]. CRN [476]. Cross [420, 815]. Cross-intersecting [420]. Crossing [417]. Crunch [618]. Cryptanalysis [348, 349, 714, 223, 184]. Cryptographic [738, 312]. Cryptographic-Key [312]. Cryptography [1402, 355]. Cryptosystem [824, 1551]. CSP [191, 29]. CSPs [96, 147]. CTRL [1549]. CTRL-PACE [1549]. Cube [782, 353]. Cubes [783]. Cubic [1356, 355]. Cultured [559]. Curve [862, 716]. Curves [1548, 1547, 906, 352, 1268]. Cut [1284, 505, 1060]. Cut-Operations [1284]. Cuts [1052, 1485, 1194]. Cutting [1345]. CWW [563]. Cyber [803]. Cyber-infrastructure [803]. Cyber-ITS [803]. Cycle [169, 1431]. Cycles [567]. Cyclic [1017, 400, 126]. Cycloid [1540]. Cyclotomic [591].

D [1144, 1297, 1298, 602, 861, 1113]. D-FLAT [1144]. D-Volume [602, 861]. Daggers [1217]. DAGs [117]. DAN [1532]. Dart [614, 1022]. DASs [1130]. Data [922, 1, 323, 1106, 1053, 106, 686, 784, 1298, 790, 783, 5, 1067, 1260, 107, 596, 597, 659, 109, 1452, 859, 618, 780, 923, 467, 1531, 4, 6, 561, 1172, 1514, 1445, 1167, 142, 477, 921, 1165, 1505, 1428, 1440, 655, 1462, 15, 8, 781, 789, 1490, 803]. Data-driven [789]. Data-Intensive [1445]. Database [707, 549, 727, 655, 68]. Databases [1354, 73, 926, 734]. Dataflow [685]. Datalog [1361]. Datasets [1154, 644, 3]. DCOPs [1346]. Deadlock [228]. Deadlockability [1409]. Dealing [1379, 877, 1165]. Decentralized [35]. Decidability [609, 946, 722, 157, 1557, 129, 610, 1463]. Deciding [794, 877]. Decipherability [256]. Decision [1417, 1484, 462, 636, 694, 791, 487, 665, 1026, 1052, 1485, 528, 695, 904, 1310, 228, 1371, 1372, 764, 1168, 401, 1038, 623, 974, 371, 139, 319, 1014, 1246, 1478, 1086, 523, 661, 984, 537, 69, 167, 648, 788, 1332, 1494, 373, 370, 620, 369]. Decision-making [523, 167]. Decision-Theoretic [623, 369, 1038, 974, 371, 370, 620]. Decisions [487, 665, 786, 370, 1480]. Declarative [99]. Decoding [824]. Decomposition [605, 783, 589, 643, 867, 739, 1268]. Decompositions [1144, 1510, 174]. Decreasing [1103]. Dedekind [1562]. Dedicated [105]. Deduction [1558]. Deep [1528, 1442]. Default [1499, 319]. Defeasible [1399, 1291]. Defects [381]. Deficiency [217, 155]. Definability [711, 391, 1373]. Definable [20, 1012, 1516]. Defining [671]. Definite [1190, 1356]. Definition [1171, 1470, 480, 869]. Degeneracy [1387]. Degenerations [798]. Degree [518, 1120]. Degrees [182, 280]. Delay [1225]. Delays [1380, 494]. Deletion [244, 1287, 950]. Deliberation [1347]. Delta [674]. Demographics [5]. Dense [1098]. Density [1230, 399, 1523]. Deontic [1062]. Dependence [1281]. Dependencies [1354, 38, 639]. Dependency [1408, 324, 11]. Dependent [92, 140, 767, 1225]. Dependently [93, 94]. Depth [695, 1236]. Derivative [1227, 1223, 1228]. Derivatives [935, 1193, 471]. Derived [154, 1531]. Deriving [1087, 1146, 988, 192, 1105]. DES-like [349]. Describing [832, 1151, 1223]. Description [105, 1152, 1362, 1126, 928, 647, 1008, 142, 877, 1059, 652, 1325, 1326, 705, 819, 691, 1387]. Descriptive [1457, 649]. Descriptor [235]. Design [1143, 1432, 749, 919, 558, 561, 1403, 14, 205, 841]. Designing [1139]. Designs [200, 180]. Desirable [1354]. Detecting [162, 1405, 76]. Detection [273, 1530, 660, 1044, 645, 166, 1525]. Detector [1433]. Determinant [697]. Determination [9]. Determined [750, 25, 934, 1036]. Determining [1548, 551, 1307]. determinism [1430, 113]. Deterministic [10, 1140, 1274, 317, 1284, 1385, 892, 1242, 204, 471, 734, 777]. Development [441]. Device [871]. Devices [902]. DFAs [252]. Diabetes [659]. Diagnosability [83, 1407]. Diagnosis [1529]. Diagnostic [640]. Diagonal [424]. Diagonalization [985]. Diagram [156, 569]. Diagrams [381, 551]. Dialogue [1412, 875, 1489, 1401, 680]. Dialogues [1401]. Diameter [999, 590]. Dicomplemented [1319, 1320]. Dictionary [1559]. Difference [403]. Differences [1241]. Differencing [440, 18]. Different [1022, 238]. Differential [1224, 349, 1435, 1056, 1228, 1225, 826]. Diffraction [1114]. Diffusible [449]. Diffusion [802]. Digital [961, 602, 861, 969, 865, 313, 1369, 1415, 1527]. Digraphs [1264]. Dilation [1464]. Dimension [1013, 798]. Dimensional [323, 1223, 1178, 612, 6, 469, 344, 345, 997, 998, 1461, 524, 10, 1495, 529, 365, 103, 1114]. Dimensions [446, 862, 1419, 646, 3]. Dining [1398]. Diophantine [239, 587]. Direct [815, 1024]. Directed [245, 733, 235]. Directional [1019]. Directions [1299, 1265, 781]. Dirichlet [1496]. Disambiguation [454]. Discourse [310, 215]. Discovering [4]. Discovery [1271, 669, 923, 813, 333, 206, 746]. Discrete [1308, 599, 595, 855, 1299, 53, 629, 1319, 1320, 1029, 905, 596, 597, 1300, 61, 995, 766, 598, 494, 1383, 594, 1313, 884, 267, 1037]. Discrete-Time [61, 1383, 1308]. Discretization [761]. Discriminative [1530, 730]. Disease [1381]. Disentangling [392]. Disjunctions [102]. Dispersion [1103]. Displacement [1432, 1436, 1438]. Display [1133]. Disruption [126]. Distance [71, 367, 795, 930]. Distance-based [930]. Distances [438, 1464, 76, 547]. Distinguishability [438, 1175]. Distortion [1260]. Distributed [583, 735, 1250, 147, 97, 464, 1408, 1398, 61, 683, 685]. Distributed-memory [685]. Distribution [896, 312, 1257]. Distributional [1137]. Distributions [417, 1367, 423]. Distributive [362]. Divides [351]. Diving [192]. Divisibility [1234]. Divisible [906]. Division [948]. Divisors [393]. DLPC [1099]. DNA [807, 1054, 1432, 953, 1433, 1436, 1438, 538, 539]. Document [1305, 793, 555]. Does [587]. Domain [658, 809, 1363, 923, 660, 921, 1165, 1505, 713, 1313, 739, 1017, 1438]. Domains [1353, 1394, 1562, 977, 1008, 955, 1325, 1326]. Dominance [362]. Dominance-based [362]. Dominating [737, 996]. Domination [1241, 1506]. Domino [424]. Double [418, 697, 269, 737, 1160]. down [309]. dP [290]. Drawn [600]. Drift [556]. Driven [226, 488, 1275, 1283, 789]. Driver [273]. Drug [1151]. DSM [1336]. DT0L [254]. Dualities [53, 176]. Duality [1555, 629, 905, 1076, 1323, 1324, 630]. Durda [1525]. Dyck [416, 426]. Dynamic [1149, 462, 1308, 1144, 342, 1046, 834, 1543, 453, 874, 35, 1349, 300, 1155, 1468, 1448, 57, 1049, 1169]. Dynamic-Logical [1349]. Dynamical [611]. Dynamics [106, 686, 363, 277, 1348, 941, 598, 50]. Dynkin [584, 986, 1356, 551, 569].

e-Passport [1549]. e-Tourism [241, 146]. Each [587]. Easy [1144, 111, 265]. EBNF [676]. ECL [1007]. Economy [697]. Ecosystems [969]. ECTL* [496]. EDAS [1246]. Edge [584, 22, 907, 982, 978, 979, 458, 1267, 138, 1108, 1109, 568, 1093, 1094, 1239]. Edge-bipartite [584, 907, 982, 978, 979, 458, 1267, 1108, 1109, 568, 1093, 1094, 1239]. Edge-Minimization [138]. Edges [22, 288]. Edit [71]. Editing [167]. edition [1089]. Editorial [1280]. eDonkey [237]. EEG [1531]. Effect [449, 849, 1403, 1146]. Effectful [92]. Effective [99, 1189]. Effectiveness [2]. Effects [444]. Efficacy [1135]. Efficiency [1088, 944, 1257]. Efficient [1453, 1456, 201, 287, 107, 288, 98, 1511, 873, 324, 1097, 726, 1208, 1476, 824, 202, 263, 1157, 533, 223, 335, 189]. Efficiently [1185, 964]. Eigenvector [1310, 1086]. Elastic [200]. Election [475]. Electrode [559]. Element [344, 345, 997, 998]. Elementary [1465, 874, 162, 515, 493, 701, 722, 1057, 1447, 114, 951, 537, 479]. Elements [1242]. Elettra [1023]. Elicitation [524]. Eliminating [1018]. Elimination [97, 1055, 621, 635, 688, 22, 505, 476]. Elite [1435, 1449]. Elitist [482]. Ellipse [346]. Elliptic [1548, 906, 352, 1268]. Embedded [422, 460]. Embedding [1260, 431, 460]. Emotional [1209, 1532]. Emotions [1532]. Empathy [1209]. Empirical [188, 703, 101, 739, 1197]. Emptiness [1541]. Empty [1474]. eMule [237]. Enclosing [372]. Encoded [1454]. Encoding [656, 1489, 173]. Encodings [101]. Encrypted [926]. Encryption [80, 805, 281, 341, 1551, 16]. ending [556]. Energy [1511, 897]. Energy-aware [1511]. Enforcing [1279, 1301]. Engine [149, 685]. Engineering [871, 64]. Engines [747]. Enhanced [1023, 1261]. Enhancement [79]. Enjoying [829]. ENL [63]. ENL-Transition [63]. Enlargement [1064]. Enriching [839]. Entailment [1399, 189]. Entropy [13]. Enumeration [1509, 1003, 475, 904]. Environment [360, 89]. Environments [871, 161, 556]. Enzymatic [953]. Epichristoffel [847]. Epidemic [1403]. Episode [674]. Epistemic [171, 296, 360, 1007, 1499, 50, 65]. Equal [1107]. Equal-Conflict [1107]. Equality [397]. Equally [1011]. Equation [1226, 1223, 587, 820, 1103]. Equational [160, 1211]. Equations [1224, 1495, 604, 1231, 667, 1228, 141, 797, 826, 403, 1423, 239, 802, 357]. Equilibrium [1122]. Equivalence [463, 966, 1276, 161, 254, 1157, 678]. Equivalence-checking [1157]. Equivalences [469, 1196]. Erosion [1464]. Erratum [241, 341]. Error [9, 76, 1551]. Error-detecting [76]. Essential [1381]. Estimate [1221]. Estimation [1367, 823, 101]. Estimations [133]. ET0L [542]. Evaluating [1183, 96, 1360]. Evaluation [349, 185, 537, 335]. Evaluations [1187]. Event [844, 267, 1542]. Events [150]. Evidence [373]. Evolution [1435, 897, 258, 1348]. Evolution-Communication [897]. Evolutionary [758, 305, 887, 110, 7, 26, 130]. Evolvability [753]. Evolving [735]. Exact [1537, 797, 1404, 911]. Example [632, 836]. Exceptions [1400]. Exchange [924, 1015]. Excluded [230]. Exclusion [1336]. Exclusive [464]. Executable [448]. Execution [682, 481, 1158, 683, 655, 1255]. Exist [587]. Existence [1522, 265, 1327, 1328]. Existential [1363, 1519, 525]. Exo [244]. Exo-Operations [244]. Expansion [500, 1510, 503]. Expansion-based [1510]. Experimental [1484, 1184, 638, 185, 1022, 598]. Experts [660]. Explicit [1007, 793]. Exploiting [1189]. Exploration [383, 524]. Explosion [185]. Exponent [1102]. Exponential [1017]. Exponents [1242]. Exposure [922]. expressible [962]. Expressing [49]. Expression [561, 233]. Expressions [457, 216, 935, 1193, 688, 1306, 645, 1533, 538, 539]. Expressive [1362]. ExpTime [877, 1155]. Extended [1011, 1258, 1408, 278, 797, 309, 137, 789]. Extending [1396, 145, 901]. Extension [74, 355, 374]. Extensionality [631]. Extensions [1003, 1513, 1465, 66, 1325, 1326, 1030]. Extensions* [1080]. Extensive [1349]. Extra [1544]. Extracting [726]. Extraction [2, 909, 730]. Extractors [1029]. Extrapolation [1341]. Extremal [1515].

Face [1533]. Facility [1023]. Factor [1456, 1451, 1346]. Factored [987]. Factoring [1311, 1242]. Factorization [1562, 1547]. Factors [449, 1381, 404]. Facts [1089]. Fair [1015]. Fairness [228, 1348]. Fallacies [875]. Families [10, 420, 308]. Family [1369, 1415, 1222]. Fast [961, 1297, 777, 1208, 930, 716]. Faster [1101]. Fatigue [273]. Faults [541]. FD [149, 1444]. Feasibility [186, 882]. Feature [325, 697, 1530, 11, 644, 77, 909, 166]. Feature-Based [166]. Features [188, 1133, 1189, 868]. Fedorov [1112]. Feeding [454]. Fences [704]. Fermat [195]. Few [857]. Fewest [40]. Fiber [1103]. Field [784, 417, 957, 716]. Fields [135]. Filter [926, 853]. Filtered [853]. Filtering [960]. Filters [305, 887, 1022]. Filtrations [535]. Final [426]. find [1113, 1539]. Finding [1230, 72, 637, 728, 1369, 1415, 938, 867]. Fine [74]. Finite [1446, 337, 528, 1394, 1470, 774, 1284, 135, 906, 181, 689, 259, 137, 138, 1447, 260, 798, 943, 344, 345, 613, 997, 998, 615, 533, 427, 587, 894, 716]. Finite-valued [337]. Finitely [1020]. Finiteness [1541, 1557]. Fire [1121]. First [588, 171, 1317, 1318, 174, 1188, 654, 233]. First-Order [171, 1317, 1318, 174, 1188, 588]. Fitness [517, 1449]. Fitness-based [1449]. Fitting [176]. fix [1212]. Fixed [1408, 50]. Fixed-Point [1408, 50]. Fixing [381]. Flat [1521, 203, 1144]. Flexible [1023, 186, 453, 1025, 445]. Flips [1018]. Floating [808]. Flooding [541]. Flow [1227, 1308, 297, 670, 674, 562]. Flower [994]. Flowgraph [662]. Flows [1445, 1448, 684]. Floyd [479]. Fluid [1227, 598]. fly [738]. Focusing [1002]. Folding [1461]. Following [1348]. Footprint [596, 597]. Forcing [51, 1317, 1318]. Forest [1121]. Forests [1446, 323]. Foreword [408, 1526]. Form [217, 218, 155, 539, 516, 587]. Formal [757, 763, 953, 1126, 180, 700, 875, 384, 270, 1245, 1186, 382, 535, 921, 622, 705, 1540]. Formalism [105]. Formalisms [1291, 916, 173]. Formalization [1556]. Format [872, 1445]. Formation [628]. Forms [458]. Formula [1487]. Formulas [1548, 531, 75, 164, 182, 280]. Formulation [1504]. Forward [815, 281, 341]. Forward-Secure [281, 341]. Foundation [1171]. Foundational [1272, 209, 627]. Foundations [1501]. Four [1432, 440, 1161]. Four-Pixel [440]. Fourier [1024, 1220]. FPGA [1154, 1170]. Fractional [1224, 1496, 1227, 1226, 1223, 1228, 1225, 1221, 802, 1222, 1220, 1495]. Fragment [171, 1202]. Fragments [1463]. Frame [532]. Frames [1493]. Framework [1200, 1074, 1272, 32, 871, 784, 1027, 1397, 783, 52, 687, 1204, 849, 726, 319, 878, 1156, 28, 880, 921, 312, 1386, 313, 1025, 568, 441, 1301, 803, 617]. Frameworks [436, 1201, 1292]. Free [390, 1211, 916, 825, 893, 17, 1070, 1288, 776, 406, 57]. Free-Choice [17]. Freeness [1420, 400]. Frequencies [1423]. Frequency [1455]. Frequent [287]. Friendly [1195]. Frontier [835, 944]. Frozen [17]. Full [348, 1458, 174, 1340]. Full-Newton [1458]. Function [1496, 1055, 1231, 520, 1413, 37, 1522, 440, 1478, 1086, 234, 975, 797]. Functional [1354, 1444, 1215, 38, 1245, 94]. Functions [966, 773, 962, 977, 350, 289, 810, 343, 12, 565, 1119, 422, 136, 1132, 1221, 607, 383, 355, 42, 933, 1480, 869]. Functors [1473]. Fundamental [1159]. Further [633]. Fusion [1371]. Fusion-Based [1371]. Futures [1366]. Fuzzy [372, 791, 1125, 1309, 1378, 1046, 822, 1033, 977, 214, 33, 273, 801, 34, 1371, 518, 1372, 1466, 37, 936, 11, 1034, 375, 1038, 1047, 578, 1008, 1037, 1379, 1039, 561, 1044, 1467, 543, 826, 1014, 1159, 1246, 1478, 1045, 1501, 456, 677, 1061, 1035, 441, 1332, 1494, 272, 215, 1036, 1243, 271, 792, 378].

G. [361]. Gabriel [1250]. Gain [396]. Gained [230]. Galois [33, 1323, 1324]. GALS [919]. Game [212, 1123, 50]. Game-Theoretic [212, 1123]. Games [1388, 529, 1213, 288, 850, 1349, 1365, 983, 1489]. Games* [1072]. Gamma [1564]. Gamma-Reduction [1564]. Gap [1069]. Gap-Order [1069]. Gate [836]. Gates [134]. Gauss [1298, 423]. Gaussian [651]. Gender [1425]. Gender-Based [1425]. Gene [1378, 110, 951, 445]. General [1027, 1397, 512, 22, 515, 1459, 28]. Generalised [1173]. Generalization [381, 1501, 586, 1095]. Generalizations [195, 179, 1203]. Generalized [297, 1411, 21, 1043, 656, 797, 1122, 1070, 1263, 1503, 805, 1162]. Generalizing [917]. Generated [248, 899, 1047, 378, 1128]. Generating [1441, 351, 422, 177, 1099, 1030]. Generation [1395, 709, 635, 736, 858, 1170, 1476, 738, 734]. Generative [276]. Generator [964]. Generators [1168, 156]. Generic [28, 29, 116, 803]. Genes [245]. Genetic [2, 482, 876, 836, 1379, 517, 878]. Genome [120]. GEO [1156]. Geodesic [1544]. Geometric [436, 1087]. Geometrically [690]. Geometries [551, 568, 569, 1239]. Geometry [961, 1124]. get [703]. girls [414]. Given [247, 72, 858, 728]. glass [239]. Global [1336, 392, 602, 1261, 88, 1155, 523, 167]. Glushkov [216]. GLV [1268]. Go [762]. Go-through [762]. Goals [52, 98]. Gödel [530]. Goppa [824]. Gower [1522]. GPU [348, 902, 880, 1024, 1025]. GPU-Like [348]. GPU-optimized [1024]. Graded [52, 801, 152, 229, 466, 532, 1155, 1147]. Gradient [1113]. Gradual [780]. Grafting [550]. Graining [446]. Gram [1093, 1094]. Grammar [284, 196, 1245, 1059, 344, 345, 233, 173]. Grammar-Based [284]. Grammars [276, 1137, 548, 763, 1139, 1289, 1414, 307, 47, 542, 1233, 27]. Grammatical [1138]. Granular [1, 780, 699, 653, 1364, 651, 650, 652, 56, 234, 1048, 883, 1173, 376]. Granularity [780]. Granulation [679, 547, 480, 1039]. Granules [928, 812, 652]. Granulometric [961]. Granulometry [1297]. Graph [1190, 249, 668, 520, 1537, 1444, 1139, 1457, 1241, 1459, 432, 344, 345, 434, 997, 998, 1527, 1100, 867, 27, 1387, 189]. Graphical [1477]. Graphs [245, 437, 342, 846, 60, 297, 1408, 1411, 21, 590, 584, 188, 863, 396, 1419, 386, 771, 433, 907, 1120, 982, 978, 979, 458, 1013, 1121, 1267, 1346, 1123, 55, 1108, 1109, 674, 562, 568, 1093, 1094, 1506, 1314, 1239, 169]. Grasp [709]. Grasping [300]. Gravitational [965, 618]. Greater [587]. Greedy [1116, 487, 665, 139]. Green [1285]. Grey [1261]. Grid [931, 402, 602]. Grill [25]. Grothendieck [1108, 1109]. Group [171, 43, 1310, 521, 1108, 1109, 1086, 1242, 1096]. Groups [755, 982, 827, 607, 301, 1094]. Growing [307]. gs [846]. gs-graphs [846]. gs-monoidal [846]. Guaranteeing [206]. Guarantees [97]. Guard [1064, 1217]. Guarded [1363, 1217]. Guide [120]. Guidelines [202].

Haar [1301]. Hadamard [414]. Hairpin [935, 257, 261]. Half [413]. Half-lattice [413]. Halting [46]. Hamacher [1332, 1494]. Hamiltonian [437, 1431, 797]. Hamming [1260]. Handling [1430, 1203, 882].

Handwritten [822]. Hankel [80]. Hard [1187, 1488, 265, 554]. Hardest [136]. Hardness [247]. Hardware [348, 459]. Hardware/Software [459]. Harju [386]. Hash [607]. Hashing [926, 166]. Hasse [156]. HBC [736]. HDT0L [397]. Head [260]. Heap [479]. Heap-Construction [479]. Heat [756]. Hedges [704]. Height [426]. Henkin [1474]. Herbrand [1202]. Herbrand-satisfiability [1202]. Hesitant [1371, 1014]. Heterogeneous [973, 880]. Heuristic [1304, 98, 188, 777, 823, 728, 453]. Heuristics [1491, 688, 729]. HIBE [223]. Hiding [477, 15]. Hierarchical [846, 325, 1412, 1286, 174, 650, 1030, 524]. Hierarchicality [1030]. Hierarchies [250, 547]. Hierarchy [672, 948, 731, 943, 1307]. High [323, 329, 6, 136, 3, 685]. High-Level [329]. Higher [1413, 1286, 826, 469, 933, 739]. Higher-order [1413, 739]. Hilbert [360]. Hilbert-space [360]. Histograms [1505]. Histological [1467]. Histories [972]. HLTL [525]. Hoare [434]. Hoare-Style [434]. Hoc [160]. HOL [1559]. Hole [4]. Holistic [820]. Holm [1223]. Holonomic [421]. Homage [1393]. Home [871]. Homogeneous [840, 1198, 1210]. Homomorphism [755]. Homomorphisms [530, 963]. Honoring [386]. hop [475]. Horizontal [110, 601, 858, 907]. Horn [142, 1188]. Hornets [701, 1057]. Hosoya [1160]. Hospital [923]. Hotspots [1518]. Huge [1053]. Hull [1468, 536]. Human [1444, 558, 1533]. Human-Computer [558]. Humanoid [300]. Hybrid [1334, 967, 1375, 896, 656, 732, 14, 684, 1551]. Hybridization [618]. Hyman [1226]. Hyper [453]. Hyper-heuristic [453]. Hyperbolic [797, 949]. Hyperbolic-function [797]. Hypercomputation [985]. Hypercuboid [1172]. Hyperelliptic [1547]. Hypergeometric [423]. Hypergraph [736]. Hypersets [1005]. Hypertension [1381].

Ideal [775, 1421]. Ideals [1562]. Idempotent [181]. Identification [1053, 1297, 875]. Identifying [36, 1189]. Identity [1021, 1207, 1101, 16, 184, 281, 341]. Identity-Based [1207, 281, 341, 1101]. if [1274, 587]. Igarashi [519]. II [979, 218, 309, 1109, 29, 345, 998, 1094]. Ikonos [79]. Ill [766]. Ill-posed [766]. Image [961, 602, 861, 727, 13, 477, 79, 1535, 166, 1313, 690, 868, 1357, 1102, 80, 805]. Images [854, 855, 219, 1467, 78, 1533, 405, 15]. Images* [1044]. Imbalanced [1167]. Imidazolium [787]. Immunity [1403]. Immunization [35]. Immunological [975]. Impact [1426, 1363, 867]. Impedance [495]. Imperative [1524]. Imperfect [1388]. Implementation [571, 784, 84, 93, 841]. Implementations [92, 204]. Implementing [145, 1204]. Implicate [36]. Implication [1036]. Implications [382, 1061]. Implicators [936]. Importance [537]. Important [1089]. Improved [74, 471, 617, 480]. Improvement [1026, 1530, 714]. Improving [1296, 96, 658, 465, 301]. Impurity [424]. Incidence [679]. Inclusion [518, 229, 810, 245]. Incomplete [1354, 1397, 790, 210, 1090, 252, 734, 235, 373, 620, 1243]. Incompletion [261]. Inconsistencies [621]. Inconsistency [1091, 796, 930]. Incremental [1271, 687, 1476, 710, 640, 1049, 792]. Incrementing [931]. Increments [996]. Independent [767, 82]. Independently [735]. Index [293, 315, 336, 358, 385, 407, 428, 450, 474, 497, 509, 526, 552, 582, 603, 625, 663, 692, 717, 740, 768, 800, 828, 851, 885, 912, 939, 970, 990, 1010, 1031, 1050, 1073, 1129, 1141, 1163, 1181, 1199, 1218, 1249, 1269, 1302, 1329, 1330, 37]. Indexed [284]. Indexing [727, 793]. Indicators [796]. Indices [1091, 810, 557]. Induced [1053, 929, 291, 75, 164, 1167, 460]. Induced/Embedded [460]. Induction [707, 638, 790, 198, 54, 578, 640, 1197]. Inequalities [1105]. Inertia [1449]. Infeasibility [882]. Infeasible [1232]. Infeasible-interior-point [1232]. Infection [444]. Inference [666, 532, 102, 1138, 118]. Inferences [174]. Infinitary [1211]. Infinite [1200, 465, 277, 23, 1563, 731, 532, 67, 406, 1234, 215]. Infinite-state [1200]. Infinite-valued [532]. Inflation [458, 1267, 1108]. Influence [1184, 733]. Influencing [1381]. Information [337, 1354, 1395, 1046, 1388, 1397, 363, 210, 632, 1371, 596, 1007, 492, 670, 928, 923, 1118, 585, 974, 192, 1122, 723, 955, 66, 674, 643, 652, 1076, 734, 79, 619, 524, 546, 788, 235, 679, 620, 1243, 1049, 480]. Informational [491]. infrastructure [803]. Inhabitation [1556]. Inheritance [433, 275]. Inhibitor [60]. Initial [713]. Inner [361, 607, 519]. Innovativeness [697]. Inpainting [477]. Input [1118]. Inputs [1432]. Insecurity [16]. Insertion [244, 950, 1287]. Insertion-Deletion [1287, 950]. Insider [1405]. Insight [563]. Inspired [1019, 83, 519, 1166, 448]. Instance [1498]. Instantiation [29]. Instead [233, 430]. Institutional [1425]. Instruction [481, 966, 1192]. Instructions [1255]. Instrument [183]. Integer [1311, 587]. Integers [401, 1242]. Integral [1224]. Integration [871]. Integrity [1543, 491, 1462]. Integro [1225, 826]. Integro-differential [1225, 826]. Intelligent [30, 1, 1240, 68]. Intensive [1445]. Intentions [52]. Interacting [64]. Interaction [1359, 105, 1372, 558, 660]. Interactions [323]. Interactive [757, 1277, 812, 699, 727, 793, 524]. Interconnection [1459]. Interface [43, 205]. Interfaces [1427]. Interference* [992]. Interior [518, 1131, 1265, 1458, 1232]. Interior-point [1265, 1458]. Interleaved [65, 470]. Interleaving [1410]. Intermediate [1445]. International [317]. Interpolation [1325, 1326]. Interpretation [170, 1303, 849, 1377, 723, 652, 514]. Interpretations [947, 1074, 655]. Interpreted [470, 1062]. Interrupt [1065]. intersecting [420]. Intersection [1555, 1556, 394, 221, 503, 504, 259, 505, 506, 507, 119, 499]. Interval [1231, 520, 33, 1466, 1538, 1520, 1038, 1288, 378, 1340, 1014, 1246, 494, 313, 1243, 1049]. Interval-Timed [1340, 494]. Interval-Valued [33, 1466, 1038, 1288, 1014, 1246, 1049]. Intervals [72]. Intra [1425]. Intra-Institutional [1425]. Intrinsic [1562]. Introducing [1501]. Introduction [1469]. Introspection [1147]. Intuitionistic [1002, 1466, 1561, 543, 501, 1401, 720]. Invariant [1074, 1281, 690, 346]. Invariants [592, 1102]. Inverse [1027]. Inversion [388]. Inverted [1061].

Invertible [610]. Investigation [209, 148]. Involutive [419]. Ion [1433]. Ions [1433]. IR [1230]. Irreducibility [135]. Irredundant [929]. Isabelle [1559]. Isabelle/HOL [1559]. Isometric [1289]. Isotropy [982, 1094]. Issue [1250, 1, 386, 1497, 1089]. Issues [447, 627, 1171, 924, 483, 211]. Italian [1023]. Iterated [32, 257]. Iterating [958]. Iteration [1370, 826]. Iterative [536]. IUML [1446]. IUML-algebras [1446]. Izhikevich [559].

JaCaMo [1359]. Jacobians [1547]. Java [227, 463, 666, 1204, 1009]. Jeffrey [32]. Jerzy [105]. Job [1511]. Johann [19]. Join [1095]. Joint [468]. JSetL [1009]. JTabWb [1204]. Jumping [1233]. Justification [1482]. Juvenile [659].

Kad [237]. Kadanoff [366]. Karp [511]. KDD [713]. Kepler [683]. Kernel [1130, 1227, 1228, 1522, 7]. Kernel-Trick [1522]. Kernelized [273]. Kernels [1234, 189]. Key [1402, 1168, 355, 312, 281, 341, 1015]. kFOIL [325]. Kite [614]. Kleene [306]. Kmeans [617]. Knowledge [599, 658, 171, 1497, 34, 558, 454, 923, 187, 1047, 142, 1166, 921, 1165, 1505, 677, 713, 178, 679, 373, 369]. Knowledge-lean [454]. Knowledge-oriented [369]. Kolmogorov [415]. kR [890]. Kripke [814, 866]. Krohn [393].

L [236]. Label [922, 442]. Labeled [795]. Labelled [1396]. Lambda [93]. Language [604, 282, 845, 1554, 856, 721, 763, 953, 231, 1059, 403, 818, 1138, 705, 706, 1545]. Languages [1417, 947, 10, 528, 1303, 1274, 248, 390, 391, 952, 548, 848, 1317, 1318, 843, 1028, 761, 762, 250, 251, 1079, 1279, 1211, 775, 386, 1126, 397, 928, 256, 1523, 899, 76, 672, 702, 522, 308, 731, 1524, 311, 1421, 535, 893, 404, 533, 799, 235, 1128, 1266]. Large [1311, 763, 906, 1154, 644, 727, 653, 3, 733]. Latency [1304, 981]. Lattice [1011, 33, 1153, 362, 421, 422, 1561, 866, 425, 1080, 642, 413]. Lattice-ordered [1080]. Lattice-valued [866]. Lattices [379, 605, 154, 1319, 1320, 1493, 269, 306, 1047, 1101, 1516, 271]. Lattices* [1076]. Law [517]. Layer [705]. layered [9]. Layers [640]. Lazy [432]. LCP [1452]. Leader [475]. Leading [1122]. lean [454]. Leap [1383]. Learn [764]. Learning [1212, 1528, 1435, 1137, 107, 763, 1412, 328, 1139, 687, 764, 1530, 556, 467, 12, 565, 1008, 1348, 560, 472, 165, 1535, 1449, 819, 1035, 788, 1529]. Least [1472]. Left [542]. Leftist [1287]. LEM2 [1170]. Lemma [1317, 1318, 770]. Length [831, 694, 418, 1454, 307, 14, 406, 1169]. Length- [406]. Length-reducing [307]. Level [416, 329, 75, 164, 1186, 460, 574]. Lexicalized [1245]. Lexicographic [389]. Lexicon [555]. Lexicon-based [555]. Leximin [1346]. Library [1513]. Life [1258]. Light [1522, 1025, 333, 495]. Lightweight [1097]. Like [348, 349, 1434]. Limit [246]. Limitations [1286]. Limited [1178, 1282, 262, 893, 1180, 1349]. Limiting [922]. Line [743, 1024]. Linear [976, 171, 1231, 349, 1216, 1452, 1253, 163, 585, 1113, 89]. Linear-time [171, 163]. Linearity [113]. Lines [40]. Lingua [901]. Linguistic [1371]. Linkage [106, 686]. Linking [1486]. Listing [737]. Literals [1509]. Live [993]. Liveness [82, 490, 1107, 1409, 299]. Loading [861]. Local [583, 388, 393, 431, 1289, 188, 54, 190, 3, 55, 865, 624, 853, 770, 690, 1222]. Local-Correlated [3]. Localised [954]. Locality [926, 1325, 1326]. Localization [862]. Location [4]. Location-Based [4]. Logarithmic [343, 1544]. Logarithms [1029]. Logic [1251, 1001, 19, 241, 1002, 711, 49, 1184, 436, 171, 296, 720, 1394, 1553, 1378, 1396, 574, 338, 150, 1360, 1470, 1398, 152, 687, 1393, 1543, 1361, 1007, 229, 1433, 1253, 279, 163, 585, 1560, 1520, 175, 1499, 231, 1561, 578, 75, 164, 751, 99, 532, 877, 1155, 1059, 177, 146, 102, 579, 42, 182, 280, 580, 508, 1401, 119, 1438, 1147]. Logic-Algebraic [229]. Logic-Based [241, 150, 146, 1398, 578, 580]. Logical [172, 391, 952, 635, 1277, 1007, 1349, 1501, 1165, 178]. Logics [337, 1273, 1055, 1482, 1362, 1008, 65, 176, 1447, 142, 1501, 1352, 1325, 1326, 819, 50]. Logs [1542]. Long [1348, 969]. Long-running [969]. Long-Term [1348]. Longest [1456, 712, 45, 1451]. Look [197]. Look-back [197]. Looking [263]. Loop [75, 164]. Loops [978, 979, 1267]. Loopy [589]. Loss [841]. Lovász [770]. Low [1260, 126, 1268]. Low-cost [1268]. Lower [736, 289, 1117, 1236, 1448, 1503]. LTL [1360]. LTLK [470]. Lucas [355]. Lucent [1542]. Łukasiewicz [905, 1214, 532, 1501].

M&M [354]. MABAC [1246]. Machine [763, 764, 309, 472, 1247, 1035, 642]. Machines [1285, 945, 7, 892, 543, 942, 264, 183, 957, 1266]. MacWilliams [395]. Made [1144, 111]. Magda [1554]. Maintenance [923, 710, 792]. Making [1310, 1371, 1372, 319, 1014, 1246, 1478, 1086, 788, 1332, 1494, 538, 523, 167]. Makowsky [19]. Malicious [1550, 1405]. Mammograms [1529]. Management [668, 8]. Manipulating [1006]. Manipulation [1126]. Manner [1521]. Many [487, 665, 176, 1020, 685]. Many-task [685]. Many-Valued [487, 665, 176]. Map [697]. Maple [1513]. Mapping [1005, 795, 75, 164, 1234, 1096]. MapReduce [617]. Maps [430, 805]. Mario [1393]. Marked [1537, 254]. Markov [1298, 1194, 198]. MAS [871]. Mashups [1427]. Mass [109, 1501]. Mass-based [1501]. Matching [1278, 1117, 703, 222, 1545, 437, 968]. Materialization [1445]. Mathematical [209, 1087, 1560, 215]. Matrices [1190, 1088, 1295, 379, 1090, 395, 439, 1418, 1097, 1018, 1262, 1134, 1314, 989]. Matrix [387, 1500, 388, 1087, 548, 1465, 584, 439, 907, 7, 1356, 1235, 551, 27, 80]. Matrix-computational [1465]. Matroid [198, 715]. Max [840]. Maximal [76, 405, 691]. Maximally [295, 1521]. Maximization [733]. Maximizing [971, 11]. Maximum [1308, 1230, 1278, 255, 100, 1448]. MaxSAT [191]. McEliece [824, 1551]. MCINTYRE [579]. McNaughton [308]. MDS [1097]. Mean [1087]. Meaningful [637]. Means [1444, 1522, 505, 561, 1260, 456]. Measurable [759]. Measure [1019, 848, 1457, 549, 1246, 1313, 1503, 240]. Measurement [668]. Measurements [237, 1020]. Measures [637, 1305, 647, 1040, 1123, 1460, 733, 1247, 234, 1197]. Mechanical [181]. Mechanism [1436]. Mechanisms [889, 1345]. Media [1426, 1527]. mediated [1382]. Medical [8, 640]. Medicine [8]. Meet [1045]. Meeting [1385]. Membership [773, 1480]. Membrane [832, 954, 835, 896, 258, 948, 838, 839, 633]. Membranes [837, 1257]. Memoriam [478]. Memory [464, 1152, 941, 943, 685]. Mercury [1433]. Mereological [1501]. Mereology [816, 1501, 752]. mereotopological [57]. Mereotopologies [1327, 1328]. Mereotopology [57]. Merging [873]. Merotopological [25]. Mesh [907, 978, 979, 239, 551, 568, 569, 1239, 978]. Message [924, 667, 927, 987]. Meta [1304, 721, 641, 883]. Meta-Actions [641]. Meta-Granular [883]. Meta-Heuristic [1304]. Meta-Theory [721]. Metaheuristic [1528, 452, 1237]. Method [446, 1495, 1151, 1055, 967, 1310, 662, 621, 743, 1223, 1131, 1458, 644, 823, 988, 1531, 440, 578, 1364, 797, 1044, 877, 344, 345, 738, 1383, 997, 998, 1423, 622, 1024, 1113, 103, 793, 167, 1268, 1243, 357]. Methodological [519, 803]. Methodologies [30]. Methods [1250, 1528, 1090, 1537, 324, 1367, 1246, 853, 654]. Metric [243, 1260, 1419, 1013, 166]. Metrics [275]. MGRSs [1040]. Micro [602]. Micro-tomography [602]. Microarray [165]. Microbleeds [1530]. Micropalaeontology [624]. Microtomography [861]. Miller [1311, 511]. Mimicking [753]. Mind [1209]. Mini [1094]. Mini-groups [1094]. Minimal [277, 1120, 63, 737, 218, 1071, 511, 1257, 742, 691, 538, 539]. Minimality [678]. Minimally [903]. Minimization [1144, 847, 891, 138, 158]. Minimum [372, 1304, 1230, 452, 255, 1118, 1391, 1113, 999]. Minimum-Enclosing-Ball [372]. Mining [325, 287, 659, 813, 4, 6, 3, 1166, 964, 921, 1165, 1505, 460, 641, 8, 657]. miRNA [645]. Mirror [405]. Misclassifications [695]. Mismatch [495]. Missing [545]. MITL [525]. Mixes [183]. Mizar [1153]. MLEM2 [54]. Mobile [160, 4]. Modal [574, 338, 1470, 576, 530, 1493, 231, 176, 1447, 1059, 205, 1080, 1401, 1173]. Modalities [1321, 1322]. Mode [903, 739]. Model [1509, 707, 1227, 1272, 296, 444, 482, 282, 1066, 1396, 1402, 338, 1298, 213, 1427, 1444, 1240, 953, 743, 366, 1375, 110, 1207, 1433, 329, 454, 1367, 973, 900, 87, 65, 1403, 1069, 232, 1390, 344, 345, 1157, 1209, 1341, 115, 1468, 559, 1489, 818, 1015, 1113, 223, 1301, 495, 706, 1410, 525, 884, 1062, 1438, 804, 211, 274]. Modeled [1381]. Modeling [1184, 782, 573, 140, 109, 521, 997, 998, 1188, 116, 448, 1366]. Modelling [1251, 960, 1376, 1378, 1351, 747, 1352, 1138, 370]. Models [1251, 756, 411, 1470, 843, 1548, 687, 455, 700, 750, 1561, 1121, 1245, 1443, 955, 238, 1503, 1243, 1542]. Modern [96, 1488]. Modification [932]. Modified [399, 1265, 559, 1437]. Modifying [696]. Modular [489, 1242]. Modules [917, 798]. Moduli [1472]. Modulo [1145, 1092]. Modulo- [1092].

Modulus [440, 587]. Moisil [905]. Moments [1102]. Monadic [49, 308, 208]. Monitoring [1258, 834, 273, 285]. Monodic [171]. Monoid [952, 256]. monoidal [846]. Monotonic [623, 1422, 1212, 845]. Monotonicity [621]. Monte [644, 579]. Monteiro [750]. Moore [847]. Mop [287]. Morgan [362, 1076]. Morphisms [292]. Morphogenesis [448]. Morphological [1464, 601]. Morphos [149]. Morsifications [584, 907, 551, 568]. Most [986, 22]. Motion [300]. MR [1044]. MRI [594]. MSO [20, 455]. MTL [884]. mTOR [1375]. Muchnik [1370]. Multi [965, 9, 243, 922, 707, 475, 5, 148, 1344, 602, 1471, 1207, 671, 1350, 65, 1039, 40, 26, 260, 1478, 204, 559, 314, 1401, 714, 301, 524, 69, 788, 804, 18, 370, 805, 1161, 513]. Multi-Agent [148, 1401, 69, 65, 370]. Multi-attribute [788]. Multi-clock [204]. Multi-Color [40]. Multi-Context [1350]. Multi-Covering [1161]. Multi-Criteria [243, 1478]. Multi-Dimensional [524]. Multi-Electrode [559]. Multi-granulation [1039]. Multi-grid [602]. Multi-Head [260]. Multi-hop [475]. Multi-Label [922]. Multi-layered [9]. Multi-Objective [26]. Multi-Party [1344, 671]. Multi-Pixel [18]. Multi-Proxy [1207, 714]. Multi-Quantifier [1471]. Multi-Receiver [314]. Multi-Relational [707]. Multi-Robot [301]. Multi-sawtooth [805]. Multi-Step [804]. Multi-Website [5]. Multi-workflow [513]. Multicast [748, 696]. Multicolor [599]. Multicounter [250, 398]. Multidimensional [782, 783, 647]. Multifaceted [786]. Multigranulation [974]. Multilayer [815]. Multimodal [1172, 1443, 274]. Multinomial [1367]. Multiobjective [1511, 896]. Multiple [947, 247, 1384, 1371, 1372, 896, 371, 1346, 78, 1075, 1332, 1494]. Multiple-category [371]. Multiplications [716]. Multiplicative [1368, 224]. Multiplicities [1070]. Multiplicity [1262]. Multiprocessor [973]. Multirelations [53]. Multiset [295, 834, 943]. Multisingular [732]. Multisource [1428]. Multispectral [79]. Multitask [325]. Multivalued [1022]. Music [647, 321]. Musical [183]. Mutex [515]. Mutual [1336].

Naive [162, 454]. Names [157, 332]. Naming [1259]. NAO [752]. Nash [1122]. Natural [763, 597, 1558, 141, 875, 680]. Near [932, 827, 562, 441, 70, 1490]. Nearness [466, 1457, 564, 827, 624, 675]. Negation [1004, 382]. Negative [492, 986]. Negotiating [1185]. Negotiation [282]. Neighborhood [1304, 981, 1131, 1232, 1265, 1263, 1373, 1479]. Neighbourhood [520, 1561]. Neighbourhoods [562]. Nelson [750]. Nervous [753]. Nested [318, 575, 490, 431]. Net [1271, 60, 297, 632, 1375, 220, 698, 329, 1058, 514, 818, 1539, 1366]. Net-Based [698]. Nets [1064, 82, 1106, 540, 1376, 1335, 226, 1407, 918, 154, 821, 327, 1378, 512, 808, 696, 1537, 490, 328, 1381, 914, 915, 747, 521, 516, 1377, 1107, 1409, 515, 61, 62, 299, 671, 745, 1379, 64, 155, 917, 744, 995, 916, 1070, 232, 732, 156, 29, 1340, 738, 879, 494, 511, 157, 332, 1030, 1307, 677, 1105, 1196, 678, 17, 820, 513].

Nets-based [738]. Network [1375, 1367, 968, 4, 35, 1443, 1386, 1535, 882]. Networks [830, 708, 1308, 1376, 758, 475, 1435, 305, 887, 1380, 1067, 107, 589, 198, 160, 698, 35, 1379, 1459, 567, 733, 815, 1515, 1386, 1448, 1527, 130, 445, 657]. Neumann [1479]. Neural [1430, 1435, 290, 898, 899, 968, 901, 292, 262, 840, 815, 1535, 1527, 1312, 1198, 841, 1210, 842]. Neuro [822]. Neuro-Fuzzy [822]. Neuroimaging [1443]. Neurons [1375, 559, 1198, 1210]. Neutral [1225]. Never [556, 1185]. Never-ending [556]. Newton [1458]. Niho [1413, 933]. NNV [797]. No [1227]. Noise [860, 1017]. Noisy [859, 1020]. Nominal [755, 666, 1260]. Nominal-continuous [1260]. Nominals [877]. Non [1183, 992, 1212, 845, 1430, 1456, 773, 952, 996, 251, 1079, 1289, 986, 726, 777, 897, 646, 423, 623, 561, 127, 113, 734, 802]. Non-coding [561]. Non-Convex [1183]. Non-cooperative [897]. Non-determinism [1430, 113]. Non-Deterministic [734, 777]. Non-dominating [996]. Non-Interference* [992]. Non-Isometric [1289]. Non-Membership [773]. Non-Monotonic [623, 1212, 845]. Non-negative [986]. Non-overlapping [1456]. Non-Parallel [726]. Non-probabilistic [802]. Non-Regular [127, 251]. Non-Returning [1079]. Non-Terminating [423]. Non-Zero [646]. Nonarchimedian [761]. Nonclassical [1130]. Nondeterminism [127, 405]. Nondeterministic [303, 259, 137, 138, 1009, 533, 235]. Nonlinear [1137, 1103]. Nonlinearities [350, 289]. Nonlinearity [1413, 933]. Nonlocal [1224]. Nonmonotonic [282]. Nonnegative [1231]. Nonterminal [1414]. Nonterminals [47]. Nonviolent [919]. Normal [804, 539]. Normalising [501]. Normality [135]. Normalization [169, 1558, 1467, 506]. Norms [1348]. Note [133, 338, 391, 51, 1413, 23, 1525, 70, 1210]. Notes [720]. nothing [1345]. Noting [66]. Notion [83, 631]. Novel [1130, 1004, 621, 1372, 440, 1386, 1535, 1437, 1313, 805, 274]. NP [764, 601, 554, 1288]. NP-complete [601]. NP-hard [554]. NTT [860]. Nulls [1519]. Number [695, 860, 1414, 1118, 1120, 1544, 1241, 47, 942, 877, 563]. Numbers [1005, 141, 425, 1368]. Numeric [652, 1239]. Numerical [1496, 826, 1325, 1326, 1128]. NURIKABE [253]. Nursing [923].

Object [782, 855, 161, 874, 62, 299, 493, 722, 275, 1369, 1415, 818, 713, 235, 1049, 792]. Object-Based [782, 855]. Object-Directed [235]. Object-Oriented [161, 275, 818]. Objective [965, 1346, 26, 880]. Objects [1191, 675, 770, 781, 1366]. Observational [170]. Observations [230, 1089]. Observed [703]. Obtained [776]. Obtaining [889, 1460]. Occasion [105, 1250]. Occurrence [540, 154, 821]. OECS [1017]. off [522]. Offs [196]. OLAP [783]. OLAPing [784]. Old [613, 742]. Omega [471]. Omega-automaton [471]. Omniscience [178]. On-line [1024]. On-the-fly [738]. Once [773, 138]. One [1495, 604, 668, 365, 124, 863, 612, 892, 402, 424, 607, 1036, 1081, 1474, 1266].

One-Catalyst [1081]. One-Dimensional [612, 1495, 365]. One-regular [863]. One-Rule [402]. One-sided [604, 1266]. One-Time [607]. One-Way [892]. Online [657]. Only [848, 996, 340]. onto [559]. Ontological [68]. Ontologies [1361, 1364, 1400]. Ontology [578]. Opacity [83, 1407, 1487, 811]. Open [514, 1025]. OpenOpt [878]. Operating [1125, 202]. Operation [1471, 1389, 1438]. Operation-Based [1471]. Operational [220, 1538, 112]. Operations [244, 1175, 1284, 1079, 394, 549, 18]. Operator [1297, 151, 427, 1036]. Operators [154, 457, 530, 518, 1372, 1466, 1493, 231, 1221, 1159, 1332, 1494, 911, 272, 1036, 1096, 271]. Opinion [1426]. Opposition [1435, 1033, 1449, 630]. Opposition-Based [1435, 1449]. Oppositions [1085]. Optimal [1484, 134, 636, 987, 944, 1341, 938, 334, 1248]. Optimality [74]. Optimisation [87, 766]. Optimiser [159]. Optimization [1491, 462, 1528, 97, 822, 452, 554, 1261, 1131, 1232, 1265, 7, 973, 26, 100, 880, 1512, 684, 713, 1449, 1169]. Optimized [1017, 1024]. Oracle [764]. Oracles [948, 184, 281, 341]. Orbits [611, 568, 569, 551]. Order [1496, 49, 171, 389, 1317, 1318, 520, 435, 906, 175, 1012, 1561, 174, 65, 1069, 825, 606, 826, 66, 1188, 933, 802, 976, 588, 350, 289, 1413, 1265, 739]. Order-based [1012]. Ordered [1371, 795, 974, 546, 1049, 1080]. Ordering [96, 1498]. Orders [1538]. Ordinal [621]. Ordinals [44]. Ore [1465]. Organization [655]. organized [449]. Organizing [697]. Orientation [1236, 1114]. Orientation-space [1114]. Oriented [85, 696, 161, 35, 275, 818, 1034, 369]. Orienteering [876]. Origin [426]. Ornaghi [1393]. Orthogonal [423, 1518]. Orthomodular [1492]. Orthopairs [1446, 213]. Other [910]. Outlier [660]. Outliers [6]. Overlap [245, 406]. Overlap-inclusion [245]. Overlapping [1028, 1456]. Overview [921]. OWL [1364].

P [1430, 1434, 1521, 901, 980]. P-Lingua [901]. p16 [1382]. p16-mediated [1382]. p70S6K [1375]. PACE [1549]. Packing [103]. PageRank [1367]. Pair [936]. Pairing [1195]. Pairing-Friendly [1195]. Pairs [431, 1042, 420]. Pairwise [1088, 1091, 1087, 1090, 796, 930, 1089, 823, 988, 1083, 1085, 1135]. Palindromes [1453]. Palindromic [712, 856]. Paper [519, 241, 341]. Paraconsistent [1344, 163]. Paradox [1337]. Parallel [295, 45, 726, 1521, 1208, 973, 187, 40, 47, 26, 903, 938, 894, 173, 1230]. Parameter [1378, 814, 598]. Parameterized [1200, 904, 1067]. Parameters* [1065]. Parametric [1308, 84, 61]. Parametrized [1258]. Parikh [1134]. Parity [1192, 983, 426]. Parsing [233, 1138]. Part [431, 309, 28, 29, 344, 345, 1209, 997, 998]. Partial [462, 430, 1176, 771, 65]. Partially [475, 161]. Particle [822, 973, 1512, 1449, 159]. Particles [417]. Partitioning [193, 1048]. Partitions [1011, 1477]. Partners [331]. Party [1344, 671]. Passages [746]. Passing [667, 927, 987, 1256]. Passport [1549]. Password [1549]. Path [709, 876, 425, 752, 1369, 1415, 867, 938]. Path-Finding [938, 867]. Path-Relinking [709]. Paths [413, 416, 249, 108, 421, 422, 426, 427]. Pathway [1382]. Pattern [412, 1303, 1078, 924, 418, 287, 1441, 650, 378, 1490]. Pattern-based [1078]. Patterning [449, 445]. Patterns [1053, 72, 659, 649, 964, 661, 559, 427, 730]. Pawlak [679]. PBW [1513]. PCFGs [1138]. PDEs [1222]. PDL [88]. Pebble [1213, 124, 23]. PEG [676, 817, 1060, 1342]. Penrose [614]. Perception [1151, 70]. Perface [1416]. Perfect [749, 1117]. Performance [1184, 747, 301, 685]. Performing [1457]. Periodic [611, 771, 401]. Permutation [438, 1021, 731, 1512]. Permutation-based [1512]. Permutations [600, 418, 1441, 419, 1018, 776]. Persistent [327, 993, 1275, 919, 119]. Personal [1560]. Perspective [1001, 1317, 1318, 507, 114, 173]. Perspectives [699, 7]. Persuasion [1344]. Perturbed [1088]. Petri [1064, 82, 1106, 1271, 1376, 1335, 1407, 918, 327, 1378, 512, 1537, 490, 328, 1381, 1375, 914, 915, 747, 220, 521, 516, 698, 1377, 329, 1107, 1409, 61, 1379, 1058, 155, 917, 916, 1070, 232, 732, 156, 29, 1340, 738, 514, 879, 494, 511, 157, 332, 818, 677, 1105, 1540, 1196, 678, 17, 820, 1539, 1366]. Phantoms [1022]. Phase [857, 610]. Phasespace [129]. Philosophers [1398]. Phylogeny [1275]. Physarum [818]. Picture [248, 758, 1332]. Piecewise [1119]. Pierce [519]. Pitfalls [314]. Pixel [440, 78, 18]. Pixel-Value [18]. Places [513]. Plan [186, 98, 1189]. Planar [1448, 1020]. Plane [1385, 1139, 40, 728]. PlanICS [298, 878, 1156]. Planner [1184]. Planning [31, 148, 188, 987, 878, 1156, 1339, 752, 1189]. Plans [880]. Played [288]. Playing [778]. Point [195, 1408, 1548, 1131, 1391, 1468, 57, 868, 50, 1232, 1265, 1458]. Point-free [57]. Point-to-point [1391]. Points [1212, 417, 40, 728, 562, 816, 1060]. Polarizationless [1257]. Policy [1341]. Polycephalum [818]. Polycube [1112]. Polygon [411, 1544]. Polygonal [195]. Polygons [600]. Polyhedra [1112]. Polyhierarchical [656]. Polymerase [953]. Polynomial [1394, 320, 22, 1160]. Polynomials [135, 423, 551]. Polyominoes [859, 601, 858]. Polyvariance [577]. Pomykala [143]. Populations [517]. posed [766]. Posets [154, 1492, 986]. Position [1459]. Positive [1190, 338, 492, 458, 1267, 1356]. Possibilistic [32, 319]. Post [1167]. Post-processing [1167]. Postselection [534]. Potts [1298]. Power [1495, 276, 887, 890, 1434, 517, 1494, 534]. Powers [1286]. PP [349]. PP-1 [349]. Pre [1212]. Pre-fix [1212]. Precise [1355, 479]. Precision [1263]. Predicates [1363, 1327, 1328]. Predicting [5]. Prediction [1382, 763, 900, 787, 661, 472, 1386, 975]. Predictor [1150, 1232, 1265]. Predictor-corrector [1265]. Predictors [1487]. Preface [1063, 91, 1552, 48, 626, 991, 1000, 1016, 1032, 1219, 1424, 886, 1174, 1182, 121, 302, 1439, 754, 608, 1450, 1343, 527, 207, 359, 1164, 940, 913, 553, 779, 59, 225, 461, 664, 693, 806, 1051, 1148, 1333, 1104, 593, 852, 1110, 1294, 1280, 199, 1136, 570, 1392, 153, 1290, 104, 443, 1374, 95, 144, 498, 1205, 741, 769, 681, 131, 1142, 1082, 347, 242, 1115, 1536, 510, 377, 1546, 785, 326, 724, 1358, 1507, 718, 1315, 1316, 168, 316, 895, 1429, 1270, 81, 294, 485, 870, 1481, 322, 429, 451, 920, 368, 268, 616, 1229, 1406]. Preference [524]. Preferences [1351, 318, 1350, 1346]. Prefix [794, 964]. Prefixes [776, 404]. Premises [801]. Preorder [788]. Preprocessing [101, 882]. Presence [541, 1403]. Preserving [1418]. Previous [1456]. Priced [744]. Prime [906, 36]. Primes [351, 591, 1195]. Primitive [194]. Principal [775, 458, 239]. Principle [1334]. Principles [1117]. Prior [599, 1298, 1301]. priori [133]. Priorities [1058]. Priority [686, 988, 1086]. Privacy [924, 1056]. Private [230]. Probabilistic [1125, 295, 528, 303, 84, 638, 790, 783, 786, 1371, 859, 1341, 579, 770, 211, 267, 802]. Probabilistically [147]. Probability [1501, 642, 274]. Problem [132, 1304, 437, 574, 1171, 529, 774, 46, 45, 108, 1311, 1511, 896, 554, 254, 279, 876, 701, 1057, 968, 745, 1459, 1237, 100, 222, 903, 729, 816, 1099, 1391, 1095, 103, 999, 648, 178, 442, 525]. Problems [1417, 599, 31, 387, 528, 1027, 1431, 904, 228, 985, 531, 453, 185, 521, 764, 1068, 554, 1176, 1420, 217, 218, 916, 138, 1488, 404, 157, 1512, 867]. Procedural [1347]. Procedure [988, 401]. Procedures [151, 1084]. Process [1271, 1125, 43, 1252, 282, 573, 687, 1475, 86, 230, 180, 813, 64, 1460, 523, 333, 167, 746, 1542]. Processes [922, 1378, 759, 1397, 512, 1483, 161, 515, 99, 1340]. Processing [337, 833, 723, 313, 79, 1527, 713, 1167]. Processor [654]. Processors [758, 305, 887, 130]. Produced [481]. Product [361, 573, 590, 516, 155, 67]. Product-form [516]. Production [573]. Productions [47]. Productivity [1425]. Products [708, 414]. Professor [1250, 386]. Profiles [335]. Program [283, 1203, 634, 479, 581]. Programming [1143, 1149, 1183, 462, 1001, 1184, 1054, 1394, 572, 1144, 1145, 85, 1396, 721, 1360, 318, 486, 1393, 1361, 1215, 1347, 1524, 99, 177, 321, 579, 1009, 1292, 1188, 818, 89, 94, 1147, 1169]. Programs [486, 1006, 197, 75, 164, 1157, 434, 204, 102]. Progressions [535]. Progressive [960, 15]. Projection [854, 859, 858, 1098, 911]. Projections [857, 601, 67]. Prolog [581]. Promoters [1434]. Promptness [983]. Proof [169, 1465, 507, 1504, 1401, 159, 1345]. Proof-graphs [169]. Proof-theoretic [507]. Proofs [777, 502]. Propagation [1223]. Propagator [468]. Proper [291]. Properties [1150, 227, 367, 283, 392, 21, 364, 669, 86, 504, 1389, 702, 988, 232, 1217, 66, 1254, 1159, 598, 753, 383, 1540, 884, 267]. Property [169]. Propositional [1252, 1561, 1155, 680]. Protein [105, 900, 1461, 1100, 1077, 975]. Proteolysis [109]. Proteomic [109]. Proto [378]. Proto-fuzzy [378]. Protocol [748, 1336, 1398, 1015]. Protocols [92, 738]. Prototypes [1412]. Provable [748]. Proving [634, 534, 680]. Proximity [675, 649, 1502]. Proxy [1207, 1390, 143, 16, 714]. Pruning [511]. Pseudo [452, 1551]. Pseudo-random [1551]. Pseudo-Triangulations [452]. Pseudopower [339]. Pseudorandom [352]. PSF [58]. PSPACE [1187]. PSPACE-Hard [1187]. PTIME [142]. Public [436, 1402, 355, 281, 341]. Public-Key [355, 281, 341]. Publications [959]. Pure [1388]. Purely [889]. Purpose [348]. Pushdown [304, 1177, 1283, 127]. Pyglaf [1508]. Pyramidal [602]. Pythagorean [1372, 1159, 1478, 1494].

QBF [1510, 187, 1187, 101]. QSAT [1187]. Quadrance [1260]. Quadratic [725, 458]. Quadraticity [353]. Qualified [877]. Qualitative [1251, 1026, 258]. Quality [97, 1427, 234, 1197]. Quality-aware [1427]. Quantification [492]. Quantified [574, 1202]. Quantifier [635, 1471, 1474]. Quantifiers [49, 21, 1043, 1321, 1322]. Quantisation [456]. Quantitative [557, 866, 1378]. Quantization [313]. Quantizer [14]. Quantum [360, 1273, 122, 848, 79]. Quarantine [1403]. Quartet [100]. Quasi [1190, 1012, 1037, 606, 1356, 1248]. Quasi-Cartan [1190, 1356]. Quasi-discrete [1037]. Quasilattices [1323, 1324]. Quasiorder [750]. Quaternary [787]. Quaternion [1102]. Queries [1452, 38]. Query [925, 880, 655, 883, 706, 1545]. Querying [783, 781]. Quests [766]. Queue [1179]. Quickly [1299]. Quillen [1465]. QVT [340]. QVT-Relations [340].

Rabin [1311]. Radial [865]. Radiation [1023]. Radio [4]. Rainbow [291]. Raising [18]. Random [323, 541, 887, 520, 860, 858, 771, 950, 644, 931, 1121, 542, 1075, 373, 184, 281, 341, 1551]. Random-Grid [931]. Randomized [193]. Randomness [1029, 1549]. Range [1452, 279]. Rank [1310, 1556, 531, 1086]. Rank-width [531]. Ranking [1084]. Rasiowa [1317, 1318]. RASP [145]. Rating [823]. Ratio [1087]. Rational [1417, 391, 1399, 1362, 568]. Rationality [1276, 977]. Ratios [1135]. ray [1023, 602, 861, 1301]. rays [1020]. RCRA [185]. Re [64, 16, 1266]. Re-encryption [16]. Re-Engineering [64]. Reachability [46, 465, 1068, 1541, 701, 1057, 1365, 879, 1072, 529, 1127]. Reaction [756, 1150, 1486, 1487, 698, 633]. Reactive [150, 116]. Read [773]. Read-Once [773]. Real [748, 1298, 646, 729, 1440, 524]. Real-Time [524, 748, 1440]. Real-world [729]. Realisability [503]. Reality [703]. Realtime [1498]. Reason [1351]. Reason-based [1351]. Reasoning [1508, 1497, 1397, 1361, 160, 1399, 585, 1291, 319, 182, 280, 473, 580, 677, 1061]. Receiver [314]. Recognition [822, 1108, 1109, 650, 1095, 730, 183, 189]. Recognizable [10]. Recognized [848]. Recommendation [647]. Recommender [243, 981]. Reconciling [846]. Reconfigurable [514]. Reconstructed [1299]. Reconstructing [1295, 856]. Reconstruction [595, 855, 1298, 662, 859, 601, 1499, 598, 1025, 1020, 882]. Reconstructions [1296]. Recovery [857]. Recurrence [1303]. Recurrent [395]. Recursion [1564, 1217]. Recursive [1183, 194, 1387]. Recut [1133]. Red [1285]. Redefinition [537]. Reduced [437, 36]. Reduced-by-matching [437]. Reducibility [502]. Reducing [556, 307]. Reduction [463, 286, 774, 596, 597, 1017, 439, 1476, 270, 623, 1047, 1564, 1242, 523, 622, 641, 374, 373, 117]. Reductions [1563, 65, 989]. Reducts [644]. Reference [1353, 283]. Refinement [1208, 513]. Reflection [1347]. Reflections [1560]. Region [931, 333]. Region-based [333]. Regions [1230, 1299, 813, 63, 57]. Register [1285]. Registers [1067]. Regression [7, 1197]. Regular [1417, 947, 1274, 1492, 935, 794, 1028, 1079, 1289, 1279, 688, 397, 1215, 256, 1523, 76, 311, 127, 1306, 142, 1059, 404, 799, 251, 863, 982, 978, 979, 1267].

Regularising [766]. Regularity [257]. Regularized [602, 1172]. Regulated [285, 127]. Regulatory [1378, 1379, 120]. Related [1548, 455, 916, 567, 425, 655]. Relating [1068, 897, 1291]. Relation [711, 1046, 794, 629, 231, 532, 1235, 1489, 789, 272, 1162]. Relation-based [1162]. Relational [707, 325, 783, 73, 928, 38, 39, 549]. Relations [540, 605, 1303, 60, 1146, 340, 1281, 750, 181, 400, 560, 378, 1501, 1041, 1477, 934, 57]. Relationships [694, 695]. Relative [1169]. Relatively [591]. Relevant [177]. Reliability [1160]. Reliable [748, 1475, 1386, 841]. Relinking [709, 876]. Remarks [1152, 135, 1472, 519, 352, 356]. Reminiscing [409]. Remote [78]. Reo [719]. Reoptimization [247]. Repeats [255]. Repetition [400]. Repetition-freeness [400]. Replanning [98]. Replication [157]. Representation [194, 757, 1497, 1319, 1320, 1007, 558, 1493, 421, 1404, 1514, 1080, 1325, 1326, 677, 555]. Representations [245, 202, 533]. Represented [67]. Representing [1292]. Reproducing [1130]. Requirements [33]. Research [633, 1425]. Reset [1116]. Resets [996]. Residual [1495, 1411, 567, 1535, 1506]. Residuated [379]. Resistance [349]. ResNet18 [1529]. Resolution [588, 1117, 102, 1535, 1442]. Resource [226, 488, 92, 1004, 673, 744]. Resource-Constrained [744]. Resource-Driven [488]. Resources [464, 585]. Response [756]. Restarting [123, 1178, 825]. Resting [1444]. Restoring [191]. Restricted [108, 1475, 264, 1468]. Restrictions [874, 877, 66, 426]. Result [925]. Results [367, 186, 398, 722, 949, 916, 1070, 1445, 425, 633, 613, 499]. Retrieval [1118, 727, 1189, 1313, 868]. Retrospective [58]. Returning [1079]. Reveals [540]. Reversal [1310, 394, 1086]. Reversibility [1107, 1409]. Reversible [219, 898, 1179, 1282, 260, 942, 477, 1255]. Reversing [1335]. Review [861]. Revised [1243]. Revision [96, 32, 908, 1400]. Revisited [548, 1310, 112, 839]. Revisited* [1108]. Rewriting [1251, 295, 686, 1500, 1078, 308, 1134, 883, 1545]. RFW [1044]. Rhodes [393]. Riccati [1228]. Riesz [1223]. Risk [699, 8]. RLE [1451]. RNA [561]. Robot [300, 301]. Robots [85, 752]. Robust [372, 1118, 561, 78, 865, 445, 739]. Robustness [1064]. Rogozhin [959, 957]. Role [1289, 1560]. Roles [449, 392, 1325, 1326]. Root [978, 586, 568, 569]. Rooted [872, 795]. Roots [1262, 239, 551, 1239]. ROSECON [619]. Rosenau [1226]. Rosser [111]. Rothe [425]. Rough [372, 791, 1309, 209, 627, 1171, 1264, 1033, 210, 632, 214, 273, 629, 1043, 1486, 810, 936, 669, 873, 1153, 362, 212, 1034, 699, 1337, 750, 549, 375, 1389, 1154, 1170, 644, 1168, 1012, 1038, 270, 623, 974, 653, 371, 1039, 561, 1044, 1172, 1467, 13, 311, 26, 606, 1166, 787, 674, 643, 675, 937, 816, 1045, 1501, 752, 1041, 456, 182, 280, 56, 473, 655, 654, 68, 1263, 1373, 234, 1502, 799, 934, 1238, 1035, 691, 789, 715, 1314, 1503, 70, 208, 628, 1173, 272, 215, 1036, 1096, 370, 620, 1480, 271, 211, 630, 369, 792, 1042, 1161, 1162, 240]. Rough-Fuzzy [561, 1044, 1467]. Rough-Granular [699]. Roughness [557, 240]. Route [662]. Routing [453, 729, 442]. Row [1295, 379]. RRWW [890]. RSDS [68, 881]. Rule [1149, 462, 1026, 32, 638, 790, 668, 1006, 54, 402, 1059, 1166, 734, 1197]. Rule-based [1006, 1166]. Rulers [749]. Rules [2, 1491, 636, 694, 1213, 1203, 1363, 324, 1434, 258, 1170, 1168, 980, 1092, 1167, 921, 1165, 536, 537, 1527, 1312, 641, 640, 842, 1169]. Run [1454]. running [969]. Runs [125, 426, 1248]. Runtime [592, 285].

S [890]. S-kR-RRWW-automata [890]. S4 [1401]. S5 [576]. SAC [1497]. Safe [721, 299, 701, 1057, 232]. Safeness [62]. Safety [744]. Salesman [1431, 896]. SAM [666]. Samples [415]. Sampling [192, 640, 1114]. Sand [239]. Sand-glass [239]. Sandpile [1066, 366]. SAT [764, 484, 1488, 470, 103, 89, 1062, 496]. SAT-Based [484, 470, 103, 1062]. SAT-solvers [1488]. Satellite [79]. Satisfaction [31, 217, 218]. Satisfiability [574, 1055, 904, 531, 191, 1447, 151, 1504, 1202]. Saving [201]. sawtooth [805]. SBVR [580]. Scalable [682, 468, 3]. Scalar [716]. Scale [763, 733, 1386]. Scaled [653]. Scattered [47, 1233, 41]. Scenarios [918, 110, 484]. Scheduler [1125]. Scheduling [200, 967, 1511, 521, 973, 1401]. Schema [566, 177]. Schemata [588]. Scheme [1207, 1390, 477, 354, 143, 714, 640, 16, 739, 80, 805]. Schemes [607, 405, 314, 356, 223]. Schnorr [1015]. Schrödinger [1495, 1103]. Schützenberger [306]. Science [19, 1560]. Scientific [683]. Scientist [120]. SCIFF [571]. Scope [1227]. Score [1478]. Scott [628]. scrambling [1418]. Search [965, 1304, 188, 747, 324, 618, 190, 193, 646, 35, 925, 1237, 1401, 911, 1131]. Searching [68, 1248]. Second [49, 435, 350, 289, 1265, 175, 1561]. Second-Order [49, 435, 175, 1561, 350, 289, 1265]. Secret [931, 356]. Secure [845, 281, 341]. Security [83, 86, 491, 1056, 738, 314]. see [703]. Seed [749]. Seeking [1305]. Segmental [795]. Segmentation [13, 1440, 690]. Seismic [1113]. Selection [1498, 1007, 11, 644, 77, 909, 1445]. Selective [850]. Selectors [1289]. Self [449, 411, 1054, 283, 284, 696, 697, 1286, 7, 968]. Self-Adaptation [7]. Self-Assembly [1054, 1286]. Self-avoiding [411]. Self-Correcting [968]. Self-Indexed [284]. Self-Modifying [696]. Self-organized [449]. Self-Organizing [697]. Self-Reference [283]. Semantic [34, 454, 955, 1504, 793, 211]. Semantical [1002]. Semantics [1354, 295, 463, 1206, 1351, 1396, 760, 1483, 310, 1043, 435, 220, 1471, 340, 1253, 1538, 503, 61, 112, 1158, 995, 177, 1293, 1256, 235, 1377]. Semantics* [1004]. Semelparous [115]. Semi [1140, 1378, 614, 950, 727, 402, 838, 1473]. Semi-conditional [950]. Semi-Deterministic [1140]. Semi-quantitative [1378]. Semi-supervised [727]. Semi-Thue [402]. Semi-totalistic [614]. Semi-Uniformity [838]. Semidefinite [1232, 1265]. Semigroup [387]. Semilinear [945, 533]. Sense [454]. Sensing [78]. Sensitive [926]. Sensitivity [1460]. Sensor [1386]. Sentences [1471, 726]. Separability [327]. Separated [1274]. Separating [40]. Separation [367, 944, 1247]. Sequence [1192, 1351, 1258, 254]. Sequences [1453, 481, 966, 996, 1017, 397, 1454, 1538, 1119, 421, 352, 730]. Sequent [1204]. Sequential [31, 973, 964, 1128]. Sequents [1401]. Seriation [1133]. Series [1495, 306, 423]. server [228]. Service [298, 1186, 1339, 206]. Services [924, 4, 331, 673, 285, 484]. Set [1143, 1183, 1145, 1202, 209, 1264, 318, 1004, 486, 801, 1043, 873, 362, 1034, 420, 1476, 644, 270, 974, 653, 1039, 728, 1166, 787, 674, 643, 321, 562, 1235, 523, 511, 332, 42, 1468, 1292, 1188, 473, 654, 68, 1373, 1502, 587, 1035, 789, 715, 1490, 215, 1243, 211, 376, 630, 369, 792, 1503, 1049, 649]. Set-based [644]. Set-pairs [420]. Set-theoretic [1202, 376]. Set-theoretical [42]. Sets [1417, 1074, 372, 246, 821, 20, 627, 1171, 1033, 790, 210, 1005, 214, 1355, 273, 1486, 1371, 1466, 466, 936, 669, 1153, 212, 397, 945, 1541, 699, 1337, 750, 141, 1389, 1154, 1170, 1168, 737, 1012, 401, 1038, 6, 623, 371, 932, 606, 1014, 624, 675, 1071, 1045, 1099, 383, 1263, 265, 934, 334, 742, 441, 691, 1314, 1516, 70, 208, 628, 272, 370, 620, 1042, 1161, 1162, 240]. Sets-based [1038]. Settings [1173, 1197]. Several [125, 1557]. Sevilla [839]. SHA [1550]. SHA-3 [1550]. Shallow [1223]. Shannon [438]. Shared [464, 151]. Sharing [931, 187, 356]. Shedding [106]. Sheet [1227]. Shelah [1370]. Sheng [478]. Shift [1143, 277, 1288]. Shift-free [1288]. Shifting [18]. Shock [756]. Shop [1511]. SHOQ [877]. Short [688, 777]. Shorter [1101]. Shortest [1369, 1415]. Shows [1187]. Shuffle [390, 672, 702]. Side [729]. Sided [935, 604, 1266]. Sieve [1311]. Sight [1349]. Sight-limited [1349]. Sign [1231]. Signal [446, 4, 313]. Signaling [1382]. Signalling [1375]. Signals [1252, 1531]. Signature [1207, 1390, 607, 354, 143, 314, 714]. Signatures [1015, 1101]. Signcryption [314, 184]. Signed [1011]. Significance [909]. Sikorski [1317, 1318]. Silver [1433]. Similarities [1241]. Similarity [1026, 810, 73, 467, 1305, 647, 646, 1237, 1246, 166, 1313, 911]. Similarity-Based [73]. Simple [1088, 976, 227, 463, 213, 1544, 1557, 1449, 158]. Simpler [1101]. Simplification [1436]. Simplified [1479]. Simplifying [437]. Simsun [418]. Simulated [1156, 559]. Simulating [902, 1254]. Simulation [1382, 201, 1444, 811, 1286, 698, 901, 1383, 594]. Simulation-based [1382]. Simulations [609, 1430]. Simulator [980]. Simultaneous [909]. Single [257]. Singular [1227, 1228]. Situations [9]. Six [1114]. Six-dimensional [1114]. Size [1192, 367, 303, 136]. Skeleton [601, 730]. Skew [1513, 396]. Skolem [1553]. SLGAD [102]. Small [956, 1120, 942, 263, 1099, 1386, 957, 1198]. Smart [871]. SMC4AC [1240]. Smirnov [415]. Smooth [1133]. Smoothness [1066]. SMT [878, 1156, 879, 1248]. SMT-Based [879, 1248]. Sniffing [1385]. SOAP [924]. SoBT [1044]. SoBT-RFW [1044]. Soccer [752]. Social [708, 924, 1426, 733, 1209, 657]. Soft [372, 282, 845, 936, 564, 1014, 1246, 1235, 1243]. Software [1023, 668, 459, 753, 301]. Sokoban [581]. Solid [857]. Soliton [1103]. Solution [1495, 1226, 320, 279, 222, 903, 50]. Solutions [1130, 1496, 247, 1231, 797, 1349, 100, 802, 587, 1103]. Solvability [1224, 46, 113]. Solve [1231, 729]. solvers [1488]. Solving [599, 31, 1184, 147, 1510, 317, 185, 1511, 896, 554, 968, 187, 1237, 826, 1488, 239, 103, 867, 617, 357, 581]. Some [411, 367, 389, 574, 209, 627, 1171, 1137, 350, 1372, 135, 188, 1472, 519, 308, 1459, 838, 66, 215]. Sorites [1337]. Sorted [1493]. Sorting [384]. Sound [666, 183]. Soundness [745, 744, 995, 499]. Source [1025]. Sources [1025]. Space [583, 1495, 759, 196, 201, 1260, 1112, 1452, 521, 343, 892, 837, 610, 374, 911, 1480, 360, 1114]. Spaced [749]. Spaces [379, 361, 635, 1355, 193, 750, 1225, 25, 1012, 963, 1037, 1039, 949, 564, 827, 675, 649, 984, 56, 524, 1173, 639]. Spanning [999]. Spans [430]. Sparse [1236]. Sparsity [1301]. Spatial [444]. Spatially [954]. Spatiotemporal [4]. Spearman [438]. Special [1250, 1, 348, 1497, 887, 386, 1089]. Special-Purpose [348]. Specialising [1253]. Specialization [465, 577]. Specialization-based [577]. Specialized [1150]. Species [1275, 115]. Species-Driven [1275]. Specific [1458]. Specification [872, 312]. Specified [308, 619]. Spectral [584, 978, 979, 568, 1093, 1094]. Spectrometry [109]. Sphere [1260]. Spike [840]. Spikes [1312, 841]. Spiking [1430, 290, 898, 899, 901, 292, 262, 840, 1312, 1198, 841, 1210, 842]. Splicing [266]. Square [776, 586, 630]. Square-free [776]. Squares [1085]. SSWSNs [1386]. Stable [246, 108, 1105, 57]. Stack [1283]. stage [1535]. Stages [713]. Standard [1402, 1207, 1390, 223]. Standpoint [208]. Star [390, 394, 221]. Star-Free [390]. State [583, 1384, 305, 1444, 1079, 465, 394, 688, 221, 1225, 259, 41, 820, 495, 706, 1200, 614]. State-dependent [1225]. Statecharts [1158]. Stateless [250, 398]. Statements [1157]. States [529, 1541, 1521]. Static [99]. Statistical [324, 1341]. Statistics [410, 739]. Steering [698]. Steganographic [440]. Steganography [219, 18]. Stenosis [1151]. Step [1272, 1458, 654, 804, 112]. Step-by-step [1272]. Stephen [236]. Steps [919]. Stereo [968]. Stereo-matching [968]. Stimulation [1442]. Stochastic [609, 1382, 444, 1206, 1388, 759, 1381, 517, 1383, 999, 841]. Stopping [482]. Storage [468, 1462]. Strand [1432, 1438]. Strands [1436]. Strategies [719, 96, 1388, 554, 1365, 951]. Strategy [200, 1006, 556, 1345, 1312, 1449]. Strategy-proof [1345]. Stream [353, 1387]. Streaming [1468]. String [1353, 71, 966, 899, 308, 1128]. String-Rewriting [308]. Strings [976, 833, 1455, 767]. Strip [103]. Strong [1558, 1013, 1086, 506]. Strongly [1421, 1180, 501]. Structural [994, 874, 166]. Structure [1011, 709, 1452, 362, 1367, 380, 218, 900, 951, 15, 1162]. Structured [1106, 908]. Structures [844, 360, 1297, 1033, 60, 297, 632, 53, 1277, 1281, 433, 814, 656, 866, 378, 1490, 215, 547]. Structuring [1194]. Stubborn [334]. Studies [1505]. Study [1484, 665, 1184, 584, 1442, 986, 1153, 35, 1040, 1306, 1022, 426, 101, 594, 546]. Study* [1109]. Studying [1487, 1423]. Stupp [1370]. Sturmian [389, 1119]. Style [434, 1256]. Subclasses [1409]. Subformula [169]. Subgraphs [291]. Subgroup [396, 1122]. Subintervals [751]. Sublinear [892]. Submodules [392]. SubRule [640]. Subsequence [712, 45]. Subset [1144, 1422]. Subsets [265]. Subshifts [615, 778]. Subspaces [6]. Substitutable [1139].

Substitution [476]. Substring [1454]. Substrings [255, 41]. Substructural [721]. Subsumption [1325, 1326]. Subtree [287]. Subtrees [460]. Subtyping [505]. Subwords [404]. Such [351]. Suffices [1474]. Sufficiently [562]. Suffixes [404]. Suitable [1440]. Sum [850, 1013, 1312]. Sums [1295]. Super [1442, 1535, 1114]. Super-resolution [1442]. Superposition [132, 1334]. Superstrings [41]. Supervised [1026, 1172, 165, 727]. Support [1087, 580, 183]. Supported [1170]. Supporting [558]. Supports [888, 38]. Sure [267]. Surveillance [662]. Survey [398, 945, 722, 309, 1443, 902, 1035, 1220]. Survival [1197]. Suslin [1465]. Swarm [822, 973, 1512, 1449, 159]. Sweep [743]. Sweep-Line [743]. Switching [396, 1018]. Syllabic [762]. Syllogisms [357]. Syllogistic [574]. SyLVaaS [1186]. Symbolic [1240, 277, 797, 1093, 1094, 1527, 1255, 1048, 1103]. Symbols [1414, 942]. Symmetric [1190, 388, 1131, 403]. Symport [1434, 980]. Symport/Antiport [1434, 980]. Symposium [1497]. Synapses [1312, 1198, 1210, 842]. Synchronising [489]. Synchronizable [513]. Synchronization [1422, 1437, 1357, 17]. Synchronizing [775, 1421]. Synchronous [200, 794, 696]. Synchrotron [1023, 1025]. Syntactic [454, 401]. Syntactocentric [173]. Synthesis [134, 1271, 719, 1273, 993, 60, 297, 1537, 516, 814, 42, 619, 1539]. Synthesize [331]. Synthesizing [486]. System [9, 682, 1224, 571, 241, 756, 666, 149, 1231, 822, 632, 558, 923, 77, 974, 1186, 78, 1339, 753, 146, 579, 1437, 68, 881, 705, 934, 69, 620, 1049]. SystemC [180]. Systems [243, 1200, 244, 831, 956, 1149, 1250, 30, 1496, 1125, 757, 83, 1150, 720, 1201, 1273, 993, 1500, 1046, 1430, 529, 981, 834, 196, 1431, 954, 1285, 228, 662, 46, 1194, 489, 1486, 431, 919, 465, 889, 927, 849, 1487, 874, 1286, 254, 290, 950, 1287, 1541, 700, 1434, 1521, 897, 1225, 978, 258, 493, 722, 898, 899, 63, 765, 973, 402, 1350, 837, 308, 65, 901, 980, 902, 1069, 459, 838, 766, 292, 222, 903, 66, 262, 674, 643, 633, 840, 1254, 470, 263, 944, 514, 1504, 997, 998, 1405, 734, 1077, 1312, 619]. Systems [1263, 1373, 234, 1134, 301, 1257, 1540, 266, 894, 648, 546, 17, 224, 235, 628, 679, 1062, 1198, 1081, 841, 267, 1210, 842, 1128]. Szilard [548].

T0L [1285]. Table [1558, 780]. Tableau [1204, 877]. Tableaux [1155]. Tabled [1184, 581]. Tables [487, 665, 363, 38, 373]. Tabu [1304]. Tabular [457]. Tackling [1512]. tag [46]. Taking [69]. Tally [952]. Tampering [166]. Tanimoto [646]. Tape [942]. Target [130]. Targeting [5]. Task [399, 685]. Tasks [1528, 973]. Tautologies [680]. Tavaxy [682]. TDOA [862]. Team [175]. Teams [565]. Technique [584, 826, 1114]. Techniques [1184, 85, 286, 197, 813, 333]. Technology [99]. Temperature [1286]. Temporal [1053, 1485, 296, 363, 152, 163, 1520, 65, 232, 560, 674]. Temporal-Epistemic [296, 65]. Ten [1554]. Tense [905]. Tensor [1527]. Term [1371, 1348, 1545]. Terminating [1204, 423]. Termination [504, 206]. Terms [1395, 389, 1563, 501]. Tero [386]. Test [1003, 415, 764, 459, 266, 882, 1387]. Testing [135, 670, 725, 1325, 1326]. Testing* [1075]. Tests [487, 353]. Tetravalent [863]. Text [1026]. Textual [467, 189]. Texture [346]. Textures [214, 1355]. Their [755, 179, 1085, 1096, 1102, 1295, 1193, 53, 1466, 986, 907, 1040, 1014, 1080, 568, 1325, 1326, 1030, 1020, 1314, 1332, 1239, 989]. theirs [42]. Them [668]. Theorem [194, 393, 1465, 306, 74, 421, 1553, 1522]. Theorems [755, 634, 1080, 550, 1325, 1326]. Theoretic [860, 212, 623, 1123, 1100, 369, 1202, 1038, 974, 371, 507, 370, 620, 376]. Theoretical [1190, 1001, 784, 42]. Theories [846, 653, 151]. Theory [1027, 172, 209, 338, 721, 1264, 952, 51, 198, 1475, 873, 1153, 1211, 1564, 1459, 932, 177, 1166, 787, 1209, 205, 1124, 1368, 1502, 57, 1035, 373, 1096, 630, 369, 657, 518, 624]. Therapeutic [659]. there [587]. these [587]. Thickening [22]. Thin [1227]. Thinking [1209]. Thorn [1411]. Thorough [169]. Threading [1100]. Three [446, 418, 638, 786, 1299, 1038, 330, 47, 344, 345, 840, 997, 998, 984, 1461, 370, 1480]. Three-way [786, 1038, 984, 370, 1480]. Threshold [320, 645, 143, 356]. Thresholds [1116]. Thue [402]. Tiered [1191]. Tight [1284, 1117, 1247]. Tightly [707]. Tightly-Coupled [707]. Tiles [807, 1054, 248, 1028, 1119]. Tiling [1112, 1098]. Tilings [614, 949, 424]. Time [1064, 1495, 82, 1397, 1216, 196, 201, 1483, 152, 521, 780, 876, 61, 1158, 995, 1069, 607, 879, 1383, 524, 158, 1196, 178, 1308, 171, 748, 163, 1440]. Time-Aware [1483]. Time-fractional [1495]. Time-independent [82]. Timed [735, 1065, 761, 809, 521, 1068, 700, 814, 1158, 995, 1340, 494, 90, 820, 884, 1248]. Timed-Arc [995]. Timestamps [1336]. Timetabling [317]. Tinput [1283]. Tinput-Driven [1283]. Tissue [222, 944, 1077]. Tiuryn [105]. Toeplitz [80]. Together [165]. Tokenflows [918]. Tokens [17]. Tolerance [1457, 1235, 1041, 789]. Tolerances [929]. Tolerant [1252]. Tomographic [1296, 857, 860]. Tomography [1019, 599, 855, 1023, 1298, 1299, 596, 597, 1017, 1300, 1098, 598, 853, 1440, 594, 1024, 1113, 1114, 1301, 602]. Tool [658, 331, 619]. Toolbox [839]. Tools [1023, 863]. Toolset [298]. Top [309]. Top-down [309]. Topic [1367]. Topics [633]. Topological [518, 455, 657, 1162]. Topologies [887]. TOPSIS [1243]. Tori [239]. Toroidal [569]. totalistic [614]. Totally [1484]. Tour [710]. Tourism [241, 146]. Toxicity [1380]. Toy [900]. Trace [391, 306, 738]. Traces [277, 515, 330, 1217, 67]. Track [1497]. Tracking [557]. Tractability [835]. Tractable [1059]. Trade [196, 522]. Trade-off [522]. Trade-Offs [196]. Traffic [399, 472]. Training [1435]. Trajectories [1518]. Trance [321]. Transactions [969]. Transducer [1140]. Transducers [286, 278, 1177, 612, 309, 958]. Transduction [446]. Transfer [110, 1529]. Transform [860, 1021, 865, 1534, 1313, 1220]. Transformation [431, 634, 433, 1458, 651, 432, 997, 998, 1301, 804]. Transformations [1427, 340, 126, 238, 1103]. Transition [1273, 60, 297, 774, 1194, 252, 700, 897, 63, 710, 1489, 224]. Transitions [1335, 857, 1058]. Translation [463, 309, 732, 525, 496]. Translucent [1542]. Transmission [15]. Transversal [736]. Traveling [896]. Travelling [1431]. Travels [1339]. Traversing [287]. Treasure [649]. Treatment [169]. Tree [830, 707, 71, 1052, 1485, 296, 1144, 1274, 97, 1510, 1137, 888, 1441, 278, 439, 737, 309, 24, 1306, 139, 238, 999, 15]. Tree-Child [830]. Tree-walking [1274]. Tree-width [24]. Trees [1484, 791, 665, 49, 695, 288, 1253, 795, 422, 964, 1160, 550, 110]. Treewidth [101]. Triangular [55, 1098, 1369, 1415]. Triangulations [452]. Trick [1522]. Tries [36]. Trim [1421]. Trio [728]. Tripartite [222]. TripICS [1339]. Trips [1339]. Trivium [353, 357]. Trivium-based [357]. Trophic [1376]. Trove [649]. True [969, 1196]. Truth [1470, 1558, 1076, 182, 280]. Trying [1342]. Tube [266]. Tubes [239, 266]. Tumor [1044]. Turbine [685]. Turing [892, 543, 942, 264, 448, 1247, 957, 1266]. Turing-Inspired [448]. Tutorial [93]. Twelve [1187]. Two [10, 1274, 1027, 935, 862, 305, 365, 1028, 986, 1223, 1493, 23, 1178, 259, 260, 613, 1468, 1085, 1535, 655, 103, 266, 223, 1438, 211, 1239]. Two-center [1468]. Two-Dimensional [1223, 1178, 10, 365, 103]. Two-domain [1438]. Two-pebble [23]. Two-Sided [935]. Two-Sorted [1493]. Two-stage [1535]. Two-State-Automata [305]. Two-Way [260, 613, 1028, 259]. Type [1226, 666, 721, 51, 1556, 584, 849, 350, 1564, 1109, 1356, 615, 118, 863]. Type-Safe [721]. Typed [666, 433, 93, 94]. Types [1496, 666, 92, 140, 1555, 986, 1215, 503, 504, 505, 506, 507, 1557, 508, 1161, 499]. Typical [249]. Typicality [1400]. Typicality-based [1400].

U [455]. Ultimately [401]. UML [381]. Unambiguous [10, 343, 1414]. Unary [122, 303, 533]. Uncertain [802]. Uncertainties [1453]. Uncertainty [213, 699, 549, 1040, 1503]. Unconventional [836]. Uncovering [323]. Undecidability [364, 1556, 751, 1474]. Undecidable [1274, 256]. Underlying [454]. Understand [1342]. Understanding [763]. Unfoldings [671]. Unified [1126, 1428, 495, 706, 869]. Uniform [222, 903]. Uniformity [838]. Unifying [653, 469]. Union [1555, 394, 221, 259, 506, 508, 499, 1539]. Union/Find [1539]. Unions [151]. Unique [255, 256]. Uniqueness [1299]. Unit [458]. Univariate [141]. Universal [956, 985, 1519, 614, 942, 840, 706, 957, 1198, 1479]. Universality [946, 1287, 949, 1254, 404, 266, 1081]. Universe [1501]. Universes [215]. Unknown [1378, 1185]. Unranked [888, 238]. Unreliable [134]. Unsatisfiability [218]. Unstable [246, 57]. Unsupervised [454, 467, 727, 557, 165, 1138]. Untrusted [1185]. Update [84, 468]. Uplinks [117]. Upper [1057, 1460, 733, 1503]. Usage [92]. Use [381, 162]. Used [213, 638]. Useful [101]. User [1427]. Using [583, 323, 599, 296, 32, 918, 348, 1378, 92, 996, 486, 5, 635, 452, 749, 863, 747, 780, 987, 646, 1347, 909, 1022, 840, 964, 661, 471, 79, 713, 374, 183, 1036, 803, 18, 369, 357, 746, 1251, 1130, 1271, 1203, 1412, 1381, 1367, 923, 1544, 671, 1476, 1058, 321, 1341, 116, 460, 868, 739, 1301]. Utility [1345]. Utilizing [657].

Valid [1105]. Validation [592]. Validity [1460]. Valuations [172]. Value [1048, 18]. Valued [487, 665, 33, 646, 176, 1041, 789, 337, 1466, 1038, 532, 1288, 866, 1014, 1246, 1049]. Values [1378, 598, 545]. Variable [96, 1304, 14, 1263, 1103]. Variable-coefficient [1103]. Variables [1396, 503, 1544]. Variance [482]. Variants [956, 108, 308]. Variation [410, 792]. Variational [826, 1440]. Variations [1188]. Varieties [1195, 311]. Various [1539]. Vector [529, 1541, 926, 1420, 622, 183, 1551]. Vector-based [622]. Vectors [1336, 996, 646, 459]. Vehicle [453, 729]. Verification [583, 1200, 1407, 918, 84, 186, 1006, 1203, 1398, 489, 577, 61, 836, 1186, 99, 434, 619]. Verify [92]. Verifying [1052, 1030]. Versatility [310]. Version [54, 770, 881]. Versus [124, 454, 1168, 878, 1447]. Vertex [291, 1241]. Vertical [601]. Vertices [1120]. Vessels [1151]. via [1508, 1509, 31, 597, 11, 653, 673, 114, 1076, 1527, 884, 373, 547]. Video [662, 1534]. View [992, 1004, 809, 198, 816, 90]. Violation [617]. Viral [444]. Virasoro [413]. Virus [1403]. Viscous [1227]. Visibly [1177]. Visits [416]. Visual [931, 675, 661, 868]. Visualizing [1532]. Vocabulary [1474]. Volume [293, 315, 336, 358, 385, 407, 428, 450, 474, 497, 509, 526, 552, 582, 603, 625, 663, 692, 717, 740, 768, 800, 828, 851, 885, 912, 939, 970, 990, 1010, 1031, 1050, 1073, 1129, 1141, 1163, 1181, 1199, 1218, 1249, 1269, 1302, 1329, 1330, 602, 861]. Volumes [854]. Voting [1201]. vs [1516].

w.r.t [88]. Walk [411]. walking [1274]. Wang [807, 1119]. Warehouses [781]. Water [1223]. Watermarking [865, 690, 739, 1357]. Watson [1285, 250, 125]. Wavelet [1223, 1044]. Waves [1223]. Way [213, 892, 260, 613, 786, 1028, 259, 1038, 984, 370, 1480]. WDBA [1478]. Weak [4, 827, 206, 639]. Weaker [838]. Weakly [420, 1241]. Web [721, 5, 298, 285, 484, 1339]. Weber [195]. Website [5]. Weight [575, 452, 75, 164, 136, 1449, 1243]. Weighted [904, 1537, 888, 1371, 278, 1278, 1409, 309, 1123, 645, 1160, 1423, 1062, 1248, 842]. Weighting [1090]. Weights [139, 788]. Wendy [331]. Weyl [551]. WFAs [216]. Where [1045]. Whether [794]. which [587]. Wide [590, 1232, 1265]. Wide-diameter [590]. Widely [213]. width [531, 24]. Wiener [1160]. Wilf [74]. Window [1295]. Windows [98, 876]. Winning [1365]. Wireless [1386]. Within [441, 10, 1227, 1228, 813, 1499, 532, 1288]. without [1451, 1228, 1262, 265, 17, 184, 281, 341]. Witnesses [1311]. Wnt [1375]. Wolf [1261]. Word [1500, 1176, 454, 257, 825, 1423, 1525]. Word-Order [825]. Words [412, 772, 856, 389, 847, 386, 771, 776, 406]. Work [297, 829]. Workflow [682, 328, 671, 745, 64, 744, 995, 1030, 1307, 513]. Workflows [683]. Working [1521, 903, 1312]. World [560, 729]. Worst [479]. Worst-Case [479].

X [1023, 602, 861, 1020, 1301]. X-ray [1023, 602, 861, 1301]. X-rays [1020]. x86 [1255]. XML [1126]. XOR [836].

Yawn [273]. Years [1554, 1187]. Yoneda [114]. Yu [478]. Yurii [959, 957].

Zadeh [362]. Zero [962, 646, 155]. Zero-expressible [962].

References

Anonymous:2010:SII

[1] Anonymous. Special issue on intelligent data analysis in granular computing. Fundamenta Informaticae, 98(1):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alcala-Fdez:2010:AEG

[2] Jesus Alcala-Fdez, Nicolo Flugy-Pape, Andrea Bonarini, and Francisco Herrera. Analysis of the effectiveness of the genetic algorithms based on extraction of association rules. Fundamenta Informaticae, 98(1):1–14, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lu:2010:SCM

[3] Kun-Che Lu and Don-Lin Yang. Scalable clustering for mining localcorrelated clusters in high dimensions and large datasets. Fundamenta Informaticae, 98(1):15–32, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lee:2010:DRS

[4] Wei-Hsun Lee, Shian-Shyong Tseng, Ching-Hung Wang, and Shen-Lung Tung. Discovering the radio signal coverage hole and weak coverage area in mobile network by spatiotemporal data mining on location-based services. Fundamenta Informaticae, 98(1):33–47, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeBock:2010:PWA

[5] Koen W. De Bock and Dirk Van den Poel. Predicting Website audience demographics for Web advertising targeting using multi-Website clickstream data. Fundamenta Informaticae, 98(1):49–70, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leng:2010:MOC

[6] Jinsong Leng and Tzung-Pei Hong. Mining outliers in correlated subspaces for high dimensional data sets. Fundamenta Informaticae, 98(1): 71–86, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kramer:2010:CMS

[7] Oliver Kramer. Covariance matrix self-adaptation and kernel regression — perspectives of evolutionary optimization in kernel machines. Fundamenta Informaticae, 98(1):87–106, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2010:RMM

[8] Shusaku Tsumoto and Shoji Hirano. Risk mining in medicine: Application of data mining to medical risk management. Fundamenta Informaticae, 98(1):107–121, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abe:2010:CED

[9] Akinori Abe, Yukio Ohsawa, Hiromi Itoh Ozaku, Kaoru Sagara, Noriaki Kuwahara, and Kiyoshi Kogure. Communication error determination system for multi-layered or chained situations. Fundamenta Informaticae, 98(1):123–142, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anselmo:2010:DUF

[10] Marcella Anselmo, Dora Giammarresi, and Maria Madonia. Deterministic and unambiguous families within recognizable two-dimensional languages. Fundamenta Informaticae, 98(2–3):143–166, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hu:2010:FSM

[11] Qinghua Hu, Pengfei Zhu, Jinfu Liu, Yongbin Yang, and Daren Yu. Feature selection via maximizing fuzzy dependency. Fundamenta Informaticae, 98(2–3):167–181, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalyanasundaram:2010:LBF

[12] Bala Kalyanasundaram and Mahe Velauthapillai. Learning behaviors of functions. Fundamenta Informaticae, 98(2–3):183–198, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Malyszko:2010:ARE

[13] Dariusz Małyszko and Jarosław Stepaniuk. Adaptive rough entropy clustering algorithms in image segmentation. Fundamenta Informaticae, 98 (2–3):199–231, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peric:2010:DHQ

[14] Zoran H. Perić, Milan R. Dinčić, and Marko D. Petković. Design of a hybrid quantizer with variable length code. Fundamenta Informaticae, 98 (2–3):233–256, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsai:2010:TSB

[15] Piyu Tsai. Tree structure based data hiding for progressive transmission images. Fundamenta Informaticae, 98(2–3):257–275, April 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2010:IIB

[16] Xu An Wang and Xiaoyuan Yang. On the insecurity of an identity based proxy re-encryption scheme. Fundamenta Informaticae, 98(2–3): 277–281, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wehler:2010:FCP

[17] Joachim Wehler. Free-choice Petri nets without frozen tokens, and bipolar synchronization systems. Fundamenta Informaticae, 98(2–3):283–320, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2010:CRS

[18] Cheng-Hsing Yang, Shiuh-Jeng Wang, and Chi-Yao Weng. Capacityraising steganography using multi-pixel differencing and pixel-value shifting operations. Fundamenta Informaticae, 98(2–3):321–336, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2010:BLC

[19] Anonymous. Bridging logic and computer science: to Johann A. Makowsky for his 60th birthday. Fundamenta Informaticae, 98(4):??, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bojanczyk:2010:BCM

[20] Mikołaj Bojańczyk, Damian Niwiński, Alexander Rabinovich, Adam Radziwończyk-Syta, and Michał Skrzypczak. On the Borel complexity of MSO definable sets of branches. Fundamenta Informaticae, 98(4): 337–349, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dawar:2010:PAA

[21] Anuj Dawar and Erich Grädel. Properties of almost all graphs and generalized quantifiers. Fundamenta Informaticae, 98(4):351–372, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoffmann:2010:MGE

[22] Christian Hoffmann. A most general edge elimination polynomial — thickening of edges. Fundamenta Informaticae, 98(4):373–378, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaminski:2010:NTP

[23] Michael Kaminski and Tony Tan. A note on two-pebble automata over infinite alphabets. Fundamenta Informaticae, 98(4):379–390, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meer:2010:TWA

[24] Klaus Meer. Tree-width in algebraic complexity. Fundamenta Informaticae, 98(4):391–409, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khare:2010:GDA

[25] Mona Khare and Surabhi Tiwari. Grill determined L-approach merotopological spaces. Fundamenta Informaticae, 99(1):1–12, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maulik:2010:ERP

[26] Ujjwal Maulik and Anasua Sarkar. Evolutionary rough parallel multiobjective optimization algorithm. Fundamenta Informaticae, 99(1):13– 27, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Velasco:2010:MGG

[27] Pedro Pablo Pérez Velasco and Juan de Lara. Matrix graph grammars with application conditions. Fundamenta Informaticae, 99(1):29–62, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orejas:2010:GACa

[28] Fernando Orejas, Hartmut Ehrig, Markus Klein, Julia Padberg, Elvira Pino, and Sonia Pérez. A generic approach to connector architectures. Part I: The general framework. Fundamenta Informaticae, 99(1):63–93, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orejas:2010:GACb

[29] Fernando Orejas, Hartmut Ehrig, Markus Klein, Julia Padberg, Elvira Pino, and Sonia Pérez. A generic approach to connector architectures. Part II: Instantiation to Petri nets and CSP. Fundamenta Informaticae, 99(1):95–124, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2010:MIS

[30] Anonymous. Methodologies for intelligent systems. Fundamenta Informaticae, 99(2):??, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bartak:2010:SSP

[31] Roman Barták and Daniel Toropila. Solving sequential planning problems via constraint satisfaction. Fundamenta Informaticae, 99(2):125– 145, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benferhat:2010:FIB

[32] Salem Benferhat, Didier Dubois, Henri Prade, and Mary-Anne AnneWilliams. A framework for iterated belief revision using possibilistic counterparts to Jeffrey’s rule. Fundamenta Informaticae, 99(2):147–168, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Djouadi:2010:IVF

[33] Yassine Djouadi and Henri Prade. Interval-valued fuzzy Galois connections: Algebraic requirements and concept lattice construction. Fundamenta Informaticae, 99(2):169–186, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esposito:2010:FCS

[34] Floriana Esposito, Claudia d’Amato, and Nicola Fanizzi. Fuzzy clustering for semantic knowledge bases. Fundamenta Informaticae, 99(2): 187–205, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2010:AOS

[35] Jiming Liu, Chao Gao, and Ning Zhong. Autonomy-oriented search in dynamic community networks: A case study in decentralized network immunization. Fundamenta Informaticae, 99(2):207–226, April 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murray:2010:IPI

[36] Neil V. Murray and Erik Rosenthal. Identifying prime implicate branches in reduced implicate tries. Fundamenta Informaticae, 99(2):227–243, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Georgescu:2010:AIF

[37] Irina Georgescu. Arrow index of a fuzzy choice function. Fundamenta Informaticae, 99(3):245–261, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jen:2010:CSC

[38] Tao-Yuan Jen, Dominique Laurent, and Nicolas Spyratos. Computing supports of conjunctive queries on relational tables with functional dependencies. Fundamenta Informaticae, 99(3):263–292, August 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2010:RCR

[39] Feng Jiang, Yuefei Sui, and Cungen Cao. Relational contexts and relational concepts. Fundamenta Informaticae, 99(3):293–314, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Majumder:2010:SMC

[40] Subhashis Majumder, Subhas C. Nandy, and Bhargab B. Bhattacharya. Separating multi-color points on a plane with fewest axis-parallel lines. Fundamenta Informaticae, 99(3):315–324, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okhotin:2010:SCS

[41] Alexander Okhotin. On the state complexity of scattered substrings and superstrings. Fundamenta Informaticae, 99(3):325–338, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rytsar:2010:STC

[42] Bohdan Rytsar, Piotr Romanowski, and Adrian Shvay. Set-theoretical constructions of Boolean functions and theirs applications in logic synthesis. Fundamenta Informaticae, 99(3):339–354, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2010:IGP

[43] Jan A. Bergstra and Cornelis A. Middelburg. An interface group for process components. Fundamenta Informaticae, 99(4):355–382, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bloom:2010:AO

[44] Stephen L. Bloom and Zoltan Ésik. Algebraic ordinals. Fundamenta Informaticae, 99(4):383–407, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deorowicz:2010:BPA

[45] Sebastian Deorowicz. Bit-parallel algorithm for the constrained longest common subsequence problem. Fundamenta Informaticae, 99(4):409– 433, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeMol:2010:SHR

[46] Liesbeth De Mol. Solvability of the halting and reachability problem for binary 2-tag systems. Fundamenta Informaticae, 99(4):435–471, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Masopust:2010:BNP

[47] Tomáš Masopust. Bounded number of parallel productions in scattered context grammars with three nonterminals. Fundamenta Informaticae, 99(4):473–480, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2010:P

[48] Anonymous. Preface. Fundamenta Informaticae, 100(1–4):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barany:2010:ECQ

[49] Vince Bárány, Łukasz Kaiser, and Alex Rabinovich. Expressing cardinality quantifiers in monadic second-order logic over trees. Fundamenta Informaticae, 100(1–4):1–17, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanBenthem:2010:GSE

[50] Johan van Benthem and Amélie Gheerbrant. Game solution, epistemic dynamics and fixed-point logics. Fundamenta Informaticae, 100(1–4): 19–41, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Coquand:2010:NFT

[51] Thierry Coquand and Guilhem Jaber. A note on forcing and type theory. Fundamenta Informaticae, 100(1–4):43–52, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dunin-Keplicz:2010:FGB

[52] Barbara Dunin-Kȩplicz, Linh Anh Nguyen, and Andrzej Szałas. A framework for graded beliefs, goals and intentions. Fundamenta Informaticae, 100(1–4):53–76, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2010:SMT

[53] Ivo Düntsch, Ewa Orłowska, and Ingrid Rewitzky. Structures with multirelations, their discrete dualities and applications. Fundamenta Informaticae, 100(1–4):77–98, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzymala-Busse:2010:LVM

[54] Jerzy W. Grzymala-Busse and Wojciech Rza̧sa. A local version of the MLEM2 algorithm for rule induction. Fundamenta Informaticae, 100(1– 4):99–116, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazurkiewicz:2010:LCT

[55] Antoni Mazurkiewicz. Local computations on triangular graphs. Fundamenta Informaticae, 100(1–4):117–140, January 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2010:ASR

[56] Andrzej Skowron, Jarosław Stepaniuk, and Roman Swiniarski. Approximation spaces in rough–granular computing. Fundamenta Informaticae, 100(1–4):141–157, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vakarelov:2010:DMP

[57] Dimiter Vakarelov. Dynamic mereotopology: A point-free theory of changing regions. I. Stable and unstable mereotopological relations. Fundamenta Informaticae, 100(1–4):159–180, January 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Veltink:2010:PR

[58] Gerrit Jan Veltink. PSF — a retrospective. Fundamenta Informaticae, 100(1–4):181–227, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2010:P

[59] Ludwik Czaja. Preface. Fundamenta Informaticae, 101(1–2):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2010:SAN

[60] Ludwik Czaja and Manfred Kudlek. Synthesis and analysis of net structures and transition graphs with inhibitor relations. Fundamenta Informaticae, 101(1–2):1–7, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Knapik:2010:BPV

[61] Michał Knapik, Wojciech Penczek, Maciej Szreter, Agata, and Półrola. Bounded parametric verification for distributed time Petri nets with discrete-time semantics. Fundamenta Informaticae, 101(1–2):9–27, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2010:SON

[62] Michael Köhler-Bußmeier and Frank Heitmann. Safeness for object nets. Fundamenta Informaticae, 101(1–2):29–43, January 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutny:2010:MRE

[63] Maciej Koutny and Marta Pietkiewicz-Koutny. Minimal regions of ENLtransition systems. Fundamenta Informaticae, 101(1–2):45–58, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2010:IWN

[64] Irina A. Lomazova. Interacting workflow nets for workflow process reengineering. Fundamenta Informaticae, 101(1–2):59–70, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomuscio:2010:POR

[65] Alessio Lomuscio, Wojciech Penczek, and Hongyang Qu. Partial order reductions for model checking temporal-epistemic logics over interleaved multi-agent systems. Fundamenta Informaticae, 101(1–2):71–90, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Palasinski:2010:CER

[66] Marek Pałasiński and Krzysztof Pancerz. On computing extensions and restrictions of information systems noting some order properties. Fundamenta Informaticae, 101(1–2):91–103, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2010:IPT

[67] Roman R. Redziejowski. Infinite product of traces represented by projections. Fundamenta Informaticae, 101(1–2):105–113, January 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2010:TIS

[68] Zbigniew Suraj and Piotr Grochowalski. Toward intelligent searching the rough set database system (RSDS): an ontological approach. Fundamenta Informaticae, 101(1–2):115–123, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wakulicz-Deja:2010:MAD

[69] Alicja Wakulicz-Deja and Małgorzata Przybyła-Kasperek. Multi-agent decision taking system. Fundamenta Informaticae, 101(1–2):125–141, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2010:PCN

[70] Marcin Wolski. Perception and classification. A note on near sets and rough sets. Fundamenta Informaticae, 101(1–2):143–155, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aratsu:2010:ATE

[71] Taku Aratsu, Kouichi Hirata, and Tetsuji Kuboyama. Approximating tree edit distance through string edit distance for binary tree codes. Fundamenta Informaticae, 101(3):157–171, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crochemore:2010:FPG

[72] Maxime Crochemore, Marcin Kubica, Tomasz Waleń, Costas S. Iliopoulos, and M. Sohel Rahman. Finding patterns in given intervals. Fundamenta Informaticae, 101(3):173–186, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honko:2010:SBC

[73] Piotr Hońko. Similarity-based classification in relational databases. Fundamenta Informaticae, 101(3):187–213, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2010:IBE

[74] Lila Kari and Shinnosuke Seki. An improved bound for an extension of fine and Wilf’s theorem and its optimality. Fundamenta Informaticae, 101 (3):215–236, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2010:LMI

[75] Guohua Liu and Jia-Huai You. Level mapping induced loop formulas for weight constraint and aggregate logic programs. Fundamenta Informaticae, 101(3):237–255, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konstantinidis:2010:CME

[76] Stavros Konstantinidis and Pedro V. Silva. Computing maximal errordetecting capabilities and distances of regular languages. Fundamenta Informaticae, 101(4):257–270, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kursa:2010:BSF

[77] Miron B. Kursa, Aleksander Jankowski, and Witold R. Rudnicki. Boruta — a system for feature selection. Fundamenta Informaticae, 101(4):271– 285, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maulik:2010:RMC

[78] Ujjwal Maulik and Debasis Chakraborty. A robust multiple classifier system for pixel classification of remote sensing images. Fundamenta Informaticae, 101(4):286–304, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shanker:2010:EMI

[79] N. R. Shanker and S. S. Ramakrishnan. Enhancement of multispectral Ikonos satellite image using quantum information processing. Fundamenta Informaticae, 101(4):305–320, December 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ye:2010:ACC

[80] Guodong Ye. Another constructed chaotic image encryption scheme based on Toeplitz matrix and Hankel matrix. Fundamenta Informaticae, 101(4):321–333, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2010:P

[81] Wojciech Penczek. Preface. Fundamenta Informaticae, 102(1):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bachmann:2010:TIL

[82] Jörg Peter Bachmann and Louchka Popova-Zeugmann. Timeindependent liveness in time Petri nets. Fundamenta Informaticae, 102 (1):1–17, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2010:NBD

[83] Roberto Barbuti, Andrea Maggiolo-Schettini, Paolo Milazzo, and Damas P. Gruska. A notion of biological diagnosability inspired by the notion of opacity in systems security. Fundamenta Informaticae, 102 (1):19–34, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Budzynska:2010:UPB

[84] Katarzyna Budzyńska, Magdalena Kacprzak, and Paweł Rembelski. Update of probabilistic beliefs: Implementation and parametric verification. Fundamenta Informaticae, 102(1):35–48, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Burkhard:2010:AOT

[85] Hans-Dieter Burkhard. Agent oriented techniques for programming autonomous robots. Fundamenta Informaticae, 102(1):49–62, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2010:PAC

[86] Damas P. Gruska. Process algebra contexts and security properties. Fundamenta Informaticae, 102(1):63–76, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomuscio:2010:MCO

[87] Alessio Lomuscio, Ben Strulo, Nigel G. Walker, and Peng Wu. Model checking optimisation based congestion control algorithms. Fundamenta Informaticae, 102(1):77–96, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2010:CCA

[88] Linh Anh Nguyen and Andrzej Szałas. Checking consistency of an ABox w.r.t. global assumptions in PDL. Fundamenta Informaticae, 102(1): 97–113, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Srebrny:2010:SPE

[89] Marian Srebrny and Lidia Stepień. SAT as a programming environment for linear algebra. Fundamenta Informaticae, 102(1):115–127, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Virbitskaite:2010:CVT

[90] Irina Virbitskaite, Natalya Gribovskaya, and Eike Best. A categorical view of timed behaviours. Fundamenta Informaticae, 102(1):129–143, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Altenkirch:2010:P

[91] Thorsten Altenkirch and Tarmo Uustalu. Preface. Fundamenta Informaticae, 102(2):??, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brady:2010:CCC

[92] Edwin Brady and Kevin Hammond. Correct-by-construction concurrency: Using dependent types to verify implementations of effectful resource usage protocols. Fundamenta Informaticae, 102(2):145–176, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Loh:2010:TID

[93] Andres Löh, Conor McBride, and Wouter Swierstra. A tutorial implementation of a dependently typed Lambda calculus. Fundamenta Informaticae, 102(2):177–207, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wilson:2010:ADT

[94] Sean Wilson, Jacques Fleuriot, and Alan Smaill. Automation for dependently typed functional programming. Fundamenta Informaticae, 102 (2):209–228, April 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gavanelli:2010:Pa

[95] Marco Gavanelli and Toni Mancini. Preface. Fundamenta Informaticae, 102(3–4):??, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balafoutis:2010:EIM

[96] Thanasis Balafoutis and Kostas Stergiou. Evaluating and improving modern variable and revision ordering strategies in CSPs. Fundamenta Informaticae, 102(3–4):229–261, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brito:2010:CTE

[97] Ismel Brito and Pedro Meseguer. Cluster tree elimination for distributed constraint optimization with quality guarantees. Fundamenta Informaticae, 102(3–4):263–286, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gerevini:2010:EPA

[98] Alfonso E. Gerevini and Ivan Serina. Efficient plan adaptation through replanning windows and heuristic goals. Fundamenta Informaticae, 102(3–4):287–323, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Montali:2010:ALP

[99] Marco Montali, Paolo Torroni, Federico Chesani, Paola Mello, Marco Alberti, and Evelina Lamma. Abductive logic programming as an effective technology for the static verification of declarative business processes. Fundamenta Informaticae, 102(3–4):325–361, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Morgado:2010:COS

[100] António Morgado and Joao Marques-Silva. Combinatorial optimization solutions for the maximum quartet consistency problem. Fundamenta Informaticae, 102(3–4):363–389, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pulina:2010:ESQ

[101] Luca Pulina and Armando Tacchella. An empirical study of QBF encodings: from treewidth estimation to useful preprocessing. Fundamenta Informaticae, 102(3–4):391–427, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Riguzzi:2010:SRI

[102] Fabrizio Riguzzi. SLGAD resolution for inference on logic programs with annotated disjunctions. Fundamenta Informaticae, 102(3–4):429–466, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Soh:2010:SBM

[103] Takehide Soh, Katsumi Inoue, Naoyuki Tamura, Mutsunori Banbara, and Hidetomo Nabeshima. A SAT-based method for solving the twodimensional strip packing problem. Fundamenta Informaticae, 102(3–4):

467–487, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gambin:2010:P

[104] Anna Gambin, Damian NiwińSki, and Paweł Urzyczyn. Preface. Fundamenta Informaticae, 103(1–4):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2010:FDP

[105] Roberto Barbuti, Andrea Maggiolo-Schettini, Angelo Troina, Mariangiola Dezani-Ciancaglini, and Paolo Milazzo. A formalism for the description of protein interaction dedicated to Jerzy Tiuryn on the occasion of his 60th birthday. Fundamenta Informaticae, 103(1–4):1–29, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2010:DLD

[106] J. A. Bergstra and C. A. Middelburg. Data linkage dynamics with shedding. Fundamenta Informaticae, 103(1–4):31–52, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dojer:2010:EAL

[107] Norbert Dojer. An efficient algorithm for learning Bayesian networks from data. Fundamenta Informaticae, 103(1–4):53–67, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Donnelly:2010:CRV

[108] Kevin Donnelly, Assaf Kfoury, and Andrei Lapets. The complexity of restricted variants of the stable paths problem. Fundamenta Informaticae, 103(1–4):69–87, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gambin:2010:MPM

[109] Anna Gambin and Bogusław Kluge. Modeling proteolysis from mass spectrometry proteomic data. Fundamenta Informaticae, 103(1–4):89– 104, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorecki:2010:TME

[110] Paweł Górecki. H-trees: a model of evolutionary scenarios with horizontal gene transfer. Fundamenta Informaticae, 103(1–4):105–128, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kozen:2010:CRM

[111] Dexter Kozen. Church–Rosser made easy. Fundamenta Informaticae, 103(1–4):129–136, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kusmierek:2010:BSO

[112] Jarosław Dominik Mateusz Kuśmierek and Viviana Bono. Big-step operational semantics revisited. Fundamenta Informaticae, 103(1–4):137–172, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pagani:2010:LND

[113] Michele Pagani and Simona Ronchi Della Rocca. Linearity, nondeterminism and solvability. Fundamenta Informaticae, 103(1–4):173– 202, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pratt:2010:CYE

[114] Vaughan Pratt. Communes via Yoneda, from an elementary perspective. Fundamenta Informaticae, 103(1–4):203–218, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudnicki:2010:AAS

[115] Ryszard Rudnicki and Radosław Wieczorek. Asymptotic analysis of a semelparous species model. Fundamenta Informaticae, 103(1–4):219–233, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Setty:2010:MBU

[116] Yaki Setty, Irun R. Cohen, and David Harel. Modeling biology using generic reactive animation. Fundamenta Informaticae, 103(1–4):235–246, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shivers:2010:BRU

[117] Olin Shivers and Mitchell Wand. Bottom-up β-reduction: Uplinks and λ-DAGs. Fundamenta Informaticae, 103(1–4):247–287, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tyszkiewicz:2010:CTI

[118] Jerzy Tyszkiewicz. Complexity of type inference. Fundamenta Informaticae, 103(1–4):289–301, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urzyczyn:2010:LPI

[119] Paweł Urzyczyn. The logic of persistent intersection. Fundamenta Informaticae, 103(1–4):303–322, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wilczynski:2010:CSG

[120] Bartek Wilczyński and Torgeir R. Hvidsten. A computer scientist’s guide to the regulatory genome. Fundamenta Informaticae, 103(1–4):323–332, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bordihn:2010:P

[121] Henning Bordihn, Mika Hirvensalo, Martin Kutrib, Rudolf Freund, Markus Holzer, and Friedrich Otto. Preface. Fundamenta Informaticae, 104(1–2):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bianchi:2010:BUQ

[122] Maria Paola Bianchi and Beatrice Palano. Behaviours of unary quantum automata. Fundamenta Informaticae, 104(1–2):1–15, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cerno:2010:CRA

[123] Peter Černo and František Mráz. Clearing restarting automata. Fundamenta Informaticae, 104(1–2):17–54, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geffert:2010:OPV

[124] Viliam Geffert, Giovanni Pighizzini, and Carlo Mereghetti. One pebble versus ϵ · log n bits. Fundamenta Informaticae, 104(1–2):55–69, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leupold:2010:WCA

[125] Peter Leupold and Benedek Nagy. 5^{→}3′ Watson–Crick automata with several runs. Fundamenta Informaticae, 104(1–2):71–91, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin:2010:LDT

[126] Gema M. Martín, Francisco J. Vico, Jüegen Dassow, and Bianca Truthe. Low disruption transformations on cyclic automata. Fundamenta Informaticae, 104(1–2):93–110, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Masopust:2010:RNP

[127] Tomáš Masopust. Regulated nondeterminism in pushdown automata: The non-regular case. Fundamenta Informaticae, 104(1–2):111–124, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nishio:2010:ACC

[128] Hidenosuke Nishio. Automorphissm classification of cellular automata. Fundamenta Informaticae, 104(1–2):125–140, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sutner:2010:CAD

[129] K. Sutner. Cellular automata, decidability and phasespace. Fundamenta Informaticae, 104(1–2):141–160, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Truthe:2010:TBA

[130] Bianca Truthe. Target based accepting networks of evolutionary processors. Fundamenta Informaticae, 104(1–2):161–183, January 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hromkovic:2010:P

[131] Juraj Hromkovič, Borislav Šuster, and Eduard Toman. Preface. Fundamenta Informaticae, 104(3):??, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ablayev:2010:CAS

[132] Farid Ablayev and Svetlana Ablayeva. A communication approach to the superposition problem. Fundamenta Informaticae, 104(3):185–200, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Albrecht:2010:NPE

[133] Andreas A. Albrecht, Alexander V. Chashkin, Costas S. Iliopoulos, Oktay M. Kasim-Zade, Georgios Lappas, and Kathleen K. Steinhöfel. A note on a priori estimations of classification circuit complexity. Fundamenta Informaticae, 104(3):201–217, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alekhina:2010:SCA

[134] Marina An. Alekhina. Synthesis and complexity of asymptotically optimal circuits with unreliable gates. Fundamenta Informaticae, 104(3): 219–225, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gashkov:2010:SRT

[135] Sergej B. Gashkov and Igor B. Gashkov. Some remarks on testing irreducibility of polynomials and normality of bases in finite fields. Fundamenta Informaticae, 104(3):227–238, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lozhkin:2010:HAA

[136] Sergei A. Lozhkin and Alexander E. Shiganov. High accuracy asymptotic bounds on the BDD size and weight of the hardest functions. Fundamenta Informaticae, 104(3):239–253, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Melnikov:2010:ENF

[137] Boris Melnikov. Extended nondeterministic finite automata. Fundamenta Informaticae, 104(3):255–265, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Melnikov:2010:OME

[138] Boris Melnikov. Once more on the edge-minimization of nondeterministic finite automata and the connected problems. Fundamenta Informaticae, 104(3):267–283, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moshkov:2010:GAW

[139] Mikhail Ju. Moshkov. Greedy algorithm with weights for decision tree construction. Fundamenta Informaticae, 104(3):285–292, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dapoigny:2010:MCD

[140] Richard Dapoigny and Patrick Barlatier. Modeling contexts with dependent types. Fundamenta Informaticae, 104(4):293–327, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jez:2010:UES

[141] Artur Jeż and Alexander Okhotin. Univariate equations over sets of natural numbers. Fundamenta Informaticae, 104(4):329–348, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2010:HKB

[142] Linh Anh Nguyen. Horn knowledge bases in regular description logics with PTIME data complexity. Fundamenta Informaticae, 104(4):349– 384, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shao:2010:BPB

[143] Zuhua Shao. Breaking Pomykala–Barabasz threshold proxy signature scheme. Fundamenta Informaticae, 104(4):385–392, December 2010. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gavanelli:2010:Pb

[144] Marco Gavanelli, Fabrizio Riguzzi, and Alberto Pettorossi. Preface. Fundamenta Informaticae, 105(1–2):??, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Costantini:2010:EIR

[145] Stefania Costantini, Andrea Formisano, and Davide Petturiti. Extending and implementing RASP. Fundamenta Informaticae, 105(1–2):1–33, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ricca:2010:LBS

[146] Francesco Ricca, Antonella Dimasi, Giovanni Grasso, Salvatore Maria Ielpa, Salvatore Iiritano, Marco Manna, and Nicola Leone. A logic-based system for e-tourism. Fundamenta Informaticae, 105(1–2):35–55, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See erratum [241].

Bistarelli:2010:SDC

[147] Stefano Bistarelli and Giorgio Gosti. Solving distributed CSPs probabilistically. Fundamenta Informaticae, 105(1–2):57–78, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dovier:2010:IMA

[148] Agostino Dovier, Andrea Formisano, and Enrico Pontelli. An investigation of multi-agent planning in CLP. Fundamenta Informaticae, 105 (1–2):79–103, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Campagna:2010:MCE

[149] Dario Campagna, Christian De Rosa, Agostino Dovier, Angelo Montanari, and Carla Piazza. Morphos configuration engine: the core of a commercial configuration system in CLP(FD). Fundamenta Informaticae, 105(1–2):105–133, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chesani:2010:LBR

[150] Federico Chesani, Paola Mello, Marco Montali, and Paolo Torroni. A logic-based, reactive calculus of events. Fundamenta Informaticae, 105(1– 2):135–161, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nicolini:2010:CSP

[151] Enrica Nicolini, Christophe Ringeissen, and Michael Rusinowitch. Combining satisfiability procedures for unions of theories with a shared counting operator. Fundamenta Informaticae, 105(1–2):163–187, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Faella:2010:GAT

[152] Marco Faella, Margherita Napoli, and Mimmo Parente. Graded alternating-time temporal logic. Fundamenta Informaticae, 105(1–2): 189–210, January 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Franceschinis:2010:P

[153] Giuliana Franceschinis, Wojciech Penczek, and Karsten Wolf. Preface. Fundamenta Informaticae, 105(3):??, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardinello:2010:COL

[154] Luca Bernardinello, Lucia Pomello, and Stefania Rombolà. Closure operators and lattices derived from concurrency in posets and occurrence nets. Fundamenta Informaticae, 105(3):211–235, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mairesse:2010:DZP

[155] Jean Mairesse and Hoang-Thach Nguyen. Deficiency zero Petri nets and product form. Fundamenta Informaticae, 105(3):237–261, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Oliveira:2010:HDG

[156] Mateus de Oliveira Oliveira. Hasse diagram generators and Petri nets. Fundamenta Informaticae, 105(3):263–289, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rosa-Velardo:2010:DPP

[157] Fernando Rosa-Velardo and David de Frutos-Escrig. Decidability problems in Petri nets with names and replication. Fundamenta Informaticae, 105(3):291–317, August 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valmari:2010:SBM

[158] Antti Valmari. Simple bisimilarity minimization in O(m log n) time. Fundamenta Informaticae, 105(3):319–339, August 2010. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vandenBergh:2010:CPP

[159] Frans van den Bergh and Andries Petrus Engelbrecht. A convergence proof for the particle swarm optimiser. Fundamenta Informaticae, 105(4): 341–374, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghassemi:2010:ERM

[160] Fatemeh Ghassemi, Wan Fokkink, and Ali Movaghar. Equational reasoning on mobile ad hoc networks. Fundamenta Informaticae, 105(4): 375–415, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grigoryan:2010:EPP

[161] Hayk Grigoryan and Arsen Shoukourian. Equivalence of processes in partially commutative object-oriented environments. Fundamenta Informaticae, 105(4):417–434, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kenaza:2010:UNB

[162] Tayeb Kenaza, Karim Tabia, and Salem Benferhat. On the use of naive Bayesian classifiers for detecting elementary and coordinated attacks.

Fundamenta Informaticae, 105(4):435–466, December 2010. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamide:2011:PLT

[163] Norihiro Kamide and Heinrich Wansing. A paraconsistent linear-time temporal logic. Fundamenta Informaticae, 106(1):1–23, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2011:LMI

[164] Guohua Liu and Jia-Huai You. Level mapping induced loop formulas for weight constraint and aggregate logic programs. Fundamenta Informaticae, 106(1):25–43, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saha:2011:USL

[165] Indrajit Saha, Ujjwal Maulik, Sanghamitra Bandyopadhyay, and Dariusz Plewczynski. Unsupervised and supervised learning approaches together for microarray analysis. Fundamenta Informaticae, 106(1):45–73, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tang:2011:SFB

[166] Zhenjun Tang, Shuozhong Wang, Xinpeng Zhang, and Weimin Wei. Structural feature-based image hashing and similarity metric for tampering detection. Fundamenta Informaticae, 106(1):75–91, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wakulicz-Deja:2011:AME

[167] Alicja Wakulicz-Deja and Małgorzata Przybyła-Kasperek. Application of the method of editing and condensing in the process of global decisionmaking. Fundamenta Informaticae, 106(1):93–117, January 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ono:2011:P

[168] Hiroakira Ono, Makoto Kanazawa, and Ruy de Queiroz. Preface. Fundamenta Informaticae, 106(2–4):??, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alves:2011:PGT

[169] Gleifer V. Alves, Anjolina G. de Oliveira, and Ruy de Queiroz. Proofgraphs: a thorough cycle treatment, normalization and subformula property. Fundamenta Informaticae, 106(2–4):119–147, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amato:2011:OCA

[170] Gianluca Amato and Francesca Scozzari. Observational completeness on abstract interpretation. Fundamenta Informaticae, 106(2–4):149–173, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Belardinelli:2011:FOL

[171] Francesco Belardinelli and Alessio Lomuscio. First-order linear-time epistemic logic with group knowledge: An axiomatisation of the monodic fragment. Fundamenta Informaticae, 106(2–4):175–190, April 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Caleiro:2011:TBA

[172] Carlos Caleiro and Ricardo Gonçalves. Towards a behavioral algebraic theory of logical valuations. Fundamenta Informaticae, 106(2–4):191– 209, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deGroote:2011:APS

[173] Philippe de Groote, Sylvain Pogodalla, and Carl Pollard. About parallel and syntactocentric formalisms: A perspective from the encoding of convergent grammar into abstract categorial grammar. Fundamenta Informaticae, 106(2–4):211–231, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Link:2011:IFF

[174] Sebastian Link. On inferences of full first-order hierarchical decompositions. Fundamenta Informaticae, 106(2–4):233–258, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kontinen:2011:TLS

[175] Juha Kontinen and Ville Nurmi. Team logic and second-order logic. Fundamenta Informaticae, 106(2–4):259–272, April 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maruyama:2011:DAF

[176] Yoshihiro Maruyama. Dualities for algebras of Fitting’s many-valued modal logics. Fundamenta Informaticae, 106(2–4):273–294, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nieves:2011:SGR

[177] Juan Carlos Nieves, Mauricio Osorio, and Claudia Zepeda. A schema for generating relevant logic programming semantics and its applications in argumentation theory. Fundamenta Informaticae, 106(2–4):295–319, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2011:KTP

[178] Ren-June Wang. Knowledge, time, and the problem of logical omniscience. Fundamenta Informaticae, 106(2–4):321–338, April 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cui:2011:CCT

[179] Bo Cui, Lila Kari, and Shinnosuke Seki. K-comma codes and their generalizations. Fundamenta Informaticae, 107(1):1–18, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hojjat:2011:FAS

[180] Hossein Hojjat, Mohammad Reza Mousavi, and Marjan Sirjani. Formal analysis of SystemC designs in process algebra. Fundamenta Informaticae, 107(1):19–42, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kammuller:2011:MAF

[181] Florian Kammüller. Mechanical analysis of finite idempotent relations. Fundamenta Informaticae, 107(1):43–65, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

She:2011:RTDa

[182] Yanhong She, Xiaoli He, and Guojun Wang. Rough truth degrees of formulas and approximate reasoning in rough logic. Fundamenta Informaticae, 107(1):67–83, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wieczorkowska:2011:ARM

[183] Alicja Wieczorkowska, Elżbieta Kubera, and Agnieszka Kubik-Komar. Analysis of recognition of a musical instrument in sound mixes using support vector machines. Fundamenta Informaticae, 107(1):85–104, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xiong:2011:CIB

[184] Hu Xiong, Zhiguang Qin, and Fagen Li. Cryptanalysis of an identity based signcryption without random oracles. Fundamenta Informaticae, 107(1):105–109, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gavanelli:2011:REE

[185] Marco Gavanelli, Toni Mancini, and Alberto Pettorossi. RCRA 2009 experimental evaluation of algorithms for solving problems with combinatorial explosion. Fundamenta Informaticae, 107(2–3):??, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cesta:2011:FPV

[186] Amedeo Cesta, Simone Fratini, Andrea Orlandini, Alberto Finzi, and Enrico Tronci. Flexible plan verification: Feasibility results. Fundamenta Informaticae, 107(2–3):111–137, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lewis:2011:PQS

[187] Matthew Lewis, Tobias Schubert, Bernd Becker, Paolo Marin, Massimo Narizzano, and Enrico Giunchiglia. Parallel QBF solving with advanced knowledge sharing. Fundamenta Informaticae, 107(2–3):139–166, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gerevini:2011:EAS

[188] Alfonso Gerevini, Alessandro Saetti, and Ivan Serina. An empirical analysis of some heuristic features for planning through local search and action graphs. Fundamenta Informaticae, 107(2–3):167–197, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zanzotto:2011:EGK

[189] Fabio Massimo Zanzotto, Lorenzo Dell’Arciprete, and Alessandro Moschitti. Efficient graph kernels for textual entailment recognition. Fundamenta Informaticae, 107(2–3):199–222, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

He:2011:ACL

[190] Jun He, Pierre Flener, and Justin Pearson. An automaton constraint for local search. Fundamenta Informaticae, 107(2–3):223–248, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lynce:2011:RCS

[191] Inês Lynce and Joao Marques-Silva. Restoring CSP satisfiability with MaxSAT. Fundamenta Informaticae, 107(2–3):249–266, April 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lombardi:2011:DIS

[192] Michele Lombardi, Michela Milano, Andrea Roli, and Alessandro Zanarini. Deriving information from sampling and diving. Fundamenta Informaticae, 107(2–3):267–287, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hyvarinen:2011:PSS

[193] Antti E. J. Hyvärinen, Tommi Junttila, and Ilkka Niemelä. Partitioning search spaces of a randomized search. Fundamenta Informaticae, 107 (2–3):289–311, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Andary:2011:RTP

[194] Philippe Andary, Bruno Patrou, and Pierre Valarcher. A representation theorem for primitive recursive algorithms. Fundamenta Informaticae, 107(4):313–330, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhattacharya:2011:FWP

[195] Bhaswar B. Bhattacharya. On the Fermat–Weber point of a polygonal chain and its generalizations. Fundamenta Informaticae, 107(4):331–343, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cojocaru:2011:TOB

[196] Liliana Cojocaru. Trade-offs between time, space, cooperation, and communication complexity for CD grammar systems. Fundamenta Informaticae, 107(4):345–378, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Faber:2011:LBT

[197] Wolfgang Faber, Nicola Leone, Marco Maratea, and Francesco Ricca. Look-back techniques for ASP programs with aggregates. Fundamenta Informaticae, 107(4):379–413, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Feng:2011:AMB

[198] Dingcheng Feng, Feng Chen, and Wenli Xu. Analysis of Markov boundary induction in Bayesian networks: A new view from matroid theory. Fundamenta Informaticae, 107(4):415–434, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Edwards:2011:P

[199] Stephen A. Edwards, Ryszard Janicki, and Walter Vogler. Preface. Fundamenta Informaticae, 108(1–2):??, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carmona:2011:SSS

[200] Josep Carmona, Jorge Júlvez, Jordi Cortadella, and Michael Kishinevsky. A scheduling strategy for synchronous elastic designs. Fundamenta Informaticae, 108(1–2):1–21, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Crafa:2011:SST

[201] Silvia Crafa, Francesco Ranzato, and Francesco Tapparo. Saving space in a time efficient simulation algorithm. Fundamenta Informaticae, 108 (1–2):23–42, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lohmann:2011:CRE

[202] Niels Lohmann and Karsten Wolf. Compact representations and efficient algorithms for operating guidelines. Fundamenta Informaticae, 108(1–2): 43–62, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mokhov:2011:FA

[203] Andrey Mokhov, Victor Khomenko, and Alex Yakovlev. Flat arbiters. Fundamenta Informaticae, 108(1–2):63–90, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Potop-Butucaru:2011:CMC

[204] Dumitru Potop-Butucaru, Yves Sorel, Robert de Simone, and Jean-Pierre Talpin. From concurrent multi-clock programs to deterministic asynchronous implementations. Fundamenta Informaticae, 108(1–2):91– 118, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Raclet:2011:MIT

[205] Jean-Baptiste Raclet, Eric Badouel, Albert Benveniste, Benoît Caillaud, Axel Legay, and Roberto Passerone. A modal interface theory for component-based design. Fundamenta Informaticae, 108(1–2):119–149, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolf:2011:GWT

[206] Karsten Wolf, Christian Stahl, Daniela Weinberg, Janine Ott, and Robert Danitz. Guaranteeing weak termination in service discovery. Fundamenta Informaticae, 108(1–2):151–180, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2011:P

[207] Davide Ciucci and Yiyu Yao. Preface. Fundamenta Informaticae, 108 (3–4):??, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2011:MAS

[208] Marcin Wolski. Monadic algebras: a standpoint on rough sets. Fundamenta Informaticae, 108(3–4):181–196, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2011:IAR

[209] Gianpiero Cattaneo. An investigation about rough set theory: Some foundational and mathematical aspects. Fundamenta Informaticae, 108 (3–4):197–221, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Couso:2011:RSC

[210] Inés Couso and Didier Dubois. Rough sets, coverings and incomplete information. Fundamenta Informaticae, 108(3–4):223–247, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2011:TSI

[211] Yiyu Yao. Two semantic issues in a probabilistic rough set model. Fundamenta Informaticae, 108(3–4):249–265, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herbert:2011:GTR

[212] Joseph P. Herbert and JingTao Yao. Game-theoretic rough sets. Fundamenta Informaticae, 108(3–4):267–286, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2011:OSW

[213] Davide Ciucci. Orthopairs: A simple and widely used way to model uncertainty. Fundamenta Informaticae, 108(3–4):287–304, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diker:2011:TFR

[214] Murat Diker. Textures and fuzzy rough sets. Fundamenta Informaticae, 108(3–4):305–336, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2011:SMS

[215] Wei-Zhi Wu. On some mathematical structures of T -fuzzy rough set algebras in infinite universes of discourse. Fundamenta Informaticae, 108(3–4):337–369, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Caron:2011:GWE

[216] Pascal Caron and Marianne Flouret. From Glushkov WFAs to Kexpressions. Fundamenta Informaticae, 109(1):1–25, January 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kullmann:2011:CSPa

[217] Oliver Kullmann. Constraint satisfaction problems in clausal form I: Autarkies and deficiency. Fundamenta Informaticae, 109(1):27–81, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kullmann:2011:CSPb

[218] Oliver Kullmann. Constraint satisfaction problems in clausal form II: Minimal unsatisfiability and conflict structure. Fundamenta Informaticae, 109(1):83–119, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2011:RSB

[219] Chin-Chen Chang, Chih-Yang Lin, and Yi-Hsuan Fan. Reversible steganography for BTC-compressed images. Fundamenta Informaticae, 109(2):121–134, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gorrieri:2011:OPN

[220] Roberto Gorrieri and Cristian Versari. An operational Petri net semantics for A^{2} CCS. Fundamenta Informaticae, 109(2):135–160, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiraskova:2011:SCS

[221] Galina Jirásková and Alexander Okhotin. On the state complexity of star of union and star of intersection. Fundamenta Informaticae, 109 (2):161–178, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niu:2011:TSB

[222] Yunyun Niu, Linqiang Pan, Mario J. Pérez-Jiménez, and Miquel Rius Font. A tissue P systems based uniform solution to tripartite matching problem. Fundamenta Informaticae, 109(2):179–188, April 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2011:CTE

[223] Xu An Wang, Xiaoyuan Yang, and Minqing Zhang. Cryptanalysis of two efficient HIBE schemes in the standard model. Fundamenta Informaticae, 109(2):189–200, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Winkowski:2011:MTS

[224] Józef Winkowski. Multiplicative transition systems. Fundamenta Informaticae, 109(2):201–222, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2011:P

[225] Ludwik Czaja. Preface. Fundamenta Informaticae, 109(3):??, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bashkin:2011:RDA

[226] Vladimir A. Bashkin and Irina A. Lomazova. Resource driven automata nets. Fundamenta Informaticae, 109(3):223–236, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2011:PJS

[227] Marco Bellia and M. Eugenia Occhiuto. Properties of Java simple closures. Fundamenta Informaticae, 109(3):237–253, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2011:DFD

[228] Ludwik Czaja. On deadlock and fairness decision problems for computations on client-server systems. Fundamenta Informaticae, 109(3): 255–264, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2011:LAA

[229] Anna Gomolińska. A logic-algebraic approach to graded inclusion. Fundamenta Informaticae, 109(3):265–279, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2011:GEP

[230] Damas P. Gruska. Gained and excluded private actions by process observations. Fundamenta Informaticae, 109(3):281–295, August 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2011:RBM

[231] Manfred Kudlek. A relation between modal logic and language closure operators. Fundamenta Informaticae, 109(3):297–304, August 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meski:2011:BBB

[232] Artur Mȩski, Wojciech Penczek, and Agata Półrola. BDD-based bounded model checking for temporal properties of 1-safe Petri nets. Fundamenta Informaticae, 109(3):305–321, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2011:BIF

[233] Roman R. Redziejowski. BITES instead of FIRST for parsing expression grammar. Fundamenta Informaticae, 109(3):323–337, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szczuka:2011:FAQ

[234] Marcin Szczuka, Andrzej Skowron, and Jarosław Stepaniuk. Function approximation and quality measures in rough–granular systems. Fundamenta Informaticae, 109(3):339–354, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2011:INI

[235] Marcin Wolski. Incomplete and nondeterministic information systems: Object-directed semantics for descriptor languages. Fundamenta Informaticae, 109(3):355–368, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esik:2011:SB

[236] Zoltán Ésik and Klaus Sutner. Stephen L. Bloom 1940–2010. Fundamenta Informaticae, 109(4):369–381, December 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Locher:2011:EEK

[237] Thomas Locher, Stefan Schmid, and Roger Wattenhofer. eDonkey & eMule’s Kad: Measurements & attacks. Fundamenta Informaticae, 109(4):383–403, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Piao:2011:TBD

[238] Xiaoxue Piao and Kai Salomaa. Transformations between different models of unranked bottom-up tree automata. Fundamenta Informaticae, 109(4):405–424, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2011:MAS

[239] Daniel Simson. Mesh algorithms for solving principal Diophantine equations, sand-glass tubes and tori of roots. Fundamenta Informaticae, 109(4):425–462, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhu:2011:AAR

[240] Ping Zhu. An axiomatic approach to the roughness measure of rough sets. Fundamenta Informaticae, 109(4):463–480, December 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2011:EPL

[241] Anonymous. Erratum to the paper: “A Logic-Based System for e-Tourism”. Fundamenta Informaticae, 109(4):481, December 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See

Karhumaki:2011:P

[242] Juhani Karhumäki, Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Preface. Fundamenta Informaticae, 110(1–4):??, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akhtarzada:2011:MCM

[243] Ali Akhtarzada, Cristian S. Calude, and John Hosking. A multi-criteria metric algorithm for recommender systems. Fundamenta Informaticae, 110(1–4):1–11, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2011:SID

[244] Artiom Alhazov, Alexander Krassovitskiy, Yurii Rogozhin, and Sergey Verlan. P systems with insertion and deletion exo-operations. Fundamenta Informaticae, 110(1–4):13–28, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Azimi:2011:DOI

[245] Sepinoud Azimi, Tero Harju, Miika Langille, Ion Petre, and Vladimir Rogojin. Directed overlap-inclusion graphs as representations of ciliate genes. Fundamenta Informaticae, 110(1–4):29–44, January 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ballier:2011:LSS

[246] Alexis Ballier, Pierre Guillon, and Jarkko Kari. Limit sets of stable and unstable cellular automata. Fundamenta Informaticae, 110(1–4):45–57, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bockenhauer:2011:HRM

[247] Hans-Joachim Böckenhauer, Juraj Hromkovič, and Andreas Sprock. On the hardness of reoptimization with multiple given solutions. Fundamenta Informaticae, 110(1–4):59–76, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bonizzoni:2011:PLG

[248] Paola Bonizzoni, Claudio Ferretti, A. Roslin Sagaya Mary, and Giancarlo Mauri. Picture languages generated by assembling tiles. Fundamenta Informaticae, 110(1–4):77–93, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cui:2011:TPG

[249] Cewei Cui, Zhe Dang, and Thomas R. Fischer. Typical paths of a graph. Fundamenta Informaticae, 110(1–4):95–109, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Egecioglu:2011:HSM

[250] Ömer Eğecioğlu, László Hegedüs, and Benedek Nagy. Hierarchies of stateless multicounter 5^{→}3′ Watson–Crick automata languages. Fundamenta Informaticae, 110(1–4):111–123, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eisman:2011:ANR

[251] Gerry Eisman and Bala Ravikumar. On approximating non-regular languages by regular languages. Fundamenta Informaticae, 110(1–4):125– 142, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gao:2011:TCI

[252] Yuan Gao, Kai Salomaa, and Sheng Yu. Transition complexity of incomplete DFAs. Fundamenta Informaticae, 110(1–4):143–158, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2011:CCN

[253] Markus Holzer, Andreas Klein, Martin Kutrib, and Oliver Ruepp. Computational complexity of NURIKABE. Fundamenta Informaticae, 110(1– 4):159–174, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:2011:SEP

[254] Juha Honkala. The sequence equivalence problem for marked DT0L systems. Fundamenta Informaticae, 110(1–4):175–182, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ilie:2011:MUS

[255] Lucian Ilie and William F. Smyth. Minimum unique substrings and maximum repeats. Fundamenta Informaticae, 110(1–4):183–195, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karhumaki:2011:UDM

[256] Juhani Karhumäki and Aleksi Saarela. The unique decipherability in the monoid of regular languages is undecidable. Fundamenta Informaticae, 110(1–4):197–200, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2011:RIH

[257] Lila Kari, Shinnosuke Seki, and Steffen Kopecki. On the regularity of iterated hairpin completion of a single word. Fundamenta Informaticae, 110(1–4):201–215, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kleijn:2011:MSQ

[258] Jetty Kleijn and Maciej Koutny. Membrane systems with qualitative evolution rules. Fundamenta Informaticae, 110(1–4):217–230, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kunc:2011:SCU

[259] Michal Kunc and Alexander Okhotin. State complexity of union and intersection for two-way nondeterministic finite automata. Fundamenta Informaticae, 110(1–4):231–239, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Morita:2011:TWR

[260] Kenichi Morita. Two-way reversible multi-head finite automata. Fundamenta Informaticae, 110(1–4):241–254, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okubo:2011:HI

[261] Fumiya Okubo and Takashi Yokomori. On the hairpin incompletion. Fundamenta Informaticae, 110(1–4):255–269, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pan:2011:LAS

[262] Linqiang Pan, Jun Wang, and Hendrik Jan Hoogeboom. Limited asynchronous spiking neural P systems. Fundamenta Informaticae, 110(1–4): 271–293, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez-Jimenez:2011:LSE

[263] Mario J. Pérez-Jiménez, Agustín Riscos-Núñez, Miquel Rius-Font, and Francisco J. Romero-Campero. Looking for small efficient P systems. Fundamenta Informaticae, 110(1–4):295–308, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Raghavan:2011:RBT

[264] Rama Raghavan, H. Ramesh, Marian Gheorghe, and Shankara Narayanan Krishna. On restricted bio-Turing machines. Fundamenta Informaticae, 110(1–4):309–320, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tran:2011:CEE

[265] Nicholas Tran. Characterizations and existence of easy sets without hard subsets. Fundamenta Informaticae, 110(1–4):321–328, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Verlan:2011:UST

[266] Sergey Verlan and Maurice Margenstern. Universality of splicing test tube systems with two tubes. Fundamenta Informaticae, 110(1–4):329– 342, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yen:2011:ASP

[267] Hsu-Chun Yen. On almost-sure properties of probabilistic discrete event systems. Fundamenta Informaticae, 110(1–4):343–359, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2011:P

[268] Wei-Zhi Wu, Dominik Ślȩzak, and Hung Son Nguyen. Preface. Fundamenta Informaticae, 111(1):??, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haruna:2011:DAC

[269] Taichi Haruna and Yukio-Pegio Gunji. Double approximation and complete lattices. Fundamenta Informaticae, 111(1):1–14, January 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2011:ARF

[270] Tong-Jun Li and Wei-Zhi Wu. Attribute reduction in formal contexts: A covering rough set approach. Fundamenta Informaticae, 111(1):15–32, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2011:CFC

[271] Yanqing Yao, Jusheng Mi, Zhoujun Li, and Bin Xie. The construction of fuzzy concept lattices based on (θ, σ)-fuzzy rough approximation operators. Fundamenta Informaticae, 111(1):33–45, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2011:AOB

[272] Zhengjiang Wu, Tianrui Li, Keyun Qin, and Da Ruan. Approximation operators, binary relation and basis algebra in L-fuzzy rough sets. Fundamenta Informaticae, 111(1):47–63, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Du:2011:KFR

[273] Yong Du, Qinghua Hu, Degang Chen, and Peijun Ma. Kernelized fuzzy rough sets based yawn detection for driver fatigue monitoring. Fundamenta Informaticae, 111(1):65–79, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2011:NMP

[274] Jian Yu, Miin-Shen Yang, and Pengwei Hao. A novel multimodal probability model for cluster analysis. Fundamenta Informaticae, 111(1):81–90, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mishra:2011:OOI

[275] Deepti Mishra and Alok Mishra. Object-oriented inheritance metrics in the context of cognitive complexity. Fundamenta Informaticae, 111 (1):91–117, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2011:GPG

[276] Zhe Chen. On the generative power of ω-grammars and ω-automata. Fundamenta Informaticae, 111(2):119–145, April 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Forys:2011:ITS

[277] Wit Foryś and Piotr Oprocha. Infinite traces and symbolic dynamics — the minimal shift case. Fundamenta Informaticae, 111(2):147–161, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fulop:2011:WET

[278] Zoltán Fülöp, Andreas Maletti, and Heiko Vogler. Weighted extended tree transducers. Fundamenta Informaticae, 111(2):163–202, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Intrigila:2011:SRP

[279] Benedetto Intrigila and Richard Statman. Solution to the range problem for combinatory logic. Fundamenta Informaticae, 111(2):203–222, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

She:2011:RTDb

[280] Yanhong She, Xiaoli He, and Guojun Wang. Rough truth degrees of formulas and approximate reasoning in rough logic. Fundamenta Informaticae, 111(2):223–239, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2011:FSI

[281] Jia Yu, Fanyu Kong, Xiangguo Cheng, Rong Hao, and Jianxi Fan. Forward-secure identity-based public-key encryption without random oracles. Fundamenta Informaticae, 111(2):241–256, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See erratum [341].

Bistarelli:2011:NSC

[282] Stefano Bistarelli and Francesco Santini. A nonmonotonic soft concurrent constraint language to model the negotiation process. Fundamenta Informaticae, 111(3):257–279, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Case:2011:PCP

[283] John Case and Samuel E. Moelius III. Properties complementary to program self-reference. Fundamenta Informaticae, 111(3):281–311, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Claude:2011:SIG

[284] Francisco Claude and Gonzalo Navarro. Self-indexed grammar-based compression. Fundamenta Informaticae, 111(3):313–337, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomuscio:2011:RMC

[285] Alessio Lomuscio, Wojciech Penczek, Monika Solanki, and Maciej Szreter. Runtime monitoring of contract regulated Web services. Fundamenta Informaticae, 111(3):339–355, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2011:RTA

[286] Jean-Marc Champarnaud, Jacques Farré, and Franck Guingne. Reduction techniques for acyclic cover transducers. Fundamenta Informaticae, 111(4):357–371, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deng:2011:MEA

[287] Zhi-Hong Deng, Ning Gao, and Xiao-Ran Xu. Mop: An efficient algorithm for mining frequent pattern with subtree traversing. Fundamenta Informaticae, 111(4):373–390, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gandhi:2011:EAG

[288] Aniruddh Gandhi, Bakhadyr Khoussainov, and Jiamou Liu. Efficient algorithms for games played on trees with back-edges. Fundamenta Informaticae, 111(4):391–412, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garg:2011:LBS

[289] Manish Garg and Sugata Gangopadhyay. A lower bound of the secondorder nonlinearities of Boolean bent functions. Fundamenta Informaticae, 111(4):413–422, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ionescu:2011:SND

[290] Mihai Ionescu, Gheorghe Păun, Mario J. Pérez-Jiménez, and Takashi Yokomori. Spiking neural dP systems. Fundamenta Informaticae, 111(4): 423–436, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kisielewicz:2011:RIS

[291] Andrzej Kisielewicz and Marek Szykuła. Rainbow induced subgraphs in proper vertex colorings. Fundamenta Informaticae, 111(4):437–451, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nishida:2011:CBM

[292] Taishin Yasunobu Nishida. Computing k-block morphisms by spiking neural P systems. Fundamenta Informaticae, 111(4):453–464, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2011:AIVa

[293] Anonymous. Author index volume 111. Fundamenta Informaticae, 111(4):465–466, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2011:P

[294] Wojciech Penczek. Preface. Fundamenta Informaticae, 112(1):??, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2011:MPP

[295] Roberto Barbuti, Francesca Levi, Paolo Milazzo, and Guido Scatena. Maximally parallel probabilistic semantics for multiset rewriting. Fundamenta Informaticae, 112(1):1–17, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Belardinelli:2011:MCT

[296] Francesco Belardinelli, Andrew V. Jones, and Alessio Lomuscio. Model checking temporal-epistemic logic using alternating tree automata. Fundamenta Informaticae, 112(1):19–37, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2011:SAG

[297] Ludwik Czaja and Manfred Kudlek. On synthesis and analysis of generalized work flow net structures and transition graphs. Fundamenta Informaticae, 112(1):39–45, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doliwa:2011:PWS

[298] Dariusz Doliwa, Wojciech Horzelski, Mariusz Jarocki, Artur Niewiadomski, Wojciech Penczek, Agata Półrola, Maciej Szreter, and Andrzej Zbrzezny. PlanICS — a Web service composition toolset. Fundamenta Informaticae, 112(1):47–71, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2011:LSO

[299] Michael Köhler-Bußmeier and Frank Heitmann. Liveness of safe object nets. Fundamenta Informaticae, 112(1):73–87, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mellmann:2011:DMC

[300] Heinrich Mellmann and Giuseppe Cotugno. Dynamic motion control: Adaptive bimanual grasping for a humanoid robot. Fundamenta Informaticae, 112(1):89–101, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Turek:2011:SAS

[301] Wojciech Turek, Krzysztof Cetnarowicz, and Wojciech Zaborowski. Software agent systems for improving performance of multi-robot groups. Fundamenta Informaticae, 112(1):103–117, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bordihn:2011:P

[302] Henning Bordihn, Rudolf Freund, Mika Hirvensalo, Markus Holzer, Martin Kutrib, and Friedrich Otto. Preface. Fundamenta Informaticae, 112 (2–3):??, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bianchi:2011:SUP

[303] Maria Paola Bianchi, Carlo Mereghetti, Beatrice Palano, and Giovanni Pighizzini. On the size of unary probabilistic and nondeterministic automata. Fundamenta Informaticae, 112(2–3):119–135, April 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2011:BPA

[304] Erzsébet Csuhaj-Varjú, Tomáš Masopust, and György Vaszil. Blackhole pushdown automata. Fundamenta Informaticae, 112(2–3):137–156, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2011:NEP

[305] Jürgen Dassow and Bianca Truthe. On networks of evolutionary processors with filters accepted by two-state-automata. Fundamenta Informaticae, 112(2–3):157–170, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Huschenbett:2011:KST

[306] Martin Huschenbett. A Kleene–Schützenberger theorem for trace series over bounded lattices. Fundamenta Informaticae, 112(2–3):171–191, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurdzinski:2011:GGL

[307] Tomasz Jurdziński. Growing grammars and length-reducing automata. Fundamenta Informaticae, 112(2–3):193–217, April 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leupold:2011:MFL

[308] Peter Leupold and Friedrich Otto. On McNaughton families of languages that are specified by some variants of monadic string-rewriting systems. Fundamenta Informaticae, 112(2–3):219–238, April 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maletti:2011:SWE

[309] Andreas Maletti. Survey: Weighted extended top-down tree transducers. Part II — application in machine translation. Fundamenta Informaticae, 112(2–3):239–261, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2011:VCD

[310] Anca Dinu. Versatility of ‘continuations’ in discourse semantics. Fundamenta Informaticae, 112(4):263–280, December 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martin:2011:RAV

[311] Gabriela Martin and Magnus Steinby. Rough approximations in varieties of regular languages. Fundamenta Informaticae, 112(4):281–303, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sabri:2011:AFS

[312] Khair Eddin Sabri and Ridha Khedri. Algebraic framework for the specification and analysis of cryptographic-key distribution. Fundamenta Informaticae, 112(4):305–335, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Santana:2011:FIQ

[313] Fabiana T. Santana, Fágner L. Santana, Ana Maria G. Guerreiro, Adrião D., Dória Neto, and Regivan H. N. Santiago. A framework for interval quantization and application to interval based algorithms in digital signal processing. Fundamenta Informaticae, 112(4):337–363, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shim:2011:SPC

[314] Kyung-Ah Shim and Young-Ran Lee. Security pitfalls of the certificateless signature and multi-receiver signcryption schemes. Fundamenta Informaticae, 112(4):365–376, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2011:AIVb

[315] Anonymous. Author index volume 112. Fundamenta Informaticae, 112(4):377–378, December 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Osorio:2011:P

[316] Mauricio Osorio and Victor W. Marek. Preface. Fundamenta Informaticae, 113(1):??, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chavez-Bosquez:2011:SIT

[317] Oscar Chávez-Bosquez, Pilar Pozos-Parra, and Florian Lengyel. Solving the international timetabling competition: a deterministic approach. Fundamenta Informaticae, 113(1):1–18, January 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Confalonieri:2011:NPA

[318] Roberto Confalonieri and Juan Carlos Nieves. Nested preferences in answer set programming. Fundamenta Informaticae, 113(1):19–39, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nieves:2011:PAD

[319] Juan Carlos Nieves and Roberto Confalonieri. A possibilistic argumentation decision making framework with default reasoning. Fundamenta Informaticae, 113(1):41–61, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeIta:2011:TPS

[320] Guillermo De Ita, J. Raymundo Marcial-Romero, and José Antonio Hernández. A threshold for a polynomial solution of #2SAT. Fundamenta Informaticae, 113(1):63–77, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez:2011:AAT

[321] Flavio Omar Everardo Pérez and Fernando Antonio Aguilera Ramírez. Armin: Automatic trance music composition using answer set programming. Fundamenta Informaticae, 113(1):79–96, January 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ramon:2011:P

[322] Jan Ramon, Fabrizio Costa, Christophe Costa Florêncio, and Joost Kok. Preface. Fundamenta Informaticae, 113(2):??, April 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arevalillo:2011:UBI

[323] Jorge M. Arevalillo and Hilario Navarro. Uncovering bivariate interactions in high dimensional data using random forests with data augmentation. Fundamenta Informaticae, 113(2):97–115, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hamalainen:2011:ESM

[324] Wilhelmiina Hämäläinen. Efficient search methods for statistical dependency rules. Fundamenta Informaticae, 113(2):117–150, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cilia:2011:RFM

[325] Elisa Cilia, Niels Landwehr, and Andrea Passerini. Relational feature mining with hierarchical multitask kFOIL. Fundamenta Informaticae, 113(2):151–177, April 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lilius:2011:P

[326] Johan Lilius and Wojciech Penczek. Preface. Fundamenta Informaticae, 113(3–4):??, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Best:2011:SPP

[327] Eike Best and Philippe Darondeau. Separability in persistent Petri nets. Fundamenta Informaticae, 113(3–4):179–203, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Esparza:2011:LWP

[328] Javier Esparza, Martin Leucker, and Maximilian Schlund. Learning workflow Petri nets. Fundamenta Informaticae, 113(3–4):205–228, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hostettler:2011:HLP

[329] Steve Hostettler, Alexis Marechal, Alban Linard, Matteo Risoldi, and Didier Buchs. High-level Petri net model checking with AlPiNA. Fundamenta Informaticae, 113(3–4):229–264, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Le:2011:TAC

[330] Dai Tri Man Lê. On three alternative characterizations of combined traces. Fundamenta Informaticae, 113(3–4):265–293, August 2011. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lohmann:2011:WTS

[331] Niels Lohmann and Daniela Weinberg. Wendy: A tool to synthesize partners for services. Fundamenta Informaticae, 113(3–4):295–311, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rosa-Velardo:2011:ACS

[332] Fernando Rosa-Velardo, María Martos-Salgado, and David de Frutos-Escrig. Accelerations for the coverability set of Petri nets with names. Fundamenta Informaticae, 113(3–4):313–341, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sole:2011:LRB

[333] Marc Solé and Josep Carmona. Light region-based techniques for process discovery. Fundamenta Informaticae, 113(3–4):343–376, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valmari:2011:CSS

[334] Antti Valmari and Henri Hansen. Can stubborn sets be optimal? Fundamenta Informaticae, 113(3–4):377–397, August 2011. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Weidlich:2011:CBP

[335] Matthias Weidlich, Artem Polyvyanyy, Jan Mendling, and Mathias Weske. Causal behavioural profiles — efficient computation, applications, and evaluation. Fundamenta Informaticae, 113(3–4):399–435, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2011:AIVc

[336] Anonymous. Author index volume 113. Fundamenta Informaticae, 113(3–4):437–438, August 2011. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Avron:2012:FVL

[337] Arnon Avron and Beata Konikowska. Finite-valued logics for information processing. Fundamenta Informaticae, 114(1):1–30, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Celani:2012:NMT

[338] Sergio Celani and Ramon Jansana. A note on the model theory for positive modal logic. Fundamenta Informaticae, 114(1):31–54, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chiniforooshan:2012:PA

[339] Ehsan Chiniforooshan, Lila Kari, and Zhi Xu. Pseudopower avoidance. Fundamenta Informaticae, 114(1):55–72, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guerra:2012:ASQ

[340] Esther Guerra and Juan de Lara. An algebraic semantics for QVTrelations check-only transformations. Fundamenta Informaticae, 114(1): 73–101, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2012:EPF

[341] Jia Yu, Fanyu, Kong, Xiangguo Cheng, Rong Hao, and Jianxi Fan. Erratum to the paper: Forward-Secure Identity-Based Public-Key Encryption without Random Oracles. Fundamenta Informaticae, 114(1):103, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). See [281].

Borowiecki:2012:DCG

[342] Piotr Borowiecki and Elżbieta Sidorowicz. Dynamic coloring of graphs. Fundamenta Informaticae, 114(2):105–128, April 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herman:2012:UFL

[343] Grzegorz Herman and Michael Soltys. Unambiguous functions in logarithmic space. Fundamenta Informaticae, 114(2):129–147, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paszynska:2012:GGMa

[344] Anna Paszyńska, Ewa Grabska, and Maciej Paszyński. A graph grammar model of the hp adaptive three dimensional finite element method. Part I. Fundamenta Informaticae, 114(2):149–182, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paszynska:2012:GGMb

[345] Anna Paszyńska, Ewa Grabska, and Maciej Paszyński. A graph grammar model of the hp adaptive three dimensional finite element method. Part II. Fundamenta Informaticae, 114(2):183–201, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2012:EIA

[346] Chih-Chia Yao and Kang Lee. Ellipse invariant algorithm for texture classification. Fundamenta Informaticae, 114(2):203–220, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jaworski:2012:P

[347] Jerzy Jaworski, Mieczysław Kula, Damian Niwiński, and Jerzy Urbanowicz. Preface. Fundamenta Informaticae, 114(3–4):??, August 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Biryukov:2012:CFA

[348] Alex Biryukov and Johann Großschädl. Cryptanalysis of the full AES using GPU-like special-purpose hardware. Fundamenta Informaticae, 114(3–4):221–237, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chmiel:2012:EPC

[349] Krzysztof Chmiel, Anna Grocholewska-Czurylo, and Janusz Stoklosa. Evaluation of PP-1 cipher resistance against differential and linear cryptanalysis in comparison to a DES-like cipher. Fundamenta Informaticae, 114(3–4):239–269, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gangopadhyay:2012:SON

[350] Sugata Gangopadhyay and Brajesh Kumar Singh. On second-order nonlinearities of some D_{0} type bent functions. Fundamenta Informaticae, 114(3–4):271–285, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzeskowiak:2012:AGP

[351] Maciej Grześkowiak. Algorithm for generating primes p and q such that q divides p_{4} ±p_{3} +p_{2} ±p+1. Fundamenta Informaticae, 114(3–4):287–299, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Merai:2012:RPB

[352] László Mérai. Remarks on pseudorandom binary sequences over elliptic curves. Fundamenta Informaticae, 114(3–4):301–308, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mroczkowski:2012:CAS

[353] Piotr Mroczkowski and Janusz Szmidt. The cube attack on stream cipher Trivium and quadraticity tests. Fundamenta Informaticae, 114(3–4): 309–318, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rjasko:2012:AMC

[354] Michal Rjaško and Martin Stanek. Attacking M&M collective signature scheme. Fundamenta Informaticae, 114(3–4):319–323, August 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Roettger:2012:PKC

[355] Eric Roettger and Hugh C. Williams. Public-key cryptography based on a cubic extension of the Lucas functions. Fundamenta Informaticae, 114(3–4):325–344, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Spiez:2012:RCT

[356] Stanisław Spież, Marian Srebrny, and Jerzy Urbanowicz. Remarks on the classical threshold secret sharing schemes. Fundamenta Informaticae, 114(3–4):345–357, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zajac:2012:STB

[357] Pavol Zajac. Solving Trivium-based Boolean equations using the method of syllogisms. Fundamenta Informaticae, 114(3–4):359–373, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVa

[358] Anonymous. Author index volume 114. Fundamenta Informaticae, 114(3–4):375–376, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2012:P

[359] Davide Ciucci, Alberto Dennunzio, and Roberto Leporini. Preface. Fundamenta Informaticae, 115(1):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beltrametti:2012:EQC

[360] Enrico Beltrametti, Maria Luisa Dalla Chiara, Roberto Giuntini, Roberto Leporini, and Giuseppe Sergioli. Epistemic quantum computational structures in a Hilbert-space environment. Fundamenta Informaticae, 115(1):1–14, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buhagiar:2012:CIP

[361] David Buhagiar, Emmanuel Chetcuti, and Anatolij Dvurečenskij. Completeness of inner product spaces and G. Cattaneo. Fundamenta Informaticae, 115(1):15–24, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Greco:2012:BCM

[362] Salvatore Greco, Benedetto Matarazzo, and Roman Słowiński. The bipolar complemented De Morgan Brouwer–Zadeh distributive lattice as an algebraic structure for the dominance-based rough set approach. Fundamenta Informaticae, 115(1):25–56, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2012:TDI

[363] Davide Ciucci. Temporal dynamics in information tables. Fundamenta Informaticae, 115(1):57–74, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DiLena:2012:UAP

[364] Pietro Di Lena and Luciano Margara. On the undecidability of attractor properties for cellular automata. Fundamenta Informaticae, 115(1):75– 85, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dennunzio:2012:ODT

[365] Alberto Dennunzio. From one-dimensional to two-dimensional cellular automata. Fundamenta Informaticae, 115(1):87–105, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Formenti:2012:CCA

[366] Enrico Formenti, Eric Goles, and Bruno Martin. Computational complexity of avalanches in the Kadanoff sandpile model. Fundamenta Informaticae, 115(1):107–124, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertoni:2012:SCD

[367] Alberto Bertoni, Massimiliano Goldwurm, Jianyi Lin, and Francesco Saccà. Size constrained distance clustering: Separation properties and some complexity results. Fundamenta Informaticae, 115(1):125–139, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2012:P

[368] Guoyin Wang, Andrzej Skowron, Yiyu Yao, and Hong Yu. Preface. Fundamenta Informaticae, 115(2–3):??, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2012:AKO

[369] Hong Yu, Shuangshuang Chu, and Dachun Yang. Autonomous knowledge-oriented clustering using decision-theoretic rough set theory. Fundamenta Informaticae, 115(2–3):141–156, April 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2012:MMA

[370] Xiaoping Yang and JingTao Yao. Modelling multi-agent three-way decisions with decision-theoretic rough sets. Fundamenta Informaticae, 115 (2–3):157–171, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2012:MCC

[371] Dun Liu, Tianrui Li, and Huaxiong Li. A multiple-category classification approach with decision-theoretic rough sets. Fundamenta Informaticae, 115(2–3):173–188, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

An:2012:SME

[372] Shuang An, Qinghua Hu, Daren Yu, and Jinfu Liu. Soft minimumenclosing-ball based robust fuzzy rough sets. Fundamenta Informaticae, 115(2–3):189–202, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2012:KRR

[373] Wei-Zhi Wu. Knowledge reduction in random incomplete decision tables via evidence theory. Fundamenta Informaticae, 115(2–3):203–218, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2012:ARU

[374] GuoYin Wang and Jun Hu. Attribute reduction using extension of covering approximation space. Fundamenta Informaticae, 115(2–3):219–232, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Joshi:2012:CFR

[375] Manish Joshi, Pawan Lingras, and C. Raghavendra Rao. Correlating fuzzy and rough clustering. Fundamenta Informaticae, 115(2–3):233– 246, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2012:STA

[376] Yiyu Yao, Nan Zhang, Duoqian Miao, and Feifei Xu. Set-theoretic approaches to granular computing. Fundamenta Informaticae, 115(2–3): 247–264, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryszkiewicz:2012:P

[377] Marzena Kryszkiewicz, Sergei Obiedkov, and Zbigniew W. Raś. Preface. Fundamenta Informaticae, 115(4):??, December 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pankratieva:2012:RBP

[378] Vera V. Pankratieva and Sergei O. Kuznetsov. Relations between protofuzzy concepts, crisply generated fuzzy concepts, and interval pattern structures. Fundamenta Informaticae, 115(4):265–277, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Belohlavek:2012:RCS

[379] Radim Belohlavek and Jan Konecny. Row and column spaces of matrices over residuated lattices. Fundamenta Informaticae, 115(4):279–295, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kridlo:2012:CCC

[380] Ondrej Krídlo, Stanislav Krajči, and Manuel Ojeda-Aciego. The category of L-Chu correspondences and the structure of L-bonds. Fundamenta Informaticae, 115(4):297–325, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dolques:2012:FGD

[381] Xavier Dolques, Marianne Huchard, Clémentine Nebut, and Philippe Reitz. Fixing generalization defects in UML use case diagrams. Fundamenta Informaticae, 115(4):327–356, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Missaoui:2012:CIN

[382] Rokia Missaoui, Lhouari Nourine, and Yoan Renaud. Computing implications with negation from a formal context. Fundamenta Informaticae, 115(4):357–375, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Revenko:2012:AEP

[383] Artem Revenko and Sergei O. Kuznetsov. Attribute exploration of properties of functions on sets. Fundamenta Informaticae, 115(4):377–394, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krajca:2012:CFC

[384] Petr Krajca, Jan Outrata, and Vilem Vychodil. Computing formal concepts by attribute sorting. Fundamenta Informaticae, 115(4):395–417, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVb

[385] Anonymous. Author index volume 115. Fundamenta Informaticae, 115(4):419–420, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:2012:WGA

[386] Vesa Halava, Juhani Karhumäki, Dirk Nowotka, and Grzegorz Rozenberg. Words, graphs, automata, and languages; special issue honoring the 60th birthday of Professor Tero Harju. Fundamenta Informaticae, 116(1–4):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bell:2012:CCM

[387] Paul C. Bell and Igor Potapov. On the computational complexity of matrix semigroup problems. Fundamenta Informaticae, 116(1–4):1–13, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brijder:2012:BSM

[388] Robert Brijder and Hendrik Jan Hoogeboom. Binary symmetric matrix inversion through local complementation. Fundamenta Informaticae, 116 (1–4):15–23, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bucci:2012:SCS

[389] Michelangelo Bucci, Alessandro De Luca, and Luca Q. Zamboni. Some characterizations of Sturmian words in terms of the lexicographic order. Fundamenta Informaticae, 116(1–4):25–33, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castiglione:2012:SSF

[390] Giusi Castiglione and Antonio Restivo. On the shuffle of star-free languages. Fundamenta Informaticae, 116(1–4):35–44, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Choffrut:2012:NLD

[391] Christian Choffrut. A note on the logical definability of rational trace languages. Fundamenta Informaticae, 116(1–4):45–50, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czeizler:2012:BAD

[392] Elena Czeizler, Andrzej Mizera, and Ion Petre. A Boolean approach for disentangling the roles of submodules to the global properties of a biomodel. Fundamenta Informaticae, 116(1–4):51–63, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diekert:2012:KRT

[393] Volker Diekert, Manfred Kufleitner, and Benjamin Steinberg. The Krohn–Rhodes theorem and local divisors. Fundamenta Informaticae, 116(1–4):65–77, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gao:2012:SCC

[394] Yuan Gao and Sheng Yu. State complexity of combined operations with union, intersection, star and reversal. Fundamenta Informaticae, 116(1– 4):79–92, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gogin:2012:RCM

[395] Nikita Gogin and Mika Hirvensalo. Recurrent construction of MacWilliams and Chebyshev matrices. Fundamenta Informaticae, 116 (1–4):93–110, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hage:2012:SSS

[396] Jurriaan Hage. Subgroup switching of skew gain graphs. Fundamenta Informaticae, 116(1–4):111–122, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honkala:2012:CRL

[397] Juha Honkala. A characterization of regular languages as equality sets of HDT0L sequences. Fundamenta Informaticae, 116(1–4):123–128, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ibarra:2012:SRS

[398] Oscar H. Ibarra and Ömer Eğecioğlu. A survey of results on stateless multicounter automata. Fundamenta Informaticae, 116(1–4):129–140, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2012:MTC

[399] Jarkko Kari and Bastien Le Gloannec. Modified traffic cellular automaton for the density classification task. Fundamenta Informaticae, 116(1– 4):141–156, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karki:2012:RFC

[400] Tomi Kärki. Repetition-freeness with cyclic relations and chain relations. Fundamenta Informaticae, 116(1–4):157–174, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lacroix:2012:SCU

[401] Anne Lacroix, Narad Rampersad, Michel Rigo, and Élise Vandomme. Syntactic complexity of ultimately periodic sets of integers and application to a decision procedure. Fundamenta Informaticae, 116(1–4): 175–187, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Latteux:2012:ORG

[402] Michel Latteux and Yves Roos. On one-rule grid semi-Thue systems. Fundamenta Informaticae, 116(1–4):189–204, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okhotin:2012:LES

[403] Alexander Okhotin. Language equations with symmetric difference. Fundamenta Informaticae, 116(1–4):205–222, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rampersad:2012:CCU

[404] Narad Rampersad, Jeffrey Shallit, and Zhi Xu. The computational complexity of universality problems for prefixes, suffixes, factors, and subwords of regular languages. Fundamenta Informaticae, 116(1–4):223–236, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salomaa:2012:MIS

[405] Arto Salomaa. Mirror images and schemes for the maximal complexity of nondeterminism. Fundamenta Informaticae, 116(1–4):237–249, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Seebold:2012:LOF

[406] Patrice Séébold. Length-k-overlap-free binary infinite words. Fundamenta Informaticae, 116(1–4):251–263, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVc

[407] Anonymous. Author index volume 116. Fundamenta Informaticae, 116 (1–4):265, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mohanty:2012:F

[408] S. G. Mohanty, S. Rinaldi, and J. Makowsky. Foreword. Fundamenta Informaticae, 117(1–4):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mohanty:2012:R

[409] Sri Gopal Mohanty. Reminiscing over. Fundamenta Informaticae, 117 (1–4):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Archibald:2012:VSC

[410] Margaret Archibald, Arnold Knopfmacher, and Toufik Mansour. Variation statistics on compositions. Fundamenta Informaticae, 117(1–4): 1–17, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beaton:2012:SNS

[411] Nicholas R. Beaton, Philippe Flajolet, Timothy M. Garoni, and Anthony J. Guttmann. Some new self-avoiding walk and polygon models. Fundamenta Informaticae, 117(1–4):19–33, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bilotta:2012:PAB

[412] Stefano Bilotta, Donatella Merlini, Elisa Pergola, and Renzo Pinzani. Pattern 1^{j+1}0^{j} avoiding binary words. Fundamenta Informaticae, 117 (1–4):35–55, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blondeau-Fournier:2012:HLP

[413] Olivier Blondeau-Fournier, Pierre Mathieu, and Trevor A. Welsh. Halflattice paths and Virasoro characters. Fundamenta Informaticae, 117 (1–4):57–83, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bodini:2012:BGB

[414] Olivier Bodini, Danièle Gardy, and Olivier Roussel. Boys-and-girls birthdays and Hadamard products. Fundamenta Informaticae, 117(1–4):85– 101, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bohm:2012:KST

[415] Walter Böhm and Kurt Hornik. A Kolmogorov–Smirnov test for r samples. Fundamenta Informaticae, 117(1–4):103–125, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brennan:2012:VLD

[416] Charlotte Brennan and Simon Mavhungu. Visits to level r by Dyck paths. Fundamenta Informaticae, 117(1–4):127–145, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Charalambides:2012:DAP

[417] Charalambos A. Charalambides. On the distributions of absorbed particles in crossing a field containing absorption points. Fundamenta Informaticae, 117(1–4):147–154, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chuang:2012:SDS

[418] Wan-Chen Chuang, Sen-Peng Eu, Tung-Shan Fu, and Yeh-Jong Pan. On simsun and double simsun permutations avoiding a pattern of length three. Fundamenta Informaticae, 117(1–4):155–177, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fusy:2012:BCI

[419] Éric Fusy. Bijective counting of involutive Baxter permutations. Fundamenta Informaticae, 117(1–4):179–188, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kiraly:2012:FWC

[420] Zoltán Király, Zoltán L. Nagy, Dömötör Pálvölgyi, and Mirkó Visontai. On families of weakly cross-intersecting set-pairs. Fundamenta Informaticae, 117(1–4):189–198, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kotek:2012:RTH

[421] Tomer Kotek and Johann A. Makowsky. A representation theorem for holonomic sequences based on counting lattice paths. Fundamenta Informaticae, 117(1–4):199–213, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuba:2012:GFE

[422] Markus Kuba. Generating functions of embedded trees and lattice paths. Fundamenta Informaticae, 117(1–4):215–227, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kyriakoussis:2012:TNT

[423] Andreas Kyriakoussis and Malvina Vamvakari. On terminating and nonterminating q-Gauss hypergeometric series distributions and the associated q-orthogonal polynomials. Fundamenta Informaticae, 117(1–4): 229–248, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nakano:2012:DTO

[424] Fumihiko Nakano and Taizo Sadahiro. Domino tilings with one diagonal impurity. Fundamenta Informaticae, 117(1–4):249–264, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Panny:2012:LPC

[425] Wolfgang Panny. A lattice path combinatorial approach to Rothe numbers and related convolution results. Fundamenta Informaticae, 117(1– 4):265–277, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Prodinger:2012:DPP

[426] Helmut Prodinger. Dyck paths with parity restrictions for the final runs to the origin: a study of the height. Fundamenta Informaticae, 117(1–4): 279–285, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sullivan:2012:BFO

[427] Shaun Sullivan. Bivariate finite operator calculus and counting ballot paths containing patterns. Fundamenta Informaticae, 117(1–4):287–309, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVd

[428] Anonymous. Author index volume 117. Fundamenta Informaticae, 117 (1–4):311–312, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rensink:2012:P

[429] Arend Rensink, Grzegorz Rozenberg, and Andy Schürr. Preface. Fundamenta Informaticae, 118(1–2):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heindel:2012:APM

[430] Tobias Heindel. Adhesivity with partial maps instead of spans. Fundamenta Informaticae, 118(1–2):1–33, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ehrig:2012:ATS

[431] Hartmut Ehrig, Ulrike Golas, Annegret Habel, Leen Lambers, and Fernando Orejas. M-adhesive transformation systems with nested application conditions. Part 2: Embedding, critical pairs and local confluence. Fundamenta Informaticae, 118(1–2):35–63, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orejas:2012:LGT

[432] Fernando Orejas and Leen Lambers. Lazy graph transformation. Fundamenta Informaticae, 118(1–2):65–96, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jurack:2012:TTC

[433] Stefan Jurack and Gabriele Taentzer. Transformation of typed composite graphs with inheritance and containment structures. Fundamenta Informaticae, 118(1–2):97–134, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Poskitt:2012:HSV

[434] Christopher M. Poskitt and Detlef Plump. Hoare-style verification of graph programs. Fundamenta Informaticae, 118(1–2):135–175, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gadducci:2012:CSS

[435] Fabio Gadducci, Alberto Lluch Lafuente, and Andrea Vandin. Counterpart semantics for a second-order μ-calculus. Fundamenta Informaticae, 118(1–2):177–205, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baskent:2012:PAL

[436] Can Başkent. Public announcement logic in geometric frameworks. Fundamenta Informaticae, 118(3):207–223, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blazewicz:2012:RMG

[437] Jacek Blazewicz and Marta Kasprzak. Reduced-by-matching graphs: Toward simplifying Hamiltonian circuit problem. Fundamenta Informaticae, 118(3):225–244, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bortolussi:2012:SPD

[438] Luca Bortolussi, Liviu P. Dinu, and Andrea Sgarro. Spearman permutation distances and Shannon’s distinguishability. Fundamenta Informaticae, 118(3):245–252, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzecza:2012:TMM

[439] Marcin Grzecza, Stanisław Kasjan, and Andrzej Mróz. Tree matrices and a matrix reduction algorithm of Belitskii. Fundamenta Informaticae, 118 (3):253–279, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liao:2012:NSM

[440] Xin Liao, Qiao yan Wen, Ze li Zhao, and Jie Zhang. A novel steganographic method with four-pixel differencing and modulus function. Fundamenta Informaticae, 118(3):281–289, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2012:DNS

[441] Lidong Wang, Yan Ren, and Xiaodong Liu. Development of near sets within the framework of axiomatic fuzzy sets. Fundamenta Informaticae, 118(3):291–304, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Widuch:2012:LCA

[442] Jacek Widuch. A label correcting algorithm for the bus routing problem. Fundamenta Informaticae, 118(3):305–326, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gambin:2012:P

[443] Anna Gambin and Anna Marciniak-Czochra. Preface. Fundamenta Informaticae, 118(4):??, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bertolusso:2012:SSE

[444] Roberto Bertolusso and Marek Kimmel. Spatial and stochastic effects in a model of viral infection. Fundamenta Informaticae, 118(4):327–343, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vakulenko:2012:FRP

[445] Sergei Vakulenko and Ovidiu Radulescu. Flexible and robust patterning by centralized gene networks. Fundamenta Informaticae, 118(4):345–369, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Archuleta:2012:ACG

[446] Michelle N. Archuleta, Jason E. McDermott, Jeremy S. Edwards, and Haluk Resat. An adaptive coarse graining method for signal Transduction in three dimensions. Fundamenta Informaticae, 118(4):371–384, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Blazewicz:2012:CIC

[447] Jacek Blazewicz and Marta Kasprzak. Complexity issues in computational biology. Fundamenta Informaticae, 118(4):385–401, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Setty:2012:EMM

[448] Yaki Setty, Irun R. Cohen, and David Harel. Executable modeling of morphogenesis: A Turing-inspired approach. Fundamenta Informaticae, 118(4):403–417, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Batmanov:2012:SOP

[449] Kirill Batmanov, Céline Kuttler, Cédric Lhoussaine, and Yasushi Saka. Self-organized patterning by diffusible factors: Roles of a community effect. Fundamenta Informaticae, 118(4):419–461, December 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVe

[450] Anonymous. Author index volume 118. Fundamenta Informaticae, 118(4):463–464, December 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Riff:2012:P

[451] Maria Cristina Riff. Preface. Fundamenta Informaticae, 119(1):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dorzan:2012:AMW

[452] Maria Gisela Dorzán, Edilma Olinda Gagliardi, Mario Guillermo Leguizamón, and Gregorio Hernández Peñalver. Approximations on minimum weight triangulations and minimum weight pseudo-triangulations using ant colony optimization metaheuristic. Fundamenta Informaticae, 119(1):1–27, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garrido:2012:FAH

[453] Pablo Garrido and Carlos Castro. A flexible and adaptive hyper-heuristic approach for (dynamic) capacitated vehicle routing problems. Fundamenta Informaticae, 119(1):29–60, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hristea:2012:FSV

[454] Florentina Hristea and Mihaela Colhon. Feeding syntactic versus semantic knowledge to a knowledge-lean unsupervised word sense disambiguation algorithm with an underlying naïve Bayes model. Fundamenta Informaticae, 119(1):61–86, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hummel:2012:TCM

[455] Szczepan Hummel and Michał Skrzypczak. The topological complexity of MSO+U and related automata models. Fundamenta Informaticae, 119 (1):87–111, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schaefer:2012:RMF

[456] Gerald Schaefer, Qinghua Hu, Huiyu Zhou, James F. Peters, and Aboul Ella Hassanien. Rough C-means and fuzzy rough C-means for colour quantisation. Fundamenta Informaticae, 119(1):113–120, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bourguiba:2012:TEO

[457] Imen Bourguiba and Belaid Moa. Tabular expressions operators. Fundamenta Informaticae, 119(2):121–147, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kosakowska:2012:IAP

[458] Justyna Kosakowska. Inflation algorithms for positive and principal edgebipartite graphs and unit quadratic forms. Fundamenta Informaticae, 119(2):149–162, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mrozek:2012:ATV

[459] Ireneusz Mrozek and Vyacheslav Yarmolik. Antirandom test vectors for BIST in hardware/software systems. Fundamenta Informaticae, 119(2): 163–185, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tan:2012:MIE

[460] Henry Tan, Fedja Hadzic, and Tharam S. Dillon. Mining induced/ embedded subtrees using the level of embedding constraint. Fundamenta Informaticae, 119(2):187–231, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2012:P

[461] Ludwik Czaja. Preface. Fundamenta Informaticae, 119(3–4):??, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amin:2012:DPA

[462] Talha Amin, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Dynamic programming approach for partial decision rule optimization. Fundamenta Informaticae, 119(3–4):233–248, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2012:ERT

[463] Marco Bellia and M. Eugenia Occhiuto. The equivalence of reduction and translation semantics of Java simple closures. Fundamenta Informaticae, 119(3–4):249–264, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2012:EAR

[464] Ludwik Czaja. Exclusive access to resources in distributed shared memory architecture. Fundamenta Informaticae, 119(3–4):265–280, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fioravanti:2012:IRA

[465] Fabio Fioravanti, Alberto Pettorossi, Maurizio Proietti, and Valerio Senni. Improving reachability analysis of infinite state systems by specialization. Fundamenta Informaticae, 119(3–4):281–300, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2012:GNS

[466] Anna Gomolińska and Marcin Wolski. On graded nearness of sets. Fundamenta Informaticae, 119(3–4):301–317, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janusz:2012:USL

[467] Andrzej Janusz, Dominik Ślȩzak, and Hung Son Nguyen. Unsupervised similarity learning from textual data. Fundamenta Informaticae, 119(3– 4):319–336, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leszczynski:2012:UPJ

[468] Paweł Leszczyński and Krzysztof Stencel. Update propagator for joint scalable storage. Fundamenta Informaticae, 119(3–4):337–355, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Oshevskaya:2012:UEH

[469] Elena Oshevskaya, Irina Virbitskaite, and Eike Best. Unifying equivalences for higher dimensional automata. Fundamenta Informaticae, 119(3–4):357–372, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2012:TSB

[470] Wojciech Penczek, Bożena Woźna-Szcześniak, and Andrzej Zbrzezny. Towards SAT-based BMC for LTLK over interleaved interpreted systems. Fundamenta Informaticae, 119(3–4):373–392, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2012:ICD

[471] Roman R. Redziejowski. An improved construction of deterministic Omega-automaton using derivatives. Fundamenta Informaticae, 119(3– 4):393–406, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rzeszotko:2012:MLT

[472] Jarosław Rzeszótko and Sinh Hoa Nguyen. Machine learning for traffic prediction. Fundamenta Informaticae, 119(3–4):407–420, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skowron:2012:RSB

[473] Andrzej Skowron, Jarosław Stepaniuk, Andrzej Jankowski, Jan G. Bazan, and Roman Swiniarski. Rough set based reasoning about changes. Fundamenta Informaticae, 119(3–4):421–437, August 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVf

[474] Anonymous. Author index volume 119. Fundamenta Informaticae, 119(3–4):439–440, August 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chalopin:2012:ELE

[475] Jérémie Chalopin, Yves Métivier, and Thomas Morsellino. Enumeration and leader election in partially anonymous and multi-hop broadcast networks. Fundamenta Informaticae, 120(1):1–27, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazzanti:2012:CES

[476] Stefano Mazzanti. CRN elimination and substitution bases for complexity classes. Fundamenta Informaticae, 120(1):29–58, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qin:2012:RDH

[477] Chuan Qin, Zhi-Hui Wang, Chin-Chen Chang, and KuoNan Chen. Reversible data hiding scheme based on image inpainting. Fundamenta Informaticae, 120(1):59–70, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salomaa:2012:SYM

[478] Arto Salomaa, Kai Salomaa, and Andrew L. Szilard. Sheng Yu (1950– 2012) in memoriam. Fundamenta Informaticae, 120(1):71–74, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suchenek:2012:EYP

[479] Marek A. Suchenek. Elementary yet precise worst-case analysis of Floyd’s heap-construction program. Fundamenta Informaticae, 120(1):75–92, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhu:2012:IAD

[480] Ping Zhu. An improved axiomatic definition of information granulation. Fundamenta Informaticae, 120(1):93–109, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2012:BPI

[481] J. A. Bergstra and C. A. Middelburg. On the behaviours produced by instruction sequences under execution. Fundamenta Informaticae, 120 (2):111–144, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bhandari:2012:VSC

[482] Dinabandhu Bhandari, C. A. Murthy, and Sankar K. Pal. Variance as a stopping criterion for genetic algorithms with elitist model. Fundamenta Informaticae, 120(2):145–164, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dennunzio:2012:CIA

[483] Alberto Dennunzio, Enrico Formenti, and Luca Manzoni. Computing issues of asynchronous CA. Fundamenta Informaticae, 120(2):165–180, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2012:TAC

[484] Artur Niewiadomski, Wojciech Penczek, Agata, Półrola, Maciej Szreter, and Andrzej Zbrzezny. Towards automatic composition of Web services: SAT-based concretisation of abstract scenarios. Fundamenta Informaticae, 120(2):181–203, April 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2012:P

[485] Wojciech Penczek. Preface. Fundamenta Informaticae, 120(3–4):??, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeAngelis:2012:SCP

[486] Emanuele De Angelis, Alberto Pettorossi, and Maurizio Proietti. Synthesizing concurrent programs using answer set programming. Fundamenta Informaticae, 120(3–4):205–229, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Azad:2012:GAC

[487] Mohammad Azad, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Greedy algorithms for construction of approximate tests for decision tables with many-valued decisions. Fundamenta Informaticae, 120(3–4): 231–242, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bashkin:2012:CRD

[488] Vladimir A. Bashkin and Irina A. Lomazova. Cellular resource-driven automata. Fundamenta Informaticae, 120(3–4):243–257, July 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drabik:2012:CMV

[489] Peter Drábik, Andrea Maggiolo-Schettini, and Paolo Milazzo. On conditions for modular verification in systems of synchronising components. Fundamenta Informaticae, 120(3–4):259–274, July 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dworzanski:2012:CBL

[490] Leonid W. Dworzański and Irina A. Lomazova. On compositionality of boundedness and liveness for nested Petri nets. Fundamenta Informaticae, 120(3–4):275–293, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2012:IAS

[491] Damas P. Gruska. Informational analysis of security and integrity. Fundamenta Informaticae, 120(3–4):295–309, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2012:QPN

[492] Damas P. Gruska. Quantification of positive and negative Attacker’s information. Fundamenta Informaticae, 120(3–4):311–324, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2012:CEO

[493] Michael Köhler-Bußmeier and Frank Heitmann. Conservative elementary object systems. Fundamenta Informaticae, 120(3–4):325–339, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Popova-Zeugmann:2012:ACI

[494] Louchka Popova-Zeugmann and Elisabeth Pelz. Algebraical characterisation of interval-timed Petri nets with discrete delays. Fundamenta Informaticae, 120(3–4):341–357, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wisniewski:2012:IML

[495] Piotr Wiśniewski, Marta Burzańska, and Krzysztof Stencel. The impedance mismatch in light of the unified state model. Fundamenta Informaticae, 120(3–4):359–374, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zbrzezny:2012:NTE

[496] Andrzej Zbrzezny. A new translation from ECTL* to SAT. Fundamenta Informaticae, 120(3–4):375–395, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVg

[497] Anonymous. Author index volume 120. Fundamenta Informaticae, 120 (3–4):397–398, July 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghilezan:2012:P

[498] Silvia Ghilezan and Luca Paolini. Preface. Fundamenta Informaticae, 121(1–4):??, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanBakel:2012:CSR

[499] Steffen van Bakel. Completeness and soundness results for X with intersection and union types. Fundamenta Informaticae, 121(1–4):1–41, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carlier:2012:AE

[500] Sébastien Carlier and J. B. Wells. The algebra of expansion. Fundamenta Informaticae, 121(1–4):43–82, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Santo:2012:CSN

[501] José Espírito Santo, Jelena Ivetić, and Silvia Likavec. Characterising strongly normalising intuitionistic terms. Fundamenta Informaticae, 121 (1–4):83–120, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamareddine:2012:RPC

[502] Fairouz Kamareddine, Vincent Rahli, and J. B. Wells. Reducibility proofs in the λ-calculus. Fundamenta Informaticae, 121(1–4):121–152, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamareddine:2012:RSI

[503] Fairouz Kamareddine, Karim Nour, Vincent Rahli, and J. B. Wells. On realisability semantics for intersection types with expansion variables. Fundamenta Informaticae, 121(1–4):153–184, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koletsos:2012:ITT

[504] George Koletsos. Intersection types and termination properties. Fundamenta Informaticae, 121(1–4):185–202, January 2012. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laurent:2012:ITS

[505] Olivier Laurent. Intersection types with subtyping by means of cut elimination. Fundamenta Informaticae, 121(1–4):203–226, January 2012. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Piccolo:2012:SNC

[506] Mauro Piccolo. Strong normalization in the π-calculus with intersection and union types. Fundamenta Informaticae, 121(1–4):227–252, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pimentel:2012:ITP

[507] Elaine Pimentel, Simona Ronchi Della Rocca, and Luca Roversi. Intersection types from a proof-theoretic perspective. Fundamenta Informaticae, 121(1–4):253–274, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stavrinos:2012:TLU

[508] Yiorgos Stavrinos and Anastasia Veneti. Towards a logic for union types. Fundamenta Informaticae, 121(1–4):275–302, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2012:AIVh

[509] Anonymous. Author index volume 121. Fundamenta Informaticae, 121 (1–4):303, January 2012. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kristensen:2013:P

[510] Lars Michael Kristensen, Wojciech Penczek, and Laure Petrucci. Preface. Fundamenta Informaticae, 122(1–2):??, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Reynier:2013:MCS

[511] Pierre-Alain Reynier and Frédéric Servais. Minimal coverability set for Petri nets: Karp and Miller algorithm with pruning. Fundamenta Informaticae, 122(1–2):1–30, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Couvreur:2013:BPG

[512] Jean-Michel Couvreur, Denis Poitrenaud, and Pascal Weil. Branching processes of general Petri nets. Fundamenta Informaticae, 122(1–2):31– 58, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanHee:2013:RSP

[513] Kees M. van Hee, Natalia Sidorova, and Jan Martijn van der Werf. Refinement of synchronizable places with multi-workflow nets. Fundamenta Informaticae, 122(1–2):59–83, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peschanski:2013:PNI

[514] Frédéric Peschanski, Hanna Klaudel, and Raymond Devillers. A Petri net interpretation of open reconfigurable systems. Fundamenta Informaticae, 122(1–2):85–117, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kleijn:2013:MCP

[515] Jetty Kleijn and Maciej Koutny. Mutex causality in processes and traces of general elementary nets. Fundamenta Informaticae, 122(1–2):119–146, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haddad:2013:SAP

[516] Serge Haddad, Jean Mairesse, and Hoang-Thach Nguyen. Synthesis and analysis of product-form Petri nets. Fundamenta Informaticae, 122(1– 2):147–172, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ma:2013:SPP

[517] Zhanshan Sam Ma. Stochastic populations, power law and fitness aggregation in genetic algorithms. Fundamenta Informaticae, 122(3):173–206, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fashandi:2013:CFT

[518] Homa Fashandi and James F. Peters. Crisp and fuzzy topological interior and closure operators with inclusion degree. Theory and applications. Fundamenta Informaticae, 122(3):207–225, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Langmaack:2013:SMR

[519] Hans Langmaack and Andrzej Salwicki. Some methodological remarks inspired by the paper “On inner classes” by A. Igarashi and B. Pierce. Fundamenta Informaticae, 122(3):227–274, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic). Cited paper in Lecture Notes in Computer Science, 1850 129–153 (2000).

Daubner:2013:NCO

[520] Jakub Daubner and Eduard Toman. Neighbourhood of constant order in the interval graph of a random Boolean function. Fundamenta Informaticae, 122(4):275–295, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Graff:2013:MGS

[521] Daniel Graff, Jan Richling, and Matthias Werner. Modeling group scheduling problems in space and time by timed Petri nets. Fundamenta Informaticae, 122(4):297–313, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuppusamy:2013:TBA

[522] Lakshmanan Kuppusamy, Anand Mahendran, and Kamala Krithivasan. On the trade-off between ambiguity and complexity in contextual languages. Fundamenta Informaticae, 122(4):315–326, October 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przybyla-Kasperek:2013:ARS

[523] Małgorzata Przybyła-Kasperek and Alicja Wakulicz-Deja. Application of reduction of the set of conditional attributes in the process of global decision-making. Fundamenta Informaticae, 122(4):327–355, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tzitzikas:2013:IEM

[524] Yannis Tzitzikas and Panagiotis Papadakos. Interactive exploration of multi-dimensional and hierarchical information spaces with real-time preference elicitation. Fundamenta Informaticae, 122(4):357–399, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna-Szczesniak:2013:TEM

[525] Bożena Woźna-Szcześniak and Andrzej Zbrzezny. A translation of the existential model checking problem from MITL to HLTL. Fundamenta Informaticae, 122(4):401–420, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVa

[526] Anonymous. Author index volume 122. Fundamenta Informaticae, 122 (4):421, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciabattoni:2013:P

[527] Agata Ciabattoni, Rusin Freivalds, Antonín Kučera, Igor Potapov, and Stefan Szeider. Preface. Fundamenta Informaticae, 123(1):??, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bell:2013:DPP

[528] Paul C. Bell, Vesa Halava, and Mika Hirvensalo. Decision problems for probabilistic finite automata on bounded languages. Fundamenta Informaticae, 123(1):1–14, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chaloupka:2013:RPG

[529] Jakub Chaloupka. Z-reachability problem for games on 2-dimensional vector addition systems with states is in P. Fundamenta Informaticae, 123(1):15–42, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fasching:2013:GHG

[530] Oliver Fasching. Gödel homomorphisms as Gödel modal operators. Fundamenta Informaticae, 123(1):43–57, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ganian:2013:BAS

[531] Robert Ganian, Petr Hliněný, and Jan Obdržálek. Better algorithms for satisfiability problems for formulas of bounded rank-width. Fundamenta Informaticae, 123(1):59–76, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Muino:2013:CRG

[532] David Picado Muiño. A consequence relation for graded inference within the frame of infinite-valued Łukasiewicz logic. Fundamenta Informaticae, 123(1):77–95, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sawa:2013:ECS

[533] Zdeněk Sawa. Efficient construction of semilinear representations of languages accepted by unary nondeterministic finite automata. Fundamenta Informaticae, 123(1):97–106, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yakaryilmaz:2013:PPP

[534] Abuzer Yakaryılmaz and A. C. Cem Say. Proving the power of postselection. Fundamenta Informaticae, 123(1):107–134, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mousavi:2013:FFL

[535] Hamoon Mousavi and Jeffrey Shallit. Filtrations of formal languages by arithmetic progressions. Fundamenta Informaticae, 123(2):135–142, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sikora:2013:CCH

[536] Marek Sikora and Adam Gudyś. CHIRA — convex hull based iterative algorithm of rules aggregation. Fundamenta Informaticae, 123(2):143– 170, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sikora:2013:RDR

[537] Marek Sikora. Redefinition of decision rules based on the importance of elementary conditions evaluation. Fundamenta Informaticae, 123(2): 171–197, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanVliet:2013:MDE

[538] Rudy van Vliet and Hendrik Jan Hoogeboom. Making DNA expressions minimal. Fundamenta Informaticae, 123(2):199–226, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanVliet:2013:MNF

[539] Rudy van Vliet and Hendrik Jan Hoogeboom. A minimal normal form for DNA expressions. Fundamenta Informaticae, 123(2):227–243, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balaguer:2013:BON

[540] Sandie Balaguer, Thomas Chatain, and Stefan Haar. Building occurrence nets from reveals relations. Fundamenta Informaticae, 123(3):245–272, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cichon:2013:FPR

[541] Jacek Cichoń and Marek Klonowski. On flooding in the presence of random faults. Fundamenta Informaticae, 123(3):273–287, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meduna:2013:LRC

[542] Alexander Meduna and Petr Zemek. Left random context ET0L grammars. Fundamenta Informaticae, 123(3):289–304, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moniri:2013:FIF

[543] Morteza Moniri. Fuzzy and intuitionistic fuzzy Turing machines. Fundamenta Informaticae, 123(3):305–315, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mroz:2013:CCB

[544] Andrzej Mróz. On the computational complexity of Bongartz’s algorithm. Fundamenta Informaticae, 123(3):317–329, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Siminski:2013:CMV

[545] Krzysztof Simiński. Clustering with missing values. Fundamenta Informaticae, 123(3):331–350, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2013:CSO

[546] Changzhong Wang, Degang Chen, Qiang He, and Qinhua Hu. A comparative study of ordered and covering information systems. Fundamenta Informaticae, 123(3):351–363, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2013:CHG

[547] Xibei Yang, Yuhua Qian, and Jingyu Yang. On characterizing hierarchies of granulation structures via distances. Fundamenta Informaticae, 123 (3):365–380, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cojocaru:2013:CSL

[548] Liliana Cojocaru and Erkki Mäkinen. The complexity of Szilard languages of matrix grammars revisited. Fundamenta Informaticae, 123(4): 381–399, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2013:ROU

[549] Feng Jiang, Xiaoyan Wan, Yuefei Sui, Cungen Cao, and Junwei Du. Relational operations and uncertainty measure in rough relational database. Fundamenta Informaticae, 123(4):401–416, October 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schmitt:2013:ATC

[550] Dominique Schmitt and Jean-Claude Spehner. Araucaria trees: Construction and grafting theorems. Fundamenta Informaticae, 123(4):417– 445, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2013:ADM

[551] Daniel Simson. Algorithms determining matrix morsifications, Weyl orbits, Coxeter polynomials and mesh geometries of roots for Dynkin diagrams. Fundamenta Informaticae, 123(4):447–490, October 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVb

[552] Anonymous. Author index volume 123. Fundamenta Informaticae, 123 (4):491–492, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cui:2013:P

[553] Zhihua Cui, Sheela Ramanna, James F. Peters, and Sankar K. Pal. Preface. Fundamenta Informaticae, 124(1–2):??, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hidalgo-Herrero:2013:CPS

[554] Mercedes Hidalgo-Herrero, Pablo Rabanal, Ismael Rodríguez, and Fernando Rubio. Comparing problem solving strategies for NP-hard optimization problems. Fundamenta Informaticae, 124(1–2):1–25, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Virginia:2013:LBD

[555] Gloria Virginia and Hung Son Nguyen. Lexicon-based document representation. Fundamenta Informaticae, 124(1–2):27–46, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hruschka:2013:CSR

[556] E. R. Hruschka, Jr., M. C. Duarte, and M. C. Nicoletti. Coupling as strategy for reducing concept-drift in never-ending learning environments. Fundamenta Informaticae, 124(1–2):47–61, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pal:2013:UTR

[557] Sankar K. Pal and Debarati Chakraborty. Unsupervised tracking, roughness and quantitative indices. Fundamenta Informaticae, 124(1–2):63–90, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabska:2013:KRH

[558] Ewa Grabska, Grażyna Ślusarczyk, and Szymon Gajek. Knowledge representation for human-computer interaction in a system supporting conceptual design. Fundamenta Informaticae, 124(1–2):91–110, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saito:2013:SAP

[559] J. H. Saito, J.-F. Mari, E. Pedrino, J. B. Destro-Filho, and M. C. Nicoletti. Simulated activation patterns of biological neurons cultured onto a multi-electrode array based on a modified Izhikevich’s model. Fundamenta Informaticae, 124(1–2):111–132, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nicoletti:2013:ALT

[560] M. C. Nicoletti, F. O. S. S. Lisboa, and E. R. Hruschka, Jr. Automatic learning of temporal relations under the closed world assumption. Fundamenta Informaticae, 124(1–2):133–151, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2013:RRF

[561] Pradipta Maji and Sushmita Paul. Robust rough-fuzzy C-means algorithm: Design and applications in coding and non-coding RNA expression data clustering. Fundamenta Informaticae, 124(1–2):153–174, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2013:SNN

[562] James F. Peters and Doungrat Chitcharoen. Sufficiently near neighbourhoods of points in flow graphs. A near set approach. Fundamenta Informaticae, 124(1–2):175–196, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pal:2013:INA

[563] Sankar K. Pal, Romi Banerjee, Soumitra Dutta, and Samar Sen Sarma. An insight into the Z-number approach to CWW. Fundamenta Informaticae, 124(1–2):197–229, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ozturk:2013:SNA

[564] Mehmet Ali Öztürk and Ebubekir İnan. Soft nearness approximation spaces. Fundamenta Informaticae, 124(1–2):231–250, January 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalyanasundaram:2013:LBF

[565] Bala Kalyanasundaram and Mahe Velauthapillai. Learning behaviors of functions with teams. Fundamenta Informaticae, 124(3):251–270, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ma:2013:CCS

[566] Hui Ma, René Noack, Klaus-Dieter Schewe, Bernhard Thalheim, and Qing Wang. Complete conceptual schema algebras. Fundamenta Informaticae, 124(3):271–295, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Odabas:2013:RCC

[567] Zeynep Nihan Odabaş and Aysun Aytaç. Residual closeness in cycles and related networks. Fundamenta Informaticae, 124(3):297–307, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2013:FCS

[568] Daniel Simson. A framework for Coxeter spectral analysis of edgebipartite graphs, their rational morsifications and mesh geometries of root orbits. Fundamenta Informaticae, 124(3):309–338, July 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2013:TAM

[569] Daniel Simson. Toroidal algorithms for mesh geometries of root orbits of the Dynkin diagram D_{4}. Fundamenta Informaticae, 124(3):339–364, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fioravanti:2013:P

[570] Fabio Fioravanti, Alberto Pettorossi, and Gianfranco Rossi. Preface. Fundamenta Informaticae, 124(4):??, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alberti:2013:CBI

[571] Marco Alberti, Marco Gavanelli, and Evelina Lamma. The CHR-based implementation of the SCIFF abductive system. Fundamenta Informaticae, 124(4):365–381, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bistarelli:2013:CAA

[572] Stefano Bistarelli and Francesco Santini. Coalitions of arguments: An approach with constraint programming. Fundamenta Informaticae, 124 (4):383–401, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Campagna:2013:PPP

[573] Dario Campagna and Andrea Formisano. Product and production process modeling and configuration. Fundamenta Informaticae, 124(4):403– 425, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cantone:2013:SPL

[574] Domenico Cantone and Marianna Nicolosi Asmundo. On the satisfiability problem for a 4-level quantified syllogistic and some applications to modal logic. Fundamenta Informaticae, 124(4):427–448, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Costantini:2013:NWC

[575] Stefania Costantini and Andrea Formisano. Nested weight constraints in ASP. Fundamenta Informaticae, 124(4):449–464, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DAgostino:2013:MCS

[576] Giovanna D’Agostino and Giacomo Lenzi. On modal μ-calculus in S5 and applications. Fundamenta Informaticae, 124(4):465–482, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fioravanti:2013:CPS

[577] Fabio Fioravanti, Alberto Pettorossi, Maurizio Proietti, and Valerio Senni. Controlling polyvariance for specialization-based verification. Fundamenta Informaticae, 124(4):483–502, October 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lisi:2013:LBC

[578] Francesca A. Lisi and Umberto Straccia. A logic-based computational method for the automated induction of fuzzy ontology axioms. Fundamenta Informaticae, 124(4):503–519, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Riguzzi:2013:MMC

[579] Fabrizio Riguzzi. MCINTYRE: A Monte Carlo system for probabilistic logic programming. Fundamenta Informaticae, 124(4):521–541, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Solomakhin:2013:LBR

[580] Dmitry Solomakhin, Enrico Franconi, and Alessandro Mosca. Logicbased reasoning support for SBVR. Fundamenta Informaticae, 124(4): 543–560, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhou:2013:TPP

[581] Neng-Fa Zhou and Agostino Dovier. A tabled Prolog program for solving Sokoban. Fundamenta Informaticae, 124(4):561–575, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVc

[582] Anonymous. Author index volume 124. Fundamenta Informaticae, 124 (4):577–578, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abid:2013:LVU

[583] Chiheb Ameur Abid and Belhassen Zouari. Local verification using a distributed state space. Fundamenta Informaticae, 125(1):1–20, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Felisiak:2013:CAT

[584] Mariusz Felisiak. Computer algebra technique for Coxeter spectral study of edge-bipartite graphs and matrix morsifications of Dynkin type A_{n}. Fundamenta Informaticae, 125(1):21–49, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kamide:2013:RAR

[585] Norihiro Kamide and Ken Kaneiwa. Reasoning about resources and information: A linear logic approach. Fundamenta Informaticae, 125 (1):51–70, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rotaru:2013:CGA

[586] Armand Stefan Rotaru and Sorin Iftene. A complete generalization of Atkin’s square root algorithm. Fundamenta Informaticae, 125(1):71–94, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tyszka:2013:DTE

[587] Apoloniusz Tyszka. Does there exist an algorithm which to each Diophantine equation assigns an integer which is greater than the modulus of integer solutions, if these solutions form a finite set? Fundamenta Informaticae, 125(1):95–99, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aravantinos:2013:RCF

[588] Vincent Aravantinos, Mnacho Echenim, and Nicolas Peltier. A resolution calculus for first-order schemata. Fundamenta Informaticae, 125(2):101– 133, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dong:2013:DAL

[589] Jianwu Dong, Feng Chen, Yanyan Huo, and Hong Liu. Decomposition and approximation of loopy Bayesian networks. Fundamenta Informaticae, 125(2):135–152, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Erves:2013:WDP

[590] Rija Erveš and Janez Žerovnik. Wide-diameter of product graphs. Fundamenta Informaticae, 125(2):153–160, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzeskowiak:2013:ARC

[591] Maciej Grześkowiak. Algorithms for relatively cyclotomic primes. Fundamenta Informaticae, 125(2):161–181, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kosiuczenko:2013:VIR

[592] Piotr Kosiuczenko. On the validation of invariants at runtime. Fundamenta Informaticae, 125(2):183–222, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2013:P

[593] P. Dulio, A. Frosini, and G. Rozenberg. Preface. Fundamenta Informaticae, 125(3–4):??, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Segers:2013:DTM

[594] Hilde Segers, Willem Jan Palenstijn, Kees Joost Batenburg, and Jan Sijbers. Discrete tomography in MRI: a simulation study. Fundamenta Informaticae, 125(3–4):223–237, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Batenburg:2013:ADR

[595] K. Joost Batenburg, Wagner Fortes, and Robert Tijdeman. Approximate discrete reconstruction algorithm. Fundamenta Informaticae, 125(3–4): 239–259, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fiorini:2013:DTDa

[596] Rodolfo A. Fiorini and Giuseppe Laguteta. Discrete tomography data footprint reduction by information conservation. Fundamenta Informaticae, 125(3–4):261–272, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fiorini:2013:DTDb

[597] Rodolfo A. Fiorini, Andrea Condorelli, and Giuseppe Laguteta. Discrete tomography data footprint reduction via natural compression. Fundamenta Informaticae, 125(3–4):273–284, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Petra:2013:CPV

[598] Stefania Petra, Christoph Schnörr, and Andreas Schröder. Critical parameter values and reconstruction properties of discrete tomography: Application to experimental fluid dynamics. Fundamenta Informaticae, 125(3–4):285–312, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barcucci:2013:SMD

[599] Elena Barcucci and Stefano Brocchi. Solving multicolor discrete tomography problems by using prior knowledge. Fundamenta Informaticae, 125 (3–4):313–328, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bilotta:2013:PDP

[600] Stefano Bilotta, Simone Rinaldi, and Samanta Socci. Polygons drawn from permutations. Fundamenta Informaticae, 125(3–4):329–342, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hantos:2013:RPH

[601] Norbert Hantos and Péter Balázs. The reconstruction of polyominoes from horizontal and vertical projections and morphological skeleton is NP-complete. Fundamenta Informaticae, 125(3–4):343–359, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fedele:2013:RPM

[602] Roberto Fedele, Antonia Ciani, Luca Galantucci, Matteo Bettuzzi, and Luca Andena. A regularized, pyramidal multi-grid approach to global 3D-volume digital image correlation based on X-ray micro-tomography. Fundamenta Informaticae, 125(3–4):361–376, July 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVd

[603] Anonymous. Author index volume 125. Fundamenta Informaticae, 125 (3–4):377–378, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baader:2013:LEO

[604] Franz Baader and Alexander Okhotin. On language equations with onesided concatenation. Fundamenta Informaticae, 126(1):1–35, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berghammer:2013:DRC

[605] Rudolf Berghammer and Michael Winter. Decomposition of relations and concept lattices. Fundamenta Informaticae, 126(1):37–82, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nagarajan:2013:ARS

[606] E. K. R. Nagarajan and D. Umadevi. Algebra of rough sets based on quasi order. Fundamenta Informaticae, 126(1):83–101, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pan:2013:CHF

[607] Ping Pan, Licheng Wang, Yixian Yang, Yuanju Gan, Lihua Wang, and Chengqian Xu. Chameleon hash functions and one-time signature schemes from inner automorphism groups. Fundamenta Informaticae, 126(1):103–119, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cervelle:2013:P

[608] Julien Cervelle, Alberto Dennunzio, Enrico Formenti, and Andrzej Skowron. Preface. Fundamenta Informaticae, 126(2–3):??, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arrighi:2013:SCA

[609] Pablo Arrighi, Nicolas Schabanel, and Guillaume Theyssier. Stochastic cellular automata: Correlations, decidability and simulations. Fundamenta Informaticae, 126(2–3):121–156, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wacker:2013:CDP

[610] Simon Wacker and Thomas Worsch. On completeness and decidability of phase space invertible asynchronous cellular automata. Fundamenta Informaticae, 126(2–3):157–181, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dennunzio:2013:POD

[611] Alberto Dennunzio, Pietro Di Lena, Enrico Formenti, and Luciano Margara. Periodic orbits and dynamical complexity in cellular automata. Fundamenta Informaticae, 126(2–3):183–199, April 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2013:ODC

[612] Martin Kutrib and Andreas Malcher. One-dimensional cellular automaton transducers. Fundamenta Informaticae, 126(2–3):201–224, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pighizzini:2013:TWF

[613] Giovanni Pighizzini. Two-way finite automata: Old and recent results. Fundamenta Informaticae, 126(2–3):225–246, April 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Imai:2013:SUS

[614] Katsunobu Imai, Takahiro Hatsuda, Victor Poupet, and Kota Sato. A 6-state universal semi-totalistic cellular automaton on kite and dart Penrose tilings. Fundamenta Informaticae, 126(2–3):247–261, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salo:2013:CCS

[615] Ville Salo and Ilkka Törmä. Constructions with countable subshifts of finite type. Fundamenta Informaticae, 126(2–3):263–300, April 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2013:P

[616] JingTao Yao, Andrzej Skowron, Guoyin Wang, and Hung Son Nguyen. Preface. Fundamenta Informaticae, 126(4):??, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2013:ICK

[617] Yan Yang, Tonny Rutayisire, Chao Lin, Tianrui Li, and Fei Teng. An improved Cop–Kmeans clustering for solving constraint violation based on MapReduce framework. Fundamenta Informaticae, 126(4):301–318, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hatamlou:2013:HGS

[618] Abdolreza Hatamlou and Masoumeh Hatamlou. Hybridization of the gravitational search algorithm and big bang–big crunch algorithm for data clustering. Fundamenta Informaticae, 126(4):319–333, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2013:RCT

[619] Zbigniew Suraj and Krzysztof Pancerz. ROSECON: a computer tool for synthesis and verification of concurrent systems specified by information systems. Fundamenta Informaticae, 126(4):335–351, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2013:DTR

[620] Xiaoping Yang, Zunjing Lu, and Tong-Jun Li. Decision-theoretic rough sets in incomplete information system. Fundamenta Informaticae, 126 (4):353–375, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deng:2013:NME

[621] Weibin Deng, Feng Hu, Guoyin Wang, Jerzy Błaszczyński, Roman Słowiński, and Marcin Szela̧g. A novel method for elimination of inconsistencies in ordinal classification with monotonicity constraints. Fundamenta Informaticae, 126(4):377–395, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shao:2013:VBA

[622] Mingwen Shao, Min Liu, and Li Guo. Vector-based attribute reduction method for formal contexts. Fundamenta Informaticae, 126(4):397–414, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2013:NMA

[623] Huaxiong Li, Xianzhong Zhou, Jiabao Zhao, and Dun Liu. Nonmonotonic attribute reduction in decision-theoretic rough sets. Fundamenta Informaticae, 126(4):415–432, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2013:NSL

[624] James F. Peters. Nearness of sets in local admissible covers. Theory and application in micropalaeontology. Fundamenta Informaticae, 126 (4):433–444, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVe

[625] Anonymous. Author index volume 126. Fundamenta Informaticae, 126 (4):445–446, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:P

[626] Anonymous. Preface. Fundamenta Informaticae, 127(1–4):??, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:2013:RSS

[627] Mihir K. Chakraborty and Mohua Banerjee. Rough sets: Some foundational issues. Fundamenta Informaticae, 127(1–4):1–15, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2013:CFR

[628] Marcin Wolski and Anna Gomolińska. Concept formation: Rough sets and Scott systems. Fundamenta Informaticae, 127(1–4):17–33, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2013:DDR

[629] Ivo Düntsch and Ewa Orłowska. Discrete duality for rough relation algebras. Fundamenta Informaticae, 127(1–4):35–47, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yao:2013:DRS

[630] Yiyu Yao. Duality in rough set theory based on the square of opposition. Fundamenta Informaticae, 127(1–4):49–64, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Semeniuk-Polkowska:2013:NEA

[631] Maria Semeniuk-Polkowska and Lech Polkowski. On a notion of extensionality for artifacts. Fundamenta Informaticae, 127(1–4):65–80, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2013:RNS

[632] Ludwik Czaja. Rough net structures — example of information system. Fundamenta Informaticae, 127(1–4):81–97, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2013:BMR

[633] Gheorghe Păun, Mario J. Pérez-Jiménez, and Grzegorz Rozenberg. Bridging membrane and reaction systems — further results and research topics. Fundamenta Informaticae, 127(1–4):99–114, January 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fioravanti:2013:PTP

[634] Fabio Fioravanti, Alberto Pettorossi, Maurizio Proietti, and Valerio Senni. Proving theorems by program transformation. Fundamenta Informaticae, 127(1–4):115–134, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doherty:2013:AGL

[635] Patrick Doherty and Andrzej Szałas. Automated generation of logical constraints on approximation spaces using quantifier elimination. Fundamenta Informaticae, 127(1–4):135–149, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amin:2013:CBO

[636] Talha Amin, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Classifiers based on optimal decision rules. Fundamenta Informaticae, 127(1– 4):151–160, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Greco:2013:FMB

[637] Salvatore Greco, Roman Słowiński, and Izabela Szczȩch. Finding meaningful Bayesian confirmation measures. Fundamenta Informaticae, 127 (1–4):161–176, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:2013:ECT

[638] Patrick G. Clark and Jerzy W. Grzymala-Busse. An experimental comparison of three probabilistic approximations used for rule induction. Fundamenta Informaticae, 127(1–4):177–191, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ziarko:2013:WDA

[639] Wojciech Ziarko and Xugunag Chen. Weak dependencies in approximation spaces. Fundamenta Informaticae, 127(1–4):193–207, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tsumoto:2013:IIM

[640] Shusaku Tsumoto and Shoji Hirano. Incremental induction of medical diagnostic rules based on incremental sampling scheme and SubRule layers. Fundamenta Informaticae, 127(1–4):209–223, January 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Touati:2013:MMA

[641] Hakim Touati and Zbigniew W. Ras. Mining meta-actions for action rules reduction. Fundamenta Informaticae, 127(1–4):225–240, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2013:LMC

[642] Hui Wang, Ivo Düntsch, and Luis Trindade. Lattice machine classification based on contextual probability. Fundamenta Informaticae, 127(1– 4):241–256, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pancerz:2013:RSA

[643] Krzysztof Pancerz and Zbigniew Suraj. A rough set approach to information systems decomposition. Fundamenta Informaticae, 127(1–4): 257–272, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kruczyk:2013:RRM

[644] Marcin Kruczyk, Nicholas Baltzer, Jakub Mieczkowski, Michał Dramiński, Jacek Koronacki, and Jan Komorowski. Random reducts: A Monte Carlo rough set-based method for feature selection in large datasets. Fundamenta Informaticae, 127(1–4):273–288, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pal:2013:WTD

[645] Jayanta Kumar Pal, Shubhra Sankar Ray, and Sankar K. Pal. A weighted threshold for detection of cancerous miRNA expressions. Fundamenta Informaticae, 127(1–4):289–305, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryszkiewicz:2013:UNZ

[646] Marzena Kryszkiewicz. Using non-zero dimensions for the cosine and Tanimoto similarity search among real valued vectors. Fundamenta Informaticae, 127(1–4):307–323, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kostek:2013:MRB

[647] Bozena Kostek and Andrzej Kaczmarek. Music recommendation based on multidimensional description and similarity measures. Fundamenta Informaticae, 127(1–4):325–340, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wakulicz-Deja:2013:CDS

[648] Alicja Wakulicz-Deja, Agnieszka Nowak-Brzezińska, and Małgorzata Przybyła-Kasperek. Complex decision systems and conflicts analysis problem. Fundamenta Informaticae, 127(1–4):341–356, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2013:TTB

[649] James F. Peters and Sheela Ramanna. Treasure trove at Banacha. Set patterns in descriptive proximity spaces. Fundamenta Informaticae, 127(1–4):357–367, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2013:APR

[650] Sinh Hoa Nguyen, Tuan Trung Nguyen, Marcin Szczuka, and Hung Son Nguyen. An approach to pattern recognition based on hierarchical granular computing. Fundamenta Informaticae, 127(1–4):369–384, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2013:GCB

[651] Yuchao Liu, Deyi Li, Wen He, and Guoyin Wang. Granular computing based on Gaussian cloud transformation. Fundamenta Informaticae, 127(1–4):385–398, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pedrycz:2013:NGD

[652] Witold Pedrycz. From numeric to granular description and interpretation of information granules. Fundamenta Informaticae, 127(1–4):399–412, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lin:2013:URS

[653] Tsau Young Lin, Yong Liu, and Wenliang Huang. Unifying rough set theories via large scaled granular computing. Fundamenta Informaticae, 127(1–4):413–428, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stepaniuk:2013:FST

[654] Jaroslaw Stepaniuk, Maciej Kopczynski, and Tomasz Grzes. The first step toward processor for rough set methods. Fundamenta Informaticae, 127(1–4):429–443, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slezak:2013:TDR

[655] Dominik Ślȩzak, Piotr Synak, Arkadiusz Wojna, and Jakub Wróblewski. Two database related interpretations of rough approximations: Data organization and query execution. Fundamenta Informaticae, 127(1–4): 445–459, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lewandowski:2013:GHE

[656] Jacek Lewandowski and Henryk Rybiński. Generalized hybrid encoding of polyhierarchical structures. Fundamenta Informaticae, 127(1–4):461– 477, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2013:UBA

[657] Hao Lan Zhang, Jiming Liu, and Yanchun Zhang. Utilizing BDI agents and a topological theory for mining online social networks. Fundamenta Informaticae, 127(1–4):479–494, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2013:DKT

[658] Jan G. Bazan, Sylwia Buregwa-Czuma, and Andrzej W. Jankowski. A domain knowledge as a tool for improving classifiers. Fundamenta Informaticae, 127(1–4):495–511, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Froelich:2013:MTP

[659] Wojciech Froelich, Rafał Deja, and Grażyna Deja. Mining therapeutic patterns from clinical data for juvenile diabetes. Fundamenta Informaticae, 127(1–4):513–528, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krasuski:2013:ODI

[660] Adam Krasuski and Piotr Wasilewski. Outlier detection by interaction with domain experts. Fundamenta Informaticae, 127(1–4):529–544, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qin:2013:DPU

[661] Linchan Qin, Ning Zhong, Shengfu Lu, and Mi Li. Decision prediction using visual patterns. Fundamenta Informaticae, 127(1–4):545–560, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czyzewski:2013:AMA

[662] Andrzej Czyżewski and Karol Lisowski. Adaptive method of adjusting flowgraph for route reconstruction in video surveillance systems. Fundamenta Informaticae, 127(1–4):561–576, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVf

[663] Anonymous. Author index volume 127. Fundamenta Informaticae, 127 (1–4):577–579, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2013:P

[664] Ludwik Czaja. Preface. Fundamenta Informaticae, 128(1–2):??, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Azad:2013:GAC

[665] Mohammad Azad, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. A greedy algorithm for construction of decision trees for tables with many-valued decisions — a comparative study. Fundamenta Informaticae, 128(1–2):1–15, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2013:JST

[666] Marco Bellia and M. Eugenia Occhiuto. Java SAM typed closures: A sound and complete type inference system for nominal types. Fundamenta Informaticae, 128(1–2):17–33, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2013:EAM

[667] Ludwik Czaja. Equations for asynchronous message passing. Fundamenta Informaticae, 128(1–2):35–45, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dabrowski:2013:OGR

[668] Robert Da̧browski, Grzegorz Timoszuk, and Krzysztof Stencel. One graph to rule them all software measurement and management. Fundamenta Informaticae, 128(1–2):47–63, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2013:ADP

[669] Adam Grabowski. Automated discovery of properties of rough sets. Fundamenta Informaticae, 128(1–2):65–79, January 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2013:IFT

[670] Damas P. Gruska. Information flow testing. Fundamenta Informaticae, 128(1–2):81–95, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2013:DMP

[671] Michael Köhler-Bußmeier. Defining multi-party compromises using unfoldings of workflow nets. Fundamenta Informaticae, 128(1–2):97–111, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2013:HLC

[672] Manfred Kudlek and Nils Erik Flick. A hierarchy of languages with catenation and shuffle. Fundamenta Informaticae, 128(1–2):113–128, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2013:ACS

[673] Irina A. Lomazova and Ivan V. Romanov. Analyzing compatibility of services via resource conformance. Fundamenta Informaticae, 128(1–2): 129–141, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pancerz:2013:ABC

[674] Krzysztof Pancerz, Arkadiusz Lewicki, Ryszard Tadeusiewicz, and Jan Warchoł. Ant-based clustering in delta episode information systems based on temporal rough set flow graphs. Fundamenta Informaticae, 128(1–2):143–158, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2013:NVO

[675] James F. Peters, Andrzej Skowron, and Jaroslaw Stepaniuk. Nearness of visual objects. application of rough sets in proximity spaces. Fundamenta Informaticae, 128(1–2):159–176, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2013:EP

[676] Roman R. Redziejowski. From EBNF to PEG. Fundamenta Informaticae, 128(1–2):177–191, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2013:NCF

[677] Zbigniew Suraj. A new class of fuzzy Petri nets for knowledge representation and reasoning. Fundamenta Informaticae, 128(1–2):193–207, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wagler:2013:MEP

[678] Annegret K. Wagler and Jan-Thierry Wegener. On minimality and equivalence of Petri nets. Fundamenta Informaticae, 128(1–2):209–222, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2013:IAA

[679] Marcin Wolski and Anna Gomolińska. An incidence algebra approach to knowledge granulation in Pawlak information systems. Fundamenta Informaticae, 128(1–2):223–238, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yaskorska:2013:PPT

[680] Olena Yaskorska, Katarzyna Budzynska, and Magdalena Kacprzak. Proving propositional tautologies in a natural dialogue. Fundamenta Informaticae, 128(1–2):239–253, January 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hidders:2013:P

[681] Jan Hidders, Paolo Missier, Jacek Sroka, and Jan Van den Bussche. Preface. Fundamenta Informaticae, 128(3):??, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abouelhoda:2013:TSC

[682] Mohamed Abouelhoda, Shady Issa, and Moustafa Ghanem. Towards scalable and cost-aware bioinformatics workflow execution in the cloud — recent advances to the Tavaxy workflow system. Fundamenta Informaticae, 128(3):255–280, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plociennik:2013:ADE

[683] Marcin Płóciennik, Tomasz Żok, Ilkay Altintas, Jianwu Wang, Daniel Crawl, David Abramson, Frederic Imbeaux, Bernard Guillerminet, Marcos Lopez-Caniego, Isabel Campos Plasencia, Wojciech Pych, Pawel Cieciela̧g, Bartek Palak, Michał Owsiak, and Yann Frauel. Approaches to distributed execution of scientific workflows in Kepler. Fundamenta Informaticae, 128(3):281–302, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simitsis:2013:HAF

[684] Alkis Simitsis, Kevin Wilkinson, and Umeshwar Dayal. Hybrid analytic flows — the case for optimization. Fundamenta Informaticae, 128(3): 303–335, July 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozniak:2013:TDM

[685] Justin M. Wozniak, Timothy G. Armstrong, Ketan Maheshwari, Ewing L. Lusk, Daniel S. Katz, Michael Wilde, and Ian T. Foster. Turbine:

A distributed-memory dataflow engine for high performance many-task applications. Fundamenta Informaticae, 128(3):337–366, July 2013. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2013:DLA

[686] J. A. Bergstra and C. A. Middelburg. Data linkage algebra, data linkage dynamics, and priority rewriting. Fundamenta Informaticae, 128(4):367– 412, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferilli:2013:LFI

[687] Stefano Ferilli and Floriana Esposito. A logic framework for incremental learning of process models. Fundamenta Informaticae, 128(4):413–443, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2013:SEH

[688] Yo-Sub Han. State elimination heuristics for short regular expressions. Fundamenta Informaticae, 128(4):445–462, October 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krasinski:2013:AAB

[689] Tadeusz Krasiński, Sebastian Sakowski, Jacek Waldmajer, and Tomasz Popławski. Arithmetical analysis of biomolecular finite automaton. Fundamenta Informaticae, 128(4):463–474, October 2013. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2013:BSB

[690] Xiangyang Wang, Hongying Yang, Jing Wang, Lili Chen, and Panpan Niu. Bayesian segmentation based local geometrically invariant image watermarking. Fundamenta Informaticae, 128(4):475–501, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2013:MDM

[691] Zhaohao Wang, Lan Shu, and Xiuyong Ding. Minimal description and maximal description in covering-based rough sets. Fundamenta Informaticae, 128(4):503–526, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2013:AIVg

[692] Anonymous. Author index volume 128. Fundamenta Informaticae, 128 (4):527–528, October 2013. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2014:Pa

[693] Ludwik Czaja. Preface. Fundamenta Informaticae, 129(1–2):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amin:2014:RBL

[694] Talha Amin, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Relationships between length and coverage of decision rules. Fundamenta Informaticae, 129(1–2):1–13, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chikalov:2014:RBA

[695] Igor Chikalov, Shahid Hussain, and Mikhail Moshkov. Relationships between average depth and number of misclassifications for decision trees. Fundamenta Informaticae, 129(1–2):15–26, January 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2014:SMN

[696] Ludwik Czaja. Self-modifying nets for synchronous, connection-oriented, multicast communication. Fundamenta Informaticae, 129(1–2):27–36, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czyzewska:2014:TAI

[697] Marta Czyżewska, Jarosław Szkoła, and Krzysztof Pancerz. Towards assessment of innovativeness economy determinant correlation: the double self-organizing feature map approach. Fundamenta Informaticae, 129(1– 2):37–48, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herajy:2014:PNB

[698] Mostafa Herajy and Monika Heiner. Petri net-based collaborative simulation and steering of biochemical reaction networks. Fundamenta Informaticae, 129(1–2):49–67, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jankowski:2014:PUR

[699] Andrzej Jankowski, Andrzej Skowron, and Roman W. Swiniarski. Perspectives on uncertainty and risk in rough sets and interactive roughgranular computing. Fundamenta Informaticae, 129(1–2):69–84, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jard:2014:BTF

[700] Claude Jard, Didier Lime, and Olivier H. Roux. Blending timed formal models with clock transition systems. Fundamenta Informaticae, 129(1– 2):85–100, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2014:CRP

[701] Michael Köhler-Bußmeier. On the complexity of the reachability problem for safe, elementary hornets. Fundamenta Informaticae, 129(1–2):101– 115, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kudlek:2014:PLC

[702] Manfred Kudlek and Nils Erik Flick. Properties of languages with catenation and shuffle. Fundamenta Informaticae, 129(1–2):117–132, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurbalija:2014:MOE

[703] Vladimir Kurbalija, Mirjana Ivanović, Charlotte von Bernstorff, Jens Nachtwei, and Hans-Dieter Burkhard. Matching observed with empirical reality — what you see is what you get? Fundamenta Informaticae, 129(1–2):133–147, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2014:BBF

[704] Lech Polkowski and Maria Semeniuk-Polkowska. Boundaries, borders, fences, hedges. Fundamenta Informaticae, 129(1–2):149–159, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szpyrka:2014:FDA

[705] Marcin Szpyrka, Piotr Matyasik, Rafał Mrówka, and Leszek Kotulski. Formal description of Alvis language with α^{0} system layer. Fundamenta Informaticae, 129(1–2):161–176, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wisniewski:2014:UQL

[706] Piotr Wiśniewski and Krzysztof Stencel. Universal query language for unified state model. Fundamenta Informaticae, 129(1–2):177–192, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Appice:2014:MRM

[707] Annalisa Appice, Michelangelo Ceci, and Donato Malerba. Multirelational model tree induction tightly-coupled with a relational database. Fundamenta Informaticae, 129(3):193–224, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Apt:2014:SNC

[708] Krzysztof R. Apt and Evangelos Markakis. Social networks with competing products. Fundamenta Informaticae, 129(3):225–250, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DiMauro:2014:GPR

[709] Nicola Di Mauro, Teresa M. A. Basile, Stefano Ferilli, and Floriana Esposito. Grasp and path-relinking for coalition structure generation. Fundamenta Informaticae, 129(3):251–277, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nemeth:2014:IMT

[710] Gábor Árpád Németh and Zoltán Pap. The incremental maintenance of transition tour. Fundamenta Informaticae, 129(3):279–300, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balbiani:2014:DCB

[711] Philippe Balbiani and Tinko Tinchev. Definability and canonicity for Boolean logic with a binary relation. Fundamenta Informaticae, 129(4): 301–327, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chowdhury:2014:CLC

[712] Shihabur Rahman Chowdhury, Md. Mahbubul Hasan, Sumaiya Iqbal, and M. Sohel Rahman. Computing a longest common palindromic subsequence. Fundamenta Informaticae, 129(4):329–340, October 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szczuka:2014:UDK

[713] Marcin Szczuka, Łukasz Sosnowski, Adam Krasuski, and Karol Kreński. Using domain knowledge in initial stages of KDD: Optimization of compound object processing. Fundamenta Informaticae, 129(4):341–364, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tian:2014:CIC

[714] Miaomiao Tian, Wei Yang, and Liusheng Huang. Cryptanalysis and improvement of a certificateless multi-proxy signature scheme. Fundamenta Informaticae, 129(4):365–375, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2014:RSC

[715] Shiping Wang, Qingxin Zhu, William Zhu, and Fan Min. Rough set characterization for 2-circuit matroid. Fundamenta Informaticae, 129 (4):377–393, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

You:2014:FSM

[716] Lin You, Yilin Yang, Shuhong Gao, and Yongxuan Sang. Fast scalar multiplications on the curve v^{2} = u^{p} − au − b over the finite field of characteristic p. Fundamenta Informaticae, 129(4):395–412, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVa

[717] Anonymous. Author index volume 129. Fundamenta Informaticae, 129 (4):413–414, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Muller:2014:P

[718] Berndt Müller and Melvin Fitting. Preface. Fundamenta Informaticae, 130(1):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baier:2014:SRC

[719] Christel Baier, Joachim Klein, and Sascha Klüppelholz. Synthesis of Reo connectors for strategies and controllers. Fundamenta Informaticae, 130 (1):1–20, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellin:2014:CCI

[720] Gianluigi Bellin and Alessandro Menti. On the π-calculus and cointuitionistic logic. Notes on logic for concurrency and λP systems. Fundamenta Informaticae, 130(1):21–65, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cervesato:2014:SMT

[721] Iliano Cervesato and Thierry Sans. Substructural meta-theory of a typesafe language for Web programming. Fundamenta Informaticae, 130(1): 67–97, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2014:SDR

[722] Michael Köhler-Bußmeier. A survey of decidability results for elementary object systems. Fundamenta Informaticae, 130(1):99–123, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mendler:2014:CIC

[723] Michael Mendler and Stephan Scheele. On the computational interpretation of CK_{n} for contextual information processing. Fundamenta Informaticae, 130(1):125–162, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Luong:2014:P

[724] Chi Mai Luong, Vincenzo Piuri, and Hung Son Nguyen. Preface. Fundamenta Informaticae, 130(2):??, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2014:QAT

[725] Nguyen Dinh Han, Ho Ngoc Vinh, Dang Quyet Thang, and Phan Trung Huy. Quadratic algorithms for testing of codes and ⋄-codes. Fundamenta Informaticae, 130(2):163–177, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoang:2014:EFE

[726] Cuong Hoang, Anh-Cuong Le, Phuong-Thai Nguyen, Son Bao Pham, and Tu Bao Ho. An efficient framework for extracting parallel sentences from non-parallel corpora. Fundamenta Informaticae, 130(2):179–199, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lai:2014:UIS

[727] Hien Phuong Lai, Muriel Visani, Alain Boucher, and Jean-Marc Ogier. Unsupervised and interactive semi-supervised clustering for large image database indexing and retrieval. Fundamenta Informaticae, 130(2):201– 218, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2014:BHA

[728] Ngoc Trung Nguyen and Anh-Duc Duong. A better heuristic algorithm for finding the closest trio of 3-colored points from a given set of 3-colored points on a plane. Fundamenta Informaticae, 130(2):219–229, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pham:2014:HSR

[729] San Pham, Viviane Gascon, and Tien Dinh. Heuristics to solve a realworld asymmetric vehicle routing problem with side constraints. Fundamenta Informaticae, 130(2):231–246, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Thanh:2014:EDP

[730] Tran Thang Thanh, Fan Chen, Kazunori Kotani, and Bac Le. Extraction of discriminative patterns from skeleton sequences for accurate action recognition. Fundamenta Informaticae, 130(2):247–261, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Madejski:2014:IHP

[731] Grzegorz Madejski. Infinite hierarchy of permutation languages. Fundamenta Informaticae, 130(3):263–274, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Motallebi:2014:TMH

[732] Hassan Motallebi and Mohammad Abdollahi Azgomi. Translation from multisingular hybrid Petri nets to multisingular hybrid automata. Fundamenta Informaticae, 130(3):275–315, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pal:2014:CMU

[733] Sankar K. Pal, Suman Kundu, and C. A. Murthy. Centrality measures, upper bound, and influence maximization in large scale directed social networks. Fundamenta Informaticae, 130(3):317–342, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sakai:2014:ABR

[734] Hiroshi Sakai, Mao Wu, and Michinori Nakata. Apriori-based rule generation in incomplete information databases and non-deterministic information systems. Fundamenta Informaticae, 130(3):343–376, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akshay:2014:DTA

[735] S. Akshay, Benedikt Bollig, Paul Gastin, Madhavan Mukund, and K. Narayan Kumar. Distributed timed automata with independently evolving clocks. Fundamenta Informaticae, 130(4):377–407, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Elbassioni:2014:LBH

[736] Khaled Elbassioni, Matthias Hagen, and Imran Rauf. A lower bound for the HBC transversal hypergraph generation. Fundamenta Informaticae, 130(4):409–414, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krzywkowski:2014:ALA

[737] Marcin Krzywkowski. An algorithm for listing all minimal double dominating sets of a tree. Fundamenta Informaticae, 130(4):415–421, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Permpoontanalarp:2014:FTG

[738] Yongyuth Permpoontanalarp and Panupong Sornkhom. On-the-fly trace generation approach to the security analysis of cryptographic protocols: Coloured Petri nets-based method. Fundamenta Informaticae, 130(4): 423–466, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2014:RAW

[739] Xian gyang Wang, Pan pan Niu, Hong ying Yang, Yan Zhang, and Tian xiao Ma. A robust audio watermarking scheme using higher-order statistics in empirical mode decomposition domain. Fundamenta Informaticae, 130(4):467–490, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVb

[740] Anonymous. Author index volume 130. Fundamenta Informaticae, 130 (4):491–492, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haddad:2014:P

[741] Serge Haddad, Jetty Kleijn, and Lucia Pomello. Preface. Fundamenta Informaticae, 131(1):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valmari:2014:ONA

[742] Antti Valmari and Henri Hansen. Old and new algorithms for minimal coverability sets. Fundamenta Informaticae, 131(1):1–25, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Evangelista:2014:SLM

[743] Sami Evangelista and Lars Michael Kristensen. A sweep-line method for Büchi automata-based model checking. Fundamenta Informaticae, 131 (1):27–53, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martos-Salgado:2014:SSP

[744] María Martos-Salgado and Fernando Rosa-Velardo. Safety and soundness for priced resource-constrained workflow nets. Fundamenta Informaticae, 131(1):55–80, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2014:CSP

[745] GuanJun Liu, Jun Sun, Yang Liu, and JinSong Dong. Complexity of the soundness problem of workflow nets. Fundamenta Informaticae, 131 (1):81–101, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderAalst:2014:PDC

[746] W. M. P. van der Aalst and H. M. W. Verbeek. Process discovery and conformance checking using passages. Fundamenta Informaticae, 131(1): 103–138, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gil-Costa:2014:MSE

[747] Veronica Gil-Costa, Mauricio Marin, Alonso Inostrosa-Psijas, Jair Lobos, and Carolina Bonacic. Modelling search engines performance using coloured Petri nets. Fundamenta Informaticae, 131(1):139–166, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2014:RRT

[748] Chin-Chen Chang and Ting-Fang Cheng. A reliable real-time multicast authentication protocol with provable accuracy. Fundamenta Informaticae, 131(2):167–186, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Egidi:2014:SSD

[749] Lavinia Egidi and Giovanni Manzini. Spaced seed design using perfect rulers. Fundamenta Informaticae, 131(2):187–203, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2014:MSR

[750] Jouni Järvinen and Sándor Radeleczki. Monteiro spaces and rough sets determined by quasiorder relations: Models for Nelson algebras. Fundamenta Informaticae, 131(2):205–215, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marcinkowski:2014:ULS

[751] Jerzy Marcinkowski and Jakub Michaliszyn. The undecidability of the logic of subintervals. Fundamenta Informaticae, 131(2):217–240, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przytula:2014:NSR

[752] Łukasz Przytuła. NAO soccer robots path planning based on rough mereology. Fundamenta Informaticae, 131(2):241–251, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rataj:2014:EMC

[753] Artur Rataj. Evolvability by mimicking common properties of a nervous system and computer software. Fundamenta Informaticae, 131(2):253– 278, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:2014:P

[754] Cristian S. Calude and Marian Gheorghe. Preface. Fundamenta Informaticae, 131(3–4):??, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alexandru:2014:NGT

[755] Andrei Alexandru and Gabriel Ciobanu. Nominal groups and their homomorphism theorems. Fundamenta Informaticae, 131(3–4):279–298, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Azimi:2014:RSM

[756] Sepinoud Azimi, Bogdan Iancu, and Ion Petre. Reaction system models for the heat shock response. Fundamenta Informaticae, 131(3–4):299– 312, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Banu-Demergian:2014:TFR

[757] Iulia Teodora Banu-Demergian and Gheorghe Stefanescu. Towards a formal representation of interactive systems. Fundamenta Informaticae, 131 (3–4):313–336, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bottoni:2014:ANE

[758] Paolo Bottoni, Anna Labella, and Victor Mitrana. Accepting networks of evolutionary picture processors. Fundamenta Informaticae, 131(3–4): 337–349, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cardelli:2014:MSS

[759] Luca Cardelli and Radu Mardare. The measurable space of stochastic processes. Fundamenta Informaticae, 131(3–4):351–371, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2014:CSA

[760] Gabriel Ciobanu and Eneia Nicolae Todoran. Continuation semantics for asynchronous concurrency. Fundamenta Informaticae, 131(3–4):373– 388, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dima:2014:NDT

[761] Cătălin Dima. A nonarchimedian discretization for timed languages. Fundamenta Informaticae, 131(3–4):389–407, July 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dinu:2014:SLG

[762] Liviu P. Dinu, Radu Gramatovici, and Florin Manea. Syllabic languages and go-through automata. Fundamenta Informaticae, 131(3–4):409–424, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duta:2014:NLU

[763] Nicolae Duta. Natural language understanding and prediction: from formal grammars to large scale machine learning. Fundamenta Informaticae, 131(3–4):425–440, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grozea:2014:CML

[764] Cristian Grozea and Marius Popescu. Can machine learning learn a decision oracle for NP problems? A test on SAT. Fundamenta Informaticae, 131(3–4):441–450, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krithivasan:2014:CS

[765] Kamala Krithivasan, Gheorghe Păun, and Ajeesh Ramanujan. On controlled P systems. Fundamenta Informaticae, 131(3–4):451–464, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nicolescu:2014:RIP

[766] Radu Nicolescu, Georgy Gimel’farb, John Morris, Patrice Delmas, and Rui Gong. Regularising ill-posed discrete optimisation: Quests with P systems. Fundamenta Informaticae, 131(3–4):465–483, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zimand:2014:CDI

[767] Marius Zimand. Counting dependent and independent strings. Fundamenta Informaticae, 131(3–4):485–497, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVc

[768] Anonymous. Author index volume 131. Fundamenta Informaticae, 131 (3–4):499–500, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:2014:P

[769] Vesa Halava, Juhani Karhumäki, and Yuri Matiyasevich. Preface. Fundamenta Informaticae, 132(1):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rumyantsev:2014:PCC

[770] Andrei Rumyantsev and Alexander Shen. Probabilistic constructions of computable objects and a computable version of Lovász local lemma. Fundamenta Informaticae, 132(1):1–14, January 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Idiatulina:2014:PPW

[771] Lidia A. Idiatulina and Arseny M. Shur. Periodic partial words and random bipartite graphs. Fundamenta Informaticae, 132(1):15–31, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brinda:2014:BBW

[772] Karel Břinda, Edita Pelantová, and Ondřej Turek. Balances of m-bonacci words. Fundamenta Informaticae, 132(1):33–61, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chistikov:2014:CNM

[773] Dmitry Chistikov, Valentina Fedorova, and Andrey Voronenko. Certificates of non-membership for classes of read-once functions. Fundamenta Informaticae, 132(1):63–77, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dahmoune:2014:TRP

[774] Mohamed Dahmoune, El Houcein El Abdalaoui, and Djelloul Ziadi. On the transition reduction problem for finite automata. Fundamenta Informaticae, 132(1):79–94, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gusev:2014:PIL

[775] Vladimir V. Gusev, Marina I. Maslennikova, and Elena V. Pribavkina. Principal ideal languages and synchronizing automata. Fundamenta Informaticae, 132(1):95–108, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ochem:2014:MSF

[776] Pascal Ochem. More on square-free words obtained from prefixes by permutations. Fundamenta Informaticae, 132(1):109–112, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Itsykson:2014:FHN

[777] Dmitry Itsykson and Dmitry Sokolov. On fast heuristic non-deterministic algorithms and short heuristic proofs. Fundamenta Informaticae, 132(1): 113–129, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salo:2014:PS

[778] Ville Salo and Ilkka Törmä. Playing with subshifts. Fundamenta Informaticae, 132(1):131–152, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cuzzocrea:2014:P

[779] Alfredo Cuzzocrea and Zbigniew W. Ras. Preface. Fundamenta Informaticae, 132(2):??, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Iftikhar:2014:UTG

[780] Nadeem Iftikhar and Torben Bach Pedersen. Using a time granularity table for gradual granular data aggregation. Fundamenta Informaticae, 132(2):153–176, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Viswanathan:2014:QCD

[781] Ganesh Viswanathan and Markus Schneider. Querying cardinal directions between complex objects in data warehouses. Fundamenta Informaticae, 132(2):177–202, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Boukraa:2014:COB

[782] Doulkifli Boukraa, Omar Boussaid, and Fadila Bentayeb. Complex object-based multidimensional modeling and cube construction. Fundamenta Informaticae, 132(2):203–238, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cuzzocrea:2014:DFC

[783] Alfredo Cuzzocrea and Dimitrios Gunopulos. A decomposition framework for computing and querying multidimensional OLAP data cubes over probabilistic relational data. Fundamenta Informaticae, 132(2):239– 266, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bimonte:2014:OFD

[784] Sandro Bimonte, Myoung-Ah Kang, Luca Paolino, Monica Sebillo, Mehdi Zaamoune, and Giuliana Vitiello. OLAPing field data: a theoretical and implementation framework. Fundamenta Informaticae, 132(2):267– 290, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2014:P

[785] Tianrui Li, Hongmei Chen, JingTao Yao, and Hung Son Nguyen. Preface. Fundamenta Informaticae, 132(3):??, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Deng:2014:MAP

[786] Xiaofei Deng and Yiyu Yao. A multifaceted analysis of probabilistic three-way decisions. Fundamenta Informaticae, 132(3):291–313, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Palkowski:2014:ARS

[787] Łukasz Pałkowski, Jerzy Krysiński, Jerzy Błaszczyński, Roman Słowiński, Andrzej Skrzypczak, Jan Błaszczak, Eugenia Gospodarek, and Joanna Wróblewska. Application of rough set theory to prediction of antimicrobial activity of bis-quaternary imidazolium chlorides. Fundamenta Informaticae, 132(3):315–330, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2014:PIB

[788] Baoli Wang, Jiye Liang, and Yuhua Qian. Preorder information based attributes’ weights learning in multi-attribute decision making. Fundamenta Informaticae, 132(3):331–347, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2014:DDV

[789] Guoyin Wang, Lihe Guan, Weizhi Wu, and Feng Hu. Data-driven valued tolerance relation based on the extended rough set. Fundamenta Informaticae, 132(3):349–363, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:2014:APA

[790] Patrick G. Clark, Jerzy W. Grzymala-Busse, and Zdzislaw S. Hippe. An analysis of probabilistic approximations for rule induction from incomplete data sets. Fundamenta Informaticae, 132(3):365–379, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

An:2014:FRD

[791] Shuang An, Hong Shi, Qinghua Hu, and Jianwu Dang. Fuzzy rough decision trees. Fundamenta Informaticae, 132(3):381–399, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zeng:2014:IMR

[792] Anping Zeng, Tianrui Li, Junbo Zhang, and Hongmei Chen. Incremental maintenance of rough fuzzy set approximations under the variation of object set. Fundamenta Informaticae, 132(3):401–422, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Swieboda:2014:IMS

[793] Wojciech Świeboda, Adam Krasuski, Hung Son Nguyen, and Andrzej Janusz. Interactive method for semantic document indexing based on explicit semantic analysis. Fundamenta Informaticae, 132(3):423–438, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Choffrut:2014:DWS

[794] Christian Choffrut. Deciding whether or not a synchronous relation is regular prefix. Fundamenta Informaticae, 132(4):439–459, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kan:2014:SMD

[795] Tomohiro Kan, Shoichi Higuchi, and Kouichi Hirata. Segmental mapping and distance for rooted labeled ordered trees. Fundamenta Informaticae, 132(4):461–483, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koczkodaj:2014:AII

[796] Waldemar W. Koczkodaj and Ryszard Szwarc. On axiomatization of inconsistency indicators for pairwise comparisons. Fundamenta Informaticae, 132(4):485–500, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2014:EGH

[797] Jian-Guo Liu and Zhi-Fang Zeng. Extended generalized hyperbolicfunction method and new exact solutions of the generalized Hamiltonian and NNV equations by the symbolic computation. Fundamenta Informaticae, 132(4):501–517, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mroz:2014:CAC

[798] Andrzej Mróz and Grzegorz Zwara. Combinatorial algorithms for computing degenerations of modules of finite dimension. Fundamenta Informaticae, 132(4):519–532, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Torres:2014:ARA

[799] Gabriela Martín Torres. On the accuracy of rough approximations of regular languages. Fundamenta Informaticae, 132(4):533–545, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVd

[800] Anonymous. Author index volume 132. Fundamenta Informaticae, 132 (4):547–548, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dutta:2014:GCF

[801] Soma Dutta and Mihir Kr. Chakraborty. Graded consequence with fuzzy set of premises. Fundamenta Informaticae, 133(1):1–18, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tapaswini:2014:NPS

[802] Smita Tapaswini and S. Chakraverty. Non-probabilistic solutions of uncertain fractional order diffusion equations. Fundamenta Informaticae, 133(1):19–34, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xia:2014:GMF

[803] Yingjie Xia, Ting Zhang, and Shengbao Wang. A generic methodological framework for cyber-ITS: Using cyber-infrastructure in ITS data analysis cases. Fundamenta Informaticae, 133(1):35–53, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xu:2014:NMS

[804] Changlin Xu, Guoyin Wang, and Qinghua Zhang. A new multi-step backward cloud transformation algorithm based on normal cloud model. Fundamenta Informaticae, 133(1):55–85, January 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ye:2014:NIE

[805] Ruisong Ye. A novel image encryption scheme based on generalized multisawtooth maps. Fundamenta Informaticae, 133(1):87–104, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2014:Pb

[806] Ludwik Czaja. Preface. Fundamenta Informaticae, 133(2–3):??, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2014:DTW

[807] Marco Bellia and M. Eugenia Occhiuto. DNA tiles, Wang tiles and combinators. Fundamenta Informaticae, 133(2–3):105–121, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2014:FCB

[808] Ludwik Czaja. Floating channels between communicating nets. Fundamenta Informaticae, 133(2–3):123–132, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dubtsov:2014:DVT

[809] Roman Dubtsov, Elena Oshevskaya, and Irina Virbitskaite. A domain view of timed behaviors. Fundamenta Informaticae, 133(2–3):133–147, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gomolinska:2014:RIF

[810] Anna Gomolińska and Marcin Wolski. Rough inclusion functions and similarity indices. Fundamenta Informaticae, 133(2–3):149–163, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2014:SO

[811] Damas P. Gruska. Simulation opacity. Fundamenta Informaticae, 133 (2–3):165–179, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jankowski:2014:ICG

[812] Andrzej Jankowski, Andrzej Skowron, and Roman W. Swiniarski. Interactive complex granules. Fundamenta Informaticae, 133(2–3):181–196, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalenkova:2014:DCR

[813] Anna A. Kalenkova and Irina A. Lomazova. Discovery of cancellation regions within process mining techniques. Fundamenta Informaticae, 133 (2–3):197–209, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Knapik:2014:PST

[814] Michał Knapik and Wojciech Penczek. Parameter synthesis for timed Kripke structures. Fundamenta Informaticae, 133(2–3):211–226, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Placzek:2014:AMN

[815] Stanisław Płaczek and Bijaya Adhikari. Analysis of multilayer neural networks with direct and cross forward connection. Fundamenta Informaticae, 133(2–3):227–240, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2014:PBM

[816] Lech Polkowski and Maria Semeniuk-Polkowska. On the problem of boundaries from mereology and rough mereology points of view. Fundamenta Informaticae, 133(2–3):241–255, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2014:MAC

[817] Roman R. Redziejowski. More about converting BNF to PEG. Fundamenta Informaticae, 133(2–3):257–270, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schumann:2014:TOO

[818] Andrew Schumann and Krzysztof Pancerz. Towards an object-oriented programming language for Physarum Polycephalum computing: A Petri net model approach. Fundamenta Informaticae, 133(2–3):271–285, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tran:2014:BBC

[819] Thanh-Luong Tran, Quang-Thuy Ha, Thi-Lan-Giao Hoang, Linh Anh Nguyen, and Hung Son Nguyen. Bisimulation-based concept learning in description logics. Fundamenta Informaticae, 133(2–3):287–303, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Werner:2014:HSE

[820] Matthias Werner, Louchka Popova-Zeugmann, Mario Haustein, and E. Pelz. A holistic state equation for timed Petri nets. Fundamenta Informaticae, 133(2–3):305–322, April 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardinello:2014:CSO

[821] Luca Bernardinello, Carlo Ferigato, Stefan Haar, and Lucia Pomello. Closed sets in occurrence nets with conflicts. Fundamenta Informaticae, 133(4):323–344, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chang:2014:NFS

[822] Bae-Muu Chang. A neuro-fuzzy system combined with particle swarm optimization for handwritten character recognition. Fundamenta Informaticae, 133(4):345–366, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kulakowski:2014:HRE

[823] Konrad Kułakowski. Heuristic rating estimation approach to the pairwise comparisons method. Fundamenta Informaticae, 133(4):367–386, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lim:2014:EDG

[824] Seongan Lim, Hyang-Sook Lee, and Mijin Choi. An efficient decoding of Goppa codes for the McEliece cryptosystem. Fundamenta Informaticae, 133(4):387–397, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mraz:2014:FWO

[825] František Mráz, Friedrich Otto, and Martin Plátek. Free word-order and restarting automata. Fundamenta Informaticae, 133(4):399–419, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Narayanamoorthy:2014:NAV

[826] S. Narayanamoorthy and K. Murugan. A numerical algorithm and a variational iteration technique for solving higher order fuzzy integrodifferential equations. Fundamenta Informaticae, 133(4):421–431, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ozturk:2014:NGW

[827] Mehmet Ali Öztürk, Mustafa Uçkun, and Ebubekir İnan. Near groups of weak cosets on nearness approximation spaces. Fundamenta Informaticae, 133(4):433–448, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVe

[828] Anonymous. Author index volume 133. Fundamenta Informaticae, 133 (4):449–450, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gheorghe:2014:EW

[829] Marian Gheorghe, Gheorghe Păun, Agustín Riscos-Núñez, and Grzegorz Rozenberg. Enjoying to work. Fundamenta Informaticae, 134(1–2):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alcala:2014:TCC

[830] Adrià Alcalà, Mercè Llabrés, Francesc Rosselló, and Pau Rullan. Treechild cluster networks. Fundamenta Informaticae, 134(1–2):1–15, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2014:LS

[831] Artiom Alhazov, Rudolf Freund, and Sergiu Ivanov. Length P systems. Fundamenta Informaticae, 134(1–2):17–37, January 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Battyanyi:2014:DMC

[832] Péter Battyányi and György Vaszil. Describing membrane computations with a chemical calculus. Fundamenta Informaticae, 134(1–2):39–50, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cienciala:2014:CPS

[833] Luděk Cienciala, Lucie Ciencialová, and Erzsébet Csuhaj-Varjú. P colonies processing strings. Fundamenta Informaticae, 134(1–2):51–65, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciobanu:2014:MCD

[834] Gabriel Ciobanu and Dragoş Sburlan. Monitoring changes in dynamic multiset systems. Fundamenta Informaticae, 134(1–2):67–82, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Diaz-Pernil:2014:AFT

[835] Daniel Díaz-Pernil, Francisco Peña-Cantillana, Artiom Alhazov, Rudolf Freund, and Miguel A. Gutiérrez-Naranjo. Antimatter as a frontier of tractability in membrane computing. Fundamenta Informaticae, 134(1– 2):83–96, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Konur:2014:CVU

[836] Savas Konur, Marian Gheorghe, Ciprian Dragomir, Florentin Ipate, and Natalio Krasnogor. Conventional verification for unconventional computing: a genetic XOR gate example. Fundamenta Informaticae, 134(1–2): 97–110, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leporati:2014:CSS

[837] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. Constant-space P systems with active membranes. Fundamenta Informaticae, 134(1–2):111–128, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murphy:2014:UWT

[838] Niall Murphy and Damien Woods. Uniformity is weaker than semiuniformity for some membrane systems. Fundamenta Informaticae, 134 (1–2):129–152, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Orellana-Martin:2014:SCR

[839] David Orellana-Martín, Carmen Graciani, Luis-Felipe Macías-Ramos, Miguel Ángel Martínez del Amor, Agustín Riscos-Núñez, Álvaro Romero-Jiménez, and Luis Valencia-Cabrera. Sevilla carpets revisited: Enriching the membrane computing toolbox. Fundamenta Informaticae, 134(1–2): 153–166, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2014:TUH

[840] Andrei Păun and Petr Sosík. Three universal homogeneous spiking neural P systems using max spike. Fundamenta Informaticae, 134(1–2):167–182, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xu:2014:SLS

[841] Zihan Xu, Matteo Cavaliere, Pei An, Sarma Vrudhula, and Yu Cao. The stochastic loss of spikes in spiking neural P systems: Design and implementation of reliable arithmetic circuits. Fundamenta Informaticae, 134(1–2):183–200, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2014:WSN

[842] Xingyi Zhang, Xiangxiaing Zeng, and Linqiang Pan. Weighted spiking neural P systems with rules on synapses. Fundamenta Informaticae, 134(1–2):201–218, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Degano:2014:ML

[843] Pierpaolo Degano, Juhani Karhumäki, and Paolo Massazza. From models to languages. Fundamenta Informaticae, 134(3–4):??, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bartoletti:2014:CCE

[844] Massimo Bartoletti, Tiziana Cimoli, G. Michele Pinna, and Roberto Zunino. Circular causality in event structures. Fundamenta Informaticae, 134(3–4):219–259, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bistarelli:2014:SNM

[845] Stefano Bistarelli and Francesco Santini. A secure non-monotonic soft concurrent constraint language. Fundamenta Informaticae, 134(3–4): 261–285, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bruni:2014:HGR

[846] Roberto Bruni, Ugo Montanari, Gordon Plotkin, and Daniele Terreni. On hierarchical graphs: Reconciling bigraphs, gs-monoidal theories and gs-graphs. Fundamenta Informaticae, 134(3–4):287–317, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castiglione:2014:EWM

[847] Giusi Castiglione and Marinella Sciortino. Epichristoffel words and minimization of Moore automata. Fundamenta Informaticae, 134(3–4):319– 333, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Comin:2014:ACC

[848] Carlo Comin. Algebraic characterization of the class of languages recognized by measure only quantum automata. Fundamenta Informaticae, 134(3–4):335–353, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Galletta:2014:AIF

[849] Letterio Galletta. An abstract interpretation framework for type and effect systems. Fundamenta Informaticae, 134(3–4):355–393, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honsell:2014:CCG

[850] Furio Honsell, Marina Lenisa, and Daniel Pellarini. Categories of coalgebraic games with selective sum. Fundamenta Informaticae, 134(3–4): 395–414, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVf

[851] Anonymous. Author index volume 134. Fundamenta Informaticae, 134 (3–4):415–416, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2014:P

[852] Paolo Dulio, Andrea Frosini, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 135(1–2):??, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Plantagie:2014:AAT

[853] Linda Plantagie and K. Joost Batenburg. Approximating algebraic tomography methods by filtered backprojection: A local filter approach. Fundamenta Informaticae, 135(1–2):1–19, January 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bleichrodt:2014:API

[854] Folkert Bleichrodt, Jan De Beenhouwer, Jan Sijbers, and Kees Joost Batenburg. Aligning projection images from binary volumes. Fundamenta Informaticae, 135(1–2):21–42, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brocchi:2014:NAR

[855] Stefano Brocchi. A new approach for the reconstruction of object-based images in discrete tomography. Fundamenta Informaticae, 135(1–2):43– 57, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brlek:2014:RWP

[856] Srečko Brlek and Nadia Lafreniére. Reconstructing words from a σpalindromic language. Fundamenta Informaticae, 135(1–2):59–72, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Denitiu:2014:PTC

[857] Andreea Deniţiu, Stefania Petra, Claudius Schnörr, and Christoph Schnörr. Phase transitions and cosparse tomographic recovery of compound solid bodies from few projections. Fundamenta Informaticae, 135 (1–2):73–102, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hantos:2014:RGC

[858] Norbert Hantos and Péter Balázs. Random generation of hv-convex polyominoes with given horizontal projection. Fundamenta Informaticae, 135(1–2):103–115, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Goupy:2014:PRC

[859] Alexandre Goupy and Silvia M. C. Pagani. Probabilistic reconstruction of hv-convex polyominoes from noisy projection data. Fundamenta Informaticae, 135(1–2):117–134, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fiorini:2014:HRY

[860] Rodolfo A. Fiorini. How random is your tomographic noise? A number theoretic transform (NTT) approach. Fundamenta Informaticae, 135(1– 2):135–170, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fedele:2014:XRM

[861] Roberto Fedele, Antonia Ciani, and Fabrizio Fiori. X-ray microtomography under loading and 3D-volume digital image correlation. A review. Fundamenta Informaticae, 135(1–2):171–197, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Compagnoni:2014:TBL

[862] Marco Compagnoni and Roberto Notari. TDOA–based localization in two dimensions: the bifurcation curve. Fundamenta Informaticae, 135(1– 2):199–210, January 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghasemi:2014:ATC

[863] Mohsen Ghasemi. An algorithmic-type classification of tetravalent oneregular graphs using computer algebra tools. Fundamenta Informaticae, 135(3):211–228, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kostanyan:2014:CBA

[864] Armen Kostanyan and Paruyr Khnkoyan. Computations in Boolean algebra with approximation. Fundamenta Informaticae, 135(3):229–241, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niu:2014:RDW

[865] Pan pan Niu, Xiang yang Wang, Hong ying Yang, Pei Wang, and Ai long Wang. A robust digital watermarking based on local complex angular radial transform. Fundamenta Informaticae, 135(3):243–268, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pan:2014:QAL

[866] Haiyu Pan, Min Zhang, Hengyang Wu, and Yixiang Chen. Quantitative analysis of lattice-valued Kripke structures. Fundamenta Informaticae, 135(3):269–293, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Surynek:2014:IBC

[867] Pavel Surynek, Petra Surynková, and Miloš Chromý. The impact of a bi-connected graph decomposition on solving cooperative path-finding problems. Fundamenta Informaticae, 135(3):295–308, July 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2014:CBI

[868] Xiang-Yang Wang, Yong-Wei Li, Pan-Pan Niu, Hong-Ying Yang, and Dong-Ming Li. Content-based image retrieval using visual attention point features. Fundamenta Informaticae, 135(3):309–329, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhu:2014:UDC

[869] Ping Zhu, Huiyang Xie, and Qiaoyan Wen. A unified definition of consistent functions. Fundamenta Informaticae, 135(3):331–340, July 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2014:P

[870] Wojciech Penczek and Ludwik Czaja. Preface. Fundamenta Informaticae, 135(4):??, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Betts:2014:EDI

[871] Jack Betts and Berndt Müller. Engineering MAS — a device integration framework for smart home environments. Fundamenta Informaticae, 135 (4):341–353, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castiglioni:2014:SFR

[872] Valentina Castiglioni, Ruggero Lanotte, and Simone Tini. A specification format for rooted branching bisimulation. Fundamenta Informaticae, 135 (4):355–369, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2014:ERS

[873] Adam Grabowski. Efficient rough set theory merging. Fundamenta Informaticae, 135(4):371–385, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Heitmann:2014:SDR

[874] Frank Heitmann and Michael Köhler-Bußmeier. Structural and dynamic restrictions of elementary object systems. Fundamenta Informaticae, 135 (4):387–401, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kacprzak:2014:IFF

[875] Magdalena Kacprzak and Anna Sawicka. Identification of formal fallacies in a natural dialogue. Fundamenta Informaticae, 135(4):403–417, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karbowska-Chilinska:2014:GAP

[876] Joanna Karbowska-Chilinska and Pawel Zabielski. Genetic algorithm with path relinking for the orienteering problem with time windows. Fundamenta Informaticae, 135(4):419–431, October 2014. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2014:ETM

[877] Linh Anh Nguyen and Joanna Golińska-Pilarek. An ExpTime tableau method for dealing with nominals and qualified number restrictions in deciding the description logic SHOQ. Fundamenta Informaticae, 135(4): 433–449, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2014:SVG

[878] Artur Niewiadomski, Jaroslaw Skaruz, Wojciech Penczek, Maciej Szreter, and Mariusz Jarocki. SMT versus genetic and OpenOpt algorithms: Concrete planning in the PlanICS framework. Fundamenta Informaticae, 135 (4):451–466, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polrola:2014:SBR

[879] Agata Półrola, Piotr Cybula, and Artur Mȩski. SMT-based reachability checking for bounded time Petri nets. Fundamenta Informaticae, 135 (4):467–482, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przymus:2014:BOO

[880] Piotr Przymus, Krzysztof Kaczmarski, and Krzysztof Stencel. A biobjective optimization framework for heterogeneous CPU/GPU query plans. Fundamenta Informaticae, 135(4):483–501, October 2014. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2014:ANV

[881] Zbigniew Suraj and Piotr Grochowalski. About new version of RSDS system. Fundamenta Informaticae, 135(4):503–519, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wagler:2014:PNR

[882] Annegret K. Wagler and Jan-Thierry Wegener. Preprocessing for network reconstruction: Feasibility test and handling infeasibility. Fundamenta Informaticae, 135(4):521–535, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wisniewski:2014:QRB

[883] Piotr Wiśniewski and Krzysztof Stencel. Query rewriting based on metagranular aggregation. Fundamenta Informaticae, 135(4):537–551, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna-Szczesniak:2014:CMP

[884] Bożena Woźna-Szcześniak and Andrzej Zbrzezny. Checking MTL properties of discrete timed automata via bounded model checking. Fundamenta Informaticae, 135(4):553–568, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2014:AIVg

[885] Anonymous. Author index volume 135. Fundamenta Informaticae, 135 (4):569–570, October 2014. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bensch:2015:P

[886] Suna Bensch, Frank Drewes, Mika Hirvensalo, and Friedrich Otto. Preface. Fundamenta Informaticae, 136(1–2):??, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dassow:2015:PAN

[887] Jürgen Dassow, Florin Manea, and Bianca Truthe. On the power of accepting networks of evolutionary processors with special topologies and random context filters. Fundamenta Informaticae, 136(1–2):1–35, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Droste:2015:SWU

[888] Manfred Droste and Doreen Heusel. The supports of weighted unranked tree automata. Fundamenta Informaticae, 136(1–2):37–58, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Freund:2015:CPC

[889] Rudolf Freund, Marion Oswald, and Gheorghe Păun. Catalytic and purely catalytic P systems and P automata: Control mechanisms for obtaining computational completeness. Fundamenta Informaticae, 136 (1–2):59–84, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hoffmann:2015:PKR

[890] Petr Hoffmann. Power of S-kR-RRWW-automata. Fundamenta Informaticae, 136(1–2):85–112, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2015:MCB

[891] Markus Holzer and Sebastian Jakobi. Minimization and characterizations for biautomata. Fundamenta Informaticae, 136(1–2):113–137, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2015:DOW

[892] Martin Kutrib, Julien Provillard, György Vaszil, and Matthias Wendlandt. Deterministic one-way Turing machines with sublinear space. Fundamenta Informaticae, 136(1–2):139–155, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pighizzini:2015:LAC

[893] Giovanni Pighizzini and Andrea Pisoni. Limited automata and contextfree languages. Fundamenta Informaticae, 136(1–2):157–176, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vollweiler:2015:ASP

[894] Marcel Vollweiler. Asynchronous systems of parallel communicating finite automata. Fundamenta Informaticae, 136(1–2):177–197, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pan:2015:P

[895] Linqiang Pan, Gheorghe Păun, and Mario J. Pérez-Jiménez. Preface. Fundamenta Informaticae, 136(3):??, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

He:2015:HDA

[896] Juanjuan He and Kai Zhang. A hybrid distribution algorithm based on membrane computing for solving the multiobjective multiple traveling salesman problem. Fundamenta Informaticae, 136(3):199–208, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Juayong:2015:RCN

[897] Richelle Ann B. Juayong and Henry N. Adorna. Relating computations in non-cooperative transition P systems and evolution-communication P systems with energy. Fundamenta Informaticae, 136(3):209–217, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kong:2015:RSN

[898] Yuan Kong, Xiaolong Shi, Jinbang Xu, and Xinquan Huang. Reversible spiking neural P systems with astrocytes. Fundamenta Informaticae, 136(3):219–230, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kong:2015:SLG

[899] Yuan Kong, Zheng Zhang, and Yang Liu. On string languages generated by spiking neural P systems with astrocytes. Fundamenta Informaticae, 136(3):231–240, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2015:ABC

[900] Yanzhang Li, Changjun Zhou, and Xuedong Zheng. Artificial bee colony algorithm for the protein structure prediction based on the toy model. Fundamenta Informaticae, 136(3):241–252, July 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Macias-Ramos:2015:ESA

[901] Luis F. Macías-Ramos, Mario J. Pérez-Jiménez, Tao Song, and Linqiang Pan. Extending simulation of asynchronous spiking neural P systems in P-Lingua. Fundamenta Informaticae, 136(3):253–267, July 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martinez-del-Amor:2015:SSG

[902] Miguel A. Martínez del Amor, Manuel García-Quismondo, Luis F. Macías-Ramos, Luis Valencia-Cabrera, Agustin Riscos-Núñez, and Mario J. Pérez-Jiménez. Simulating P systems on GPU devices: A survey. Fundamenta Informaticae, 136(3):269–284, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niu:2015:USC

[903] Yunyun Niu, Ibrahim Venkat, Ahamad Tajudin Khader, and K. G. Subramanian. Uniform solution to common algorithmic problem by P systems working in the minimally parallel mode. Fundamenta Informaticae, 136(3):285–296, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Creignou:2015:PCW

[904] Nadia Creignou and Heribert Vollmer. Parameterized complexity of weighted satisfiability problems: Decision, enumeration, counting. Fundamenta Informaticae, 136(4):297–316, October 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Figallo:2015:DDT

[905] Aldo V. Figallo and Gustavo Pelaitay. Discrete duality for tense Łukasiewicz–Moisil algebras. Fundamenta Informaticae, 136(4):317–329, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzeskowiak:2015:ACF

[906] Maciej Grześkowiak. An algorithmic construction of finite elliptic curves of order divisible by a large prime. Fundamenta Informaticae, 136(4): 331–343, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kaniecki:2015:HMA

[907] Mariusz Kaniecki, Justyna Kosakowska, Piotr Malicki, and Grzegorz Marczak. A horizontal mesh algorithm for a class of edge-bipartite graphs and their matrix morsifications. Fundamenta Informaticae, 136(4):345– 379, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Korpusik:2015:CBR

[908] Michał Korpusik, Witold Łukaszewicz, and Ewa Madalińska-Bugaj. Consistency-based revision of structured belief bases. Fundamenta Informaticae, 136(4):381–404, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2015:SFS

[909] Pradipta Maji and Partha Garai. Simultaneous feature selection and extraction using feature significance. Fundamenta Informaticae, 136(4): 405–431, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazzanti:2015:BAO

[910] Stefano Mazzanti. Bases for AC^{0} and other complexity classes. Fundamenta Informaticae, 136(4):433–460, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wichert:2015:PBO

[911] Andreas Wichert and Catarina Moreira. On projection based operators in l^{p} space for exact similarity search. Fundamenta Informaticae, 136 (4):461–474, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVa

[912] Anonymous. Author index volume 136. Fundamenta Informaticae, 136 (4):475–476, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Colom:2015:P

[913] José Manuel Colom, Jörg Desel, and Jetty Kleijn. Preface. Fundamenta Informaticae, 137(1):??, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fraca:2015:CAC

[914] Estíbaliz Fraca and Serge Haddad. Complexity analysis of continuous Petri nets. Fundamenta Informaticae, 137(1):1–28, January 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geeraerts:2015:PNA

[915] Gilles Geeraerts, Alexander Heußner, M. Praveen, and Jean-François Raskin. ω-Petri nets: Algorithms and complexity. Fundamenta Informaticae, 137(1):29–60, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mayr:2015:CRP

[916] Ernst W. Mayr and Jeremias Weihmann. Complexity results for problems of communication-free Petri nets and related formalisms. Fundamenta Informaticae, 137(1):61–86, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marechal:2015:GCP

[917] Alexis Marechal and Didier Buchs. Generalizing the compositions of Petri nets modules. Fundamenta Informaticae, 137(1):87–116, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenthum:2015:VSP

[918] Robin Bergenthum and Robert Lorenz. Verification of scenarios in Petri nets using compact tokenflows. Fundamenta Informaticae, 137(1):117– 142, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernandes:2015:PNS

[919] Johnson Fernandes, Maciej Koutny, Łukasz Mikulski, Marta Pietkiewicz-Koutny, Danil Sokolov, and Alex Yakovlev. Persistent and nonviolent steps and the design of GALS systems. Fundamenta Informaticae, 137 (1):143–170, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Slezak:2015:P

[920] Dominik Ślȩzak, Benjamin C. M. Fung, and William K. Cheung. Preface. Fundamenta Informaticae, 137(2):??, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rauch:2015:FFD

[921] Jan Rauch. Formal framework for data mining with association rules and domain knowledge — overview of an approach. Fundamenta Informaticae, 137(2):171–217, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anciaux:2015:LDE

[922] Nicolas Anciaux, Danae Boutara, Benjamin Nguyen, and Michalis Vazirgiannis. Limiting data exposure in multi-label classification processes. Fundamenta Informaticae, 137(2):219–236, April 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Iwata:2015:MDD

[923] Haruko Iwata, Shoji Hirano, and Shusaku Tsumoto. Maintenance and discovery of domain knowledge for nursing care using data in hospital information system. Fundamenta Informaticae, 137(2):237–252, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2015:PIS

[924] Wuhui Chen, Incheon Paik, and Patrick C. K. Hung. Privacy issues in SOAP message exchange pattern for social services. Fundamenta Informaticae, 137(2):253–271, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meina:2015:SRC

[925] Michał Meina and Hung Son Nguyen. Search result clustering based on query context. Fundamenta Informaticae, 137(2):273–290, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kawamoto:2015:LSH

[926] Junpei Kawamoto. A locality sensitive hashing filter for encrypted vector databases. Fundamenta Informaticae, 137(2):291–304, April 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fusco:2015:CCC

[927] Emanuele G. Fusco and Andrzej Pelc. Communication complexity of consensus in anonymous message passing systems. Fundamenta Informaticae, 137(3):305–322, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Honko:2015:DLR

[928] Piotr Hońko. Description languages for relational information granules. Fundamenta Informaticae, 137(3):323–340, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jarvinen:2015:TII

[929] Jouni Järvinen and Sándor Radeleczki. Tolerances induced by irredundant coverings. Fundamenta Informaticae, 137(3):341–353, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koczkodaj:2015:FCD

[930] Waldemar W. Koczkodaj, Marek Kosiek, Jacek Szybowski, and Ding Xu. Fast convergence of distance-based inconsistency in pairwise comparisons. Fundamenta Informaticae, 137(3):355–367, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kumar:2015:RGB

[931] Sachin Kumar and Rajendra Kumar Sharma. Random-grid based region incrementing visual secret sharing. Fundamenta Informaticae, 137(3): 369–386, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marei:2015:MNS

[932] Emad A. Marei, Mohamed E. Abd El-Monsef, and Hassan M. Abu-Donia. Modification of near sets theory. Fundamenta Informaticae, 137 (3):387–402, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sun:2015:HON

[933] Guanghong Sun and Chuankun Wu. Higher order nonlinearity of Niho functions. Fundamenta Informaticae, 137(3):403–412, July 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umadevi:2015:CRS

[934] Dwaraganathan Umadevi. On the completion of rough sets system determined by arbitrary binary relations. Fundamenta Informaticae, 137 (3):413–424, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2015:TSD

[935] Jean-Marc Champarnaud, Jean-Philippe Dubernard, Hadrien Jeanne, and Ludovic Mignot. Two-sided derivatives for regular expressions and for hairpin expressions. Fundamenta Informaticae, 137(4):425–455, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gong:2015:CFS

[936] Zengtai Gong and Xiaoxia Zhang. On characterization of fuzzy soft rough sets based on a pair of border implicators. Fundamenta Informaticae, 137 (4):457–491, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peters:2015:ARC

[937] Georg Peters. Assessing rough classifiers. Fundamenta Informaticae, 137 (4):493–515, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Surynek:2015:COP

[938] Pavel Surynek. On the complexity of optimal parallel cooperative pathfinding. Fundamenta Informaticae, 137(4):517–548, October 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVb

[939] Anonymous. Author index volume 137. Fundamenta Informaticae, 137 (4):549–550, October 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cojocaru:2015:P

[940] Svetlana Cojocaru, Maurice Margenstern, Gheorghe Păun, and Sergey Verlan. Preface. Fundamenta Informaticae, 138(1–2):??, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martinez:2015:DCA

[941] Genaro J. Martínez, Andrew Adamatzky, and Ramon Alonso-Sanz. On the dynamics of cellular automata with memory. Fundamenta Informaticae, 138(1–2):1–16, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Morita:2015:URT

[942] Kenichi Morita. Universal reversible Turing machines with a small number of tape symbols. Fundamenta Informaticae, 138(1–2):17–29, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Okubo:2015:FAM

[943] Fumiya Okubo and Takashi Yokomori. Finite automata with multiset memory: A new characterization of Chomsky hierarchy. Fundamenta Informaticae, 138(1–2):31–44, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez-Jimenez:2015:OFE

[944] Mario J. Pérez-Jiménez and Petr Sosík. An optimal frontier of the efficiency of tissue P systems with cell separation. Fundamenta Informaticae, 138(1–2):45–60, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ibarra:2015:SSC

[945] Oscar H. Ibarra and Shinnosuke Seki. Semilinear sets and counter machines: a brief survey. Fundamenta Informaticae, 138(1–2):61–76, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calude:2015:UAD

[946] Cristian S. Calude and Damien Desfontaines. Universality and almost decidability. Fundamenta Informaticae, 138(1–2):77–84, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alarcon:2015:AMI

[947] Pedro Pablo Alarcón, Fernando Arroyo, Henning Bordihn, Victor Mitrana, and Mike Müller. Ambiguity of the multiple interpretations on regular languages. Fundamenta Informaticae, 138(1–2):85–95, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leporati:2015:MDO

[948] Alberto Leporati, Luca Manzoni, Giancarlo Mauri, Antonio E. Porreca, and Claudio Zandron. Membrane division, oracles, and the counting hierarchy. Fundamenta Informaticae, 138(1–2):97–111, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Margenstern:2015:AAT

[949] Maurice Margenstern. An algorithmic approach to tilings of hyperbolic spaces: Universality results. Fundamenta Informaticae, 138(1–2):113– 125, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ivanov:2015:RCS

[950] Sergiu Ivanov and Sergey Verlan. Random context and semi-conditional insertion-deletion systems. Fundamenta Informaticae, 138(1–2):127–144, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rogojin:2015:SES

[951] Vladimir Rogojin and Ion Petre. The structure of elementary strategies for gene assembly in ciliates. Fundamenta Informaticae, 138(1–2):145– 158, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Choffrut:2015:LTM

[952] Christian Choffrut and Serge Grigorieff. Logical theory of the monoid of languages over a non tally alphabet. Fundamenta Informaticae, 138(1– 2):159–177, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Enaganti:2015:FLM

[953] Srujan Kumar Enaganti, Lila Kari, and Steffen Kopecki. A formal language model of DNA polymerase enzymatic activity. Fundamenta Informaticae, 138(1–2):179–192, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2015:SLM

[954] Erzsébet Csuhaj-Varjú, Marian Gheorghe, Mike Stannett, and György Vaszil. Spatially localised membrane systems. Fundamenta Informaticae, 138(1–2):193–205, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Muravitsky:2015:DMS

[955] Alexei Y. Muravitsky. Domains as models for semantic information. Fundamenta Informaticae, 138(1–2):207–225, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhazov:2015:VSU

[956] Artiom Alhazov and Rudolf Freund. Variants of small universal P systems with catalysts. Fundamenta Informaticae, 138(1–2):227–250, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Woods:2015:YRC

[957] Damien Woods and Turlough Neary. Yurii Rogozhin’s contributions to the field of small universal Turing machines. Fundamenta Informaticae, 138(1–2):251–258, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sutner:2015:IT

[958] K. Sutner. Iterating transducers. Fundamenta Informaticae, 138(1–2): 259–272, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:YRB

[959] Anonymous. Yurii Rogozhin’s biography and publications. Fundamenta Informaticae, 138(1–2):273–284, January 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Armano:2015:MPF

[960] Giuliano Armano. Modelling progressive filtering. Fundamenta Informaticae, 138(3):285–320, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bera:2015:FAG

[961] Sahadev Bera, Arindam Biswas, and Bhargab B. Bhattacharya. A fast and automated granulometric image analysis based on digital geometry. Fundamenta Informaticae, 138(3):321–338, July 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dabrowski:2015:CCZ

[962] Robert Da̧browski and Wojciech Plandowski. Complete characterization of zero-expressible functions. Fundamenta Informaticae, 138(3):339–350, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lang:2015:HBC

[963] Guangming Lang, Qingguo Li, and Lankun Guo. Homomorphisms between covering approximation spaces. Fundamenta Informaticae, 138 (3):351–371, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pham:2015:EMS

[964] Thi-Thiet Pham. Efficiently mining sequential generator patterns using prefix trees. Fundamenta Informaticae, 138(3):373–386, July 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abbasian:2015:CBA

[965] Mohammad Amir Abbasian, Hossein Nezamabadi-pour, and Maryam Amoozegar. A clustering based archive multi objective gravitational search algorithm. Fundamenta Informaticae, 138(4):387–409, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2015:AEI

[966] J. A. Bergstra and C. A. Middelburg. On algorithmic equivalence of instruction sequences for computing bit string functions. Fundamenta Informaticae, 138(4):411–434, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chunlin:2015:HCS

[967] Li Chunlin and Li LaYuan. Hybrid cloud scheduling method for cloud bursting. Fundamenta Informaticae, 138(4):435–455, ???? 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Laskowski:2015:SCN

[968] Lukasz Laskowski and Jerzy Jelonkiewicz. Self-correcting neural network for stereo-matching problem solving. Fundamenta Informaticae, 138(4): 457–482, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Moschoyiannis:2015:TCL

[969] Sotiris Moschoyiannis and Paul J. Krause. True concurrency in longrunning transactions for digital ecosystems. Fundamenta Informaticae, 138(4):483–514, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVc

[970] Anonymous. Author index volume 138. Fundamenta Informaticae, 138 (4):515–516, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:2015:MC

[971] Gregory Clark and Jason Teutsch. Maximizing T -complexity. Fundamenta Informaticae, 139(1):1–19, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2015:CCH

[972] Ryszard Janicki, Jetty Kleijn, Maciej Koutny, and Łukasz Mikulski. Characterising concurrent histories. Fundamenta Informaticae, 139(1): 21–42, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kumar:2015:PSO

[973] Paulraj Ranjith Kumar, P. Babu, and Sankaran Palani. Particle swarm optimization based sequential and parallel tasks scheduling model for heterogeneous multiprocessor systems. Fundamenta Informaticae, 139 (1):43–65, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2015:MDT

[974] Wentao Li and Weihua Xu. Multigranulation decision-theoretic rough set in ordered information system. Fundamenta Informaticae, 139(1): 67–89, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zekri:2015:ICP

[975] Meriem Zekri, Karima Alem, and Labiba Souici-Meslati. Immunological computation for protein function prediction. Fundamenta Informaticae, 139(1):91–114, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alatabbi:2015:SLC

[976] Ali Alatabbi, Jacqueline W. Daykin, M. Sohel Rahman, and William F. Smyth. Simple linear comparison of strings in V -order. Fundamenta Informaticae, 139(2):115–126, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Desai:2015:RFC

[977] Santosh S. Desai and Shrikant R. Chaudhari. On Q-rationality of fuzzy choice functions on base domains. Fundamenta Informaticae, 139(2):127– 151, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kasjan:2015:MACa

[978] Stanisław Kasjan and Daniel Simson. Mesh algorithms for Coxeter spectral classification of Cox-regular edge-bipartite graphs with loops, I. Mesh root systems. Fundamenta Informaticae, 139(2):153–184, ???? 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kasjan:2015:MACb

[979] Stanisław Kasjan and Daniel Simson. Mesh algorithms for Coxeter spectral classification of Cox-regular edge-bipartite graphs with loops, II. application to Coxeter spectral analysis. Fundamenta Informaticae, 139 (2):185–209, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Macias-Ramos:2015:PBS

[980] Luis F. Macías-Ramos, Luis Valencia-Cabrera, Bosheng Song, Tao Song, Linqiang Pan, and Mario J. Pérez-Jiménez. A P_{Lingua} based simulator for P systems with symport/antiport rules. Fundamenta Informaticae, 139(2):211–227, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chojnacki:2015:LNB

[981] Szymon Chojnacki and Mieczysław Kłopotek. Latency of neighborhood based recommender systems. Fundamenta Informaticae, 139(3):229–248, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kasjan:2015:AIG

[982] Stanisław Kasjan and Daniel Simson. Algorithms for isotropy groups of Cox-regular edge-bipartite graphs. Fundamenta Informaticae, 139(3): 249–275, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mogavero:2015:PPG

[983] Fabio Mogavero, Aniello Murano, and Loredana Sorrentino. On promptness in parity games. Fundamenta Informaticae, 139(3):277–305, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shakiba:2015:AST

[984] Ali Shakiba and Mohammad R. Hooshmandasl. S-approximation spaces: A three-way decision approach. Fundamenta Informaticae, 139(3):307– 328, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eberbach:2015:HUD

[985] Eugene Eberbach. On hypercomputation, universal and diagonalization complete problems. Fundamenta Informaticae, 139(4):329–346, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gasiorek:2015:ASN

[986] Marcin Ga̧siorek and Katarzyna Zaja̧c. On algorithmic study of nonnegative posets of corank at most two and their Coxeter–Dynkin types. Fundamenta Informaticae, 139(4):347–367, ???? 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jezequel:2015:FCO

[987] Loïg Jezequel and Eric Fabre. Factored cost-optimal planning using message passing algorithms*. Fundamenta Informaticae, 139(4):369–401, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kulakowski:2015:PPD

[988] Konrad Kułakowski. On the properties of the priority deriving procedure in the pairwise comparisons method. Fundamenta Informaticae, 139(4): 403–419, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhu:2015:BMT

[989] Kai Zhu, Guilong Liu, and Yanbin Feng. Boolean matrices and their applications to covering reductions. Fundamenta Informaticae, 139(4): 421–433, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVd

[990] Anonymous. Author index volume 139. Fundamenta Informaticae, 139 (4):435–436, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:Pa

[991] Anonymous. Preface. Fundamenta Informaticae, 140(1):??, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baldan:2015:CVN

[992] Paolo Baldan and Alberto Carraro. A causal view on non-interference*. Fundamenta Informaticae, 140(1):1–38, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Best:2015:SLB

[993] Eike Best and Raymond Devillers. Synthesis of live and bounded persistent systems. Fundamenta Informaticae, 140(1):39–59, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Avellaneda:2015:CSB

[994] Florent Avellaneda and Rémi Morin. Catching a structural bug with a flower. Fundamenta Informaticae, 140(1):61–87, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mateo:2015:STA

[995] José Antonio Mateo, Jiří Srba, and Mathias Grund Sørensen. Soundness of timed-arc workflow nets in discrete and continuous-time semantics. Fundamenta Informaticae, 140(1):89–121, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czerwinski:2015:NDS

[996] Wojciech Czerwiński, Tomasz Gogacz, and Eryk Kopczyński. Nondominating sequences of vectors using only resets and increments. Fundamenta Informaticae, 140(2):123–127, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ryszka:2015:GTSa

[997] Iwona Ryszka, Anna Paszyńska, Ewa Grabska, Marcin Sieniek, and Maciej Paszyński. Graph transformation systems for modeling three dimensional finite element method. Part I. Fundamenta Informaticae, 140 (2):129–172, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ryszka:2015:GTSb

[998] Iwona Ryszka, Anna Paszyńska, Ewa Grabska, Marcin Sieniek, and Maciej Paszyński. Graph transformation systems for modeling three dimensional finite element method. Part II. Fundamenta Informaticae, 140 (2):173–203, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Torkestani:2015:SBD

[999] Javad Akbari Torkestani. Stochastic bounded diameter minimum spanning tree problem. Fundamenta Informaticae, 140(2):205–219, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:Pb

[1000] Anonymous. Preface. Fundamenta Informaticae, 140(3–4):??, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ancona:2015:TPC

[1001] Davide Ancona and Agostino Dovier. A theoretical perspective of coinductive logic programming. Fundamenta Informaticae, 140(3–4):221– 246, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Avellone:2015:SAF

[1002] Alessandro Avellone, Camillo Fiorentini, and Alberto Momigliano. A semantical analysis of focusing and contraction in intuitionistic logic. Fundamenta Informaticae, 140(3–4):247–262, ???? 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bistarelli:2015:CTE

[1003] Stefano Bistarelli, Fabio Rossi, and Francesco Santini. A comparative test on the enumeration of extensions in abstract argumentation*. Fundamenta Informaticae, 140(3–4):263–278, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Costantini:2015:NRN

[1004] Stefania Costantini and Andrea Formisano. Negation as a resource: a novel view on answer set semantics*. Fundamenta Informaticae, 140(3– 4):279–305, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DAgostino:2015:MSH

[1005] Giovanna D’Agostino, Eugenio G. Omodeo, Alberto Policriti, and Alexandru I. Tomescu. Mapping sets and hypersets into numbers. Fundamenta Informaticae, 140(3–4):307–328, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeAngelis:2015:RBV

[1006] Emanuele De Angelis, Fabio Fioravanti, Alberto Pettorossi, and Maurizio Proietti. A rule-based verification strategy for array manipulating programs. Fundamenta Informaticae, 140(3–4):329–355, ???? 2015. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gentilini:2015:ECL

[1007] Paolo Gentilini, Maurizio Martelli, and Giuseppe Rosolini. Explicit constructive logic ECL: a new representation of construction and selection of logical information by an epistemic agent. Fundamenta Informaticae, 140(3–4):357–372, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lisi:2015:LDL

[1008] Francesca A. Lisi and Umberto Straccia. Learning in description logics with fuzzy concrete domains. Fundamenta Informaticae, 140(3–4):373– 391, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rossi:2015:NPJ

[1009] Gianfranco Rossi and Federico Bergenti. Nondeterministic programming in Java with JSetL. Fundamenta Informaticae, 140(3–4):393–412, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVe

[1010] Anonymous. Author index volume 140. Fundamenta Informaticae, 140 (3–4):413–414, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cattaneo:2015:LSE

[1011] Gianpiero Cattaneo, Giampiero Chiaselotti, Tommaso Gentile, and Paolo A. Oliverio. The lattice structure of equally extended signed partitions. Fundamenta Informaticae, 141(1):1–36, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kumar:2015:ADR

[1012] Arun Kumar and Mohua Banerjee. Algebras of definable and rough sets in quasi order-based approximation spaces. Fundamenta Informaticae, 141(1):37–55, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kuziak:2015:SMD

[1013] Dorota Kuziak, Ismael G. Yero, and Juan A. Rodríguez-Velázquez. On the strong metric dimension of Cartesian sum graphs. Fundamenta Informaticae, 141(1):57–69, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peng:2015:IVH

[1014] Xindong Peng and Yong Yang. Interval-valued hesitant fuzzy soft sets and their application in decision making. Fundamenta Informaticae, 141 (1):71–93, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shao:2015:CBF

[1015] Zuhua Shao and Yipeng Gao. Certificate-based fair exchange protocol of Schnorr signatures in chosen-key model. Fundamenta Informaticae, 141(1):95–114, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:Pc

[1016] Anonymous. Preface. Fundamenta Informaticae, 141(2–3):??, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fiorini:2015:CTN

[1017] Rodolfo A. Fiorini. Computerized tomography noise reduction by CICT optimized exponential cyclic sequences (OECS) co-domain. Fundamenta Informaticae, 141(2–3):115–134, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hantos:2015:ESC

[1018] Norbert Hantos and Péter Balázs. Eliminating switching components in binary matrices by 0–1 flips and column permutations. Fundamenta Informaticae, 141(2–3):135–150, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Balazs:2015:MDC

[1019] Péter Balázs, Zoltán Ozsvár, Tamás S. Tasi, and László G. Nyúl. A measure of directional convexity inspired by binary tomography. Fundamenta Informaticae, 141(2–3):151–167, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vincze:2015:ARC

[1020] Csaba Vincze and Ábris Nagy. An algorithm for the reconstruction of hv-convex planar bodies by finitely many and noisy measurements of their coordinate X-rays. Fundamenta Informaticae, 141(2–3):169–189, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Frosini:2015:ITP

[1021] Andrea Frosini, Daniela Battaglino, Simone Rinaldi, and Samanta Socci. The identity transform of a permutation and its applications. Fundamenta Informaticae, 141(2–3):191–205, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nagy:2015:ESM

[1022] Antal Nagy. Experimental study on multivalued phantoms using different filters in the DART algorithm. Fundamenta Informaticae, 141(2–3):207– 231, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brun:2015:EFS

[1023] Francesco Brun, Serena Pacilè, Agostino Accardo, George Kourousias, Diego Dreossi, Lucia Mancini, Giuliana Tromba, and Roberto Pugliese. Enhanced and flexible software tools for X-ray computed tomography at the Italian Synchrotron Radiation Facility Elettra. Fundamenta Informaticae, 141(2–3):233–243, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shkarin:2015:GOD

[1024] Roman Shkarin, Evelina Ametova, Suren Chilingaryan, Timo Dritschler, Andreas Kopmann, Alessandro Mirone, Andrei Shkarin, Matthias Vogelgesang, and Sergey Tsapko. GPU-optimized direct Fourier method for on-line tomography. Fundamenta Informaticae, 141(2–3):245–258, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shkarin:2015:OSG

[1025] Andrei Shkarin, Evelina Ametova, Suren Chilingaryan, Timo Dritschler, Andreas Kopmann, Matthias Vogelgesang, Roman Shkarin, and Sergey Tsapko. An open source GPU accelerated framework for flexible algebraic reconstruction at synchrotron light sources. Fundamenta Informaticae, 141(2–3):259–274, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Basu:2015:SBS

[1026] Tanmay Basu and C. A. Murthy. A similarity based supervised decision rule for qualitative improvement of text categorization. Fundamenta Informaticae, 141(4):275–295, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bortolussi:2015:CTG

[1027] Luca Bortolussi, Liviu P. Dinu, Laura Franzoi, and Andrea Sgarro. Coding theory: A general framework and two inverse problems. Fundamenta Informaticae, 141(4):297–310, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dicky:2015:TWA

[1028] Anne Dicky and David Janin. Two-way automata and regular languages of overlapping tiles. Fundamenta Informaticae, 141(4):311–343, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Durnoga:2015:CDL

[1029] Konrad Durnoga and Bartosz Źrałek. On computing discrete logarithms in bulk and randomness extractors. Fundamenta Informaticae, 141(4): 345–366, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sroka:2015:GHW

[1030] Jacek Sroka, Piotr Chrza̧stowski-Wachtel, and Jan Hidders. On generating hierarchical workflow nets and their extensions and verifying hierarchicality. Fundamenta Informaticae, 141(4):367–398, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVf

[1031] Anonymous. Author index volume 141. Fundamenta Informaticae, 141 (4):399–400, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:Pd

[1032] Anonymous. Preface. Fundamenta Informaticae, 142(1–4):??, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2015:SOF

[1033] Davide Ciucci, Didier Dubois, and Henri Prade. Structures of opposition in fuzzy rough sets. Fundamenta Informaticae, 142(1–4):1–19, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Inuiguchi:2015:AOF

[1034] Masahiro Inuiguchi. Approximation-oriented fuzzy rough set approaches. Fundamenta Informaticae, 142(1–4):21–51, ???? 2015. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vluymans:2015:AFR

[1035] Sarah Vluymans, Lynn D’eer, Yvan Saeys, and Chris Cornelis. Applications of fuzzy rough set theory in machine learning: a survey. Fundamenta Informaticae, 142(1–4):53–86, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2015:UOA

[1036] Wei-Zhi Wu, Tong-Jun Li, and Shen-Ming Gu. Using one axiom to characterize fuzzy rough approximation operators determined by a fuzzy implication operator. Fundamenta Informaticae, 142(1–4):87–104, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2015:QDF

[1037] Guilong Liu. On quasi-discrete fuzzy closure spaces. Fundamenta Informaticae, 142(1–4):105–115, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lang:2015:DTR

[1038] Guangming Lang and Tian Yang. Decision-theoretic rough sets-based three-way approximations of interval-valued fuzzy sets. Fundamenta Informaticae, 142(1–4):117–143, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ma:2015:RSA

[1039] Jian-Min Ma and Yiyu Yao. Rough set approximations in multigranulation fuzzy approximation spaces. Fundamenta Informaticae, 142 (1–4):145–160, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ma:2015:CSM

[1040] Zhou-Ming Ma and Ju-Sheng Mi. A comparative study of MGRSs and their uncertainty measures. Fundamenta Informaticae, 142(1–4):161– 181, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qin:2015:RAB

[1041] Keyun Qin, Junfang Luo, and Zheng Pei. Rough approximations based on valued tolerance relations. Fundamenta Informaticae, 142(1–4):183– 194, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2015:RAP

[1042] Xiaohong Zhang, Yanning Zhang, Zhanao Xue, and Yingcang Ma. T - rough approximation pairs and covering based rough sets. Fundamenta Informaticae, 142(1–4):195–212, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dutta:2015:GQC

[1043] Soma Dutta and Andrzej Skowron. Generalized quantifiers in the context of rough set semantics. Fundamenta Informaticae, 142(1–4):213– 236, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2015:SRR

[1044] Pradipta Maji and Shaswati Roy. SoBT-RFW: Rough-fuzzy computing and wavelet analysis based automatic brain tumor detection method from MR images*. Fundamenta Informaticae, 142(1–4):237–267, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2015:WRS

[1045] Lech Polkowski and Maria Semeniuk-Polkowska. Where rough sets and fuzzy sets meet. Fundamenta Informaticae, 142(1–4):269–284, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cai:2015:CDF

[1046] Mingjie Cai and Qingguo Li. Compression of dynamic fuzzy relation information systems. Fundamenta Informaticae, 142(1–4):285–306, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Li:2015:KRC

[1047] Mei-Zheng Li and Guo-Yin Wang. Knowledge reduction in crisply generated fuzzy concept lattices. Fundamenta Informaticae, 142(1–4):307–335, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wen:2015:GCA

[1048] Liu-Ying Wen and Fan Min. A granular computing approach to symbolic value partitioning. Fundamenta Informaticae, 142(1–4):337–371, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2015:ICA

[1049] Jianhang Yu and Weihua Xu. Incremental computing approximations with the dynamic object set in interval-valued ordered information system. Fundamenta Informaticae, 142(1–4):373–397, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2015:AIVg

[1050] Anonymous. Author index volume 142. Fundamenta Informaticae, 142 (1–4):399–400, ???? 2015. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2016:Pa

[1051] Ludwik Czaja, Wojciech Penczek, and Krzysztof Stencel. Preface. Fundamenta Informaticae, 143(1–2):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2016:CBD

[1052] Jan G. Bazan, Stanislawa Bazan-Socha, Sylwia Buregwa-Czuma, Lukasz Dydo, Wojciech Rza̧sa, and Andrzej Skowron. A classifier based on a decision tree with verifying cuts. Fundamenta Informaticae, 143(1–2): 1–18, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2016:CBP

[1053] Jan G. Bazan, Marcin Szpyrka, Adam Szczur, Lukasz Dydo, and Hubert Wojtowicz. Classifiers for behavioral patterns identification induced from huge temporal data. Fundamenta Informaticae, 143(1–2):19–34, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellia:2016:PSA

[1054] Marco Bellia and M. Eugenia Occhiuto. Programming self-assembly of DNA tiles. Fundamenta Informaticae, 143(1–2):35–49, ???? 2016. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Castiglioni:2016:FEM

[1055] Valentina Castiglioni, Ruggero Lanotte, and Simone Tini. A function elimination method for checking satisfiability of arithmetical logics. Fundamenta Informaticae, 143(1–2):51–71, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gruska:2016:DPS

[1056] Damas P. Gruska. Differential privacy and security. Fundamenta Informaticae, 143(1–2):73–87, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kohler-Bussmeier:2016:UBR

[1057] Michael Köhler-Bußmeier and Frank Heitmann. An upper bound for the reachability problem of safe, elementary hornets. Fundamenta Informaticae, 143(1–2):89–100, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lomazova:2016:CPN

[1058] Irina A. Lomazova and Louchka Popova-Zeugmann. Controlling Petri net behavior using priorities for transitions. Fundamenta Informaticae, 143 (1–2):101–112, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2016:TRL

[1059] Linh Anh Nguyen. A tractable rule language in the modal and description logic that combines CPDL with regular grammar logic. Fundamenta Informaticae, 143(1–2):113–139, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2016:CPP

[1060] Roman R. Redziejowski. Cut points in PEG. Fundamenta Informaticae, 143(1–2):141–149, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Suraj:2016:IFI

[1061] Zbigniew Suraj, Agnieszka Lasek, and Piotr Lasek. Inverted fuzzy implications in approximate reasoning. Fundamenta Informaticae, 143(1–2):

151–171, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna-Szczesniak:2016:SBB

[1062] Bożena Woźna-Szcześniak. SAT-based bounded model checking for weighted Deontic interpreted systems. Fundamenta Informaticae, 143 (1–2):173–205, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abdulla:2016:P

[1063] Parosh Aziz Abdulla, Stéphane Demri, Alain Finkel, Jérôme Leroux, and Igor Potapov. Preface. Fundamenta Informaticae, 143(3–4):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akshay:2016:RTP

[1064] S. Akshay, Loïc Hélouët, Claude Jard, and Pierre-Alain Reynier. Robustness of time Petri nets under guard enlargement. Fundamenta Informaticae, 143(3–4):207–234, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berard:2016:ITA

[1065] Beatrice Bérard, Serge Haddad, Aleksandra Jovanović, and Didier Lime. Interrupt timed automata with auxiliary clocks and parameters*. Fundamenta Informaticae, 143(3–4):235–259, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brocchi:2016:NSM

[1066] Stefano Brocchi and Paolo Massazza. A new sandpile model with smoothness assumptions. Fundamenta Informaticae, 143(3–4):261–286, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delzanno:2016:ADR

[1067] Giorgio Delzanno, Arnaud Sangnier, and Riccardo Traverso. Adding data registers to parameterized networks with broadcast. Fundamenta Informaticae, 143(3–4):287–316, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Haase:2016:RRP

[1068] Christoph Haase, Joël Ouaknine, and James Worrell. Relating reachability problems in timed and counter automata. Fundamenta Informaticae, 143(3–4):317–338, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mayr:2016:BTM

[1069] Richard Mayr and Patrick Totzke. Branching-time model checking gaporder constraint systems. Fundamenta Informaticae, 143(3–4):339–353, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mayr:2016:CRG

[1070] Ernst W. Mayr and Jeremias Weihmann. Completeness results for generalized communication-free Petri nets with arbitrary arc multiplicities. Fundamenta Informaticae, 143(3–4):355–391, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Piipponen:2016:CMC

[1071] Artturi Piipponen and Antti Valmari. Constructing minimal coverability sets. Fundamenta Informaticae, 143(3–4):393–414, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Reichert:2016:CCR

[1072] Julien Reichert. On the complexity of counter reachability games*. Fundamenta Informaticae, 143(3–4):415–436, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVa

[1073] Anonymous. Author index volume 143. Fundamenta Informaticae, 143 (3–4):437–438, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alexandru:2016:AIF

[1074] Andrei Alexandru and Gabriel Ciobanu. Abstract interpretations in the framework of invariant sets. Fundamenta Informaticae, 144(1):1– 22, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mrozek:2016:MCR

[1075] Ireneusz Mrozek and Vyacheslav Yarmolik. Multiple controlled random testing*. Fundamenta Informaticae, 144(1):23–43, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Radzikowska:2016:DTI

[1076] Anna Maria Radzikowska. Duality via truth for information algebras based on De Morgan lattices*. Fundamenta Informaticae, 144(1):45– 75, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Song:2016:TSP

[1077] Bosheng Song, Linqiang Pan, and Mario J. Pérez-Jiménez. Tissue P systems with protein on cells. Fundamenta Informaticae, 144(1):77–107, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bottoni:2016:PBR

[1078] Paolo Bottoni, Esther Guerra, and Juan de Lara. Pattern-based rewriting through abstraction. Fundamenta Informaticae, 144(2):109–160, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eom:2016:SCB

[1079] Hae-Sung Eom, Yo-Sub Han, and Galina Jirásková. State complexity of basic operations on non-returning regular languages. Fundamenta Informaticae, 144(2):161–182, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Radzikowska:2016:RTL

[1080] Anna Maria Radzikowska. Representation theorems for lattice-ordered modal algebras and their axiomatic extensions*. Fundamenta Informaticae, 144(2):183–203, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2016:UCO

[1081] Tingfang Wu, Zhiqiang Zhang, Gheorghe Păun, and Linqiang Pan. On the universality of colored one-catalyst P systems. Fundamenta Informaticae, 144(2):205–212, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2016:P

[1082] Ryszard Janicki and Konrad Kułakowski. Preface. Fundamenta Informaticae, 144(3–4):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saaty:2016:CPC

[1083] Thomas L. Saaty. Continuous pairwise comparisons. Fundamenta Informaticae, 144(3–4):213–221, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sandrasagra:2016:CRP

[1084] Barbara Sandrasagra and Michael Soltys. Complex ranking procedures. Fundamenta Informaticae, 144(3–4):223–240, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schumann:2016:TSO

[1085] Andrew Schumann and Jan Woleński. Two squares of oppositions and their applications in pairwise comparisons analysis. Fundamenta Informaticae, 144(3–4):241–254, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez:2016:EPF

[1086] Joaquín Pérez and Ethel Mokotoff. Eigenvector priority function causes strong rank reversal in group decision making. Fundamenta Informaticae, 144(3–4):255–261, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Choo:2016:MSG

[1087] Eng U. Choo, William C. Wedley, and Diederik J. D. Wijnmalen. Mathematical support for the geometric mean when deriving a consistent matrix from a pairwise ratio matrix. Fundamenta Informaticae, 144(3–4): 263–278, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abele-Nagy:2016:EAS

[1088] Kristóf Ábele-Nagy and Sándor Bozóki. Efficiency analysis of simple perturbed pairwise comparison matrices. Fundamenta Informaticae, 144 (3–4):279–289, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koczkodaj:2016:IFO

[1089] Waldemar W. Koczkodaj, Ludmil Mikhailov, Grzegorz Redlarski, Michael Soltys, Jacek Szybowski, Gaik Tamazian, Eliza Wajch, and Kevin Kam Fung Yuen. Important facts and observations about pairwise comparisons (the special issue edition). Fundamenta Informaticae, 144 (3–4):291–307, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csato:2016:IPC

[1090] László Csató and Lajos Rónyai. Incomplete pairwise comparison matrices and weighting methods. Fundamenta Informaticae, 144(3–4):309– 320, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brunelli:2016:RAI

[1091] Matteo Brunelli. Recent advances on inconsistency indices for pairwise comparisons — a commentary. Fundamenta Informaticae, 144(3–4):321– 332, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Martins:2016:CMC

[1092] Claudio L. M. Martins and Pedro P. B. de Oliveira. Computing modulo-n by composing cellular automata rules. Fundamenta Informaticae, 145(1): 1–17, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2016:SACa

[1093] Daniel Simson. Symbolic algorithms computing Gram congruences in the Coxeter spectral classification of edge-bipartite graphs, I. A Gram classification. Fundamenta Informaticae, 145(1):19–48, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Simson:2016:SACb

[1094] Daniel Simson. Symbolic algorithms computing Gram congruences in the Coxeter spectral classification of edge-bipartite graphs, II. Isotropy minigroups. Fundamenta Informaticae, 145(1):49–80, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shalu:2016:GJA

[1095] M. A. Shalu and S. Devi Yamini. A generalization of join and an algorithmic recognition problem. Fundamenta Informaticae, 145(1):81–91, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xu:2016:RAO

[1096] Jiuping Xu, Rui Qiu, and Zhimiao Tao. Rough approximation operators in group mapping and their applications to coding theory. Fundamenta Informaticae, 145(1):93–109, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Han:2016:CEM

[1097] Huiting Han, Chunming Tang, Yu Lou, and Maozhi Xu. Construction of efficient MDS matrices based on block circulant matrices for lightweight application. Fundamenta Informaticae, 145(2):111–124, ???? 2016. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nagy:2016:DPT

[1098] Benedek Nagy and Tibor Lukić. Dense projection tomography on the triangular tiling. Fundamenta Informaticae, 145(2):125–141, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pomykala:2016:SGS

[1099] Jacek Pomykała. Small generating sets and DLPC problem. Fundamenta Informaticae, 145(2):143–150, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Song:2016:NGT

[1100] Yinglei Song, Junfeng Qu, Ying Xu, and Liming Cai. A new graph theoretic approach for protein threading. Fundamenta Informaticae, 145 (2):151–170, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tian:2016:IBS

[1101] Miaomiao Tian and Liusheng Huang. Identity-based signatures from lattices: Simpler, faster, shorter. Fundamenta Informaticae, 145(2):171– 187, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2016:QEM

[1102] Hong ying Yang, Lin lin Liang, Yong wei Li, and Xiang yang Wang. Quaternion exponent moments and their invariants for color image. Fundamenta Informaticae, 145(2):189–205, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zeng:2016:TSS

[1103] Zhi-Fang Zeng, Jian-Guo Liu, Yan Jiang, and Bin Nie. Transformations and soliton solutions for a variable-coefficient nonlinear Schrödinger equation in the dispersion decreasing fiber with symbolic computation. Fundamenta Informaticae, 145(2):207–219, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2016:P

[1104] Raymond Devillers, Antti Valmari, and Wojciech Penczek. Preface. Fundamenta Informaticae, 146(1):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Triebel:2016:CSD

[1105] Marvin Triebel and Jan Sürmeli. Characterizing stable and deriving valid inequalities of Petri nets. Fundamenta Informaticae, 146(1):1–34, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Badouel:2016:PNS

[1106] Eric Badouel, Loïc Hélouët, and Christophe Morvan. Petri nets with structured data. Fundamenta Informaticae, 146(1):35–82, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hujsa:2016:LRE

[1107] Thomas Hujsa, Jean-Marc Delosme, and Alix Munier-Kordon. On liveness and reversibility of equal-conflict Petri nets. Fundamenta Informaticae, 146(1):83–119, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mroz:2016:CEBa

[1108] Andrzej Mróz. Congruences of edge-bipartite graphs with applications to Grothendieck group recognition I. Inflation algorithm revisited*. Fundamenta Informaticae, 146(2):121–144, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mroz:2016:CEBb

[1109] Andrzej Mróz. Congruences of edge-bipartite graphs with applications to Grothendieck group recognition II. Coxeter type study*. Fundamenta Informaticae, 146(2):145–177, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2016:P

[1110] Paolo Dulio, Andrea Frosini, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 146(2):179–183, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brunetti:2016:ABA

[1111] Sara Brunetti and Carla Peri. On J-additivity and bounded additivity. Fundamenta Informaticae, 146(2):185–195, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gambini:2016:TSP

[1112] Ian Gambini and Laurent Vuillon. Tiling the space by polycube analogues of Fedorov’s polyhedra. Fundamenta Informaticae, 146(2):197– 209, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Smaglichenko:2016:MFM

[1113] Tatyana A. Smaglichenko, Ingi Th. Bjarnason, Alexander V. Smaglichenko, and Wolfgang R. Jacoby. Method to find the minimum 1D linear gradient model for seismic tomography. Fundamenta Informaticae, 146(2):211– 217, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vigano:2016:OSS

[1114] Nicola Viganò, Kees Joost Batenburg, and Wolfgang Ludwig. An orientation-space super sampling technique for six-dimensional diffraction contrast tomography. Fundamenta Informaticae, 146(2):219–230, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karhumaki:2016:P

[1115] J. Karhumäki, V. Mazalov, and Yu. Matiyasevich. Preface. Fundamenta Informaticae, 145(3):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ananichev:2016:ART

[1116] Dimitry S. Ananichev and Vladimir V. Gusev. Approximation of reset thresholds with greedy algorithms. Fundamenta Informaticae, 145(3): 221–227, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Itsykson:2016:TLB

[1117] Dmitry Itsykson, Vsevolod Oparin, Mikhail Slabodkin, and Dmitry Sokolov. Tight lower bounds on the resolution complexity of perfect matching principles. Fundamenta Informaticae, 145(3):229–242, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Junnila:2016:MNI

[1118] Ville Junnila and Tero Laihonen. Minimum number of input clues in robust information retrieval. Fundamenta Informaticae, 145(3):243–256, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kari:2016:PAF

[1119] Jarkko Kari. Piecewise affine functions, Sturmian sequences and Wang tiles. Fundamenta Informaticae, 145(3):257–277, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Karpov:2016:MCG

[1120] Dmitri V. Karpov. Minimal k-connected graphs with small number of vertices of degree k. Fundamenta Informaticae, 145(3):279–312, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leri:2016:FFM

[1121] Marina Leri and Yuri Pavlov. Forest fire models on configuration random graphs. Fundamenta Informaticae, 145(3):313–322, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Matsuhisa:2016:CLS

[1122] Takashi Matsuhisa. Communication leading to subgroup Nash equilibrium for generalized information. Fundamenta Informaticae, 145(3):323– 340, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mazalov:2016:GTC

[1123] V. V. Mazalov, K. E. Avrachenkov, L. I. Trukhina, and B. T. Tsynguev. Game-theoretic centrality measures for weighted graphs. Fundamenta Informaticae, 145(3):341–358, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Raigorodskii:2016:CGC

[1124] A. M. Raigorodskii. Combinatorial geometry and coding theory. Fundamenta Informaticae, 145(3):359–369, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bagchi:2016:PFP

[1125] Susmit Bagchi. Probabilistic and fuzzy process classifiers for operating systems scheduler. Fundamenta Informaticae, 145(4):405–427, ????

2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hidders:2016:FUD

[1126] Jan Hidders and Jan Paredaens. A formal and unified description of XML manipulation languages. Fundamenta Informaticae, 145(4):429– 470, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Komarath:2016:CR

[1127] Balagopal Komarath, Jayalal Sarma, and K. S. Sunil. On the complexity of L-reachability. Fundamenta Informaticae, 145(4):471–483, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2016:SLG

[1128] Zhiqiang Zhang, Tingfang Wu, and Linqiang Pan. On string languages generated by sequential numerical P systems. Fundamenta Informaticae, 145(4):485–509, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVb

[1129] Anonymous. Author index volume 145. Fundamenta Informaticae, 145 (4):511–512, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arqub:2016:ASD

[1130] Omar Abu Arqub. Approximate solutions of DASs with nonclassical boundary conditions using novel reproducing kernel algorithm. Fundamenta Informaticae, 146(3):231–254, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2016:ASI

[1131] Behrouz Kheirfam. An arc-search interior point method in the ????_{∞} neighborhood for symmetric optimization. Fundamenta Informaticae, 146(3):255–269, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mandal:2016:ACB

[1132] Bimal Mandal, Pantelimon Stănică, Sugata Gangopadhyay, and Enes Pasalic. An analysis of the C class of bent functions. Fundamenta Informaticae, 146(3):271–292, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Qi:2016:RSA

[1133] Ji Qi and Yukio Ohsawa. Recut: A seriation algorithm balancing smooth display and aggregated features. Fundamenta Informaticae, 146(3):293– 304, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Teh:2016:PMP

[1134] Wen Chean Teh. Parikh matrices and Parikh rewriting systems. Fundamenta Informaticae, 146(3):305–320, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wedley:2016:EAR

[1135] William C. Wedley, Eng U. Choo, and Diederik J. D. Wijnmalen. Efficacy analysis of ratios from pairwise comparisons. Fundamenta Informaticae, 146(3):321–338, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eyraud:2016:P

[1136] Rémi Eyraud, Colin de la Higuera, Makoto Kanazawa, and Ryo Yoshinaka. Preface. Fundamenta Informaticae, 146(4):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Clark:2016:DLS

[1137] Alexander Clark, Makoto Kanazawa, Gregory M. Kobele, and Ryo Yoshinaka. Distributional learning of some nonlinear tree grammars. Fundamenta Informaticae, 146(4):339–377, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Scicluna:2016:GIP

[1138] James Scicluna and Colin de la Higuera. Grammatical inference of PCFGs applied to language modelling and unsupervised parsing. Fundamenta Informaticae, 146(4):379–402, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Eyraud:2016:DLS

[1139] Rémi Eyraud, Jean-Christophe Janodet, Tim Oates, and Frédéric Papadopoulos. Designing and learning substitutable plane graph grammars. Fundamenta Informaticae, 146(4):403–430, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Beros:2016:CSD

[1140] Achilles A. Beros and Colin de la Higuera. A canonical semi-deterministic transducer. Fundamenta Informaticae, 146(4):431–459, ???? 2016. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVc

[1141] Anonymous. Author index volume 146. Fundamenta Informaticae, 146 (4):461–462, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Inclezan:2016:P

[1142] Daniela Inclezan, Marco Maratea, and Victor Marek. Preface. Fundamenta Informaticae, 147(1):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abseher:2016:SDA

[1143] Michael Abseher, Martin Gebser, Nysret Musliu, Torsten Schaub, and Stefan Woltran. Shift design with answer set programming. Fundamenta Informaticae, 147(1):1–25, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bliem:2016:DFS

[1144] Bernhard Bliem, Günther Charwat, Markus Hecher, and Stefan Woltran. D-FLAT^{2}: Subset minimization in dynamic programming on tree decompositions made easy. Fundamenta Informaticae, 147(1):27–61, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bomanson:2016:ASP

[1145] Jori Bomanson, Martin Gebser, Tomi Janhunen, Benjamin Kaufmann, and Torsten Schaub. Answer set programming modulo acyclicity. Fundamenta Informaticae, 147(1):63–91, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fandinno:2016:TDC

[1146] Jorge Fandinno. Towards deriving conclusions from cause-effect relations. Fundamenta Informaticae, 147(1):93–131, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2016:LPG

[1147] Zhizheng Zhang, Bin Wang, and Shutao Zhang. Logic programming with graded introspection. Fundamenta Informaticae, 147(1):133–158, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2016:Pb

[1148] Ludwik Czaja, Wojciech Penczek, and Krzysztof Stencel. Preface. Fundamenta Informaticae, 147(2–3):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alsolami:2016:DPA

[1149] Fawaz Alsolami, Talha Amin, Igor Chikalov, Mikhail Moshkov, and Beata Zielosko. Dynamic programming approach for construction of association rule systems. Fundamenta Informaticae, 147(2–3):159–171, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barbuti:2016:SPR

[1150] Roberto Barbuti, Roberta Gori, Francesca Levi, and Paolo Milazzo. Specialized predictor for reaction systems with context properties. Fundamenta Informaticae, 147(2–3):173–191, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Buregwa-Czuma:2016:MDC

[1151] Sylwia Buregwa-Czuma, Jan G. Bazan, Lech Zareba, Stanislawa Bazan-Socha, Barbara Rewerska, Przemyslaw Pardel, and Lukasz Dydo. The method for describing changes in the perception of stenosis in blood vessels caused by an additional drug. Fundamenta Informaticae, 147 (2–3):193–207, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2016:RMC

[1152] Ludwik Czaja. Remarks on memory consistency description. Fundamenta Informaticae, 147(2–3):209–221, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2016:LTR

[1153] Adam Grabowski. Lattice theory for rough sets – a case study with Mizar. Fundamenta Informaticae, 147(2–3):223–240, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kopczynski:2016:CLD

[1154] Maciej Kopczyński, Tomasz Grześ, and Jarosław Stepaniuk. Core for large datasets: Rough sets on FPGA. Fundamenta Informaticae, 147 (2–3):241–259, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2016:ETG

[1155] Linh Anh Nguyen. ExpTime tableaux with global caching for graded propositional dynamic logic. Fundamenta Informaticae, 147(2–3):261– 288, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2016:CPP

[1156] Artur Niewiadomski, Jaroslaw Skaruz, Piotr Switalski, and Wojciech Penczek. Concrete planning in PlanICS framework by combining SMT with GEO and simulated annealing. Fundamenta Informaticae, 147(2– 3):289–313, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Podymov:2016:EEC

[1157] Vladislav Podymov. An efficient equivalence-checking algorithm for a model of programs with commutative and absorptive statements. Fundamenta Informaticae, 147(2–3):315–336, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Magott:2016:SET

[1158] Jan Magott and Paweł Skrobanek. Semantics and execution time of new timed statecharts. Fundamenta Informaticae, 147(4):387–414, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peng:2016:FPP

[1159] Xindong Peng and Huiyong Yuan. Fundamental properties of Pythagorean fuzzy aggregation operators. Fundamenta Informaticae, 147(4):415–446, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Poklukar:2016:RHW

[1160] Darja Rupnik Poklukar and Janez Žerovnik. Reliability Hosoya–Wiener polynomial of double weighted trees. Fundamenta Informaticae, 147(4): 447–456, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2016:FTM

[1161] Xiawei Zhang and Qingzhao Kong. On four types of multi-covering rough sets. Fundamenta Informaticae, 147(4):457–476, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2016:TSR

[1162] Yan-Lan Zhang, Jinjin Li, and Changqing Li. Topological structure of relation-based generalized rough sets. Fundamenta Informaticae, 147(4): 477–491, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVd

[1163] Anonymous. Author index volume 147. Fundamenta Informaticae, 147 (4):493–494, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ciucci:2016:P

[1164] Davide Ciucci, Dominik Ślȩzak, and Marcin Wolski. Preface. Fundamenta Informaticae, 148(1–2):??, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rauch:2016:LAD

[1165] Jan Rauch. Logical aspects of dealing with domain knowledge in data mining with association rules. Fundamenta Informaticae, 148(1–2):1– 33, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nowak-Brzezinska:2016:MRB

[1166] Agnieszka Nowak-Brzezińska. Mining rule-based knowledge bases inspired by rough set theory. Fundamenta Informaticae, 148(1–2):35–50, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Napierala:2016:PPB

[1167] Krystyna Napierala and Jerzy Stefanowski. Post-processing of BRACID rules induced from imbalanced data. Fundamenta Informaticae, 148(1– 2):51–64, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kryszkiewicz:2016:AAA

[1168] Marzena Kryszkiewicz. ACBC-adequate association and decision rules versus key generators and rough sets approximations. Fundamenta Informaticae, 148(1–2):65–85, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zielosko:2016:ADP

[1169] Beata Zielosko. Application of dynamic programming approach to optimization of association rules relative to coverage and length. Fundamenta Informaticae, 148(1–2):87–105, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kopczynski:2016:RSB

[1170] Maciej Kopczyński, Tomasz Grześ, and Jarosław Stepaniuk. Rough sets based LEM2 rules generation supported by FPGA. Fundamenta Informaticae, 148(1–2):107–122, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraborty:2016:SIF

[1171] Mihir K. Chakraborty. On some issues in the foundation of rough sets: the problem of definition. Fundamenta Informaticae, 148(1–2):123–132, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2016:RHB

[1172] Pradipta Maji and Ankita Mandal. Rough hypercuboid based supervised regularized canonical correlation for multimodal data analysis*. Fundamenta Informaticae, 148(1–2):133–155, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2016:RGC

[1173] Marcin Wolski and Anna Gomolińska. Rough granular computing in modal settings: Generalised approximation spaces. Fundamenta Informaticae, 148(1–2):157–172, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bensch:2016:P

[1174] Suna Bensch, Rudolf Freund, Mika Hirvensalo, and Friedrich Otto. Preface. Fundamenta Informaticae, 148(3–4):??, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Campeanu:2016:DOC

[1175] Cezar Câmpeanu, Nelma Moreira, and Rogério Reis. Distinguishability operations and closures. Fundamenta Informaticae, 148(3–4):243–266, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Holzer:2016:CCP

[1176] Markus Holzer, Sebastian Jakobi, and Matthias Wendlandt. On the computational complexity of partial word automata problems. Fundamenta Informaticae, 148(3–4):267–289, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ibarra:2016:VPA

[1177] Oscar H. Ibarra. Visibly pushdown automata and transducers with counters. Fundamenta Informaticae, 148(3–4):291–308, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Krtek:2016:TDL

[1178] Lukáš Krtek and František Mráz. Two-dimensional limited context restarting automata. Fundamenta Informaticae, 148(3–4):309–340, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2016:RQA

[1179] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Reversible queue automata. Fundamenta Informaticae, 148(3–4):341–368, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pighizzini:2016:SLA

[1180] Giovanni Pighizzini. Strongly limited automata. Fundamenta Informaticae, 148(3–4):369–392, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVe

[1181] Anonymous. Author index volume 148. Fundamenta Informaticae, 148 (3–4):393–394, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bistarelli:2016:P

[1182] Stefano Bistarelli, Andrea Formisano, Marco Maratea, and Paolo Torroni. Preface. Fundamenta Informaticae, 149(1–2):??, ???? 2016. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alviano:2016:EAS

[1183] Mario Alviano. Evaluating answer set programming with non-convex recursive aggregates. Fundamenta Informaticae, 149(1–2):1–34, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bartak:2016:ESI

[1184] Roman Barták and Jindřich Vodrážka. An experimental study of influence of modeling and solving techniques on performance of a tabled logic programming planner. Fundamenta Informaticae, 149(1–2):35–60, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mancini:2016:NNN

[1185] Toni Mancini. Now or never: Negotiating efficiently with unknown or untrusted counterparts. Fundamenta Informaticae, 149(1–2):61–100, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mancini:2016:SSL

[1186] Toni Mancini, Federico Mari, Annalisa Massini, Igor Melatti, and Enrico Tronci. SyLVaaS: System level formal verification as a service. Fundamenta Informaticae, 149(1–2):101–132, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marin:2016:TYQ

[1187] Paolo Marin, Massimo Narizzano, Luca Pulina, Armando Tacchella, and Enrico Giunchiglia. Twelve years of QBF evaluations: QSAT is PSPACEhard and it shows. Fundamenta Informaticae, 149(1–2):133–158, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schuller:2016:MVF

[1188] Peter Schüller. Modeling variations of first-order Horn abduction in answer set programming. Fundamenta Informaticae, 149(1–2):159–207, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vallati:2016:IEF

[1189] Mauro Vallati, Ivan Serina, Alessandro Saetti, and Alfonso Emilio Gerevini. Identifying and exploiting features for effective plan retrieval in case-based planning. Fundamenta Informaticae, 149(1–2):209–240, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Abarca:2016:GTA

[1190] M. Abarca and D. Rivera. Graph theoretical and algorithmic characterizations of positive definite symmetric quasi-Cartan matrices. Fundamenta Informaticae, 149(3):241–261, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alessi:2016:TO

[1191] Fabio Alessi and Felice Cardone. Tiered objects. Fundamenta Informaticae, 149(3):263–295, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2016:ISS

[1192] J. A. Bergstra and C. A. Middelburg. Instruction sequence size complexity of parity. Fundamenta Informaticae, 149(3):297–309, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Champarnaud:2016:CET

[1193] Jean-Marc Champarnaud, Ludovic Mignot, and Florent Nicart. Constrained expressions and their derivatives. Fundamenta Informaticae, 149(3):311–361, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Doberkat:2016:BCS

[1194] Ernst-Erich Doberkat. Bisimulation cuts for structuring Markov transition systems. Fundamenta Informaticae, 149(4):363–383, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grzeskowiak:2016:PFP

[1195] Maciej Grześkowiak. Pairing-friendly primes for Abelian varieties. Fundamenta Informaticae, 149(4):385–400, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Virbitskaite:2016:TCE

[1196] Irina Virbitskaite, Dmitry Bushin, and Eike Best. True concurrent equivalences in time Petri nets. Fundamenta Informaticae, 149(4):401–418, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wrobel:2016:RQM

[1197] Łukasz Wróbel, Marek Sikora, and Marcin Michalak. Rule quality measures settings in classification, regression and survival rule induction — an empirical approach. Fundamenta Informaticae, 149(4):419–449, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wu:2016:SUS

[1198] Tingfang Wu, Yanfeng Wang, Suxia Jiang, and Xiaolong Shi. Small universal spiking neural P systems with homogeneous neurons and synapses. Fundamenta Informaticae, 149(4):451–470, ???? 2016. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2016:AIVf

[1199] Anonymous. Author index volume 149. Fundamenta Informaticae, 149 (4):471–472, ???? 2016. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alberti:2017:FVP

[1200] Francesco Alberti, Silvio Ghilardi, and Natasha Sharygina. A framework for the verification of parameterized infinite-state systems. Fundamenta Informaticae, 150(1):1–24, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Benedetti:2017:AFV

[1201] Irene Benedetti and Stefano Bistarelli. From argumentation frameworks to voting systems and back. Fundamenta Informaticae, 150(1):25–48, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cantone:2017:HSQ

[1202] Domenico Cantone, Cristiano Longo, and Marianna Nicolosi-Asmundo. Herbrand-satisfiability of a quantified set-theoretic fragment. Fundamenta Informaticae, 150(1):49–71, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeAngelis:2017:PVU

[1203] Emanuele De Angelis, Fabio Fioravanti, Alberto Pettorossi, and Maurizio Proietti. Program verification using constraint handling rules and array constraint generalizations. Fundamenta Informaticae, 150(1):73– 117, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferrari:2017:JJF

[1204] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. JTabWb: a Java framework for implementing terminating sequent and tableau calculi. Fundamenta Informaticae, 150(1):119–142, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordano:2017:P

[1205] Laura Giordano, Valentina Gliozzi, Alberto Pettorossi, and Gian Luca Pozzato. Preface. Fundamenta Informaticae, 150(1):??, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bozapalidis:2017:SS

[1206] Symeon Bozapalidis and George Rahonis. Stochastic semantics. Fundamenta Informaticae, 150(2):143–177, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gu:2017:IBM

[1207] Ke Gu, Weijia Jia, and Jianming Zhang. Identity-based multi-proxy signature scheme in the standard model. Fundamenta Informaticae, 150 (2):179–210, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Juneam:2017:FEP

[1208] Nopadon Juneam and Sanpawat Kantabutra. Fast and efficient parallel coarsest refinement. Fundamenta Informaticae, 150(2):211–220, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Przybyszewski:2017:TME

[1209] Andrzej W. Przybyszewski and Lech T. Polkowski. Theory of mind and empathy. Part I — model of social emotional thinking. Fundamenta Informaticae, 150(2):221–230, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2017:NSN

[1210] Yang Yu, Tingfang Wu, Jinbang Xu, Yanfeng Wang, and Juanjuan He. A note on spiking neural P systems with homogeneous neurons and synapses. Fundamenta Informaticae, 150(2):231–240, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grathwohl:2017:IAE

[1211] Niels Bjørn Bugge Grathwohl, Fritz Henglein, and Dexter Kozen. Infinitary axiomatization of the equational theory of context-free languages. Fundamenta Informaticae, 150(3–4):241–257, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berardi:2017:NMP

[1212] Stefano Berardi and Ugo de’ Liguoro. Non-monotonic pre-fix points and learning. Fundamenta Informaticae, 150(3–4):259–280, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dawar:2017:PGA

[1213] Anuj Dawar and Bjarki Holm. Pebble games with algebraic rules. Fundamenta Informaticae, 150(3–4):281–316, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mio:2017:LC

[1214] Matteo Mio and Alex Simpson. Łukasiewicz μ-calculus. Fundamenta Informaticae, 150(3–4):317–346, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jeannin:2017:CFP

[1215] Jean-Baptiste Jeannin, Dexter Kozen, and Alexandra Silva. CoCaml: Functional programming with regular coinductive types. Fundamenta Informaticae, 150(3–4):347–377, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cirstea:2017:BLT

[1216] Corina Cîrstea. From branching to linear time, coalgebraically. Fundamenta Informaticae, 150(3–4):379–406, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Milius:2017:GYD

[1217] Stefan Milius and Tadeusz Litak. Guard your daggers and traces: Properties of guarded (co-)recursion. Fundamenta Informaticae, 150(3–4): 407–449, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVa

[1218] Anonymous. Author index volume 150. Fundamenta Informaticae, 150 (3–4):451–452, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baelde:2017:P

[1219] David Baelde, Arnaud Carayol, Ralph Matthes, and Igor Walukiewicz. Preface. Fundamenta Informaticae, 150(3–4):??, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2017:CSF

[1220] Yudong Zhang, Shuihua Wang, Jian-Fei Yang, Zheng Zhang, Preetha Phillips, Ping Sun, and Jie Yan. A comprehensive survey on fractional Fourier transform. Fundamenta Informaticae, 151(1–4):1–48, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Murugusundaramoorthy:2017:CEB

[1221] G. Murugusundaramoorthy, T. Janani, and Sunil Dutt Purohit. Coefficient estimate of bi-Bazilevič functions associated with fractional q-calculus operators. Fundamenta Informaticae, 151(1–4):49–62, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2017:NFL

[1222] Xiao-Jun Yang, J. A. Tenreiro Machado, and Juan J. Nieto. A new family of the local fractional PDEs. Fundamenta Informaticae, 151(1–4):63– 75, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gupta:2017:TDC

[1223] A. K. Gupta and S. Saha Ray. Two-dimensional Chebyshev wavelet method for Camassa–Holm equation with Riesz fractional derivative describing propagation of shallow water waves. Fundamenta Informaticae, 151(1–4):77–89, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ahmad:2017:SCS

[1224] Bashir Ahmad, Sotiris K. Ntouyas, Ahmed Alsaedi, and Aatef Hobiny. Solvability of a coupled system of fractional differential equations with nonlocal and integral boundary conditions. Fundamenta Informaticae, 151(1–4):91–108, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kailasavalli:2017:FNI

[1225] S. Kailasavalli, S. Suganya, and M. Mallika Arjunan. On fractional neutral integro-differential systems with state-dependent delay in Banach spaces. Fundamenta Informaticae, 151(1–4):109–133, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Babaei:2017:AAS

[1226] Afshin Babaei. On analytical approximate solution of the fractional type Rosenau–Hyman equation. Fundamenta Informaticae, 151(1–4):135–143, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Atangana:2017:MTV

[1227] Abdon Atangana and Ilknur Koca. Model of thin viscous fluid sheet flow within the scope of fractional calculus: Fractional derivative with and no singular kernel. Fundamenta Informaticae, 151(1–4):145–159, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jafari:2017:ARD

[1228] Hossein Jafari, Atena Lia, Haleh Tejadodi, and Dumitru Baleanu. Analysis of Riccati differential equations within a new fractional derivative without singular kernel. Fundamenta Informaticae, 151(1–4):161–171, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zhang:2017:P

[1229] Yudong Zhang, Xiao-Jun Yang, Carlo Cattani, Zhengchao Dong, Ti-Fei Yuan, and Liang-Xiu Han. Preface. Fundamenta Informaticae, 151(1–4): ??, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Basu:2017:FMM

[1230] Nilanjana G. Basu, Subhashis Majumder, and Wing Kai Hon. On finding the maximum and minimum density axes-parallel regions in IR^{d}. Fundamenta Informaticae, 152(1):1–12, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chakraverty:2017:SFA

[1231] Snehashish Chakraverty, Milan Hladík, and Nisha Rani Mahato. A sign function approach to solve algebraically interval system of linear equations for nonnegative solutions. Fundamenta Informaticae, 152(1):13–31, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2017:PCI

[1232] Behrouz Kheirfam. A predictor–corrector infeasible-interior-point algorithm for semidefinite optimization in a wide neighborhood. Fundamenta Informaticae, 152(1):33–50, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Meduna:2017:JSC

[1233] Alexander Meduna and Ondřej Soukup. Jumping scattered context grammars. Fundamenta Informaticae, 152(1):51–86, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shin:2017:IDC

[1234] Kilho Shin. On infinite divisibility of convolution and mapping kernels. Fundamenta Informaticae, 152(1):87–105, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polat:2017:TSS

[1235] Nazan Çakmak Polat, Gözde Yaylalı, and Bekir Tanay. Tolerance soft set relation on a soft set and its matrix applications. Fundamenta Informaticae, 152(2):107–122, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koroth:2017:DLB

[1236] Sajin Koroth and Jayalal Sarma. Depth lower bounds against circuits with sparse orientation. Fundamenta Informaticae, 152(2):123–144, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mirhosseini:2017:MSA

[1237] Mina Mirhosseini and Hossein Nezamabadi-pour. Metaheuristic search algorithms in solving the n-similarity problem. Fundamenta Informaticae, 152(2):145–166, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Umadevi:2017:CRA

[1238] Dwaraganathan Umadevi. Characterizing rough algebras. Fundamenta Informaticae, 152(2):167–183, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zajac:2017:NAC

[1239] Katarzyna Zaja̧c. Numeric algorithms for corank two edge-bipartite graphs and their mesh geometries of roots. Fundamenta Informaticae, 152(2):185–222, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

ElKholy:2017:SNS

[1240] Warda El Kholy, Jamal Bentahar, Mohamed El Menshawy, Hongyang Qu, and Rachida Dssouli. SMC4AC: A new symbolic model checker for intelligent agent communication. Fundamenta Informaticae, 152(3):223– 271, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lemanska:2017:SDB

[1241] Magdalena Lemańska, Juan Alberto Rodríguez-Velázquez, and Rolando Trujillo-Rasua. Similarities and differences between the vertex cover number and the weakly connected domination number of a graph. Fundamenta Informaticae, 152(3):273–287, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pomykala:2017:DRF

[1242] Jacek Pomykała. On deterministic reduction of factoring integers to computing the exponents of elements in modular group. Fundamenta Informaticae, 152(3):289–295, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2017:RTM

[1243] Yong Yang and Xindong Peng. A revised TOPSIS method based on interval fuzzy soft set models with incomplete weight information. Fundamenta Informaticae, 152(3):297–321, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kopczynski:2017:CCB

[1244] Eryk Kopczyński. Computational complexity on the blackboard. Fundamenta Informaticae, 152(4):323–339, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Loukanova:2017:AFF

[1245] Roussanka Loukanova. An approach to functional formal models of constraint-based lexicalized grammar. Fundamenta Informaticae, 152 (4):341–372, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peng:2017:IVF

[1246] Xindong Peng, Jingguo Dai, and Huiyong Yuan. Interval-valued fuzzy soft decision making methods based on MABAC, similarity measure and EDAS. Fundamenta Informaticae, 152(4):373–396, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sima:2017:TSB

[1247] Jiří Šíma and Stanislav Žák. On tight separation for Blum measures applied to Turing machine buffer complexity. Fundamenta Informaticae, 152(4):397–409, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wozna-Szczesniak:2017:SBS

[1248] Bożena Woźna-Szcześniak, Agnieszka M. Zbrzezny, and Andrzej Zbrzezny. SMT-based searching for k-quasi-optimal runs in weighted timed automata. Fundamenta Informaticae, 152(4):411–433, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVb

[1249] Anonymous. Author index volume 152. Fundamenta Informaticae, 152 (4):435–436, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aman:2017:MDC

[1250] Bogdan Aman, Jetty Kleijn, Maciej Koutny, and Dorel Lucanu. Methods for distributed and concurrent systems: Special issue on the occasion of the 60th birthday of Professor Gabriel Ciobanu. Fundamenta Informaticae, 153(1–2):v–vi, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alhumaidan:2017:MAQ

[1251] Abdullah Alhumaidan and Jason Steggles. Modelling and analysing qualitative biological models using rewriting logic. Fundamenta Informaticae, 153(1–2):1–28, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergstra:2017:CTP

[1252] J. A. Bergstra and C. A. Middelburg. Contradiction-tolerant process algebra with propositional signals. Fundamenta Informaticae, 153(1–2): 29–55, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Horne:2017:SSA

[1253] Ross Horne, Sjouke Mauw, and Alwen Tiu. Semantics for specialising attack trees based on linear logic. Fundamenta Informaticae, 153(1–2): 57–86, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2017:SSB

[1254] Andrei Păun and Alfonso Rodríguez-Patón. P systems simulating bacterial conjugation: Universality and properties. Fundamenta Informaticae, 153(1–2):87–103, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Stoenescu:2017:BAB

[1255] Teodor Stoenescu, Alin Stefanescu, Sorina Predut, and Florentin Ipate. Binary analysis based on symbolic execution and reversible x86 instructions. Fundamenta Informaticae, 153(1–2):105–124, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Todoran:2017:CSC

[1256] Eneia Nicolae Todoran and Nikolaos Papaspyrou. Concurrency semantics in continuation-passing style. Fundamenta Informaticae, 153(1–2):125– 146, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valencia-Cabrera:2017:CEM

[1257] Luis Valencia-Cabrera, David Orellana-Martín, Miguel A. Martínez del Amor, Agustín Riscos-Núñez, and Mario J. Pérez-Jiménez. Computational efficiency of minimal cooperation and distribution in polarizationless P systems with active membranes. Fundamenta Informaticae, 153 (1–2):147–172, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chai:2017:MPE

[1258] Ming Chai and Bernd-Holger Schlingloff. Monitoring with parametrized extended life sequence charts. Fundamenta Informaticae, 153(3):173– 198, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chlebus:2017:NCB

[1259] Bogdan S. Chlebus, Gianluca De Marco, and Muhammed Talo. Naming a channel with beeps. Fundamenta Informaticae, 153(3):199–219, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Denisiuk:2017:LDE

[1260] Aleksander Denisiuk and Michał Grabowski. Low distortion embedding of the Hamming space into a sphere with quadrance metric and k-means clustering of nominal-continuous data. Fundamenta Informaticae, 153 (3):221–233, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Joshi:2017:EGW

[1261] Himani Joshi and Sankalap Arora. Enhanced grey wolf optimization algorithm for global optimization. Fundamenta Informaticae, 153(3): 235–264, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koprowski:2017:RMC

[1262] Przemysław Koprowski. Roots multiplicity without companion matrices. Fundamenta Informaticae, 153(3):265–270, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Syau:2017:NSV

[1263] Yu-Ru Syau, En-Bing Lin, and Churn-Jung Liau. Neighborhood systems and variable precision generalized rough sets. Fundamenta Informaticae, 153(3):271–290, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chiaselotti:2017:RST

[1264] Giampiero Chiaselotti, Davide Ciucci, Tommaso Gentile, and Federico Infusino. Rough set theory and digraphs. Fundamenta Informaticae, 153 (4):291–325, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2017:WNS

[1265] Behrouz Kheirfam and Mohaddeseh Mohamadi-Sangachin. A wide neighborhood second-order predictor-corrector interior-point algorithm for semidefinite optimization with modified corrector directions. Fundamenta Informaticae, 153(4):327–346, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanLeeuwen:2017:TMO

[1266] Jan van Leeuwen and Jiří Wiedermann. Turing machines with one-sided advice and acceptance of the co-RE languages. Fundamenta Informaticae, 153(4):347–366, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Makuracki:2017:IAC

[1267] Bartosz Makuracki, Daniel Simson, and Błażej Zyglarski. Inflation algorithm for Cox-regular positive edge-bipartite graphs with loops. Fundamenta Informaticae, 153(4):367–398, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wronski:2017:CEC

[1268] Michał Wroński, Robert Dryło, Tomasz Kijko, and Piotr Bora. Constructing elliptic curves for the GLV method with low-cost decomposition. Fundamenta Informaticae, 153(4):399–413, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVc

[1269] Anonymous. Author index volume 153. Fundamenta Informaticae, 153 (4):415–416, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Paun:2017:P

[1270] Gheorghe Păun, Grzegorz Rozenberg, and Arto Salomaa. Preface. Fundamenta Informaticae, 154(1–4):vii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Badouel:2017:IPD

[1271] Eric Badouel and Uli Schlachter. Incremental process discovery using Petri net synthesis. Fundamenta Informaticae, 154(1–4):1–13, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barash:2017:FFS

[1272] Mikhail Barash and Ion Petre. A foundational framework for stepby-step model construction. Fundamenta Informaticae, 154(1–4):15–24, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bernardinello:2017:STS

[1273] Luca Bernardinello, Carlo Ferigato, Lucia Pomello, and Adrián Puerto Aubel. Synthesis of transition systems from quantum logics. Fundamenta Informaticae, 154(1–4):25–36, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bojanczyk:2017:IUI

[1274] Mikołaj Bojańczyk. It is undecidable if two regular tree languages can be separated by a deterministic tree-walking automaton. Fundamenta Informaticae, 154(1–4):37–46, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bonizzoni:2017:SDP

[1275] Paola Bonizzoni, Anna Paola Carrieri, Gianluca Della Vedova, Raffaella Rizzi, and Gabriella Trucco. Species-driven persistent phylogeny. Fundamenta Informaticae, 154(1–4):47–63, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cassaigne:2017:AER

[1276] Julien Cassaigne, Juhani Karhumäki, Svetlana Puzynina, and Markus A. Whiteland. k-Abelian equivalence and rationality. Fundamenta Informaticae, 154(1–4):65–94, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dutta:2017:ILS

[1277] Soma Dutta, Andrzej Jankowski, and Andrzej Skowron. Interactive logical structures. Fundamenta Informaticae, 154(1–4):95–108, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gabow:2017:WMA

[1278] Harold N. Gabow. The weighted matching approach to maximum cardinality matching. Fundamenta Informaticae, 154(1–4):109–130, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Genova:2017:ERL

[1279] Daniela Genova and Hendrik Jan Hoogeboom. Enforcing regular languages. Fundamenta Informaticae, 154(1–4):131–144, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Durand-Lose:2017:PE

[1280] Jérôme Durand-Lose, Jarkko Kari, and Benedek Nagy. Preface / editorial. Fundamenta Informaticae, 155(1–2):v–vii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2017:ISD

[1281] Ryszard Janicki, Jetty Kleijn, Maciej Koutny, and Łukasz Mikulski. Invariant structures and dependence relations. Fundamenta Informaticae, 155(1–2):1–29, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2017:RLA

[1282] Martin Kutrib and Matthias Wendlandt. Reversible limited automata. Fundamenta Informaticae, 155(1–2):31–58, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kutrib:2017:TDP

[1283] Martin Kutrib, Andreas Malcher, and Matthias Wendlandt. Tinputdriven pushdown, counter, and stack automata. Fundamenta Informaticae, 155(1–2):59–88, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drewes:2017:TBC

[1284] Frank Drewes, Markus Holzer, Sebastian Jakobi, and Brink van der Merwe. Tight bounds for cut-operations on deterministic finite automata. Fundamenta Informaticae, 155(1–2):89–110, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csuhaj-Varju:2017:WCT

[1285] Erzsébet Csuhaj-Varjú, Rudolf Freund, and György Vaszil. Watson– Crick T0L systems and red–green register machines. Fundamenta Informaticae, 155(1–2):111–129, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hendricks:2017:SPL

[1286] Jacob Hendricks, Matthew J. Patitz, and Trent A. Rogers. The simulation powers and limitations of higher temperature hierarchical selfassembly systems. Fundamenta Informaticae, 155(1–2):131–162, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ivanov:2017:UCC

[1287] Sergiu Ivanov and Sergey Verlan. Universality and computational completeness of controlled leftist insertion-deletion systems. Fundamenta Informaticae, 155(1–2):163–185, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nagy:2017:SFC

[1288] Benedek Nagy and Sándor Vályi. A shift-free characterization of NP within interval-valued computing. Fundamenta Informaticae, 155(1–2): 187–207, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fernau:2017:NIC

[1289] Henning Fernau, Rudolf Freund, Rani Siromoney, and K. G. Subramanian. Non-isometric contextual array grammars and the role of regular control and local selectors. Fundamenta Informaticae, 155(1–2):209–232, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gaggl:2017:P

[1290] Sarah Alice Gaggl, Juan Carlos Nieves, Hannes Strass, and Paolo Torroni. Preface. Fundamenta Informaticae, 155(3):i–iii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maher:2017:RCD

[1291] Michael J. Maher. Relating concrete defeasible reasoning formalisms and abstract argumentation. Fundamenta Informaticae, 155(3):233–260, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sakama:2017:RAF

[1292] Chiaki Sakama and Tjitze Rienstra. Representing argumentation frameworks in answer set programming. Fundamenta Informaticae, 155(3): 261–292, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Osorio:2017:ASA

[1293] Mauricio Osorio, José Luis Carballido, and Claudia Zepeda. Abducible semantics and argumentation. Fundamenta Informaticae, 155(3):293– 319, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2017:P

[1294] Paolo Dulio, Andrea Frosini, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 155(4):i–vii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alpers:2017:RBM

[1295] Andreas Alpers and Peter Gritzmann. Reconstructing binary matrices under window constraints from their row and column sums. Fundamenta Informaticae, 155(4):321–340, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arcadu:2017:IAT

[1296] Filippo Arcadu, Jakob Vogel, Marco Stampanoni, and Federica Marone. Improving analytical tomographic reconstructions through consistency conditions. Fundamenta Informaticae, 155(4):341–361, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brun:2017:FGO

[1297] Emmanuel Brun, Claudio Ferrero, and Jerome Vicente. Fast granulometry operator for the 3D identification of cell structures. Fundamenta Informaticae, 155(4):363–372, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chapdelaine:2017:BCT

[1298] Camille Chapdelaine, Ali Mohammad-Djafari, Nicolas Gac, and Estelle Parra. A 3D Bayesian computed tomography reconstruction algorithm with Gauss–Markov–Potts prior model and its application to real data. Fundamenta Informaticae, 155(4):373–405, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2017:RUQ

[1299] Paolo Dulio, Silvia M. C. Pagani, and Andrea Frosini. Regions of uniqueness quickly reconstructed by three directions in discrete tomography. Fundamenta Informaticae, 155(4):407–423, ???? 2017. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hajdu:2017:CCD

[1300] Lajos Hajdu and Rob Tijdeman. Consistency conditions for discrete tomography. Fundamenta Informaticae, 155(4):425–447, ???? 2017. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2017:XRC

[1301] Li Wang, Ali Mohammad-Djafari, and Nicolas Gac. X-ray computed tomography using a sparsity enforcing prior model based on Haar transformation in a Bayesian framework. Fundamenta Informaticae, 155(4): 449–480, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVd

[1302] Anonymous. Author index volume 155. Fundamenta Informaticae, 155 (4):481–482, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bilotta:2017:PAL

[1303] Stefano Bilotta, Elisabetta Grazzini, and Elisa Pergola. Pattern avoiding languages and recurrence relations interpretation. Fundamenta Informaticae, 156(1):1–19, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ban:2017:MHA

[1304] Ha Bang Ban and Duc Nghia Nguyen. A meta-heuristic algorithm combining between tabu and variable neighborhood search for the minimum latency problem. Fundamenta Informaticae, 156(1):21–41, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klopotek:2017:SCB

[1305] Mieczysław Kłopotek. On seeking consensus between document similarity measures. Fundamenta Informaticae, 156(1):43–68, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mignot:2017:TAC

[1306] Ludovic Mignot, Nadia Ouali Sebti, and Djelloul Ziadi. Tree automata constructions from regular expressions: a comparative study. Fundamenta Informaticae, 156(1):69–94, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sroka:2017:DHW

[1307] Jacek Sroka and Jan Hidders. On determining the AND–OR hierarchy in workflow nets. Fundamenta Informaticae, 156(1):95–123, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Avesalon:2017:MPF

[1308] Nicoleta Avesalon, Eleonor Ciurea, and Mircea Parpalea. The maximum parametric flow in discrete-time dynamic networks. Fundamenta Informaticae, 156(2):125–139, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bartl:2017:RFC

[1309] Eduard Bartl and Jan Konecny. Rough fuzzy concept analysis. Fundamenta Informaticae, 156(2):141–168, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Csato:2017:EMR

[1310] László Csató. Eigenvector method and rank reversal in group decision making revisited. Fundamenta Informaticae, 156(2):169–178, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Durnoga:2017:LSM

[1311] Konrad Durnoga and Jacek Pomykała. Large sieve, Miller–Rabin compositeness witnesses and integer factoring problem. Fundamenta Informaticae, 156(2):179–185, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Su:2017:SNS

[1312] Yansen Su, Tingfang Wu, Fei Xu, and Andrei Păun. Spiking neural P systems with rules on synapses working in sum spikes consumption strategy. Fundamenta Informaticae, 156(2):187–208, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Varish:2017:NSM

[1313] Naushad Varish, Sumit Kumar, and Arup Kumar Pal. A novel similarity measure for content based image retrieval in discrete cosine transform domain. Fundamenta Informaticae, 156(2):209–235, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2017:ABG

[1314] Jingqian Wang and William Zhu. Applications of bipartite graphs and their adjacency matrices to covering-based rough sets. Fundamenta Informaticae, 156(2):237–254, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niwinski:2017:Pa

[1315] Damian Niwiński and Ewa Orłowska. Preface. Fundamenta Informaticae, 156(3–4):i–ii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niwinski:2017:Pb

[1316] Damian Niwiński and Ewa Orłowska. Preface. Fundamenta Informaticae, 156(3–4):i–ii, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czelakowski:2017:FFOa

[1317] Janusz Czelakowski. Forcing for first-order languages from the perspective of Rasiowa–Sikorski lemma. Fundamenta Informaticae, 156(3–4): 255–280, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czelakowski:2017:FFOb

[1318] Janusz Czelakowski. Forcing for first-order languages from the perspective of Rasiowa–Sikorski lemma. Fundamenta Informaticae, 156(3–4): 255–280, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2017:DRDa

[1319] Ivo Düntsch, Léonard Kwuida, and Ewa Orłowska. A discrete representation for dicomplemented lattices. Fundamenta Informaticae, 156 (3–4):281–295, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duntsch:2017:DRDb

[1320] Ivo Düntsch, Léonard Kwuida, and Ewa Orłowska. A discrete representation for dicomplemented lattices. Fundamenta Informaticae, 156 (3–4):281–295, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:2017:MQa

[1321] Melvin Fitting. On modalities and quantifiers. Fundamenta Informaticae, 156(3–4):297–330, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:2017:MQb

[1322] Melvin Fitting. On modalities and quantifiers. Fundamenta Informaticae, 156(3–4):297–330, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romanowska:2017:DQGa

[1323] Anna B. Romanowska and Jonathan D. H. Smith. Duality for quasilattices and Galois connections. Fundamenta Informaticae, 156(3–4):331– 359, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Romanowska:2017:DQGb

[1324] Anna B. Romanowska and Jonathan D. H. Smith. Duality for quasilattices and Galois connections. Fundamenta Informaticae, 156(3–4):331– 359, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sofronie-Stokkermans:2017:RTLa

[1325] Viorica Sofronie-Stokkermans. Representation theorems and locality for subsumption testing and interpolation in the description logics ϵL, ϵL^{+} and their extensions with n-ary roles and numerical domains. Fundamenta Informaticae, 156(3–4):361–411, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sofronie-Stokkermans:2017:RTLb

[1326] Viorica Sofronie-Stokkermans. Representation theorems and locality for subsumption testing and interpolation in the description logics ϵL, ϵL^{+} and their extensions with n-ary roles and numerical domains. Fundamenta Informaticae, 156(3–4):361–411, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vakarelov:2017:MPAa

[1327] Dimiter Vakarelov. Mereotopologies with predicates of actual existence and actual contact. Fundamenta Informaticae, 156(3–4):413–432, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Vakarelov:2017:MPAb

[1328] Dimiter Vakarelov. Mereotopologies with predicates of actual existence and actual contact. Fundamenta Informaticae, 156(3–4):413–432, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVe

[1329] Anonymous. Author index volume 156. Fundamenta Informaticae, 156 (3–4):433–434, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Anonymous:2017:AIVf

[1330] Anonymous. Author index volume 156. Fundamenta Informaticae, 156 (3–4):433–434, ???? 2017. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grahne:2018:AC

[1331] Gösta Grahne and Adrian Onet. Anatomy of the chase. Fundamenta Informaticae, 157(3):221–270, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wei:2018:PFH

[1332] Guiwu Wei. Picture fuzzy hamacher aggregation operators and their application to multiple attribute decision making. Fundamenta Informaticae, 157(3):271–320, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2018:P

[1333] Ludwik Czaja, Wojciech Penczek, Holger Schlingloff, and Nguyen Hung Son. Preface. Fundamenta Informaticae, 157(4):i–ii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Akhundov:2018:SPC

[1334] Jafar Akhundov, Peter Tröger, and Matthias Werner. Superposition principle in composable hybrid automata. Fundamenta Informaticae, 157(4):321–339, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Barylska:2018:RTB

[1335] Kamila Barylska, Evgeny Erofeev, Maciej Koutny, Łukasz Mikulski, and Marcin Pia̧tkowski. Reversing transitions in bounded Petri nets. Fundamenta Informaticae, 157(4):341–357, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czaja:2018:PME

[1336] Ludwik Czaja. A protocol of mutual exclusion for DSM based on vectors of global timestamps. Fundamenta Informaticae, 157(4):359–370, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jankowski:2018:RSS

[1337] Andrzej Jankowski, Andrzej Skowron, and Piotr Wasilewski. Rough sets and sorites paradox. Fundamenta Informaticae, 157(4):371–384, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Nguyen:2018:CBB

[1338] Linh Anh Nguyen. Computing bisimulation-based comparisons. Fundamenta Informaticae, 157(4):385–401, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2018:TWS

[1339] Artur Niewiadomski, Piotr Switalski, Marcin Kowalczyk, and Wojciech Penczek. TripICS — a Web service composition system for planning trips and travels. Fundamenta Informaticae, 157(4):403–425, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pelz:2018:FAT

[1340] Elisabeth Pelz. Full axiomatisation of timed processes of interval-timed Petri nets. Fundamenta Informaticae, 157(4):427–442, ???? 2018. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rataj:2018:EOP

[1341] Artur Rataj and Bożena Woźna-Szcześniak. Extrapolation of an optimal policy using statistical probabilistic model checking. Fundamenta Informaticae, 157(4):443–461, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Redziejowski:2018:TUP

[1342] Roman R. Redziejowski. Trying to understand PEG. Fundamenta Informaticae, 157(4):463–475, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chen:2018:P

[1343] Qingliang Chen, Paolo Torroni, and Serena Villata. Preface. Fundamenta Informaticae, 158(1–3):v–vii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dunin-Keplicz:2018:MPP

[1344] Barbara Dunin-Kȩplicz and Alina Powała. Multi-party persuasion: a paraconsistent approach. Fundamenta Informaticae, 158(1–3):1–39, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ihara:2018:SPC

[1345] Takamasa Ihara, Shunsuke Tsuruta, Taiki Todo, Yuko Sakurai, and Makoto Yokoo. Strategy-proof cake cutting mechanisms for all-or-nothing utility. Fundamenta Informaticae, 158(1–3):41–61, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Matsui:2018:LMO

[1346] Toshihiro Matsui, Marius Silaghi, Tenda Okimoto, Katsutoshi Hirayama, Makoto Yokoo, and Hiroshi Matsuo. Leximin multiple objective DCOPs on factor graphs for preferences of agents. Fundamenta Informaticae, 158(1–3):63–91, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Leask:2018:PAD

[1347] Sam Leask and Brian Logan. Programming agent deliberation using procedural reflection. Fundamenta Informaticae, 158(1–3):93–120, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lorini:2018:LTB

[1348] Emiliano Lorini and Roland Mühlenbernd. The long-term benefits of following fairness norms under dynamics of learning and evolution. Fundamenta Informaticae, 158(1–3):121–148, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2018:DLC

[1349] Chanjuan Liu, Fenrong Liu, and Kaile Su. A dynamic-logical characterization of solutions to sight-limited extensive games. Fundamenta Informaticae, 158(1–3):149–169, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Le:2018:MCS

[1350] Tiep Le, Tran Cao Son, and Enrico Pontelli. Multi-context systems with preferences. Fundamenta Informaticae, 158(1–3):171–216, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calardo:2018:SSM

[1351] Erica Calardo, Guido Governatori, and Antonino Rotolo. Sequence semantics for modelling reason-based preferences. Fundamenta Informaticae, 158(1–3):217–238, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Porello:2018:LMC

[1352] Daniele Porello. Logics for modelling collective attitudes. Fundamenta Informaticae, 158(1–3):239–275, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Amadini:2018:RAD

[1353] Roberto Amadini, Graeme Gange, François Gauthier, Alexander Jordan, Peter Schachte, Harald Søndergaard, Peter J. Stuckey, and Chenyi Zhang. Reference abstract domains and applications to string analysis. Fundamenta Informaticae, 158(4):297–326, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Badia:2018:DSF

[1354] Antonio Badia and Daniel Lemire. On desirable semantics of functional dependencies over databases with incomplete information. Fundamenta Informaticae, 158(4):327–352, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dost:2018:PSA

[1355] Şenol Dost. Precise sets in approximation spaces and textures. Fundamenta Informaticae, 158(4):353–368, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Perez:2018:CAC

[1356] Claudia Pérez, Mario Abarca, and Daniel Rivera. Cubic algorithm to compute the Dynkin type of a positive definite quasi-Cartan matrix. Fundamenta Informaticae, 158(4):369–384, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2018:CIW

[1357] Xiang yang Wang, Huan Xu, Si yu Zhang, Lin lin Liang, Pan pan Niu, and Hong ying Yang. A color image watermarking approach based on synchronization correction. Fundamenta Informaticae, 158(4):385–407, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maratea:2018:P

[1358] Marco Maratea, Viviana Mascardi, Davide Ancona, and Alberto Pettorossi. Preface. Fundamenta Informaticae, 159(1–2):v–vi, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baldoni:2018:CBA

[1359] Matteo Baldoni, Cristina Baroglio, Federico Capuzzimati, and Roberto Micalizio. Commitment-based agent interaction in JaCaMo+. Fundamenta Informaticae, 159(1–2):1–33, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chesani:2018:ECL

[1360] Federico Chesani, Marco Gavanelli, Evelina Lamma, Paola Mello, and Marco Montali. Evaluating compliance: From LTL to abductive logic programming. Fundamenta Informaticae, 159(1–2):35–63, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gavanelli:2018:RDO

[1361] Marco Gavanelli, Evelina Lamma, Fabrizio Riguzzi, Elena Bellodi, Riccardo Zese, and Giuseppe Cota. Reasoning on Datalog^{±} ontologies with abductive logic programming. Fundamenta Informaticae, 159(1–2):65– 93, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordano:2018:TRC

[1362] Laura Giordano, Valentina Gliozzi, and Nicola Olivetti. Towards a rational closure for expressive description logics: the case of S⟨⟩⨿. Fundamenta Informaticae, 159(1–2):95–122, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gottlob:2018:IAD

[1363] Georg Gottlob, Andreas Pieris, and Mantas Šimkus. The impact of active domain predicates on guarded existential rules. Fundamenta Informaticae, 159(1–2):123–146, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Lisi:2018:GCM

[1364] Francesca A. Lisi and Corrado Mencar. A granular computing method for OWL ontologies. Fundamenta Informaticae, 159(1–2):147–174, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Malvone:2018:AWS

[1365] Vadim Malvone, Aniello Murano, and Loredana Sorrentino. Additional winning strategies in reachability games. Fundamenta Informaticae, 159 (1–2):175–195, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

deBoer:2018:PNB

[1366] Frank S. de Boer, Mario Bravetti, Matias D. Lee, and Gianluigi Zavattaro. A Petri net based modeling of active objects and futures. Fundamenta Informaticae, 159(3):197–256, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ikegami:2018:PTM

[1367] Kenshin Ikegami and Yukio Ohsawa. PageRank topic model: Estimation of multinomial distributions using network structure analysis methods. Fundamenta Informaticae, 159(3):257–277, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Salehi:2018:AMT

[1368] Saeed Salehi. On axiomatizability of the multiplicative theory of numbers. Fundamenta Informaticae, 159(3):279–296, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sarkar:2018:FSTa

[1369] Apurba Sarkar, Arindam Biswas, Mousumi Dutt, and Shouvick Mondal. Finding shortest triangular path and its family inside a digital object. Fundamenta Informaticae, 159(3):297–325, ???? 2018. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Caucal:2018:SSI

[1370] Didier Caucal and Teodor Knapik. Shelah–Stupp’s iteration and Muchnik’s iteration. Fundamenta Informaticae, 159(4):327–359, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Farhadinia:2018:OWH

[1371] Bahram Farhadinia and Zeshui Xu. Ordered weighted hesitant fuzzy information fusion-based approach to multiple attribute decision making with probabilistic linguistic term sets. Fundamenta Informaticae, 159 (4):361–383, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gao:2018:SNP

[1372] Hui Gao, Mao Lu, Guiwu Wei, and Yu Wei. Some novel Pythagorean fuzzy interaction aggregation operators in multiple attribute decision making. Fundamenta Informaticae, 159(4):385–428, ???? 2018. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Syau:2018:NSR

[1373] Yu-Ru Syau, En-Bing Lin, and Churn-Jung Liau. Neighborhood systems: Rough set approximations and definability. Fundamenta Informaticae, 159(4):429–450, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gambin:2018:P

[1374] Anna Gambin and Monika Heiner. Preface. Fundamenta Informaticae, 160(1–2):v–vii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fortin:2018:HPN

[1375] Mathieu Pagé Fortin and Simon V. Hardy. A hybrid Petri net model of the Akt-Wnt-mTOR-p70S6K signalling network in neurons. Fundamenta Informaticae, 160(1–2):1–25, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Baldan:2018:PNM

[1376] Paolo Baldan, Martina Bocci, Daniele Brigolin, Nicoletta Cocco, Monika Heiner, and Marta Simeoni. Petri nets for modelling and analysing trophic networks. Fundamenta Informaticae, 160(1–2):27–52, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Herajy:2018:ABS

[1377] Mostafa Herajy and Monika Heiner. Adaptive and bio-semantics of continuous Petri nets: Choosing the appropriate interpretation. Fundamenta Informaticae, 160(1–2):53–80, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bordon:2018:SQM

[1378] Jure Bordon, Miha Moškon, Nikolaj Zimic, and Miha Mraz. Semiquantitative modelling of gene regulatory processes with unknown parameter values using fuzzy logic and Petri nets. Fundamenta Informaticae, 160(1–2):81–100, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Liu:2018:CFP

[1379] Fei Liu and Siyuan Chen. Colored fuzzy Petri nets for dealing with genetic regulatory networks. Fundamenta Informaticae, 160(1–2):101– 118, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delaplace:2018:AND

[1380] Franck Delaplace, Cinzia Di Giusto, Jean-Louis Giavitto, Hanna Klaudel, and Antoine Spicher. Activity networks with delays an application to toxicity analysis. Fundamenta Informaticae, 160(1–2):119–142, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Formanowicz:2018:FIE

[1381] Dorota Formanowicz, Agnieszka Rybarczyk, and Piotr Formanowicz. Factors influencing essential hypertension and cardiovascular disease modeled and analyzed using stochastic Petri nets. Fundamenta Informaticae, 160(1–2):143–165, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bashirov:2018:SSB

[1382] Rza Bashirov and Nimet İlke Akçay. Stochastic simulation-based prediction of the behavior of the p16-mediated signaling pathway. Fundamenta Informaticae, 160(1–2):167–179, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rohr:2018:DTL

[1383] Christian Rohr. Discrete-time leap method for stochastic simulation. Fundamenta Informaticae, 160(1–2):181–198, ???? 2018. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Caron:2018:SCM

[1384] Pascal Caron, Jean-Gabriel Luque, and Bruno Patrou. State complexity of multiple catenations. Fundamenta Informaticae, 160(3):255–279, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Elouasbi:2018:DMS

[1385] Samir Elouasbi and Andrzej Pelc. Deterministic meeting of sniffing agents in the plane. Fundamenta Informaticae, 160(3):281–301, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sandhu:2018:NFR

[1386] Jasminder Kaur Sandhu, Anil Kumar Verma, and Prashant Singh Rana. A novel framework for reliable network prediction of small scale wireless sensor networks (SSWSNs). Fundamenta Informaticae, 160(3):303–341, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2018:GAD

[1387] Lin Wang and Zhi Hu. On graph algorithms for degeneracy test and recursive description of stream ciphers. Fundamenta Informaticae, 160 (3):343–359, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carayol:2018:PSI

[1388] Arnaud Carayol, Christof Löding, and Olivier Serre. Pure strategies in imperfect information stochastic games. Fundamenta Informaticae, 160 (4):361–384, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kong:2018:OPA

[1389] Qingzhao Kong and Weihua Xu. Operation properties and algebraic application of covering rough sets. Fundamenta Informaticae, 160(4): 385–408, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ming:2018:CPS

[1390] Yang Ming and Yumin Wang. Certificateless proxy signature scheme in the standard model. Fundamenta Informaticae, 160(4):409–445, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Roayaei:2018:MCP

[1391] Mehdy Roayaei and MohammadReza Razzazi. Minimum cardinality point-to-point connectivity augmentation problem. Fundamenta Informaticae, 160(4):447–463, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fiorentini:2018:P

[1392] Camillo Fiorentini, Alberto Momigliano, and Alberto Pettorossi. Preface. Fundamenta Informaticae, 161(1–2):v–vii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ferrari:2018:CLP

[1393] Mauro Ferrari, Camillo Fiorentini, and Alberto Momigliano. From constructivism to logic programming: an homage to Mario Ornaghi. Fundamenta Informaticae, 161(1–2):1–7, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bergenti:2018:CLP

[1394] Federico Bergenti, Stefania Monica, and Gianfranco Rossi. Constraint logic programming with polynomial constraints over finite domains. Fundamenta Informaticae, 161(1–2):9–27, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bozzato:2018:ABG

[1395] Loris Bozzato. ASP based generation of information terms for constructive ϵL. Fundamenta Informaticae, 161(1–2):29–51, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Calegari:2018:ELP

[1396] Roberta Calegari, Enrico Denti, Agostino Dovier, and Andrea Omicini. Extending logic programming with labelled variables: Model and semantics. Fundamenta Informaticae, 161(1–2):53–74, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chesani:2018:CBP

[1397] Federico Chesani, Paola Mello, Riccardo De Masellis, Chiara Di Francescomarino, Chiara Ghidini, Marco Montali, and Sergio Tessaris. Compliance in business processes with incomplete information and time constraints: a general framework based on abductive reasoning. Fundamenta Informaticae, 161(1–2):75–111, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Delzanno:2018:LBV

[1398] Giorgio Delzanno. Logic-based verification of the distributed dining philosophers protocol. Fundamenta Informaticae, 161(1–2):113–133, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Giordano:2018:DRR

[1399] Laura Giordano and Daniele Theseider Dupré. Defeasible reasoning in P ROEL: from rational entailment to rational closure. Fundamenta Informaticae, 161(1–2):135–161, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Micalizio:2018:ROA

[1400] Roberto Micalizio and Gian Luca Pozzato. Revision of ontologies to accommodate exceptions: a typicality-based approach. Fundamenta Informaticae, 161(1–2):163–189, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sticht:2018:MAD

[1401] Martin Sticht. Multi-agent dialogues and dialogue sequents for proof search and scheduling in intuitionistic logic and the modal logic S4. Fundamenta Informaticae, 161(1–2):191–218, ???? 2018. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Canard:2018:CPK

[1402] Sébastien Canard and Viet Cuong Trinh. Certificateless public key cryptography in the standard model. Fundamenta Informaticae, 161(3):219– 248, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Masood:2018:DEC

[1403] Zaheer Masood, Khalid Majeed, Raza Samar, and Muhammad Asif Zahoor Raja. Design of epidemic computer virus model with effect of quarantine in the presence of immunity. Fundamenta Informaticae, 161(3): 249–273, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalak:2018:BRE

[1404] Marcin Michalak and Dominik Ślȩzak. Boolean representation for exact biclustering. Fundamenta Informaticae, 161(3):275–297, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

S:2018:DIM

[1405] Smrithy G. S, Alfredo Cuzzocrea, and Ramadoss Balakrishnan. Detecting insider malicious activities in cloud collaboration systems. Fundamenta Informaticae, 161(3):299–316, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderAalst:2018:P

[1406] Wil van der Aalst, Eike Best, and Wojciech Penczek. Preface. Fundamenta Informaticae, 161(4):i–ii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Berard:2018:CDO

[1407] Béatrice Bérard, Stefan Haar, Sylvain Schmitz, and Stefan Schwoon. The complexity of diagnosability and opacity verification for Petri nets. Fundamenta Informaticae, 161(4):317–349, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dalsgaard:2018:DFP

[1408] Andreas E. Dalsgaard, Søren Enevoldsen, Peter Fogh, Lasse S. Jensen, Peter G. Jensen, Tobias S. Jepsen, Isabella Kaufmann, Kim G. Larsen, Søren M. Nielsen, Mads Chr. Olesen, Samuel Pastva, and Jiří Srba. A distributed fixed-point algorithm for extended dependency graphs. Fundamenta Informaticae, 161(4):351–381, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hujsa:2018:DLR

[1409] Thomas Hujsa and Raymond Devillers. On deadlockability, liveness and reversibility in subclasses of weighted Petri nets. Fundamenta Informaticae, 161(4):383–421, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolf:2018:IBM

[1410] Karsten Wolf. Interleaving based model checking of concurrency and causality. Fundamenta Informaticae, 161(4):423–445, ???? 2018. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dangalchev:2018:RCG

[1411] Chavdar Dangalchev. Residual closeness of generalized thorn graphs. Fundamenta Informaticae, 162(1):1–15, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dutta:2018:DHL

[1412] Soma Dutta and Piotr Wasilewski. Dialogue in hierarchical learning of concept using prototypes and counterexamples. Fundamenta Informaticae, 162(1):17–36, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garg:2018:NHO

[1413] Manish Garg. A note on the higher-order nonlinearity of Niho function. Fundamenta Informaticae, 162(1):37–42, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jez:2018:NNS

[1414] Artur Jeż and Alexander Okhotin. On the number of nonterminal symbols in unambiguous conjunctive grammars. Fundamenta Informaticae, 162(1):43–72, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sarkar:2018:FSTb

[1415] Apurba Sarkar, Arindam Biswas, Mousumi Dutt, and Shouvick Mondal. Finding shortest triangular path and its family inside a digital object. Fundamenta Informaticae, 162(1):73–100, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Halava:2018:P

[1416] Vesa Halava, Juhani Karhumäki, Yuri Matiyasevich, and Mikhail Volkov. Perface. Fundamenta Informaticae, 162(2–3):i, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Afonin:2018:DPA

[1417] Sergey Afonin. Decision problems and applications of rational sets of regular languages. Fundamenta Informaticae, 162(2–3):101–118, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guterman:2018:PSM

[1418] Alexander E. Guterman and Artem M. Maksaev. Preserving λscrambling matrices. Fundamenta Informaticae, 162(2–3):119–141, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hakanen:2018:MDG

[1419] Anni Hakanen and Tero Laihonen. On ℓ-metric dimensions in graphs. Fundamenta Informaticae, 162(2–3):143–160, ???? 2018. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ko:2018:VAF

[1420] Sang-Ki Ko and Igor Potapov. Vector ambiguity and freeness problems in SL(2, Z). Fundamenta Informaticae, 162(2–3):161–182, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maslennikova:2018:TSC

[1421] Marina Maslennikova and Emanuele Rodaro. Trim strongly connected synchronizing automata and ideal languages. Fundamenta Informaticae, 162(2–3):183–203, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ryzhikov:2018:SSM

[1422] Andrew Ryzhikov and Anton Shemyakov. Subset synchronization in monotonic automata. Fundamenta Informaticae, 162(2–3):205–221, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Saarela:2018:SWE

[1423] Aleksi Saarela. Studying word equations by a method of weighted frequencies. Fundamenta Informaticae, 162(2–3):223–235, ???? 2018. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bellatreche:2018:P

[1424] Ladjel Bellatreche and Zbigniew W. Ras. Preface. Fundamenta Informaticae, 162(4):i–iii, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Savic:2018:GBA

[1425] Miloš Savić, Mirjana Ivanović, Miloš Radovanović, and Bojana Dimić Surla. Gender-based analysis of intra-institutional research productivity and collaboration. Fundamenta Informaticae, 162(4):237–258, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ennaji:2018:ICO

[1426] Fatima Zohra Ennaji, Lobna Azaza, Zakaria Maamar, Abdelaziz El Fazziki, Marinette Savonnet, Mohamed Sadgal, Eric Leclercq, Idir Amine Amarouche, and Djamal Benslimane. Impact of credibility on opinion analysis in social media. Fundamenta Informaticae, 162(4):259–281, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Criado:2018:QAA

[1427] Javier Criado, Silverio Martínez-Fernández, David Ameller, Luis Iribarne, Nicolás Padilla, and Andreas Jedlitschka. Quality-aware architectural model transformations in adaptive mashups user interfaces. Fundamenta Informaticae, 162(4):283–309, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ravat:2018:UAM

[1428] Franck Ravat and Jiefu Song. A unified approach to multisource data analyses. Fundamenta Informaticae, 162(4):311–359, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pan:2019:P

[1429] Linqiang Pan, Mario J. Pérez-Jiménez, and Gexiang Zhang. Preface. Fundamenta Informaticae, 164(2–3):i–ii, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Carandang:2019:HND

[1430] Jym Paul Carandang, Francis George C. Cabarle, Henry Natividad Adorna, Nestine Hope S. Hernandez, and Miguel Ángel Martínez del Amor. Handling non-determinism in spiking neural P systems: Algorithms and simulations. Fundamenta Informaticae, 164(2–3):139–155, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cooper:2019:HCT

[1431] James Cooper and Radu Nicolescu. The Hamiltonian cycle and travelling salesman problems in cP systems. Fundamenta Informaticae, 164(2–3):

157–180, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Cui:2019:CCC

[1432] Guangzhao Cui, Yangyang Jiao, Jianxia Liu, Jixiang Li, Xuncai Zhang, and Zhonghua Sun. Complex complement circuit design of four inputs based on DNA strand displacement. Fundamenta Informaticae, 164(2– 3):181–194, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Guo:2019:DID

[1433] Bingjie Guo, Luhui Wang, Tao Wu, and Yafei Dong. DNA ion detector and logic circulation amplification model based on mercury and silver ions. Fundamenta Informaticae, 164(2–3):195–205, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2019:CPC

[1434] Suxia Jiang, Yanfeng Wang, Jinbang Xu, and Fei Xu. The computational power of cell-like P systems with symport/antiport rules and promoters. Fundamenta Informaticae, 164(2–3):207–225, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Choi:2019:ADE

[1435] Tae Jong Choi, Jong-Hyun Lee, Hee Yong Youn, and Chang Wook Ahn. Adaptive differential evolution with elite opposition-based learning and its application to training artificial neural networks. Fundamenta Informaticae, 164(2–3):227–242, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niu:2019:CSM

[1436] Ying Niu, Feng Han, Xuncai Zhang, and Zheng Zhou. A circuit simplification mechanism based on DNA combinatorial strands displacement. Fundamenta Informaticae, 164(2–3):243–257, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sun:2019:NCS

[1437] Junwei Sun, Nan Li, Yanfeng Wang, and Wei Wang. A novel chaotic system and its modified compound synchronization. Fundamenta Informaticae, 164(2–3):259–275, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xie:2019:LOM

[1438] Wendan Xie, Changjun Zhou, Hui Lv, and Qiang Zhang. Logic operation model of the complementer based on two-domain DNA strand displacement. Fundamenta Informaticae, 164(2–3):277–288, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Brunetti:2018:P

[1439] Sara Brunetti, Paolo Dulio, Andrea Frosini, and Grzegorz Rozenberg. Preface. Fundamenta Informaticae, 163(1):v–xiv, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sarkissian:2018:DCV

[1440] Henri Der Sarkissian, Nicola Viganò, and Kees Joost Batenburg. A data consistent variational segmentation approach suitable for real-time tomography. Fundamenta Informaticae, 163(1):1–20, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Duchi:2018:GTP

[1441] Enrica Duchi, Veronica Guerrini, and Simone Rinaldi. A generating tree for permutations avoiding the pattern 122^{+}3. Fundamenta Informaticae, 163(1):21–39, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Feschet:2018:SRC

[1442] Fabien Feschet and Jean-Jacques Lemaire. Super-resolution in clinical conditions: Deep brain stimulation case study. Fundamenta Informaticae, 163(1):41–62, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mancini:2018:NMN

[1443] Matteo Mancini and Mara Cercignani. Network models in neuroimaging: A survey of multimodal applications. Fundamenta Informaticae, 163(1): 63–91, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dulio:2018:GMS

[1444] Paolo Dulio, Paolo Finotelli, Andrea Frosini, Elisa Pergola, and Alice Presenti. Graph model simulation of human Brain’s functional activity at resting state by means of the FD model. Fundamenta Informaticae, 163(1):93–109, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Munir:2018:IRM

[1445] Rana Faisal Munir, Sergi Nadal, Oscar Romero, Alberto Abelló, Petar Jovanovic, Maik Thiele, and Wolfgang Lehner. Intermediate results materialization selection and format for data-intensive flows. Fundamenta Informaticae, 163(2):111–138, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aguzzoli:2018:FIA

[1446] Stefano Aguzzoli, Stefania Boffa, Davide Ciucci, and Brunella Gerla. Finite IUML-algebras, finite forests and orthopairs. Fundamenta Informaticae, 163(2):139–163, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michaliszyn:2018:SVF

[1447] Jakub Michaliszyn, Jan Otop, and Piotr Witkowski. Satisfiability versus finite satisfiability in elementary modal logics. Fundamenta Informaticae, 163(2):165–188, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Schiopu:2018:MFP

[1448] Camelia Schiopu and Eleonor Ciurea. Maximum flows in planar dynamic networks with lower bounds. Fundamenta Informaticae, 163(2):189–204, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tong:2018:SBP

[1449] Lyuyang Tong, Minggang Dong, Bing Ai, and Chao Jing. A simple butterfly particle swarm optimization algorithm with the fitness-based adaptive inertia weight and the opposition-based learning average elite strategy. Fundamenta Informaticae, 163(2):205–223, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Charalampopoulos:2018:P

[1450] Panagiotis Charalampopoulos, Maxime Crochemore, and Solon P. Pissis. Preface. Fundamenta Informaticae, 163(3):i, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grabowski:2018:ALC

[1451] Szymon Grabowski, Tomasz Kociumaka, and Jakub Radoszewski. On Abelian longest common factor with and without RLE. Fundamenta Informaticae, 163(3):225–244, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ganguly:2018:LSD

[1452] Arnab Ganguly, Manish Patil, Rahul Shah, and Sharma V. Thankachan. A linear space data structure for range LCP queries. Fundamenta Informaticae, 163(3):245–251, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alzamel:2018:ECP

[1453] Mai Alzamel, Jia Gao, Costas S. Iliopoulos, and Chang Liu. Efficient computation of palindromes in sequences with uncertainties. Fundamenta Informaticae, 163(3):253–266, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hooshmand:2018:CAC

[1454] Sahar Hooshmand, Neda Tavakoli, Paniz Abedin, and Sharma V. Thankachan. On computing average common substring over run length encoded sequences. Fundamenta Informaticae, 163(3):267–273, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mhaskar:2018:FCS

[1455] Neerja Mhaskar and W. F. Smyth. Frequency covers for strings. Fundamenta Informaticae, 163(3):275–289, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Chairungsee:2018:EAC

[1456] Supaporn Chairungsee. Efficient approaches to compute longest previous non-overlapping factor array. Fundamenta Informaticae, 163(3):291–304, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Henry:2018:DTN

[1457] Christopher J. Henry and Syed Aqeel Awais. A descriptive tolerance nearness measure for performing graph comparison. Fundamenta Informaticae, 163(4):305–324, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kheirfam:2018:FNS

[1458] Behrouz Kheirfam and Afsaneh Nasrollahi. A full-Newton step interiorpoint method based on a class of specific algebra transformation. Fundamenta Informaticae, 163(4):325–337, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Manuel:2018:GTG

[1459] Paul Manuel and Sandi Klavžar. The graph theory general position problem on some interconnection networks. Fundamenta Informaticae, 163(4):339–350, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Mishra:2018:TOU

[1460] Sumit Mishra, Samrat Mondal, and Sriparna Saha. Towards obtaining upper bound on sensitivity computation process for cluster validity measures. Fundamenta Informaticae, 163(4):351–374, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shaw:2018:AAT

[1461] Dipan Lal Shaw, A. S. M. Shohidull Islam, Shuvasish Karmaker, and M. Sohel Rahman. Approximation algorithms for three dimensional protein folding. Fundamenta Informaticae, 163(4):375–393, ???? 2018. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tian:2018:ABD

[1462] Miaomiao Tian, Lingyan Wang, Hong Zhong, and Jie Chen. Attributebased data integrity checking for cloud storage. Fundamenta Informaticae, 163(4):395–411, ???? 2018. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bes:2019:CDF

[1463] Alexis Bès and Christian Choffrut. Complexity and (un)decidability of fragments of ⟨ωωλ; ×⟩. Fundamenta Informaticae, 164(1):1–15, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drazic:2019:AMD

[1464] Slobodan Dražić. Advanced morphological distances based on dilation and erosion. Fundamenta Informaticae, 164(1):17–39, ???? 2019. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fajardo:2019:EMC

[1465] William Fajardo and Oswaldo Lezama. Elementary matrix-computational proof of Quillen–Suslin theorem for Ore extensions. Fundamenta Informaticae, 164(1):41–59, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Garg:2019:CIV

[1466] Harish Garg and Dimple Rani. Complex interval-valued intuitionistic fuzzy sets and their aggregation operators. Fundamenta Informaticae, 164(1):61–101, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maji:2019:RFC

[1467] Pradipta Maji and Suman Mahapatra. Rough-fuzzy circular clustering for color normalization of histological images. Fundamenta Informaticae, 164(1):103–117, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sadhu:2019:TCC

[1468] Sanjib Sadhu, Sasanka Roy, Soumen Nandi, Anil Maheshwari, and Subhas C. Nandy. Two-center of the convex hull of a point set: Dynamic model, and restricted streaming model. Fundamenta Informaticae, 164 (1):119–138, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Szymanik:2019:I

[1469] Jakub Szymanik. Introduction. Fundamenta Informaticae, 164(4):i–iii, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Czarnecki:2019:MLT

[1470] Marek Czarnecki and Konrad Zdanowski. A modal logic of a truth definition for finite models. Fundamenta Informaticae, 164(4):299–325, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grudzinska:2019:CSM

[1471] Justyna Grudzińska and Marek Zawadowski. Continuation semantics for multi-quantifier sentences: Operation-based approaches. Fundamenta Informaticae, 164(4):327–344, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kalocinski:2019:SRL

[1472] Dariusz Kalociński. Some remarks on least moduli. Fundamenta Informaticae, 164(4):345–358, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zawadowski:2019:CSA

[1473] Marek Zawadowski. co-semi-analytic functors. Fundamenta Informaticae, 164(4):359–373, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zdanowski:2019:OHQ

[1474] Konrad Zdanowski. One Henkin quantifier in the empty vocabulary suffices for undecidability. Fundamenta Informaticae, 164(4):375–386, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ghassemi:2019:RRP

[1475] Fatemeh Ghassemi and Wan Fokkink. Reliable restricted process theory. Fundamenta Informaticae, 165(1):1–41, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kovacs:2019:AUC

[1476] László Kovács. An algorithm using context reduction for efficient incremental generation of concept set. Fundamenta Informaticae, 165(1): 43–73, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Shaheen:2019:GPG

[1477] Tanzeela Shaheen and John G. Stell. Graphical partitions and graphical relations. Fundamenta Informaticae, 165(1):75–98, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Peng:2019:APF

[1478] Xindong Peng. Algorithm for Pythagorean fuzzy multi-criteria decision making based on WDBA with new score function. Fundamenta Informaticae, 165(2):99–137, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Xu:2019:UCS

[1479] Wen-Li Xu, Jia Lee, Hui-Hui Chen, and Teijiro Isokawa. Universal computation in a simplified Brownian cellular automaton with von Neumann neighborhood. Fundamenta Informaticae, 165(2):139–156, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yang:2019:TWD

[1480] Bin Yang, Bao Qing Hu, and Junsheng Qiao. Three-way decisions with rough membership functions in covering approximation space. Fundamenta Informaticae, 165(2):157–191, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Penczek:2019:P

[1481] Wojciech Penczek, Holger Schlingloff, and Piotr Wasilewski. Preface. Fundamenta Informaticae, 165(3–4):i–iii, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitting:2019:WJL

[1482] Melvin Fitting. What are justification logics? Fundamenta Informaticae, 165(3–4):193–203, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

DeAngelis:2019:SCT

[1483] Emanuele De Angelis, Fabio Fioravanti, Maria Chiara Meo, Alberto Pettorossi, and Maurizio Proietti. Semantics and controllability of timeaware business processes. Fundamenta Informaticae, 165(3–4):205–244, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Aldilaijan:2019:EST

[1484] Abdulla Aldilaijan, Mohammad Azad, and Mikhail Moshkov. Experimental study of totally optimal decision trees. Fundamenta Informaticae, 165(3–4):245–261, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazan:2019:CBD

[1485] Jan G. Bazan, Adam Szczur, Andrzej Skowron, Marian Rzepko, Paweł Król, Wojciech Bajorek, and Wojciech Czarny. A classifier based on a decision tree with temporal cuts. Fundamenta Informaticae, 165(3–4): 263–281, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dutta:2019:LRS

[1486] Soma Dutta, Andrzej Jankowski, Grzegorz Rozenberg, and Andrzej Skowron. Linking reaction systems with rough sets. Fundamenta Informaticae, 165(3–4):283–302, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gori:2019:SOR

[1487] Roberta Gori, Damas Gruska, and Paolo Milazzo. Studying opacity of reaction systems through formula based predictors. Fundamenta Informaticae, 165(3–4):303–319, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Niewiadomski:2019:AMS

[1488] Artur Niewiadomski, Piotr Switalski, Teofil Sidoruk, and Wojciech Penczek. Applying modern SAT-solvers to solving hard problems. Fundamenta Informaticae, 165(3–4):321–344, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sawicka:2019:TET

[1489] Anna Sawicka, Magdalena Kacprzak, and Andrzej Zbrzezny. Towards encoding of the transition relation in dialogue games model checking. Fundamenta Informaticae, 165(3–4):345–361, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolski:2019:DPS

[1490] Marcin Wolski and Anna Gomolińska. From data to pattern structures: Near set approach. Fundamenta Informaticae, 165(3–4):363–380, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alsolami:2019:CHO

[1491] Fawaz Alsolami, Talha Amin, Mikhail Moshkov, Beata Zielosko, and Krzysztof Żabiński. Comparison of heuristics for optimization of association rules. Fundamenta Informaticae, 166(1):1–14, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Borzyszkowski:2019:ROP

[1492] Andrzej M. Borzyszkowski. Regular orthomodular posets. Fundamenta Informaticae, 166(1):15–28, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hartonas:2019:RLM

[1493] Chrysafis Hartonas and Ewa Orłowska. Representation of lattices with modal operators in two-sorted frames. Fundamenta Informaticae, 166 (1):29–56, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wei:2019:PFH

[1494] Gui-Wu Wei. Pythagorean fuzzy hamacher power aggregation operators in multiple attribute decision making. Fundamenta Informaticae, 166(1): 57–85, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arqub:2019:ARP

[1495] Omar Abu Arqub. Application of residual power series method for the solution of time-fractional Schrödinger equations in one-dimensional space. Fundamenta Informaticae, 166(2):87–110, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Arqub:2019:NAS

[1496] Omar Abu Arqub. Numerical algorithm for the solutions of fractional order systems of Dirichlet function types with comparative analysis. Fundamenta Informaticae, 166(2):111–137, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bistarelli:2019:SIK

[1497] Stefano Bistarelli, Francesco Santini, and Paolo Torroni. Special issue of the knowledge representation and reasoning track, Symposium on Applied Computing (SAC) 2017. Fundamenta Informaticae, 166(2):139– 140, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fitzgerald:2019:CSI

[1498] Tadhg Fitzgerald and Barry O’Sullivan. Candidate selection and instance ordering for realtime algorithm configuration. Fundamenta Informaticae, 166(2):141–166, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koutras:2019:RDC

[1499] Costas D. Koutras, Christos Moyzes, and Christos Rantsoudis. A reconstruction of default conditionals within epistemic logic. Fundamenta Informaticae, 166(2):167–197, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bisht:2019:WMR

[1500] Raj Kishor Bisht, Taishin Yasunobu Nishida, and Kouhei Yamamoto. Word matrix rewriting systems. Fundamenta Informaticae, 166(3):199– 226, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Polkowski:2019:IMB

[1501] Lech Polkowski. Introducing mass-based rough mereology in a mereological universe with relations to fuzzy logics and a generalization of the Łukasiewicz logical foundations of probability. Fundamenta Informaticae, 166(3):227–249, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tiwari:2019:APR

[1502] Surabhi Tiwari and Pankaj Kumar Singh. An approach of proximity in rough set theory. Fundamenta Informaticae, 166(3):251–271, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wang:2019:UMB

[1503] Zhaohao Wang, Huifang Yue, and Jianping Deng. An uncertainty measure based on lower and upper approximations for generalized rough set models. Fundamenta Informaticae, 166(3):273–296, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pombo:2019:SCA

[1504] Carlos G. Lopez Pombo, Pablo F. Castro, Nazareno M. Aguirre, and Thomas S. E. Maibaum. Satisfiability calculus: An abstract formulation of semantic proof systems. Fundamenta Informaticae, 166(4):297–347, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rauch:2019:DMH

[1505] Jan Rauch and Milan Šimunek. Data mining with histograms and domain knowledge — case studies and considerations. Fundamenta Informaticae, 166(4):349–378, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Turaci:2019:CCR

[1506] Tufan Turacı and Aysun Aytaç. Combining the concepts of residual and domination in graphs. Fundamenta Informaticae, 166(4):379–392, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Maratea:2019:P

[1507] Marco Maratea, Ivan Serina, and Paolo Torroni. Preface. Fundamenta Informaticae, 167(1–2):v–vii, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alviano:2019:ARC

[1508] Mario Alviano. Argumentation reasoning via circumscription with Pyglaf. Fundamenta Informaticae, 167(1–2):1–30, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Alviano:2019:MEA

[1509] Mario Alviano and Carmine Dodaro. Model enumeration via assumption literals. Fundamenta Informaticae, 167(1–2):31–58, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Charwat:2019:EBQ

[1510] Günther Charwat and Stefan Woltran. Expansion-based QBF solving on tree decompositions. Fundamenta Informaticae, 167(1–2):59–92, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Gonzalez:2019:EAS

[1511] Miguel A. González, Angelo Oddi, and Riccardo Rasconi. Efficient approaches for solving a multiobjective energy-aware job shop scheduling problem. Fundamenta Informaticae, 167(1–2):93–132, ???? 2019. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Santucci:2019:TPB

[1512] Valentino Santucci, Marco Baioletti, and Alfredo Milani. Tackling permutation-based optimization problems with an algebraic particle swarm optimization algorithm. Fundamenta Informaticae, 167(1–2):133– 158, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Fajardo:2019:CML

[1513] William Fajardo. A computational Maple library for skew PBW extensions. Fundamenta Informaticae, 167(3):159–191, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Michalak:2019:BRC

[1514] Marcin Michalak and Dominik Ślȩzak. On Boolean representation of continuous data biclustering. Fundamenta Informaticae, 167(3):193–217, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Poklukar:2019:NEC

[1515] Darja Rupnik Poklukar and Janez Žerovnik. Networks with extremal closeness. Fundamenta Informaticae, 167(3):219–234, ???? 2019. CO-DEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wasilewski:2019:ADS

[1516] Piotr Wasilewski. Algebras of definable sets vs. concept lattices. Fundamenta Informaticae, 167(3):235–256, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bazhenov:2019:CCA

[1517] Nikolay Bazhenov. Computable contact algebras. Fundamenta Informaticae, 167(4):257–269, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Rudi:2019:AHO

[1518] Ali Gholami Rudi. Approximate hotspots of orthogonal trajectories. Fundamenta Informaticae, 167(4):271–285, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Grahne:2019:UEN

[1519] Gösta Grahne and Ali Moallemi. Universal (and existential) nulls. Fundamenta Informaticae, 167(4):287–321, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klaudel:2019:BAI

[1520] Hanna Klaudel, Maciej Koutny, Zhenhua Duan, and Ben Moszkowski. From Box algebra to interval temporal logic. Fundamenta Informaticae, 167(4):323–354, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jiang:2019:CPS

[1521] Suxia Jiang, Yanfeng Wang, Fei Xu, and Junli Deng. Communication P systems with channel states working in flat maximally parallel manner. Fundamenta Informaticae, 168(1):1–24, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Klopotek:2019:EKF

[1522] Mieczysław A. Kłopotek. On the existence of kernel function for kerneltrick of k-means in the light of Gower Theorem. Fundamenta Informaticae, 168(1):25–43, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Koga:2019:DRL

[1523] Toshihiro Koga. On the density of regular languages. Fundamenta Informaticae, 168(1):45–49, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Marquer:2019:ACI

[1524] Yoann Marquer. Algorithmic completeness of imperative programming languages. Fundamenta Informaticae, 168(1):51–77, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Urban:2019:NDC

[1525] Roman Urban and Hubert Anisimowicz. A note on the Durda, Caron, and Buchanan word ambiguity detection algorithm. Fundamenta Informaticae, 168(1):79–88, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:2019:F

[1526] Wladyslaw Skarbek and Yu-Dong Zhang. Foreword. Fundamenta Informaticae, 168(2–4):i–ii, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Skarbek:2019:STN

[1527] Władysław Skarbek. Symbolic tensor neural networks for digital media — from tensor processing via BNF graph rules to CREAMS applications. Fundamenta Informaticae, 168(2–4):89–184, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Biedrzycki:2019:DLO

[1528] Rafał Biedrzycki, Paweł Zawistowski, and Bartłomiej Twardowski. Deep learning optimization tasks and metaheuristic methods. Fundamenta Informaticae, 168(2–4):185–218, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Yu:2019:ADM

[1529] Xiang Yu and Shui-Hua Wang. Abnormality diagnosis in mammograms by transfer learning based on ResNet18. Fundamenta Informaticae, 168 (2–4):219–230, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hong:2019:ICM

[1530] Jin Hong, Hong Cheng, Shui-Hua Wang, and Jie Liu. Improvement of cerebral microbleeds detection based on discriminative feature learning. Fundamenta Informaticae, 168(2–4):231–248, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurowski:2019:MCB

[1531] Adam Kurowski, Katarzyna Mrozik, Bozena Kostek, and Andrzej Czyzewski. Method for clustering of brain activity data derived from EEG signals. Fundamenta Informaticae, 168(2–4):249–268, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Tautkute:2019:CVE

[1532] Ivona Tautkutė and Tomasz Trzciński. Classifying and visualizing emotions with emotional DAN. Fundamenta Informaticae, 168(2–4):269–285, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pilarczyk:2019:HFE

[1533] Rafał Pilarczyk, Xin Chang, and Władysław Skarbek. Human face expressions from images. Fundamenta Informaticae, 168(2–4):287–310, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Pesko:2019:CTV

[1534] Maciej Pȩśko, Adam Svystun, Paweł Andruszkiewicz, Przemysław Rokita, and Tomasz Trzciński. Comixify: Transform video into comics. Fundamenta Informaticae, 168(2–4):311–333, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Sharma:2019:NTS

[1535] Shailza Sharma, Vivek Singh Bawa, and Vinay Kumar. A novel twostage residual learning based convolutional neural network for image super resolution. Fundamenta Informaticae, 168(2–4):335–351, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Khomenko:2019:P

[1536] Victor Khomenko, Jetty Kleijn, Wojciech Penczek, and Olivier H. Roux. Preface. Fundamenta Informaticae, 169(1–2):v–vi, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Devillers:2019:ASW

[1537] Raymond Devillers and Thomas Hujsa. Analysis and synthesis of weighted marked graph Petri nets: Exact and approximate methods. Fundamenta Informaticae, 169(1–2):1–30, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Janicki:2019:OSI

[1538] Ryszard Janicki and Maciej Koutny. Operational semantics, interval orders and sequences of antichains. Fundamenta Informaticae, 169(1–2): 31–55, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Wolf:2019:SVP

[1539] Karsten Wolf. Synthesis for various Petri net classes with Union/Find. Fundamenta Informaticae, 169(1–2):57–84, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Valk:2019:FPP

[1540] Rüdiger Valk. Formal properties of Petri’s cycloid systems. Fundamenta Informaticae, 169(1–2):85–121, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Jancar:2019:CFC

[1541] Petr Jančar, Jérôme Leroux, and Grégoire Sutre. Co-finiteness and coemptiness of reachability sets in vector addition systems with states. Fundamenta Informaticae, 169(1–2):123–150, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

vanderAalst:2019:LPM

[1542] Wil M. P. van der Aalst. Lucent process models and translucent event logs. Fundamenta Informaticae, 169(1–2):151–177, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Feuillade:2019:DLA

[1543] Guillaume Feuillade, Andreas Herzig, and Christos Rantsoudis. A dynamic logic account of active integrity constraints. Fundamenta Informaticae, 169(3):179–210, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kavand:2019:GCS

[1544] P. Kavand and A. Mohades. Geodesic center of a simple polygon using a logarithmic number of extra variables. Fundamenta Informaticae, 169 (3):211–235, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zielinski:2019:QLB

[1545] Bartosz Zieliński. A query language based on term matching and rewriting. Fundamenta Informaticae, 169(3):237–274, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kula:2019:P

[1546] Mieczysław Kula, Damian Niwiński, and Jacek Pomykała. Preface. Fundamenta Informaticae, 169(4):i, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drylo:2019:JHC

[1547] Robert Dryło and Jacek Pomykała. Jacobians of hyperelliptic curves over Z_{n} and factorization of n. Fundamenta Informaticae, 169(4):275– 283, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Drylo:2019:DFR

[1548] Robert Dryło, Tomasz Kijko, and Michał Wroński. Determining formulas related to point compression on alternative models of elliptic curves. Fundamenta Informaticae, 169(4):285–294, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hanzlik:2019:CPC

[1549] Lucjan Hanzlik, Kamil Kluczniak, and Mirosław Kutyłowski. CTRL-PACE: Controlled randomness for e-passport password authentication. Fundamenta Informaticae, 169(4):295–330, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Morawiecki:2019:MS

[1550] Paweł Morawiecki. Malicious SHA-3. Fundamenta Informaticae, 169 (4):331–343, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Zajac:2019:HEM

[1551] Pavol Zajac. Hybrid encryption from McEliece cryptosystem with pseudo-random error vector. Fundamenta Informaticae, 169(4):345–360, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Altenkirch:2019:P

[1552] Thorsten Altenkirch and Aleksy Schubert. Preface. Fundamenta Informaticae, 170(1–3):v–vi, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bezem:2019:STC

[1553] Marc Bezem and Thierry Coquand. Skolem’s Theorem in coherent logic. Fundamenta Informaticae, 170(1–3):1–14, ???? 2019. CODEN FU-MAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Bono:2019:MLT

[1554] Viviana Bono. The Magda language: Ten years after. Fundamenta Informaticae, 170(1–3):15–38, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Downen:2019:DCI

[1555] Paul Downen, Zena M. Ariola, and Silvia Ghilezan. The duality of classical intersection and union types. Fundamenta Informaticae, 170(1–3): 39–92, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Dudenhefner:2019:UIT

[1556] Andrej Dudenhefner and Jakob Rehof. Undecidability of intersection type inhabitation at rank 3 and its formalization. Fundamenta Informaticae, 170(1–3):93–110, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Santo:2019:DSC

[1557] José Espírito Santo, Ralph Matthes, and Luís Pinto. Decidability of several concepts of finiteness for simple types. Fundamenta Informaticae, 170(1–3):111–138, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Geuvers:2019:SNT

[1558] Herman Geuvers, Iris van der Giessen, and Tonny Hurkens. Strong normalization for truth table natural deduction. Fundamenta Informaticae, 170(1–3):139–176, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Hupel:2019:CDC

[1559] Lars Hupel. Certifying dictionary construction in Isabelle/HOL. Fundamenta Informaticae, 170(1–3):177–205, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kfoury:2019:PRR

[1560] Assaf Kfoury. Personal reflections on the role of mathematical logic in computer science. Fundamenta Informaticae, 170(1–3):207–221, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Kurata:2019:NLM

[1561] Toshihiko Kurata and Ken etsu Fujita. Neighbourhood and lattice models of second-order intuitionistic propositional logic. Fundamenta Informaticae, 170(1–3):223–240, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Darkey-Mensah:2019:IFI

[1562] Mawunyo Kofi Darkey-Mensah and Przemysław Koprowski. Intrinsic factorization of ideals in Dedekind domains. Fundamenta Informaticae, 170(4):325–338, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Ketema:2019:CIT

[1563] Jeroen Ketema and Jakob Grue Simonsen. Computing with infinite terms and infinite reductions. Fundamenta Informaticae, 170(4):339– 365, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).

Loukanova:2019:GRT

[1564] Roussanka Loukanova. Gamma-reduction in type theory of acyclic recursion. Fundamenta Informaticae, 170(4):367–411, ???? 2019. CODEN FUMAAJ. ISSN 0169-2968 (print), 1875-8681 (electronic).